A Complete Bibliography of Publications in Technometrics for the decade 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.02

Title word cross-reference

(≥ 2) [Sin92]. 2^{k−p} [BH97, Bis94]. 2^{k−p} × s [BS97]. 2^{n} [SWC97]. 2^{n−k} [SCF97]. 2^{n−p} [SWC97]. 3 [Ber98]. ^{∗} [DL90a]. C_{P} [Mal95, vdVtBM97]. D [AB97, DJ94, Ney94]. D_{s} [AB97]. F [Ano00, SBD98]. k [Ode90, Sin92]. k(≥ 2) [Sin83]. L_{p} [Spo90]. M [Gri97b, Ode90]. N ≠ 2^{k} [LIV96]. P [GK90, ME99, Cha93, Lin93a, Lin95a, ZM94]. ∏ [Bre91a, Bre91b, Fri91, Gu91, HT91]. X [WJ92]. Y [WJ92].

- [AB97]. -D [Ber98]. -efficient [Lin95a, Lin93a]. -Estimators [Gri97b]. -Norm [Spo90]. -Optimal [DJ94]. -Optimality-Based [Ney94]. -Optimum [AB97]. -Spline [ME99]. -Value [Cha93].

1 [Fam95, Zie98o]. 100 [Zie94a]. 101 [Zie99a]. 1750 [Zie92a]. 1930 [Zie99b]. 1984 [Zie91f, Zie91e]. 1988 [Zie90d].

2 [Ber96b, Sch93a, Zie98p, Zie99-44]. 20th [Bro90b]. 25 [Sin92]. 29 [Ste90b].

3 [Fox91, Kus93, Sar93, Zie99-45]. 31 [BT98]. 34 [HR94]. 35 [DPL^{+}94, Lin95a].

40 [Sen99b].

5 [Zie90f, Zie99n]. 5.0 [Zie96m].

84g [Sin92]. 88e [Ste90b].

9000 [Zie93n]. 90f [BT98]. ’92 [Zie95r]. 94c [DPL^{+}94]. ’96 [Zie98u]. 97 [Zie99k]. 99h [Sen99b].

Aage [Kel97]. Abarbanel [Sti97]. Aberration [BS99, HCV98, SCF97]. Abraham [Zie98-35]. Abrupt [Rie94a, Yas95]. Abuse [Zie90z]. Accelerated [BE94, DH92, EM95, ML95, MEL98, MM94, Sch91c, Spe91, Mee91b, Sam90]. Acceptance [HL95, MT92, Tan93, Wes90]. Acceptance-Sampling [MT92]. Accompany [Gun90]. Accountants [Zie95-40]. Accumulation [Hab90, HW90b, HW90a, Hir90, Nai90b, YDB90]. Accuracy [CB93, ES94, OUH93]. Accurate [AKO94]. Achieving [Dun96]. Action [BL94]. Active [VS96, VS98]. Adams [Zie95-50]. Adaptive [HP97, Har92a, ST90a, Zha94, Har91d, Por97b]. Added [Coo96a]. Added-Variable [Coo96a]. Additional [ES94, Nel93]. Additive [dV92]. Adelchi [DeV99, Lei99]. Adequacy [Zie94c]. Adhikari [Loh94]. Adjusted [Ano91f, Haw91b]. Adjustment [BK92a, BK92b, BL94, JV93, Luc98, LG99, Mac92, Tuc92, VV92, WG92, Cha93, Fal99]. Administrators [Pro93b]. Adrian [DeV99]. Adrienne [Bru94]. Advanced [Ada96, Dea95, Tho90a, Zie94d]. Advances [Wol96, Zie96-34, Tor99, Zie94e, Zie96a, Zie98c, Zie98d, Zie98e]. Advection [MS95]. Advection-Dispersion [MS95]. Adventures [Wen93]. Afifi [Zie92e, Zie97h]. Aft [Rol93, Zie99o]. after [GS92]. Aggarwal [Kha95]. Agresti [Hap91]. Ahlbom [Zie95i]. AI [Zie95c]. Aided [And90, LaL95, WC92, Zie98f, Zie92e, Zie97h]. Air [GK93]. Aitken [Zie96-42]. Aitkin [Moo91]. Akaike [Zie99-36, Zie99-36]. Alain [Cab98]. Alan [Hap91, Lee90, Sea94, Win93, Zie98r]. Albert [Eas90, Goo92a, Zie98-34, Zie98i]. Alberto [Fal99, Job99, Lei91]. Albright [Zie99z]. Alessandro [Dog99]. Alexander [Goo95, Jac95, Sha97a, Weh91]. Algebra [Zie98-30, Zie99w]. Algorithm [Kus98, MN95, QY98, RvD99]. Algorithmic [Mad96, Ngu96a, TFV93, VTFD92]. Algorithms [LW97, Was97, Jon92, IHMT95, Sar90, Wol91]. Aliasing [CHW97]. All-or-None [VV94]. Allan [Mul97, Nan94, Sar94a]. Alldredge [Zie94m]. Allen [Gra90, Bax92]. Allison [Zie97-38]. Allocation [ES94, Yin91]. Alreck [Zie99-52]. Alsultan [Zie99y]. Alsup [McL95]. Alternative [Fai94]. Alvin [Boo98, Mat96]. American [Lew92, Zie99-33]. Amitava [Bem96]. Amount [McA90]. Analyses [Bar92a, Ko95a, Sch91c, VV99, Dri92, Mee91b, Zie97y]. Analysing [Zie96b]. Analysis [ACV99, Alb93, Alt99, Ano99e, Bar95, Bas97, Bax96b, BNC87, BS97, Bla94, Bow91, Boy90, Boy96, Bro92, Bry92, Cab94, Cal94, Cam90, Car97, CHB90, Che94, CH96a, Chi91, Chu91, Cob91, Cox93, Cox96, CB94, Dea91, Deh90, Dic91, DS93, Dre91, Elt90, Emp99, Eub90, Fai91, Far97, FL97, Fot96, Fox98, FL95, Fre95, Gle91, Goo95, Got92, Gra98, Gre98, GC91, Gru90, Hab90, Hag96, HW90b, HW90a, HW91, HS93, Han98, Hap92, Har92a, Has91, Hec96, Hir90, Hir96, HS97, Jac95, Jol90, KLR91, Kia92b, Kia93b, Kii92, KTTS90, LN95, Law98, Lee99, Lin92, Lwi94, Lyn97b, Lyn99, MW90, Maz93, McA90, MEL98, MME98, Mil90, MMY93, MS95, Nai90b, NW93, Nel93, OUH93, Pap95, Per92, Per98, Pet93, Pet99, Pri98, Pro92a, Rao98]. Analysis [Ray93, Rig93, STC99, Sau98, Sch95, Sch96b, Sch93c, Sch93b, Sch91d, Sco91, Sha97c, Sim91, Sri92, Ste91, SP95, Taa91, Tam96, Tan98, Tom95, Vin92, WL97, Wat96, Weh95, Wel90, WN97, YDB90, Yas92, ZP94, Zie90-27, Zie90v, Zie90q, Zie90f, Zie90b, Zie91q, Zie91h, Zie91o, Zie92f, Zie92q, Zie93h, Zie93g, Zie93f, Zie94o, Zie94-29, Zie94p, Zie95y, Zie95j, Zie95-46, Zie95-50, Zie95-36, Zie96m, Zie96t, Zie96i, Zie96-29, Zie96-39, Zie96-38, Zie96l, Zie97-38, Zie97x, Zie97p, Zie97-36, Zie97c, Zie97j, Zie97z, Zie97-30, Zie97-40, Zie98-44, Zie98g, Zie98r, Zie98b, Zie99-38, Zie99-31, Zie99i, Zie99d, Zie99x, vN90, Bay99, Ber92a, Boo97, DeV99, Dea94, Dov93, Geo94, Geo98, Gol99, Goo92a, Gra90, Gro99, Gun99, Had92, Ham96, Han90, Han97, Hap91, Hec95, Hei90, Job99]. Analysis [Jol93, Koh94, LaL95, Lin90, Liu92b, Mat96, Mat97, Mat98, Nai94, Nan94, Not92b, Por93, Pri93, Ruk97a, Sam99b, Sau94, Sch96a, Sch96c, Sen94b, Ter92, Tim95, Ver90, Web90, Win90, Wol96, Zel94, Zha92, Zie90o, Zie90c, Zie92k, Zie92e, Zie93-27, Zie94u, Zie94f, Zie94x, Zie94-36, Zie95-48, Zie95v, Zie95l, Zie96-50, Zie96a, Zie96c, Zie97h, Zie97r, Zie97l, Zie97f, Zie97-36, Zie97-33, Zie97t, Zie97-37, Zie98-39, Zie98m, Zie99-34, Zie99c, Zie99-35, Zin93, Car95, Cra94, Kia93a, Len93, Sti97, Tsa91, Zie91q, Zie95d, vT91]. Analysis-of-Means-Type [WN97]. AnalysisPart [Sha95]. Analysts [LeP90b]. Analytic [BR95]. Analytical [Hil95, LaL95, RL95, Zie93l]. Analyzing [CHW97, Pat91, Gro96]. and/or [CR91]. Anders [Kel97, Zie92a, Zie99b]. Andersen [Kle94, Sau98, SG99]. Anderson [Moo91, Gre94, Hap92, Lee90, Zie94-36]. Andor [Sea91]. Andre [Zie94d, Zie97y, Zie98t]. Andreas [Zie98-46]. Andrej [KB94]. Andrew [Ang92, Cox96, Ens95, Han98, Lar95, Pri93, Zie97-33]. Andrzej [Mor93b]. Ani [Loh94]. Anil [Mat98, Sar90]. Ann [Soy95, Zie99r]. Anne [Bor97b, Gun90]. Annealing [Deu96]. Annie [Bor97b]. Announcement [Mor98b]. ANOVA [BT98, BT89, Van92b, Nai96, Zie93b]. Ansell [Han97]. Anthony [Cal94, LeP90a]. Antoniadis [Zie97-43]. Anxiety [Cly94]. Any [Zie99a]. Anyone [God90]. Aoki [Zie98f, Zie92p]. APC [HP92].

Application [Ano98e, BCL98a, BCL98b, CR98, DP97, ET96, Jen92, Kaf94, Lam92, LT96, Oma91, Rob98, Sch98a, SP95, TH93, VTFD92, Bor98, Der90, Rie94a, Zie98v]. Applications [Ber96a, Boo90, Bow91, Bru93a, Cal94, CR96, Cox96, CC95, Dea97, Gla98, Hag90, HR92, Hea94, HL97, Kan95, LaL92, LPR92, LW97, LLP^{+}99, LN96, McC94, Nan94, OM96a, OM96b, Pig91, Rie94b, Smi93, Spe96b, Ste91, Swe96, Tho90a, Tuc96, Weh91, Won98, Wyc96, Zie92a, Zie94d, Zie94-27, Zie96-37, Zie99w, HR94, Zie95-50, Zie98s, Bar92b, Bar95, Bax92, Bax96a, Boo94b, Boo98, Bro90a, Chi91, Dea95, Dob96, Fai91, Gri99, Gue90, Jac95, Jon91a, Kah94, Kir92, LuV90, Mas99, Mil90, Mor93a, Not92a, Sar94b, Sin97, Tor95, Van92a, Was97, Zie91d, Zie92m, Zie95o, Zie95y, Zie98d, Zie99e, Jon90, Zie98f, Zie98g]. Applied [Bil91, Dre91, Goo91, Gra94, Jam96, Liu92a, Mat98, Mor93b, Nel95, NO96, Rie97, Sau97, Sen94b, Ste96a, Ste96b, Ver90, WVYB97, Zie91j, Zie91b, Zie94t, dVU96, Zie97-44, Ale95, Cuc92, DeV99, Dri92, Elt90, Eub93, Gas94, Gol99, Gra98, Gri97a, Gun99, Han98, Hec95, Jan94, Kia90, Kia92b, Kia93b, LC98, Lyn97b, Mar97, Nai94, Rag91, Ste98, Tho99, Wat90, Win90, Woz91, Zie90b, Zie90c, Zie90a, Zie92b, Zie93c, Zie95e, Zie96d, Zie96c, Zie97-44, Zie98h, Zie99c, Zie99d, vN90]. Apply [Pro92a]. Applying [Che98]. Approach [Cam98, CHW97, CM93b, CM93a, EH96, FL97, Gri96, Gun93, HO93, Hah93, Hec96, HLD98, LST93, ME99, Mar99, Mil94, Mor92, Ngu96a, Pet93, Sam99a, Ste99, Tho90b, USM99, VS96, VM91, Won92, Zie93-34, Zie94j, Zie95b, Ada91, Bal93, Bor97c, Boy90, Bul96, Deh90, Edg92, Fre92, Gre94, Haw90, Lar95, Liu92a, Lyn95, Mad96, MA92, Qiu95, Sar94a, Stu94, Vin92, Wil98, Zie91j, Zie93-29, Zie94t, Zie94s, Zie94z, Zie95-29]. Approaches [Ale98, Sim97]. Appropriate [HC91]. Approximate [CDKS92, DS93]. Approximation [KM99, Jol90]. ARCH [Zie99e]. Archer [Zie99-48]. Archetypal [CB94]. Architecture [Ara92]. Argentesi [War90]. Arising [MS95]. Arlene [Zie96-47, Zie99-52]. Arlinghaus [Zie98-33]. Armed [Yin91]. Armstrong [Zie97n, Zie90i, Zie92h]. Arnljot [Van96b]. Arnold [Bax92, Far93, Zie99-37, Wol92]. Arrays [HS99, Ngu96b, WW92]. Art [Gut92]. Arthur [Gil92, Law98, Spo90, Zie98f, Fai91]. Articles [Bor97a]. Artificial [Zie95f]. Ashish [Chi91]. Aspects [MS90, O’D97, Zie98-41, Zie94-28, Bas99, Zie95g]. ASQC [Zie97o, Zie97-31]. Assembled [Bis97]. Assessing [TC90]. Assessment [Dow94, GD94, Got94c, Got94b, Jou94, Liu99, Rut94, Tor94, Wad99, WB94, Zie95n, Zie99-28, Pla96, Zie91n, Zie95m, Zie98y, Zie99j]. Assessments [Zie90y]. Assignments [Zie91b]. Association [Jol97]. Assumed [DJ94]. Assumption [Kun97]. Assumptions [Col92, St.94]. Assurance [Lew92, Spa99, Zie93l, Zie99-33, Bru93b, Gan97, Jai95, Sea94, Sha92]. Astrom [Har91d]. Astronomy [Zie98g, Coa94]. Astrostatistics [Sin98]. Asymptotic [Ang90b, Ano90f, SS89]. Asymptotically [Sch92b]. Asymptotics [SP95]. Athanasios [Gro92]. Atkinson [Whe93]. Atmospheric [Ram96, Zie96-38]. Attributes [Car93, NW93]. Auditing [Zie96q]. Augmenting [RH95]. Autocorrelations [ZP94]. Automated [Kir93, Run90]. Availability [Pla96]. Avenhaus [Dav99, War90]. Average [Bax90, DP91, FHT90, Gan95, Hes95, Hun90b, LWCR92, LS90b, LS90c, MH90, Van96a, WM90]. Averill [Dea94]. Award [Zie96-46]. Ayton [Zie95-42]. Ayyub [Bry99]. Azzalini [DeV99, Lei99].

B [Alb93, Bar92b, Ber97, Bod97, Bru95, Bzi97, Cox96, Cra94, Die92, Dow95, Edg92, Ful97, Gri90, Hap92, Har91d, Hin97, Jam96, Jen92, Jen93, Koe92, Lee99, Mac90a, MA92, Ord96, Pri93, Qiu95, Run90, Rut93, Sau97, Sau98, Smi91, Spa99, Tsa91, Zie91c, Zie92t, Zie93j, Zie94u, Zie94e, Zie94-38, Zie96q, Zie96-37, Zie96-40, Zie96g, Zie96s, Zie96z, Zie97p, Zie98-36, Zie98g, Zie98b, Zie98w, Zie99-52]. Baafi [Zie98u]. Babu [Sin98, Coa94]. Back [Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d]. Backwards [Ano90e, Sim88]. Backwards-Stepping [Ano90e, Sim88]. Badavas [Zie94-34]. Badoux [Zie96e]. Bailar [Zie98-42]. Bain [Bry92]. Bajaria [Pro93a]. Bakeman [Ber92a]. Balakrishnan [Far93, Sin97, Ber96b, Fam95, Glo93, Pos98, Sen97, Zie96-34, Zie98d, Zie98e]. Balanced [AD96, BT89, Van92b, Zie98-45, BT98]. Baldeo [Joh91]. Baldrige [Zie96-46]. Band [Luc98, LG99]. Bandit [Yin91]. Bands [Dal90, Spu93]. Banks [Mil91, Zie98o, Zie98p]. Barbara [Bel95]. Bare [Kal93, Kal93]. Barker [Edg92, Zie95-30]. Barkman [Pla90]. Barlow [Gru95, Hes91, McC99]. Barnard [Die92]. Barndorff [Zie94-28, Har91c, Zie97-42]. Barndorff-Nielsen [Har91c, Zie97-42]. Barndorff-Nielson [Zie94-28]. Barnes [Boy90]. Barnett [Bre95b, Cam98, Zie94-42, Zie99-45, Zie99-44, Min93]. Barrentine [Jen93]. Barrett [Bru92a]. Barrie [Col93]. Barry [Bor97c, Far93]. Bartell [Zie94q]. Bartoszynski [Cha98]. Based [Ano91f, Boy90, CWWR97, CH96b, DH92, EMSD91, EM99, GK90, Glo95, Gri96, Haw91b, Kle94, Lei99, MS90, Mee90a, ML95, Ney94, ST90a, Sen98, Sin83, YE94, Zie99-31, Sen99b, Sin92, Cha93]. Basford [Raz91]. BASIC [Tho90a, Mil94, Moo96, Zie94-45, LC94, Nem92, Zie94h, Zie94g, Zie97b]. Basics [Zie98-46, Zie91h]. Basilevsky [Jac95]. Basseville [Rie94a]. Basu [Sin97]. Batch [MB98, NM95]. Batches [RH92]. Bates [Mil90]. Batteries [Wal92b]. Battery [Tho94]. Bayes [Ang91a, Ruk97a, MKA96, Ruk97a]. Bayesian [Deh90, Fox98, Rie97, Sam99b, Yak97, Zie96h, Zie98l, Zie99-27, CH96a, CHW97, CDKS92, DJ94, EMS98, HS93, MW90, MMY93, MT92, USM99, Ber92b, Fox91, Gru90, Gue90, Mar99, Ruk98, Wel90, Wes92, Zie96e, Zie97e, Zie97c, Zie97d, Zie98j, Zie98i]. Beating [Smi91]. Beauregard [Bel95]. Bebe [Kal93]. Bechhofer [Tam96]. Becker [Sco90]. Beckman [Ano90g]. Beebe [Tob99]. before [Zie92a]. Beginning [Lee94]. Behavior [ST90b]. Behnken [Bor95a, DDPG94]. Belief [Ter92]. Bell [Law98]. Belle [Zie94i]. Belsley [Gol93]. Bemowski [Zie99a]. Benchmarking [Zie96f]. Bendat [Nan94, Taa91]. Benjamin [Fre95]. Bennett [Beg90]. Benson [Pit92]. Bent [Bor95b, Sen99a]. Benzecri [Zie93g]. Beran [Glo97]. Berenson [Wat90, Zie96-52]. Berestizhevsky [Zie99-31]. Berger [Ang91c, Zie93v, Zie96-40]. Berk [Zie97j, Zie99i]. Bernard [Cob91, Geo98, Har91b, Lal99, Zie90x, Zie97-35]. Bernardo [Fox91]. Bernd [Bay92]. Bernhard [Haw90]. Berry [Har91b, Zie91r, Zie97-35, Zie97c, Zie97d, St.94]. Berthouex [Gat96]. Bertram [Maz93]. Best [Pro92b, Kin91, PBR92]. Besterfield [McC91]. Beta [Roc93, Zie91c]. Beth [Tob99]. Bethea [Bzi97, Jan94]. Betteley [Pan96]. Better [Alm95, Bre95a, CR93, Kel97]. Betty [Kim91]. between [Arm90, Atk96, Box90, DL90a, God90, Hah90, HK90, Hoo90, Hun90a, KMP95, Mac90b, Moo90, Sch91d, SW95, Tri90, Wan90, Zie95y]. beyond [Kra90]. Bhattacharya [Kir92]. Bhattacharyya [Boo93]. Bhote [Hoc94]. Bible [Car96, Spe96a]. Biemer [Zie90-28, Zie93r, Zie98-43]. Bilal [Bry99]. Bilateral [God95]. Billard [KR93]. Bimal [Roh91]. Binary [ET96, Tho93, Zie91q]. Binder [Zie96g]. Binomial [EMSD91, MW90]. Bioavailability [Zie94p]. Bioequivalence [Zie94p]. Biological [Zie93-28, Zie94-37, Zie96t]. Biology [Tho90a, Zie98-42, Nik93, Zie98-36]. Biometry [Zie94-27]. Biostatistical [Tso92, Zie95-27]. Biostatistics [Zie94-30, Zie96v, Zie97d, Zie94i, Zie95i, Zie95h]. Biotechnology [Jen91]. Biplot [SC93]. Biplots [Tro97]. Birkes [Fai94]. Birnbaum [RN91]. Birolini [Dog99]. Bishop [Web90]. Bissell [Boo94a]. Bivariate [GI92, Not92a, Not92c, Wel93]. Bjørn [SG99]. Blaisdell [Lyn94]. Blank [Hem99]. Block [Hol98, Khu92, Ngu94]. Blocked [Bis94, SCF97]. Blocking [Ros99, SWC97]. Blocks [DPL^{+}93, Hol98, DPL^{+}94]. Bluman [Sar94a]. Bob [Zie95w, Zie98-31]. Bolasco [Zie92k]. Bond [Zie90-27]. Bonett [Gil92]. Book [Aba98, Acr97, Ada91, Ada92, Ada93, Ada94, Ada96, Ajm99, Alb93, Ale91, Ale95, Ale98, Ale99, All96, Alt96, Alt99, Amm92, And90, Ang90a, Ang90b, Ang91a, Ang91b, Ang91c, Ang92, Ara92, Arm92, Arm93, Arm97, Arn90, Aye98, Bab98, Baj98, Bak98, Bal93, BR95, Bar91, Bar92b, Bar92a, Bar95, Bas90, Bas97, Bas99, Bat99, Bax92, Bax95, Bax96b, Bax96a, Bay92, Bay95, Bay99, Beg90, Bel95, Bel99, Bem96, Ben93, Ber91, Ber92a, Ber96b, Ber97, Ber92b, Bie92, Bil91, Bil92, Bil94, Bla94, Bla92, Bli91, Bod97, Bog92, Boo90, Boo95, Boo92, Boo93, Boo94b, Boo94a, Boo97, Boo98, Boo99, Bor97a, Bor94, Bor97b, Bor95b, Bor97c, Bor98, Bos92, Bow91, Boy90, Bre95b, Bre97, Bre90, Bre96, Bri91, Bri96, Bro90a, Bro90b, Bro92, Bru91, Bru92a, Bru92b]. Book [Bru93b, Bru93a, Bru94, Bru95, Bru96, Bru97, Bry92, Bry99, Buc90, Bul91, Bul96, Bur91, Bzi97, Cab94, Cab98, Cal94, Cal98, Cam90, Cam96, Cam98, Car90a, Car90b, Car92, Car93, Car95, Car96, Car97, Cha95, Cha92, Cha98, Cha93, Che92, Che94, Che98, Che91, Chi91, Chr97, Chu91, Cly94, Coa94, Cob91, Coh92, Col92, Col93, Coo96b, Cox93, Cox96, Cra94, Cro94, Cro99, Cry90, Cuc92, Czi92, Das93, Dav99, DeV99, Dea91, Dea94, Dea95, Dea96, Dea97, Deh90, Der90, Dic91, Dic97, Die92, Dob96, Dog93, Dog99, Dom90, Don90, Dov93, Dow95, Dre91, Dri91, Dri92, Dun90, Dun96, Dun97, Eas90, Eas99, Edg92, Edg98, Elt90, Emp92, Emp93, Emp97, Emp99, Ens95, Eub90, Eub93, Fai91, Fai94, Fal99, Fam95, Fam97, Fam99, Far90]. Book [Far92, Far93, Fot95, Fot96, Fox91, Fox94, Fox97, Fox98, Fre92, Fre99, Fre93, Fre94, Fre95, Fre97, Fri97, Ful95, Ful97, Gal95, Gal92, Gal93, Gan97, Gas97, Gas94, Gat96, GR97, Geo94, Geo98, Gho93, Gho94, Gib91, Gil92, Gla98, Gle91, Gle96, Glo93, Glo95, Glo97, Glo99, God93, God95, Goi96, Gol99, Gol95, Gol93, Gol98, Gol90, Goo91, Goo92a, Goo95, Goo92b, Got92, Got94a, Gov92, Gra90, Gra94, Gra98, Gre94, Gre98, Gri90, Gri96, Gri97a, Gri99, Gri98, Gri95, Gro92, Gro96, Gro99, Gru90, Gru95, Gru97, Gue90, Gun90, Gun96a, Gun99, Gun94b, Gun94a, Gun96b, Gut90, Gut92, Hab92, Had92, Hag90, Hag91, Hag96, Hal95, Hal98, Hal94, HH96, Ham96, Han90, Han93, Han97, Han98, Hap91, Hap92, Har91a, Har91b, Har91d, Har91c]. Book [Har92a, Has91, Haw90, Hea94, Hec90, Hec95, Hec96, Hec99, Hei90, Hem99, Hes91, Hes93, Hes97, Hey92, Hey98, Hey93, Hig98, Hil95, Hin97, Hoc94, Hof95, Hor98, Hul92, Huz95, Hwa92, IHMT95, Ing92, Iye98, Iye90, Ize90, Ize91, Jac95, Jai95, Jam96, Jan94, Jen90, Jen91, Jen92, Jen93, Job99, Joh91, Jol90, Jol93, Jol96, Jol97, Jon90, Jon91b, Jon91a, Jon92, Jon93, Kah94, Kal93, Kal98, Kan95, Kat91, KR93, KB94, Kat98, Ked96, Ked98, Kel97, Kha95, Kia90, Kia92a, Kia92b, Kia93a, Kia93b, Kia94, Kia95a, Kia95b, Kim91, Kin91, Kin93, Kin97, Kir92, Kir93, Kle94, Ko95b, Ko95a, Koe92, Koh90, Koh94, Koo90, KY91, Kot92, Kra90, Kra92, Kus92, Kus98, Kus93, LaL92, LaL95, Lal99, Lam90, Lar95, Law95, Law92, Law98]. Book [Law96, LC94, LC98, LeP90a, LeP90b, Lee90, Lee93, Lee94, Lee99, Lei91, Lei92, Lei93, Lei97, Lei99, Len90, Len93, Len95, Lew92, Lew93, LH97, Lin99, Lin92, Lin90, Lit91, Liu92a, Liu92b, Liu99, Loh92, Loh94, Lor92, Lor93, Lor99, LuV90, Lyn94, Lyn95, Lyn97a, Lyn97b, Lyn99, Mac90a, Mad96, Mad98, Mag90, Mag96, Mar97, Mar94, MA92, Mar99, Mas99, Mat96, Mat97, Mat98, Maz93, Maz98, McA90, McC90, McC91, McC98, McC94, McC99, McG94, McL95, Mee91a, Mee91b, Men91, Mi97, Mi99, Mil91, Mil94, Mil99, Mil90, Mil98, Min93, Moo91, Moo96, Moo99, Mor93a, Mor99a, Mor92, Mor99b, Mor97a, Mor93b, Mul97, Mye91, Mye96, Nai94, Nai96, Nan94, Nar93, Nay98, Nel98, Nem92, Nem95, Nem97, Neu99, Nik93, Nik95, Nip92, Not92a, Not92c, Not92b, O’D97, O’D99]. Book [Oli97, Oli98, Ord96, Osb97, Ott90, Owe91, Owe96, Pan92, Pan96, Pap95, Pap96, Par92, Pat91, Ped98, Per92, Per98, Per97, Pet99, Pfa92, Pfa99, Pie90b, Pie90a, Pig91, Pit92, Pit96, Pla90, Pla96, Pol93, Por93, Por97b, Por97a, Pos98, Pot93, Pri93, Pri98, Pro90, Pro92b, Pro92a, Pro93a, Pro93b, Pro94, Pro96, Qiu95, Rab92, Rag91, Rag93, Rag95, Ram99, Ram93, Ram91, Ram94, Ram96, Rao98, Ray93, Raz91, Rea93, Rie94b, Rie94a, Rie97, Rig92b, Rig92a, Rig93, Rig99, Rob90, Roh91, Rol93, Roz93, Ruk97a, Ruk97b, Ruk98, Run90, Rus94, Rut93, Sac93, Sai96, Sam90, Sam99b, Sam99a, San98, San92, San99, Sar90, Sar93, Sar94b, Sar94a, Sau92, Sau94, Sau97, Sau98, Sch91a, Sch92a, Sch96a, Sch91b, Sch95, Sch92c, Sch93a, Sch94a]. Book [Sch96b, Sch92d, Sch93c, Sch93b, Sch94b, Sch96c, Sco90, Sco91, Sco92, Sea90, Sea91, Sea93, Sea94, Sen99a, Sen94b, Sen94a, Sen95, Sen97, Sha92, Sha93, Sha95, Sha97c, Sha97b, Sha97a, Sha98, She90a, She90b, Sho90, Sim97, Sim91, Sin90a, Sin90b, Sin97, Sin98, Smi91, Smi93, Soy95, Spa99, SN95, Spe96a, Spe91, Spo90, Spr91, Sri92, Sri93, St.94, Ste90a, Ste91, Ste90c, Ste98, SG99, Ste99, Sti97, Sto96, Sto92, Str96, Stu94, Sys99, Taa91, Taa99, Tam96, Tan98, Tan93, Ter92, Thi91, Tho99, Tho90a, Tho93, Tho90b, Tho92b, Tim95, Tim96, Tim99, Tob99, Tom95, Tor95, Tor99, Tro97, Tsa91, Tso92, Van92a, Van91, Van96b, Ver90, Vin92, Voe94, Vuk98, Wad99, Wal92a, Wal92b, Wan94, War90, Was91, Was97, Wat90, Wat96]. Book [Web90, Weh91, Weh93, Weh95, Wei90, Wei94, Wel90, Wel93, Wen93, Wen94, Wes92, Wha95, Whe93, Wil90a, Wil90b, Wil98, Win90, Win93, Wit94, Wol91, Wol96, Wol92, Won94, Won98, Woo91, Woo92, Woz91, Wu99, Wyc96, Yak97, Yan96, Yan97, Yeh93, Yin91, Zei94, Zel94, Zha92, Zie90l, Zie90p, Zie90-28, Zie90x, Zie90-27, Zie90i, Zie90m, Zie90o, Zie90v, Zie90h, Zie90q, Zie90j, Zie90z, Zie90e, Zie90y, Zie90s, Zie90d, Zie90f, Zie90n, Zie90t, Zie90r, Zie90w, Zie90u, Zie90b, Zie90c, Zie90a, Zie90k, Zie90g, Zie91g, Zie91n, Zie91i, Zie91m, Zie91q, Zie91r, Zie91p, Zie91l, Zie91k, Zie91f, Zie91e, Zie91a, Zie91c, Zie91j, Zie91d, Zie91b, Zie91h, Zie91o, Zie92b, Zie92j, Zie92f, Zie92a, Zie92t, Zie92h, Zie92g, Zie92d, Zie92q, Zie92k, Zie92e, Zie92c, Zie92x, Zie92i]. Book [Zie92r, Zie92w, Zie92v, Zie92p, Zie92m, Zie92l, Zie92n, Zie92o, Zie92u, Zie92s, Zie93-30, Zie93e, Zie93c, Zie93j, Zie93t, Zie93b, Zie93-28, Zie93-31, Zie93o, Zie93w, Zie93l, Zie93-37, Zie93p, Zie93v, Zie93h, Zie93d, Zie93x, Zie93-33, Zie93r, Zie93-29, Zie93m, Zie93z, Zie93-27, Zie93-38, Zie93a, Zie93u, Zie93q, Zie93k, Zie93g, Zie93f, Zie93i, Zie93s, Zie93-32, Zie93-36, Zie93n, Zie93-34, Zie93-35, Zie93y, Zie94o, Zie94u, Zie94e, Zie94h, Zie94f, Zie94k, Zie94-42, Zie94d, Zie94w, Zie94r, Zie94t, Zie94x, Zie94-28, Zie94-29, Zie94j, Zie94a, Zie94v, Zie94-30, Zie94-41, Zie94-39, Zie94y, Zie94-31, Zie94-45, Zie94g, Zie94l, Zie94-44, Zie94b, Zie94s, Zie94-32, Zie94-40, Zie94-38, Zie94-43, Zie94-33, Zie94-36, Zie94-37, Zie94-27, Zie94z, Zie94i, Zie94p, Zie94c, Zie94m, Zie94n, Zie94-34, Zie94q, Zie94-35, Zie95q]. Book [Zie95-48, Zie95p, Zie95-30, Zie95u, Zie95r, Zie95-33, Zie95g, Zie95-42, Zie95-29, Zie95d, Zie95-41, Zie95i, Zie95t, Zie95-35, Zie95-45, Zie95-31, Zie95o, Zie95-38, Zie95e, Zie95-49, Zie95-32, Zie95b, Zie95h, Zie95-40, Zie95z, Zie95s, Zie95y, Zie95v, Zie95f, Zie95l, Zie95n, Zie95j, Zie95x, Zie95w, Zie95-37, Zie95-47, Zie95-46, Zie95-28, Zie95-50, Zie95m, Zie95-39, Zie95-36, Zie95-27, Zie95-34, Zie95k, Zie95c, Zie95a, Zie95-44, Zie95-43, Zie96m, Zie96-47, Zie96-34, Zie96-32, Zie96o, Zie96q, Zie96b, Zie96-28, Zie96t, Zie96d, Zie96-45, Zie96w, Zie96-30, Zie96n, Zie96-37, Zie96-41, Zie96e, Zie96-31, Zie96-44, Zie96-50, Zie96j, Zie96-43, Zie96i, Zie96h, Zie96-40, Zie96-27, Zie96a, Zie96x, Zie96-33, Zie96-35, Zie96k, Zie96-51, Zie96-52, Zie96g, Zie96-48, Zie96-29, Zie96-39, Zie96-42, Zie96p, Zie96-38, Zie96r, Zie96s, Zie96c, Zie96u, Zie96l]. Book [Zie96-46, Zie96f, Zie96z, Zie96-49, Zie96-36, Zie96v, Zie96y, Zie97-39, Zie97b, Zie97h, Zie97-32, Zie97-27, Zie97-38, Zie97m, Zie97-29, Zie97s, Zie97r, Zie97-34, Zie97-28, Zie97l, Zie97e, Zie97u, Zie97f, Zie97x, Zie97p, Zie97-36, Zie97-33, Zie97-42, Zie97-35, Zie97o, Zie97v, Zie97c, Zie97d, Zie97t, Zie97-37, Zie97-43, Zie97q, Zie97j, Zie97-41, Zie97z, Zie97g, Zie97-30, Zie97a, Zie97w, Zie97-31, Zie97k, Zie97i, Zie97y, Zie97-44, Zie97-40, Zie97n, Zie98o, Zie98-42, Zie98-45, Zie98-48, Zie98-39, Zie98h, Zie98-34, Zie98-41, Zie98-32, Zie98-27, Zie98a, Zie98-38, Zie98t, Zie98s, Zie98y, Zie98j, Zie98n, Zie98-35, Zie98q, Zie98p, Zie98-29, Zie98-47, Zie98v, Zie98-31, Zie98-33, Zie98-46, Zie98c, Zie98f, Zie98-43, Zie98l, Zie98m, Zie98-36, Zie98-44, Zie98-28, Zie98-40, Zie98g, Zie98-37, Zie98-30, Zie98i, Zie98r, Zie98x, Zie98d, Zie98e]. Book [Zie98k, Zie98z, Zie98b, Zie98w, Zie98u, Zie99-27, Zie99l, Zie99-42, Zie99k, Zie99-38, Zie99-31, Zie99-47, Zie99w, Zie99f, Zie99-45, Zie99-46, Zie99z, Zie99-49, Zie99-43, Zie99u, Zie99i, Zie99-28, Zie99-34, Zie99v, Zie99-29, Zie99-33, Zie99-37, Zie99e, Zie99-39, Zie99j, Zie99-36, Zie99s, Zie99c, Zie99-51, Zie99d, Zie99t, Zie99p, Zie99q, Zie99-52, Zie99r, Zie99-35, Zie99-50, Zie99a, Zie99b, Zie99o, Zie99y, Zie99n, Zie99-44, Zie99m, Zie99g, Zie99-48, Zie99h, Zie99-30, Zie99-41, Zie99-32, Zie99-40, Zie99x, Zin93, dV92, dV93, dC99, vN90, vT91, vT95]. Boostrapping [Zie94j]. Booth [Zie99q]. Booth-Kewley [Zie99q]. Bootstrap [LPR92, Zie98-36, GJMS^{+}95, Gri95, Hes97, KR93, Len95]. Bootstrapping [JR99]. Borgan [Kle94]. Boris [Cro94, Hig98, Pap96]. Boros [Sea91]. Borowiak [Iye90]. Bose [Rie94b]. Bothe [McC98]. Bouleau [SN95]. Boullion [Bzi97]. Bounds [BBT90, JD92]. Bouvier [Bor97b]. Bovas [Zie98-35]. Bowen [Beg90]. Bowerman [Zie91j, Zie94t]. Bowman [DeV99, Sin90a]. Box [Die92, Fal99, Zie95-46, Zie98q, Bor95a, DDPG94, GJMS^{+}95, KSJ96]. Box-Behnken [Bor95a]. Box-Jenkins [GJMS^{+}95]. Boxplot [GI92]. Bozdogan [Zie95-29]. Brad [Zie99a]. Bradley [Gri95, Ruk97a, Zie96y]. Branham [Bro92]. Brashers [Nai96]. Braun [Zie95-43]. Breakdown [HL97, MSH94]. Breakdown-Voltage [HL97]. Breakthrough [Zie94-45, Zie94-45]. Breakthroughs [Zie98k]. Brecht [Gil92]. Breckling [Ste91]. Breen [Fox97]. Brereton [Zie92c, Zie94k, Nem95, Smi93]. Bretthorst [Gru90]. Brettner [Dog93]. Brewer [Gan97]. Breyfogle [Pol93]. Brian [Gre98, Ko95a, Koh94, Moo91, Nai94, Sha93, Zie95q]. Bridging [Zie95y]. Brillinger [Zie99t]. British [Jon90]. Britt [Kel97]. Bro [Wal92b]. Brock [Zie98y]. Brockwell [Sha97b, Zie92v]. Brooker [Zie94b]. Brown [Gun90, Che91, Col93, Gat96, Gle96]. Bruce [Han98, Zie91j, Zie94t]. Brush [Wil90a]. Bryan [Car96, Nem97, Nik93, Zie95-48]. Bryk [Cal94]. Bucklew [Par92]. Buehler [JD92]. Bugaj [Cha98]. Building [Pie90a, Bel99]. Buja [Zie93e]. Bunday [Sha93]. Bunke [Zie91l, Zie91l]. Burch [Zie95-32]. Burdick [Pot93, Ste98]. Burger [Bla92]. Burke [Zie94-43]. Burman [LD92]. Burn [Sam99a]. Burn-In [Sam99a]. Burr [Owe91, Pro94]. Bus [MSF95]. Bus-Motor-Failure [MSF95]. Business [Elt90, HH98a, HH98b, MTV98, Nem98, Zie90-27, Zie95q, Zie95-50, Zie97m, Zie97v, Zie96-52, Zie96-52, SG99, Zie96g, Zie96f]. Butler [Zie96-49]. Buydens [Zie98w, BR95]. Byron [Gal93, Zie91q].

C [Aba98, Ale95, Ang92, Bak98, Bar92b, Bar92a, Ber91, Bie92, Bod97, Boo95, Boo98, Bor98, Chr97, Cro94, Dic91, Dow95, Dre91, Eas99, Ens95, Far93, Ful95, Gat96, GR97, Gra94, Gun96a, Gun94a, Hei90, Hil95, Hul92, Jen90, Ked98, Kir92, KY91, LaL92, LuV90, Mad96, Mar94, Mat96, Mat97, Maz98, Nar93, Not92a, Not92c, Ord96, Per97, Por93, Pro92a, Rag93, Rie94b, Sam90, Sar90, Sau92, Sch91b, Sch93c, Sch93b, Sha95, Sha97c, Sin90b, Spr91, Sri92, Sys99, Tim95, Tim99, Tro97, Wal92b, Wei94, Wel90, Whe93, Wol92, Yan96, Yin91, Zha92, Zie90h, Zie90e, Zie90u, Zie90k, Zie91k, Zie91a, Zie92f, Zie92g, Zie92c, Zie93-28, Zie94k, Zie94j, Zie94v, Zie94g, Zie95-41, Zie95s, Zie95n, Zie95k, Zie96t, Zie96-37, Zie96-41, Zie96-31, Zie96i, Zie96h]. C [Zie96a, Zie96-33, Zie96-42, Zie96p, Zie96-49, Zie97g, Zie97-44, Zie98o, Zie98-39, Zie98-32, Zie98p, Zie98-43, Zie98x, Zie99l, Zie99-35, dV93, vT95]. Cable [Jen92]. Calculus [Zie94d]. Calibration [Ano90f, CV97, JR99, MER91, ME96, Oma91, Sea94, SS89, Tho91, Tho92a, Gle96, Hag91]. California [GK93]. Cameron [Pet99]. Camil [Mas99]. Camp [Zie96f]. Campbell [Ber97]. Canavos [Zie95-50]. Canonical [Kii92]. Cantor [Zie98r]. Canty [Dav99]. Capability [Cha95, Sto92, McC98]. Captive [GK93]. Captive-Air [GK93]. Carbon [Ano98e, DP97]. Cardoso [Zie98-38]. Care [God90]. Carey [Zie97j, Zie99i]. Carl [Beg90, Hec95]. Carley [Zie92d]. Carlin [Ruk97a]. Carlo [Gas97, Gru95, Nik93, Was97, Zie96-27, Zie98-32, Zie98-36]. Carlson [Emp97, Zie97f]. Carpenter [Zie96-33]. Carr [Hey93, Wat96]. Carroll [Arm97, Mat98, Wei90, Ano90g]. Cary [Hei90]. Case [BH93, CPJ92, Kaf94, Nem95, RH95, Zie90i, Zie98-40, Zie99-37, Zie96h, Zie98l, Zie99f]. Case-Deletion [CPJ92]. Casella [Ang91c, Gho93, Zie99-49]. Cases [Kra92, Zie97f]. Castillo [Ang90a]. Catalog [Tür94]. Categorical [Cal98, CH96a, Cob91, Hap92, Kia93b, KTTS90, Sau98, Hap91, Zie96i]. Categorized [CR98]. Categorized-Components [CR98]. Caudill [Zie96-49]. Caulcott [Zie91h, Zie92u]. Caulcutt [Dun96]. Causal [OUH93, Zel94]. Cause [AWS94, Fel94, Kva96, Luc94, WMP94a, WMP94b, Zie94-36]. Cell [Oeh94]. Cells [Ano90e, Sim88]. Censored [EM99, HW91, Mor93a, Rao98, Sin83, VB94b, Sin92]. Censoring [HLS98]. Centered [Zie96-35]. central [Bor95a]. Century [Zie96r]. Cepstral [Ano00, SBD98]. Certain [WC92]. Cervetz [Zie95k]. Chacko [Liu92a]. Chain [Bru96, Gas97]. Chakraborti [Zie93s]. Chakraborty [Zie93-28]. Challenges [Coa94, HH98a, HH98b, MTV98, Nem98]. Chaloner [Zie97c]. Chambers [Hes93, Sco90, Bat99, Zie93-38, Zie96-44]. Chamont [Das93]. Chance [Ara92]. Chang [Kin93]. Change [TW96]. Changes [Rie94a, Yas95]. Changing [GSH98]. Chaos [Zie94-28, Voe94]. Chaotic [Sti97]. Chap [Gol99]. Characteristics [AAK99, EMSD91, Viv91].

Characterization [Li97, Cab94]. Characterizing [CCC90]. Charles [Bre97, Gho93, Hey98, Ize90, Kal98, Mag90, Pit96, Vin92, Zie97-33, dC99]. Chart [LWCR92, Pap95, Zha98, Zie90h]. Charts [AWS94, BC95, Emp93, Fel94, FB94, Gan95, HD95, Hea90, Lau94, LS90a, Luc94, NM95, PD95, RAA90a, RAA90b, RD95a, RD95b, ST90b, SGW96, SWG95, Tat96, Tat97, WMP94a, WMP94b, Wes90, Woo90, Ada96, Kel97]. Chase [Gun90]. Chatfield [Dic91, Don90, Zie96-31, Zie97-40]. Chatterjee [Maz93]. Chaturvedi [Mat98]. Chaudhuri [Zie93-32]. Cheeseman [Zie95-33]. Chemical [Ano97f, AB97, Emp97, Koh90, Lew92, Ram94, Sch95, Wol93, Zie91h, Zie92d, Zie93l, Zie94o, Zie97-31, Zie98y, Zie99-33]. Chemistry [CDKS92, LaL95, Rig92a, RL95, Tho90a, Zie96j, BR95, Bla94, Hil95, Mor93b]. Chemists [McC94, Wol91]. Chemometric [McA90, Zie94s, Bay99]. Chemometrics [Bur91, FF93b, FF93a, HM93, Nem95, Zie98w, Bay95, LaL95, Mac90a, Smi93, Tob99, Zie92c, Zie93d, Zie94k, Zie96j, Zie98m]. Chen [Joh91]. Chhikara [Bas90]. Chi [Kin97]. Chi-Squared [Kin97]. Chien [Sam99a]. Chinnappa [Zie96g]. Choe [Zel94]. Choi [Ale98]. Choice [Emp93]. Choose [Wil90a]. Choosing [Ste90a]. Chow [Zie94p, Zie96-39, Dea95, Nay98, Zie99-40]. Chris [Zie95-34, Zie97-40]. Christensen [Hab92, Nip92, Zie98-27]. Christian [Cab98, Zie99z, Zie99e, Zie99j]. Christianson [Zie96g]. Christine [Pri98]. Christopher [Bor98, Don90]. Chui [Ize90]. Chung [Dea95]. Circuit [HNF97, TH93]. Circular [WL97, Weh95]. Claims [KLR91, Woo91]. Clark [Neu99, Zie92e, Zie97h]. Clarke [Por97a]. Clarotti [Gru95, Sam90]. Class [BE94, CR93, EMS98, Gol98, Hoc94, Lin93b, Rea93]. Classical [Zie91d]. Classification [Hig98, Hol98, Sha97c, Sim97, Wad99, Ful95, Mag90]. Clayton [Zie95-39, Zie98t]. Clemen [Liu92b]. Clements [Bru92a, Kra90, Rut93]. Cleveland [Buc90, Gun94b, Zie97-41]. Clifford [All96, Kim91, Mor93a, vT95]. Climatology [Mi97]. Clinical [Zie94n, Zie96b, Zie99g]. Clive [Gun96b]. Clogg [vT95]. Cluster [Had92, Koh94]. Clustered [HNF97]. Clustering [Koh90, Rou95, Sar90, STR99, Hig98]. Clyde [Alt96]. Clymer [Dea91]. Coal [BH93]. Coating [CR98]. Cobb [Lyn99]. Cody [Zie96-36, Zie98h]. Coefficient [Gal95]. Cohen [Kim91, Zie99t, Bli91, Mor93a, Zie96-32]. Cohn [Woo91]. Coleman [Aba98]. Colin [Pet99, Zie91f, Zie91e]. Collaní [Tor99]. Collected [Zie95-43, Zie91f, Zie91e]. Collection [Wen94]. Collections [Mor98a]. Colledge [Zie96g]. Collett [Tho93, Zie95x]. Collinearities [CGG92]. Collinearity [Gol93]. Collins [Zie98-43]. Color [VD94]. Column [GD91, Tür94]. Columnwise [LW97]. Columnwise-Pairwise [LW97]. Combinations [CL93]. Combining [Viv91, Sri93]. Comment [BZM96, PBR92, PGT95]. Commentary [Arm90, Box90, God90, Hah90, Hoo90, Hun90a, Mac90b, Moo90, Sne90, Tri90, Wan90]. Comments [Ano93f, FMQ92, HSPF92, Lau94, LW98, LC90, LLW^{+}96, Mal95, MLV92, Sch98b, SDP^{+}95, WKWL95, YWMP95, vdVtBM97]. Committee [Zie97-31]. Common [Kva96, Mee90b, Zie99g]. Communication [Zie99h].

Communications [Arm90, Box90, Hah90, HK90, Hoo90, Hun90a, Mac90b, Moo90, Tri90, Wan90, God90]. Companion [Wit94, Rut93]. Comparative [Ang91b, YE94]. Comparing [Spu92]. Comparison [HL95, KM99, ME96, RH92, Spu93, Vin92, ZZ91]. Comparisons [Chr97, DS93, MME98, SE93, TI99, Tür94, Zie94-27, Tam96, Zie95-43]. Competition [Smi91]. Competitive [Zad95]. Complementary [Zad95]. Complete [Tür94]. Complex [CHW97, FL97, MW90, Zie96-29, Ano97e, LH97, Weh93]. Compliance [Dav99]. Component [EMSD91, Rag98, RM95, RL95, Tat96, Viv91, Jol90]. Component-Test [EMSD91]. Components [CR98, DPL^{+}93, KM94, MW90, SS97, Yas94, DPL^{+}94, ZM94, Gho93, Lei93, Pot93]. composite [Bor95a]. Composites [Ano98e, DP97]. Compound [AB97, HS99, Ros96, Ros99]. Compression [JS99]. COMPSTAT [Zie90d]. Computation [Zie98i, Zie92d]. Computational [Mor91, Sch96a, Wei94, Zie98-34, Zie98-41, Mi97, Zie97g]. Computationally [Bro90b]. Computations [EMS98, Thi91, Sco91]. Computer [And90, Bax92, Bla92, Boy90, Hal95, LLW^{+}96, MMY93, Mor97a, Ngu94, WBS^{+}92, Zie95-49, Zie95-50, Zie95c, Zie97z, Zie98f, Zie90e, Len90, Len95, Zie92e, Zie97h, Zin93]. Computer-Aided [Zie92e, Zie97h]. Computer-Based [Boy90]. Computer-Interactive [Zin93]. Computing [Amm92, Bro90b, Gri93, Hin97, KM99, Tho90a, Wal92a, Zie96-27, Zie93e]. Comrey [Pri93]. Concepts [Bar92b, Bar95, Cro99, Pit96, VTFD92, Was97, Iye98, Kat91, Zie96o, Huz95, Jen93]. Conceptual [Wil98]. Concise [Sch91a, Zie91g]. Conditional [Dow94, GD94, Got94c, Got94b, HR92, HR94, Jou94, Rut94, Tor94, WB94, Won92]. Conditioning [Gol93]. Condra [Fox94]. Conduct [Zie99q, Zie99-52]. Conference [Zie95-29]. Confidence [BBT90, CCC90, Dal90, DS93, JD92, Len93, Nel95, Pet93, Spu93, Pot93]. Configurations [Mil97]. Confounding [Fed96, JD96, MSB96a, MSB96b, Tob96]. Connection [Sch91d]. Connections [DL90a]. Conrad [Zie91b]. Consciousness [Zie95q]. Considerations [GC91]. Constantine [Zie98l]. Constrained [Deu96]. Constraints [Won98]. Construct [Bru92b]. Constructing [CL93, MN95, Ngu96a]. Construction [ACV99, Jac92, Jac93, LW97, LIV96, Ngu94, Ngu96b, Wad99]. Consul [LuV90]. contain [Ode90]. Contemporary [Pig91]. Contingency [Alb93, Ano90e, Sim88, Zie99x]. Continual [Bel99, Zie93i]. Continuous [CFRH99, Gun96b, Kra92, MT91, Nem92, Not92c, Sto96, Woo92, Zie94h, Ber96b, Fam95, Not92a, Van91, Zie94l]. Contouring [Zie90e]. Contrasts [VS96, VS98]. Control [Acr97, Ada94, AWS94, Ada96, Ano91f, BR95, Bax90, Bem96, Bon95, CFRH99, Col93, Dun97, FHT90, Fal99, Fel94, Fre99, FB94, Gan95, Gho94, God95, Han93, Haw91b, Hun90b, KD97, Kal93, Kra90, Lau94, LWCR92, LS90b, LS90c, Luc94, MH90, Mas99, McC90, McG94, McL95, MME98, Mil94, Pie90b, Pla90, Ram91, Roz93, Run90, SE93, Sar94b, She90a, Sin90b, ST90b, SGW96, Tat96, Tat97, Tor95, TWN98, TFV93, Tür94, VTFD92, WVYB97, WMP94a, WMP94b, Wes90, WM90, Yas92, Yas93, Zha98, Zie90l, Zie92r, Zie94-43, Zie94-33, Zie97-32, Zie97w, Zie99n, Lew92, Pro92a, Ada93, Bos92, Car90b, Car93, Edg98, GR97, Hal94, Har91d, Kal98, Kin93, McC91, Mil91, O’D97, Pan96, Rol93, Sau92, Vuk98, Zei94, Zie90t, Zie92w, Zie93-38, Zie94-34, Zie95-46, Zie96-37]. Control [Zie96-41, Zie97o, Zie99o, Zie99y, Zie93f]. Controlled [JR99]. Controllers [Har92b]. Controversies [Woo91]. Conventional [Cox96]. Convergence [Zie99j]. Cook [Gol95, Bog92, LC90, Moo99]. Cooper [Zie95-44]. Coordinate [DGTZ97a, DGTZ97b, Hul97, MN95, PE97, WH97]. Coordinate-Exchange [MN95]. Copp [Pro93a]. Coppi [Zie92k]. Corbett [Zie99-43]. Cornell [Pro92a, Czi92, Zie92i, Zie97y]. Correction [Ano90e, Ano90f, Ano91e, Ano91f, Ano93e, Ano93f, Ano97f, Ano97e, Ano98e, Ano99e, Ano00, BT98, DPL^{+}94, Lin95a, Sen99b, Sin92, Ste90b, Tho92a]. Corrections [Ano90g, HR94]. Correlated [AWS94, AAK99, Fel94, Luc94, Mee90a, WMP94a, WMP94b, Wes90, Yas93, Zie99t]. Correlation [Gov92, MQ91, Nan94, SW95, Zha94, Zie94z]. Correspondence [Jol90, Zie93g]. Corresponding [ZZ91]. Cortada [Zie96o]. Corter [Jol97]. Cost [Cro92, MW98, Nel95]. Cost-Driven [MW98]. Costs [Coo96b, ST90a, Zie90-28]. Cothern [Zie95n]. Cottman [Zie93-36]. Council [Sri93]. Cound [Zie93a]. Count [Owe96, Pet99]. Counting [Kle94]. Couper [Zie99v]. Course [Dre91, Far93, Gra94, Gun94a, Jon90, Jon91b, Kia92a, Lal99, Lor92, Mee91a, Pri93, Ste98, Zie98a, Dri91, Wat90]. Courts [Zie90y]. Covariance [RvD99, Sha97c]. Covariates [Kia95b]. Coverage [Sea90]. Covering [DM98]. Cox [Die92, Fre93, Jol96, Weh93, Zie91q, Zie96g, Zie97-42, Jol96, KSJ96, Zie99f]. Crain [Jen90]. Creating [Kel97]. Creighton [Gun94a]. Cressie [Zie94-41, dV93]. Creveling [Alt96, Bak98]. Criteria [ME96]. Critical [Hab90, HW90b, HW90a, Hir90, Nai90b, Nel93, YDB90]. Crocker [Pro92a]. Cromwell [Dow95, Ord96]. Cross [Zie91q, Zie94m, Zie94n]. Cross-Over [Zie91q, Zie94m, Zie94n]. Crossed [Hol98]. Crossed-Classification [Hol98]. Crossings [Fre95]. Crowder [Bar92a, Lin92, Zie97t]. Cryer [Zie95-50, Zie96-52]. Cumulative [Ano98e, DP97, SGW96, Zie90h]. Current [Sen98, Yas95, Sen99b]. Curvature [Coo96a]. Curve [BB95, Wha95]. Curves [Ann99, Ano99f, JFM99, KK99, ME99, PM99, Zie96s, And90]. Customer [San98, Zie95w, Zie96y, Zie97s, Zie98-31, Zie96k, Zie97i]. Customer-Focused [Zie96k]. CUSUM [Hea90, LS90a, RAA90b, Woo90, RAA90a, Yas92, Yas93]. Cyclic [Mor97a]. Cyclical [ST90b]. Cyrus [Dob96, O’D97]. Czerminski [Mor93b]. Czitrom [Zie98-40].

D [Ale99, Arm97, Bax96b, Boo97, Bri91, Car90a, Cha92, Coa94, Cuc92, Die92, Fai91, Fot96, Fox91, Fox98, Fre93, Ful95, Ful97, Gas97, Got94a, Gov92, Gri90, Hei90, Hey92, Jam96, Jen91, Kia92b, Kia93b, Kia95a, Kin91, Kle94, Koe92, LH97, Lin92, Mac90a, Mag96, McA90, Mil94, Mye91, Not92a, Not92c, Per97, Rag93, Rag95, Sac93, Sam90, Sch96b, Sea93, Sha93, Sin98, Smi91, St.94, Sti97, Sys99, Tho93, Tro97, Ver90, Voe94, Wat90, Weh93, Wei90, Wha95, Zel94, Zha92, Zie90-28, Zie90e, Zie90d, Zie91q, Zie92c, Zie92r, Zie92s, Zie93-30, Zie93p, Zie93-29, Zie93-38, Zie93f, Zie94k, Zie94y, Zie94c, Zie94m, Zie95p, Zie95-41, Zie95f, Zie95x, Zie95-50, Zie95-39, Zie95c, Zie95a, Zie96d, Zie96-37, Zie96-51, Zie96-52, Zie96g, Zie97-38, Zie97-42, Zie97c, Zie97d]. D [Zie97t, Zie97-37, Zie97a, Zie98o, Zie98-45, Zie98p, Zie98-43, Zie98l, Zie98-40, Zie98w, Zie99-27, Zie99-42, Zie99-37, Zie99t, Zie99q, Zie99-35, Zie99-32, Zin93, Ber98]. D’Agostino [Goi96]. Dagpunar [Gol90]. Dale [Iye90, McC91, Nai96, Zie96z]. Dallas [vT91]. Daly [Zie95a]. Damage [Ano98e, DP97]. Dana [Zie93a]. Daniel [Aye98, Goi96, Har92a, Ram96, Weh91, Woz91, Zie95h]. Data [Bat99, Boy96, Bre91b, Bro92, Cal94, Cam90, CH96a, Cob91, Cox93, Dav99, DeV99, DS93, Dov93, Dre91, EMSD91, EM99, Fri91, Geo98, Gol93, Got92, Gra90, Gre98, Gri96, Gu91, GS94, Had92, HR92, HW91, Ham96, Han97, HNF97, Hap91, Has91, HT91, HL97, JS99, Kel97, KM94, Kia92b, Kia93b, Kii92, KTTS90, Kva96, LaL95, LPY97, Lwi94, MA92, Mee91b, MSF95, Nai94, Nel95, Nem95, Not92b, Oli97, Ruk97a, Sau94, Sch91b, Sch96c, SC93, SGW96, SP95, Viv91, VB94b, Wen94, Zha98, Zie90o, Zie90v, Zie92q, Zie92k, Zie93-27, Zie94f, Zie94-29, Zie94-37, Zie94z, Zie95y, Zie95v, Zie95x, Zie95-50, Zie95a, Zie96b, Zie96-50, Zie96i, Zie96a, Zie96-38, Zie97r, Zie97f, Zie97x, Zie97-33, Zie97t, Zie97z, Zie98c, Zie99-47, Zin93, vT91, Bre91a, HR94, All96]. Data [Bar92a, Bas97, Bre95b, Cra94, Eub90, Gal92, Gro96, Gun94b, Hap92, Hec99, Lee90, Lee99, Mul97, Nel98, Nip92, Owe96, Pat91, Pet99, Pfa99, Ram91, Rao98, Sar90, Sau98, Sim91, Sti97, Taa91, Tho93, Tim96, Weh95, Zie91q, Zie94-41, Zie95-33, Zie95d, Zie95-36, Zie97-41, Zie99t, dV93, Che94, Fox98, Law98, Mar94, Ram94, Vin92, Zie90-27, Zie90f, Zie90g, Zie91h, Zie92f, Zie93h, Zie94o, Zie96m, Zie96l, Zie97j, Zie98n, Zie99i]. Data-Based [Gri96]. Data-Sequential [SGW96]. David [Arn90, Bru93a, Bru97, Car97, Cro94, Dea94, Dun90, Edg98, Fot96, Gol93, Gri98, Koo90, Law98, Len93, Loh94, Mil98, Moo96, Pan96, Rag95, Ram93, San99, Sau97, Sch93c, Sch93b, Tam96, Tim96, Wad99, Zie91i, Zie93-31, Zie98-30, Zie98b, Zie99-43, Zie99-34, Zie99d, Zie99t, Zie99-41]. Davidian [Tim96]. Davidson [Oli97]. Davies [Sch95]. Davis [McC98, Zie94-43, Zie96i, Hof95, Sac93, Sha97b, Zie92v]. Days [Zie95p]. Dead [Luc98, LG99]. Dean [Zie99c, SDP^{+}95]. Debarred [CL93]. Decision [Bax95, Gru95, Han90, Liu92b, Par92, Ter92, Zie98-39, Zie98d, Deh90, Liu92a, Mag90]. Decision-Making [Liu92a]. Decisions [Liu92b, Zie98n, Zie90g, Zie96n]. Defect [GS92, Bru91]. Defects [FRY96, HNF97, Lam92]. Defensive [Hes95]. Defined [BCL98a, BCL98b, Rob98, Sch98a]. Definition [Bis94]. Deformation [CHB90]. Degradation [BE94, LM93, LPY97, MEL98]. Degrees [DDPG94]. DeGroot [Fox91]. Dehnad [Pie90b]. Deistler [Cry90]. DeJong [Zie98w]. del [Ram99]. Delaney [Pat91]. Delay [Ano00, SBD98]. Delay-Fired [Ano00, SBD98]. Deletion [CPJ92]. Dell [Zie94-36]. DelMar [Zei94]. Deming [Ada91, Mac90a, Pfa92, Pro93b, Zie94s, Pro96, Zie95p, Zie93i]. Denker [Lin99]. Dennis [Car96, Gol95, Kan95, Moo99, Zie97-32]. Density [Ram93]. Dependence [DJ94, Iye98]. Dependent [Cal98, Cox93, Gri99]. Deployment [Zie96-32]. Derek [Boo94a]. Derivatives [MMY93]. Derived [LPY97]. Derman [O’D97, Dob96]. Descriptive [Wol96]. Deshpande [Bas97]. Design [Ano97f, Bak98, BS97, Bor94, BE94, Bow91, CH96b, Der90, Edg92, Emp97, EH96, Fox94, Fre99, Hoc94, Hol98, Jen91, Job99, Kin93, LN97, Lyn99, McC94, MS90, MW98, MMY93, MS95, NAM^{+}92, O’D99, Pie90b, Pit96, Ram99, Rig99, RJ97, Ros96, Sch95, STW91, Tan98, Van91, Web90, WT98, Zie90q, Zie91o, Zie94s, Zie96-29, Zie96-39, Zie99-38, Zie91q, Alt96, And90, Gil92, Hwa92, Kia92b, Kia93a, Men91, Pro92b, Sac93, San99, Tom95, Zie93w, Zie93d, Zie95-30, Ano97f, Ano97e, Emp97, Gan97, Gle91, Gre94, Oli98, Sri92, Tam96, Tom95, Yeh93, Zie93j, Zie94p, Zie95j, Zie97k]. Designed [CHW97, CM93b, CM93a, FL95, Gre98, Gun96a, Gun93, HO93, Hah93, LST93, Pan92, Sco91, Zie96-48]. Designing [Aye98, Bis97, RM95, Pat91, Pro92b]. Designs [ACV99, Atk96, AD96, AB97, BH97, BS99, Bis94, BL97, CL93, DM98, DL90a, DL90b, DPL^{+}93, DDPG94, DJ94, Fed96, FL97, HCV98, JD96, Kun97, LW97, LIV96, LD92, Lin93b, Lin95b, MB98, MN95, MSB96a, MSB96b, Ngu94, Ngu96a, Pan99, SCF97, Spu92, SWC97, Tob96, Tür94, VM91, Zie97y, Bor95a, DPL^{+}94, Lin93a, Lin95a, Mor97a, Whe93, Zie92f]. Dessy [Zie92c, Zie94k]. Destructively [SW95]. Desu [Sea93]. Detect [Cro94]. Detecting [Ano90e, Ano00, SBD98, Sim88, ST90b, TH93]. Detection [PF91, QY98, Tat96, Rie94a]. Determinant [RvD99]. Determination [Jac92]. Determine [Hul92]. Determining [AB97, HC91, USM99]. Deterministic [JV93]. DeToro [Zie93-37, Zie97u]. Deutsch [Zie95s, Zie98t]. Dev [Sha92]. Developing [Bak98, Zie95k]. Development [Kaf94, Pol93, Sch95, Zie91r, Sam99a, Zie92u, Zie93u, Zie97w, Zie99-39]. Developments [Zie95l]. Deviation [Par92, Tat97]. Device [Bla92]. Device-Level [Bla92]. Devices [CB93]. DeVor [Kin93]. Devore [Zie92l]. Dey [Zie99-27]. Diagnostic [Ber98]. Diagnostics [Ano91e, BNC87, CS92, CPJ92, Gol93, Haw91a, KSJ96, Sch91d, TW90, TW92, Zie93k, Zie93x]. Diagrams [Ter92]. Dialogue [Bur91]. Diane [Sim91]. Diaz [Job99]. Dickey [Zie99d]. Dickinson [Gro99]. Dictionary [Zie91a]. Diday [Zie90f]. Dieck [Kah94]. Diesel [MKA96]. Differences [Gol98]. Different [Kva96]. Digby [Dri91]. Diggle [Zie95d, Zie99t, Tso92]. Digital [Gru97, Sen94a]. DiIorio [LaL92]. Dijkstra [Tsa91, Zie90p]. Dimension [Jol90, Pfa92]. Dimensional [Kii92, WVYB97]. Dimitrakopoulos [Zie96r]. Dimitri [Sch92c, Sch93a, Sch94a]. Dippo [Zie98-43]. Direction [Ste91]. Directional [Cob91, Ste91]. Directions [Zie94-29, Zie93k]. Discovery [Mul97, Zie98c]. Discrepancy [PJ90]. Discrete [Bax95, BL94, Bru94, Gol98, Luc98, Mar99, Ped98, Rag95, Sim91, TWN98, Wel93, Goo95, Pos98]. Discrete-Event [Rag95]. Discrete-Valued [Ped98]. Discretization [Zie99j]. Discriminant [Hec95, Hir96, VB94a, Ray93]. Discriminating [Atk96]. Discrimination [HS97, Iye90]. Discussion [AWS94, Alm95, Ann99, Atk96, Bax90, Ber96a, BC95, Bon95, BK92a, Che95, CH96b, CM99, CR96, Dow94, FHT90, Fed96, Fel94, Fri91, Gib99, GD94, Gil99, Gu91, Gun93, HO93, Hab90, Hah93, HD95, HT91, HM93, Hea90, Hir90, HP92, Hul97, Hun90b, JFM99, JD96, Jou94, KMP95, KTTS90, KK99, Lam98, LST93, LS90a, Luc94, MH90, Mac92, MTV98, Nai90b, NAM^{+}92, Nem98, NO96, PD95, PE97, Rob98, Rou95, Rut94, Sch98a, Sne98, Spe96b, SWG95, Swe96, Tob96, Tor94, Tuc92, Tuc96, VV92, VV94, Van99, WG92, WB94, Wol93, WH97, Woo90, WM90, YDB90, Zad95, dVU96]. Disjunctive [Cam96]. Dispelling [Cly94]. Dispersion [BH97, MS95, NLB^{+}98, Pan99, Ros96, Ros99, Sen99a, WT98]. Dispersion-Effects [Pan99]. Displays [SC93]. Distinction [KMP95]. Distributed [Mee90a, WJ92]. Distribution [DH92, Gri93, HLS98, KGK90, LM93, LPY97, RN91, Sin97, USM99, WI94, WL91, Won92, Ode90, Zie94-45, Bas90, Ber97, Glo93, Ize91, Sin90a, Ste90c]. Distribution-Free [USM99]. Distributional [Col92]. Distributions [AWS94, Ber96b, Bru97, CDKS92, DS93, Fam95, Fel94, Hes95, LuV90, Luc94, Lwi94, Not92a, ST90a, Sha95, WMP94a, WMP94b, Yas92, Bru94, Loh92, Not92c, Pos98, Wel93, Zie94-38, Sha97a]. Disturbance [Ano90f, Luc98, SS89]. Divergence [Li97]. Division [Zie97o]. Doane [Dun90]. Dobson [Emp92]. Dodge [Zie90q, Fai94, Zie92q]. Dodson [Car96, Zie95-48]. Does [God90]. Doing [Zie99k]. Dominique [SN95]. Don [Col93, Job99, Mag96]. Donald [Ada96, Bel99, Bro90b, Cox96, Dog93, Fre92, God93, Har91b, Mil90, Zei94, Zie91r, Zie92x, Zie97-35, Zie98n]. Donev [Whe93]. Dongsheng [Hes97]. Donna [Maz98, Zie98v]. Doran [Zie90b]. Dorothy [Moo91]. Doty [Ada93, Mor99a, Vuk98]. Dougherty [Wal92a]. Douglas [Ale95, GR97, Gil92, Gun96a, Jen92, Mat98, Mil90, Nar93, Por93, Pro92a, Sau92, Sri92, Zie99l, Zie99-42, Zie99f, Zie99u]. Dov [Weh91]. Dovich [Ben93, Rig92b]. Dowd [Zie97n]. Dowdy [Ing92]. down [Coo96b]. Downtime [Bog92]. Dr [Zie95p]. Drain [Edg98, Gri98]. Draper [Gun99, SDP^{+}95, Zie98q]. Drift [JV93]. Driven [MW98]. Drug [Zie91r, Zie99-39]. Dshalalow [Gla98]. Duane [Sen99b, Sen98]. Duane-based [Sen99b, Sen98]. Dubes [Sar90]. Dudewicz [Joh91, Law92]. Duffy [Sim91]. Dummy [Zie94-35]. Duncan [Bla92, Zie90-28]. Dunn [Gle91, Nai94]. Duran [Bzi97]. Duration [BCL98a, BCL98b, Rob98, Sch98a]. Duvall [Zie94j]. Dux [Zie93l]. Dykstra [Sho90]. Dynamic [Amm92, Bax95, Bil94, Cab98, Peñ91, Wel90, Wes92, Win93, ZP94, Zie97e, Buc90]. Dynamics [LG99, Zie93q].

Earth [Rus94, Zie90z]. Easy [Bre97, Zie95-35]. Ebeling [Hey98]. Eckhard [Hal95]. Ecological [Zie95j, Zie94q, Zie95m]. Econometric [Zie95e]. Econometrics [Zie90b, Zie94v, Zie97c]. Economic [Elt90, Ngu96b]. Economical [STW91]. Economics [Pro96, Zie97m, Zie90-27, Zie95-50].

ECR [Sen99b]. Eddy [Len93, Mar99]. Edge [HT92]. Edge-Preserving [HT92]. Edgington [Zie98-37]. Editor [Haw94, Kaf99, Mor97b, Mor98c, Nai90a, Nai91a, Nai91b, Nai92, Var94a, Var94b, Yas90]. Editorial [Mor98b]. Edler [Zie98x]. Edmund [Bla94]. Eduardo [Hul92]. Education [Pro96]. Edward [Baj98, Bil91, Bro90b, Joh91, Ked98, Law92, Lei93, Pan96, Ram91, Sco92, Wal92a, vT95]. Edwards [Mag96, Pro96, Zie90d, Zie99q]. Effect [CH96b, HO95, Jac92, Jac93, Sch96b]. Effect-Based [CH96b]. Effect-Saturated [HO95]. Effects [Ale98, Ano99e, BT89, BH97, CWWR97, Emp99, Hol98, Khu92, LBB96, LN97, LG99, Pan99, TH93, Van92b, VV99, WL97, WT98, BT98]. Efficiency [Fam99, TWN98, Sch95]. Efficient [KD97, Lin93a, Lin95a]. Efron [Gri95]. Eigenvalues [BA96]. Einax [Zie98m, Zie96j]. Elaboration [TFV93]. Elart [Tor99]. Electrical [Lei91]. Electronics [Dog93, Kaf94]. Elementary [Hei90, Joh91, Zie99-35, Far90, Hag90, Mag96, Sar94a, Wil90b]. Elements [Zie97-41, Tim95, Zie97l]. Elfe [Zie97v]. Elisa [Zie93-27, RJ97]. Elizabeth [Por93]. Ellen [Bor97a]. Ellis [Ram91, Zie99h]. Elsayed [She90b, Lyn97a]. Emanuel [Zie99-36]. Emergency [MKA96]. Emmanuel [Bor97b]. Emphasis [BH93]. Emphasising [Not92a]. Empirical [MKA96, Ruk97a, Yas92, Ang91a, Pie90a]. Encyclopedia [Zie91g, Zie90h, Zie92g, Zie96o, Zie98o, Zie98p]. End [Zie90h]. Enders [Zie95e]. Engelhardt [Bry92]. Engine [HLD98]. Engine-Mapping [HLD98]. Engineer [Car90b, Geo94]. Engineering [Ajm99, Alm95, Ben93, Bow91, Bru93b, Bru95, CFRH99, Car96, Che91, Cro99, Goi96, Gut90, Jan94, KY91, Lar95, LN96, Men91, Nar93, Not92c, Rag91, Sch92c, Sch93a, Sha97a, She90b, Spe96a, Wal92a, Wyc96, Zie91p, Zie92l, Zie93w, Zie93v, Zie93-36, Zie94-40, Ang90a, Ano97e, Fri97, Gla98, Hey92, Hey98, Kha95, Lei91, LH97, Lyn97a, Oli98, Pap96, Sen97, Tan98, Was91, Zie94-39, Zie97-39, Alt96, Bod97, Edg92, McC99, Nan94, Sac93, Weh91, Zie99l]. Engineers [Arm90, Bar91, Bax96b, Box90, Boy90, Bzi97, God90, Hah90, HK90, Hoo90, Hun90a, LeP90b, Mac90b, Moo90, Rob90, Tho99, Tri90, Wan90, Zie93c, Zie94-31, Zie99-43, Zie99p, Zie99-30, Ale95, Bru93a, Bry99, Gat96, Gut92, Lee93, Mil99, Spa99, Taa99, Zie95-28, Zie99h]. Engineers/ Physical [Box90, God90, Hah90, HK90, Hoo90, Hun90a, Mac90b, Moo90, Tri90, Wan90]. Enhancements [Bax90, FHT90, Hun90b, LS90b, LS90c, MH90, WM90]. Enrick [Zie94-33]. Enrique [Ang90a]. Ensuring [Soy95]. Entropy [KY91]. Environment [Amm92, Zie99-45, Zie99-44, Hin97, Zie94-42]. Environmental [Cam98, Gat96, Got92, Rus94, Zie90r, Zie96j, Zie98-42, Zie98s, Zie98m, Zie99f, Ham96, Zie95n, Zie96p, Zie99m]. Environments [KS93]. Envisioning [Sco92]. Ephraim [Tho99]. Epidemiological [Zie95-36]. Epidemiologists [Zie95i]. Epidemiology [Zie95-39, Zie96-43]. Equality [Sin83, Sin92]. Equation [Hec96, Zie97b]. Equivalency [CC99]. Eric [Coa94, Len90]. Ericson [KY91]. Erling [Hap92, Sau98]. Ernest [Lyn94]. Error [Arm97, CGG92, HA93, Hir96, JV93, RL95, Roe91, TW96, VM91, Zie93t, McA90]. Error-Rate [Hir96]. Errors [BA96, Car93, CS92, GS95, LBB96, Tho91, Tho92a, WI94, Zie90-28, Zie93r]. Errors-in-Variables [CS92, Tho91, Tho92a]. Erwin [Spr91]. Escobar [Nel98]. Essays [Zie91k]. Essential [Zie90-27]. Essentials [Hal98, Zie94r]. Estelle [Zie99t]. Estimate [Hul92, Kun97, LM93, Viv91]. Estimated [AKO94]. Estimates [Gri93, TW96]. Estimating [Ann99, Ano99f, Bre91b, BHS90, CDKS92, Fri91, GS92, Gri90, Gu91, HT91, JFM99, Kun93, KK99, PM99, PJ90, SW95, Yas95, Bre91a, Ste90a]. Estimation [CR93, Har91a, Hir96, Hor98, HLS98, KM94, Kim91, Kva96, Mag90, MKA96, MPR99, Mar96, MS90, Oma91, Ram93, Roe91, RWHH97, Sen98, Sen99b, Sen95, Tat97, Tho91, Tho92a, VB94b, WT98, YE94, Gru90, Kus92, Par92, San92, Spo90, Wha95, Zie94q, Zie99-49]. Estimations [McA90]. Estimator [RvD99]. Estimators [Ang91b, Ano90f, BH93, Gri97b, Sin90a, SS89, ZZ91]. Eureka [Coo96b]. Evaluate [ET96]. Evaluating [CS93, Tho94, VM91, Bor97a, Bri91, Tor95]. Evaluation [Ber96a, CR96, OM96a, OM96b, Spe96b, Swe96, Tuc96, Zie96-42, Sea91, Roz93]. Evans [Zie94-38, Zie94m, Bru93a, McG94]. Event [Goo95, Rag95, Zie98-28]. Events [BCL98a, BCL98b, LN95, Rob98, Sch98a]. Everitt [Lee99, Nai94, Zie97p, Alb93, Ko95a, Koh94]. Every [Ada91, Car90b, Car92, Geo94]. Everyone [Pro94]. Evidence [Baj98, Glo99, Zie90y, Zie96-42, Len93, Zie98b]. Evolutionary [Zie98q]. Evolving [Zie90y]. Exact [MN95]. Example [Cob91, Maz93, Zie96-36]. Examples [Cha93, Zie96x, Zie97m, Zie99g]. Excel [Zie96m, Zie97j, Zie98n, Zie99i]. Excel^{T M} [Zie99k]. Exchange [MN95]. Execution [Sch96b]. Executives [Rea93]. Exorcist [Cly94]. Expansible [Hol98]. Expected [BBT90, Cro92]. Experience [Bod97]. Experiment [CM93b, CM93a, CR98, FL95, Gun93, HO93, Hah93, LST93, RH95]. Experimental [All96, Atk96, AD96, BE94, Emp97, EM95, Gal93, Gil92, Gri98, Job99, Kia92b, Kia93a, McC94, MN95, Pro92b, Sac93, San99, Tom95, Whe93, Zie93d, Zie95-30, Ano97f, Jen91, Web90, Zie94s]. Experimentation [Pap95, STW91, Cha92, Zie92x]. Experimenter [Rut93]. Experiments [BS97, Bis97, Bow91, CH96a, CHW97, Der90, DH92, Emp93, Fed96, Gan97, Gre94, GC91, GK93, HW91, HL97, Hoc94, HLD98, Hol98, Joh90, JD96, JR99, LBB96, LLW^{+}96, MB98, MSB96a, MSB96b, Mor91, MMY93, Oli98, Pat91, Ram99, Ros96, Ros99, SC93, TH93, Tam96, Tob96, Tom95, WBS^{+}92, WC92, Zie92i, Aye98, Bor94, Bru92b, Fox94, Gre98, Gun96a, Hal95, Lyn99, Pan92, Pit96, Pri98, Rig99, Sch95, Sco91, Sri92, Zie90q, Zie91o, Zie94-37, Zie94m, Zie95j, Zie96-48, Zie99-38, vT91, Czi92]. Expert [Jon92, Sam90, Zie90g]. Explanations [Baj98]. Exploration [Ott90, Cam90]. Exploration-Geochemical [Cam90]. Exploratory [Gra90, Sch93c, Sch93b, Not92b]. Exploring [Coo93, Dun90, KR93, Spr91]. Explosives [ET96]. Exponential [Kun93, Rag98, Sin83, Sin97, Sin92]. Exponentially [Bax90, DP91, FHT90, Gan95, Hun90b, LWCR92, LS90b, LS90c, MH90, WM90]. Exponentials [WL91]. Exponentiated [MSF95].

Exposure [Zie98y]. Extended [Nel93]. Extending [Zie98r]. Extensions [GI92, Kus98]. Extreme [Ang90b, PF91, YE94, Ang90a, Zie95o]. Extrusion [Zie95-34].

F [And90, Ber91, Bil92, Boo95, Bow91, Bru92b, Cly94, Far90, Gan97, Jol96, Nem97, Nik93, O’D99, Por97b, Pro90, Run90, Sen94a, Sha95, Sha97c, Sho90, Ste98, War90, Zie90l, Zie90-27, Zie90w, Zie91a, Zie94-43, Zie95a, Zie96-30, Zie97x, Zie97q, Zie98n, Zie99l]. Fabio [Gru95]. Fabrication [HNF97]. Factor [CL93, CH96b, DM98, Hec96, Jac95, Kun97, Lyn97b, Pri93, RH95, Bla94]. Factor-Based [CH96b]. Factor-Covering [DM98]. Factor-Level [CL93]. Factor-Sparsity [Kun97]. Factorial [Ano99e, BS99, Bru92b, CL93, Emp99, FL97, Joh90, LBB96, LN97, Mor91, Pan99, RH95, TH93]. Factorials [HO95, Mil97]. Factors [EM95, KM99, Nai96]. Fahrmeir [Glo95, Zie94e]. Failure [Ano98e, DP97, GS94, Hul92, Kva96, LM93, LPY97, MSF95, Rag98, USM99, Sch96b]. Falk [Pap96]. Family [MSF95, Roc93]. Fang [Loh92, Str96]. Farebrother [Thi91]. Farnum [Che92, Ada94]. Fasser [Dog93]. Fast [RvD99]. Fatigue [Ann99, Ano99f, JFM99, KK99, PM99]. Fatigue-Limit [Ann99, Ano99f, JFM99, KK99, PM99]. Fausett [Zie95-49]. Fayyad [Zie98c]. Feature [DGTZ97a, DGTZ97b, Hul97, JS99, PE97, WH97]. Feature-Preserving [JS99]. Federer [Wen94]. Federov [Rig99, Zie90o, Zie90q]. Feedback [BK92a, BK92b, BL94, Luc98, LG99, Mac92, Tuc92, VV92, WG92]. Feigelson [Coa94, Sin98]. Fellers [Kra92, Pro93b]. Fergenbaum [Zie92w]. Feridun [Zie99-45, Zie99-44]. Fiber [WVYB97]. Fibers [Ano98e, DP97]. Fields [Bax96a, PH99, Woo91]. Fienberg [Zie90y]. Figures [Zie99-32]. Filter [Ang92]. Filtering [Har92a]. Financial [Zie99e]. Finding [Had92]. Fingleton [Cob91]. Finish [Bor97a]. Finite [CDKS92, Ruk98, Spu93, Zie93j]. Finite-Mixture [CDKS92]. Fink [Zie99-52, Zie96-47]. Finkelstein [Zie92t]. Finn [Yak97]. Fired [Ano00, SBD98]. First [Dre91, Far93, Gun94a, Kia92a, Lal99, Pri93, Ste98, Wat90, Zie95-29, Zie98a, Lin93a, Lin95a]. first-order [Lin93a, Lin95a]. Fisher [Zie94i, Weh95]. Fishman [Was97]. Fit [Ale99, GK90, Kin91]. Fits [MSH94]. Fitting [Ram94, Bru97]. Fitzgerald [Rie97]. Five [Sto92]. Flury [Haw90, Lal99]. FMEA [Sch96b]. Focus [HH96]. Focused [Zie96k]. Folks [Bas90]. Follow [Fed96, JD96, MSB96a, MSB96b, Tob96]. Follow-Up [Fed96, JD96, MSB96a, MSB96b, Tob96]. Forecasting [Boo94b, Che92, Wes92, Zie95-46, Zie97e, Elt90, Law98, Sha97b, Zie95n, Ang92, Win93, Zha92, Zie94t, Zie97m]. Foreman [Zie93-33]. Foremen [McC90]. Forensic [Zie96-42]. Forina [LaL95]. Formation [GK93]. Forms [RJ97]. Formulae [Zie90u]. Formulation [SDP^{+}95]. Forrest [Pol93]. Forthofer [Zie96v]. Fotheringham [Zie97-30]. Foundations [Zie94-45]. Four [DPL^{+}93, DPL^{+}94, Zie95p]. Fourth [Zie95q]. Fowler [Zie97q]. Fowlkes [Alt96]. Fox [Zie90-27, Zie93x, Emp93, Gra98]. Fractional [BS99, Bru92b, CL93, HO95, Mil97, SCF97]. Fractionated [HW91]. Francis [Bil92, Has91, Moo91, Zie94e, Zie95z, Zie96-40]. Frank [Ada91, Hal98, Kat98, LaL92, Zie96l]. Franklin [Bar95, Pot93, War90, Zie95-41]. Franz [Tor99]. Fred [Gun90, McL95, Oli97, Zie94-27]. Frederick [Sch93c, Sch93b]. Free [Ste90c, USM99, Wha95]. Freedback [Fal99]. Freedom [DDPG94]. Freeman [Loh94, Zie95g]. French [Sam99b, Han90]. Frequency [Bru97, Lwi94, Owe96]. Fresh [Fre92]. Freund [Alt99, Won94, Zie92n, Zie92o, Ale91]. Frieden [MA92]. Friedrich [Bor94]. Friendly [Zie93z]. Frigon [San99]. Froidevaux [Zie98s]. Front [Ano90h, Ano90i, Ano90j, Ano90k, Ano91g, Ano91h, Ano91i, Ano91j, Ano92e, Ano92f, Ano92g, Ano92h, Ano93g, Ano93h, Ano93i, Ano93j, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h, Ano96e, Ano96f, Ano96g, Ano96h, Ano97g, Ano97h, Ano97i, Ano97j, Ano98f, Ano98g, Ano98h, Ano98i, Ano99g, Ano99h, Ano99i, Ano99j]. Frontiers [Zie95-29, Zie95f, Gla98, Zie99n]. Fu [Zie95-49]. Fuchs [Mas99]. Fuller [Dic97]. Function [Goo95, MT92, RWHH97, Zie96-32]. Functional [Bla92, Far97, Jol90, Zie91l, Geo98]. Functions [Bre91b, CR93, Fri91, Gu91, Har92b, HT91, IHMT95, Bre91a]. Fundamentals [Zie95-49, Wan94, Bem96, Pit96, Rol93, Sau97, Sch93c, Sch93b, Sen95, Sys99, Zie96q, Zie99o]. Fung [HSPF92]. future [Ode90]. Fuzzy [Alm95, Bon95, Che95, Kan95, KMP95, LSBW95b, LSBW95a, Rou95, STR99, Zad95, Gut90].

G [Aye98, BR95, Ber91, Bod97, Car97, Coa94, Cob91, Col93, Die92, Gil92, Glo95, Gru90, Har91a, Hec99, Kan95, Kia95a, KY91, Lyn97b, Mac90a, Mar99, Mil99, Mil90, Nan94, Nem95, Pan92, Per97, Pfa99, Ram91, Rig92a, San98, Sar94a, Sha92, She90b, Sin98, Smi93, Ste90a, Was97, Zie90-28, Zie90i, Zie90w, Zie92f, Zie93-29, Zie93-35, Zie94e, Zie94a, Zie94v, Zie94-39, Zie94y, Zie94-36, Zie94i, Zie95-42, Zie95-46, Zie95-50, Zie96i, Zie96-40, Zie96-42, Zie96p, Zie96-46, Zie97-27, Zie97r, Zie97p, Zie97-43, Zie98-41, Zie98t, Zie98-29, Zie99-42, Zie99d, Zie99t]. Gaetano [God95]. Galambos [Zie95o, Ang90b]. Gale [Zie96y]. Galwey [Dri91]. Gamma [KGK90, Sin90a, Won92]. Gans [Ram94]. Gantz [Bro90b]. Gao [Han98]. Gap [Arm90, Zie95y]. Garcia [Job99, Lei91]. Garcia-Diaz [Job99]. Gardiner [Zha92]. Gardner [Zie94q, Per92]. Gareth [Boo94b]. Garrity [LC94]. Garrote [Bre95a]. Gary [Acr97, Kra92, Pro93b, Vin92, Wil90a]. Gatsonis [Zie96h, Zie98l]. Gauge [VV99]. Gaul [Zie90g]. Gaussian [Bas90, CWWR97, DH92, PH99, WL91]. Gauvreau [Zie94-30]. Gavin [Kus92, San92]. Geert [Ste99]. Geiss [Zie98m]. Geisser [Law95]. General [Gun90, Pet93, Zie99-51]. Generalised [Sch94b]. Generalized [Che98, CR98, Emp92, EH96, Glo95, Gri97b, Gri93, Mar96, ME99, PF91, Qiu95, TC90, ZM94, LuV90, Sch92d, dV92]. Generated [Mor97a, WL91]. Generating [Lin95b]. Generation [Zie95q, Gol90]. Generators [MKA96]. Genshiro [Zie99-36]. GENSTAT [Dri92, Dri91].

Gentle [LaL92, Zie99w]. Geochemical [Cam90]. Geodesy [Zie91m]. GeoENV1 [Zie98s]. Geoffrey [Kus98, Mil99, Ray93, Raz91]. Geoffroy [Ott90]. Geographical [Sch96c]. Geography [Sch94b]. Geologic [Zie90e]. Geological [Wat96]. Geology [Zie90m, Zie90j]. Geometric [And90, DGTZ97a, DGTZ97b, Hul97, Mil98, PE97, WH97, Zie93-29]. Geophysics [Zie91m]. Georg [Eub90]. George [Ale95, Ang91c, Arn90, Cam90, Fal99, Gho93, Liu92a, Lyn99, Por97b, Zei94, Zie93u, Zie97w, Zie98q, Zie99l, Zie99-49, vT91]. Geoscientists [Mye91, Zie99-43]. Geostatistical [Zie94b, Zie95s, Zie90i, Zie97n, Zie98t]. Geostatistics [Zie98s, Bil91, Cam96, Mye96, Zie96-44, Zie90j, Zie92h, Zie95r, Zie96r, Zie98u]. Geosystems [Zie99-46]. Gerald [Jon93]. Gergen [Wil98]. Get [Kun97]. Getting [Zie98v]. Geweke [Zie97c]. Ghosh [Ruk98, Zie94u, Zie97-34, Zie91o]. Gibbons [Zie93s, Gov92, Gro99, Zie95-37]. Giel [Zie97i]. Gifi [Goo92a]. Gilchrist [Zie94e]. Gilks [Gas97]. Gill [Kle94]. Gillies [Bla92]. Giltinan [Tim96]. Giordano [Zie90-27]. Girden [Bor97a, Zie93b]. Giri [Mat97]. GIS [Zie97-30]. Gitlow [Bre90, Zie95-31]. Gittins [Yin91]. Glasbey [Zie96t]. Glasserman [Rag95]. Glen [Gib91, Ruk98, Stu94]. Glenn [Zie95m, Zie99g]. Gleser [Dob96]. GLIM [Zie94e, Moo91]. Global [Mar99, STC99]. Glossary [Zie97o]. Gnedenko [Pap96]. Goble [Tor95]. Godfrey [Sch91a]. Goggin [Zie99-43]. Goldsman [Tam96]. Gomez [Zie98s]. Gomez-Hernandez [Zie98s]. Gonin [Spo90]. Good [Zie99a, Nik95]. Goodman [Zie99-37]. Goodness [GK90, Kin91]. Goodness-of-Fit [GK90]. Gordon [Pan96]. Gore [Bas97]. Gottfried [Far92]. Gottman [Ber92a]. Goupy [McC94]. Gouri [Boo93]. Gourieroux [Cab98, Zie99e]. Government [Pro96]. Gower [Tro97]. GPSS [Law92]. Grace [Gal92]. Graham [Cob91, Gle91, Mil98, Nai94, Zie94o]. Grant [Zie90t]. Graph [WC92]. Graph-Aided [WC92]. Graphical [FL95, Pit92, VM91, Goo91]. Graphics [Bla92, Buc90, Gra94, Hec99, Moo99, Pfa99, Zie93e, Amm92, Gol95, Moo99, Zie93z]. Graphing [Zie97-41, Kel97]. Graupe [Har92a]. Gravely [Zie99-48]. Graybill [Bar95, Ste98, Pot93]. Green [Mat98, Qiu95]. Greenfield [Aba98]. Greenstreet [Lit91]. Greenwood [Kin97]. Gregoire [Zie99t]. Gregory [Tim95, Tim99, Zie97l, Zie98c]. Grey [Zie99-28]. Griffith [Acr97]. Groenendaal [Lor93]. Groeneveld [Tho90b]. Grosh [Car90a]. Gross [Sys99]. Groundwater [Zie95-37]. Group [Hir96, ST90a]. Group-Testing [ST90a]. Grouped [HL97, SGW96]. Groups [ES94, Had92]. Groves [Sac93, Zie90-28, Zie93r, Zie99v]. Growth [EMS98, Sen98, Sen99b, Zie96s]. Gruber [Ang91b, Fam99]. Gruet [Bor97b]. Gruska [Zie96-46]. Gryna [Zie90l]. GSLIB [Zie95s]. Gudmund [Wil98]. Gui [Dea95]. Guide [Aba98, Bat99, Bay95, Bel95, Coo96b, Cro99, Hea94, Hes91, Jon93, Kin97, Lei93, Lit91, Nik95, Not92c, Pit92, San99, Wan94, Woo91, Zie90w, Zie99-52, Bre97, Don90, Dun96, Hem99, Tob99, Zie95s, Zie96-31, Zie98t, Tan93]. Gulezian [Han93]. Gunst [Bow91, Bru92b]. Gupta [Gut90]. Gurevitch [Zie95j]. Guttorp [Rus94]. Guyon [Bax96a]. Gy [Bil92, Zie95z, Bil94, Zie95z].

H [Arm93, Bas99, Bil91, Bre90, Bru93a, Bry99, Cha93, Dri92, Fam99, Far93, Fre92, Fre94, Gla98, Gru97, Gun96a, Gun94a, Hag90, Hag91, Hag96, Hof95, Hwa92, Iye98, Jai95, Kah94, Kan95, Kia90, Kia92a, McC91, Pat91, Per98, Pfa92, Pri98, Pro92b, Rob90, Sch96b, Sha95, Sha97c, Spo90, Ste90c, Ste98, Tan98, Vin92, Win90, Zie90o, Zie90q, Zie91l, Zie91a, Zie93-32, Zie94r, Zie94x, Zie94v, Zie94-33, Zie95-29, Zie95-31, Zie95-44, Zie95-43, Zie96-43, Zie96-27, Zie96-38, Zie97z, Zie98-29, Zie98m, Zie98i, Zie99f, Zie99-51, Zie99n, Zie99-30, Zie99-32]. H.-J [Jai95]. Haaland [Jen91, Kel97]. Hackl [Rig99]. Hahn [Jon93, Zie94-39]. Haimes [Liu99]. Haining [Got92]. Hald [Zie92a, Zie99b]. Hall [Sea90]. Hamilton [Zie90e, Bie92, Boo97, Dre91, Gra94]. Hamming [Gut92]. Hand [Zie95a, Zie97t, Lin92, Tro97, Wad99, Zie95f, Zie95c]. Handbook [Bak98, Dri92, Ko95a, LC98, Rie94b, Sch92c, Sch93a, Sin90b, Zie90-27, Zie95a, Zie95-44, Zie98-33, Has91, Sch94a, Zie90l, Zie91c, Zie93v, Zie93g, Zie96l, Zie99-52, Bar91, Ber97, Bru92a, Glo93, Gri98, Sen94a, Wei94, Zie90k, Zie91i, Zie93l, Zie94u, Zie94w, Zie94v, Zie97p, Zie98w, Zie99p]. Handle [Cro94]. Handling [Cam98, Nem95, Oli97]. Hands [Stu94]. Hands-On [Stu94]. Hannan [Cry90, Ord96]. Hans [Eub90, Haw90, Mye96, Pro93a]. Hans-Georg [Eub90]. Happen [Hoc94]. Hard [Liu92b]. Härdle [Hin97, Zie98-41, Eub93]. Hardy [Zie94-35, Zie97x]. Hariharan [Bar95]. Harnett [Zie98n]. Harold [Kia93a]. Harrell [Bre97]. Harris [Sys99, Zie91p, Zie93p, Zie98-48]. Harrison [Bar91, Zie99p, Wes92, Zie97e]. Harry [Cab94, Gun99, Rig92a, Zie90-27, Zie93h]. Hart [Pro90, Ale99, Pro90]. Harvey [Ang92, Ens95]. Harville [Zie98-30]. Hasofer [Ste90a]. Hasselblad [Len93]. Hastie [dV92, Hes93]. Hastings [Zie94-38]. Haswell [Bay95]. Hatcher [Zie95b, Hec96]. Hatzinger [Zie96-40]. Haugh [Zie99-37]. Havenner [Zie98f]. Havlicek [Jen90]. Hawkins [Ver90]. Hayes [Gru97, Zie95w, Zie98-31]. Health [Woo91, Zie94c]. Heaphy [Zie96-46]. Heat [Tho94]. Hecht [Rig92a]. Hedayat [Zie93j]. Hedges [Zie95-44]. Heiberger [Sco91]. Heinen [Gol98]. Helmut [Jol93]. Helsel [Got94a]. Henk [Mad96, Spr91]. Henri [Hal95]. Henrion [Zie95-47]. Henry [Sti97]. Hernandez [Zie98s]. Herzberg [Tho92b]. hesketh [Lee99]. Hess [Bow91]. Heterogeneity [Bil94]. Heterogeneous [AHT91, Bil94]. Heteroscedasticity [CS92]. Hettmansperger [Mi99]. Heuristic [Mar99]. Heyde [Bor98]. Hiatt [Zie96-51]. Hicks [Bod97, Pit96]. Hidden [JR91, Ped98]. Hierarchical [Cal94]. Higgins [Huz95]. High [Kii92, MSH94]. High-Breakdown [MSH94]. High-Dimensional [Kii92]. Higham [Zie94w]. Higher [Fre95]. Highly [HW91]. Hild [Per97]. Hildebrand [Koo90, Zie99-41]. Hildrum [Zie95y]. Hills [Zie95-39]. Hinde [Moo91]. Hines [Nar93]. Hinkelman [Tom95]. Hinkley [Zie92s, Zie93-30, Zie97-42]. Hirotugu [Zie99-36]. Hirsch [Got94a].

Histograms [Deu96]. Histories [Zie98-28]. History [Zie92a, Zie99b]. Hjorth [Len95]. Hoaglin [Sch93c, Sch93b, Cro94]. Hocking [Dea97]. Hodges [Zie96h, Zie98l]. Hogg [Zie93c]. Hohn [Zie90j]. Hollander [Cly94, Zie99u]. Homogenizing [Bil94]. Hong [Han98]. Hong-Ye [Han98]. Honor [Zie91k]. Hopke [Zie92c, Zie94k]. Hoppe [Zie94-27]. Horgan [Zie96t]. Horrell [Zie98n]. Hosmer [Cuc92, Zie94c]. Household [Zie99v]. How-To [Zie96-45]. Howard [Goo92b, Pri93, Zie90b]. Hows [Zie93-35]. Høyland [Van96b]. Hsiang [She90b]. Hsu [Chr97]. Hubele [Run90, Zie99l]. Huberty [Hec95]. Huet [Bor97b]. Hunt [Rag93]. Hutchinson [Not92a, Not92c]. Hy [Hal94]. Hydrology [Por97a]. Hypercube [PH99, Ste87, Ste90b]. Hypotheses [KGK90, Len90, Nik95]. Hypothesis [Hor98, VS96]. Hypothesis-Testing [VS96].

Iain [Ped98]. IBM [Cam90, Dun90]. Ideas [Moo99, Rag93, Zie99a]. Identification [Har92a, Pan99, STR99, Taa91]. Identifying [VS96, VS98]. Idle [GD91]. Iglewicz [Cro94]. Ignition [Tho94]. Igor [Gri99, Pap96, Rie94a]. II [HH96, Kia93b, Sin92, Zie94-45, Sin83, Bru97, Zie90v, Zie94k, Zie96h]. III [DL90a, Zie98l, Zie98k]. Ildiko [Zie96l]. Illustrated [Nem95]. Ilya [Mi97]. Image [Zie96t]. Images [Baj98]. Iman [Gri96]. Immunoassay [RJ97]. Impact [Pan99, Zie95c]. Implications [CH96b, Zie90p]. Importance [Hes95]. Important [WC92]. Improve [Zie99a]. Improved [CR93, KM94, Mee90a, MME98]. Improvement [AHT91, Bel95, Bre90, Bre96, Dog93, Dun96, Fox94, Gan97, Gun96b, Pan96, SG99, Zie94l, Zie97a, Zie97w, Zie98-35, Aba98, Ada94, Bel99, Bem96, Gho94, Gri98, Hal98, Lam90, LC94, McG94, Nem92, Pro90, Woo92, Zie93i, Zie93-35, Zie94h, Zie95-31, Zie98q, Zie98-40, dC99]. Improving [Cha92, Fam99, San98, Zie97q]. In-Process [Pla90]. Inc. [Zie99-50]. Incomplete [SS97]. Incorporating [LT96]. Increasing [CS93, GS94, Sne98]. Independent [STC99]. Index [Zie90h]. Indices [Cha95, Yin91]. Indirect [TW96]. Industrial [Bod97, Bog92, CH96a, CM93b, CM93a, Edg98, Fre99, Gun93, HO93, Hah93, LST93, Pan92, Rol93, Zie91o, Zie92x, Zie95t, Zie96-48, Zie98-40, Zie99o, Die92, Zie98x]. Industries [Arm93, Lew92, Zie94r, Zie97-31, Zie99-33]. Industry [HH98a, HH98b, Kaf94, KTTS90, MTV98, Nem98, Pan96, Pro96, Zie91h, Dog93, Jen92]. Inference [Boo98, CWWR97, FRY96, HO95, HP97, HL97, Koe92, Law95, Lei99, LPY97, Zie93j, Ang91c, Bas99, Cha98, Das93, Gun94a, Ruk97b, Sha95, Sha98, Sho90, Ste98, Zie92q, Zie93s, Zie94j]. Inferences [Won92]. Inflated [Lam92, LLP^{+}99]. Influence [Peñ91, Sch91d, TC90, Ter92]. Influential [SD90, STR99]. Informatics [Zie90f]. Information [Ano90l, Ano91k, Ano92i, Ano93k, Ano94i, Ano95i, Ano96i, Ano97k, Ano98j, Ano99k, JS99, Koh90, SS97, Sri93, Sco92]. Informational [Zie95-29]. Infra [Zie95y]. Infra-Red [Zie95y]. Ingram [Dob96]. Inhomogeneity [Cre97]. Inputs [GSH98]. Inspection [DGTZ97a, DGTZ97b, God95, GS95, HR92, Hul97, PE97, VV94, WH97, HR94, Car93]. Institute [Zie96x, Zie99r]. Instructional [Kal93]. Integral [Li97, TWN98]. Integrated [HNF97, TH93, Tho90b, Van96a]. Integrated-Circuit [TH93]. Integrating [HP92]. Integration [CFRH99]. Intelligence [Zie95f, Zie95-49]. Intensity [CS93]. Intensive [Bro90b, Len90, Len95, Zie97z]. Interaction [Ale98, Ano99e, Ber92a, FL97, Emp99]. Interactions [Oeh94, WC92]. Interactive [Hin97, Bla92, Zin93]. Interest [Zie97-31]. Interface [Bro90b]. Intermediate [Bal93]. Interpolation [VD94]. Interpretation [Ram91, Wen94]. Interpreting [Edg92, Lee90, Zie96u]. Intersection [Rou95]. Interval [BL94, Spu93, VB94b]. Intervals [BBT90, CR91, DS93, Hea90, Jon93, LS90a, Mee90b, MER91, Pet93, Pot93, RAA90a, RAA90b, Sch92b, WI94, Woo90, DSSU98, Ode90]. Interview [Zie99v]. Introducing [Sch94b]. Introduction [Ber92a, Bil91, Bro92, Dav99, Emp92, Fre94, Gol95, Gri95, Had92, Ked98, Ko95b, Liu92b, Mor92, Pro92b, Rig92a, Sau94, Sch96a, Tom95, Yak97, Zie94f, Zie95-50, Bax96b, Ber92b, Che94, Dic91, LaL92, Law95, Len90, Tso92, Zie98j, Zie99-42, Arn90, Bab98, Bos92, Cam96, Dic97, Dun97, Far92, GR97, Gil92, Hor98, Jol93, Jon91a, Lyn99, Por93, Rig93, Sau92, Sau98, Sea90, Sen94b, Sen97, Sha97b, Taa99, Wol92, Zei94, Zie94x, Zie94y, Zie95t, Zie96w, Zie96v, Zie98y]. Introductory [Zie95-50, Bul96, Lin99, Tho90b, Zie93m]. Inverse [Bas90, DH92, MS95, WL91]. Investigation [Gib91]. Investigations [Mee91a]. Involving [Zie96n]. Iorio [Zie97x]. Irradiation [GK93]. Irwin [Wyc96]. Isaac [Lei97]. Isaaks [Bil91]. Isaksson [Zie95y]. Isham [Fre93]. Ishikawa [Bos92]. ISO [Zie93n]. Isolating [Oeh94]. Isolation [DDPG94]. Issue [Zie91p]. Issues [DGTZ97a, DGTZ97b, Hul97, KTTS90, PE97, Sri93, WH97, Zie91r, Zie99-39, Pig91]. Iteratively [Mar96]. Iversen [Wil98]. Iwasiewicz [Mor93b]. Iyer [Ste98, Bar95].

J [Ada96, Alt99, Ang91a, Arm97, Bar92a, Bas90, Bas97, Bel95, Bel99, Ber91, Boo95, Boo92, Boy90, Bro90b, Bry92, Bul96, Cab94, Cal98, Cob91, Cra94, Cro99, Cry90, Deh90, Dob96, Dri92, Dun97, Emp92, Fai91, Fam99, Fox91, Fri97, Ful95, Gas97, Gil92, Gle96, God93, Goo91, Gov92, Gre94, Gri95, Gun94a, Had92, Han97, Har91a, Har91d, Hec95, Hes91, Hes93, Huz95, Jai95, Joh91, Jon93, Ked96, Kia90, Kia92a, Kia92b, Kia93b, Kia94, Kin91, KY91, Kot92, Kus92, Kus98, Law92, LeP90b, Lei92, Lei93, Len95, Lin92, Lyn95, Mar99, Mat98, Mi99, Mil98, Mor97a, Mul97, Nem92, Nem97, Neu99, Nik93, Oli98, Ord96, Ott90, Owe96, Pit92, Pro93a, Qiu95, Rag93, Ram99, Ray93, Raz91, Rie97, Run90, Sam99b, San92, Sch92a, Sch92d, Sco90]. J [Sha95, Sha97c, Sha97b, Sha98, Sim91, Sin98, Spa99, Ste91, Tam96, Ter92, Tob99, Tro97, Tsa91, Tso92, Wad99, War90, Wei90, Wes92, Wol96, Wol92, Won94, Yan97, Zha92, Zie90l, Zie90-28, Zie90-27, Zie90n, Zie90w, Zie91m, Zie91q, Zie92x, Zie92r, Zie92s, Zie93-30, Zie93c, Zie93t, Zie93o, Zie93l, Zie93g, Zie93s, Zie93n, Zie94r, Zie94-28, Zie94-32, Zie94p, Zie94c, Zie94m, Zie95o, Zie95-38, Zie95j, Zie95-28, Zie95-50, Zie95c, Zie96o, Zie96-37, Zie96-44, Zie96j, Zie96h, Zie96a, Zie96-51, Zie96-52, Zie96-39, Zie97e, Zie97v, Zie97c, Zie97-44, Zie98s, Zie98m, Zie98-28, Zie98w, Zie99k, Zie99-43, Zie99t, Zie99n, dV92]. Jaccard [Ale98, Emp99]. Jacek [Gho94]. Jack [Bar92b, Lor93, Zie99q]. Jackknife [Hes97]. Jackknifing [Cre97]. Jackson [Nai96, Lei93]. Jacoby [Hec99, Pfa99]. Jacqueline [Zie99-52]. Jacques [Cam96, McC94]. Jaffe [Hec90, Zie95u]. Jagdish [Ber97, Fot95]. Jain [Sar90]. Jambu [Not92b]. James [Ale98, Bal93, Boo97, Bow91, Che98, Dea96, Emp99, Gho94, Gue90, Jol97, McG94, Pap96, Par92, Pie90a, Sen94a, She90a, Wat96, Wel90, Zie93q, Zie96n, Zie98n, Zie98l, Zie98i, Zie99w]. Jan [Bay92, Glo97, Jol90, Kel97, Sch91b]. Jan-Aage [Kel97]. Jan-Bernd [Bay92]. Janacek [Boo94b]. Janos [Ang90b]. Janssen [Zie96a]. Jason [Chr97]. Jay [Hec90, Zie92l, Zie95u]. Jaya [Zie90s]. Jean [Gro99, Zie96-38]. Jeff [Dow95, Ord96]. Jeffrey [Ale99, Che94, Zie98h]. Jenkins [Zie95-46, GJMS^{+}95]. Jennrich [Sch96a]. Jensen [Zie94-28, Zie99-43, Tor99, Yak97]. Jerry [Goo92b, Mil91, Zie99-43]. Jersy [Zie99k]. Jessica [Fam97]. Jewell [Zie97r]. Jewgeni [Gla98]. Jill [Zie99-29]. Jim [Geo98]. Jiri [LaL95]. JMP [Zie99r]. Jobe [Spa99]. Jobson [Kia92b, Kia93b]. Joe [Iye98]. Joel [Zie99-29]. Jogesh [Coa94]. John [Bat99, Bay95, Ber92a, Bro90b, Bul91, Car95, Che91, Czi92, Dav99, Dea91, Dom90, Dov93, Gol90, Gra98, Gro96, Hag96, Hes93, Kin93, Mag96, Moo91, Mor97a, Owe91, Pro92a, Pro94, Rao98, Roz93, Yin91, Zie91f, Zie91e, Zie92b, Zie92i, Zie93x, Zie95v, Zie95-43, Zie96-35, Zie97y, Zie98-42, Zie99d, Zie99r, vN90, Bru93b, SDP^{+}95, Zie99-38]. Johnson [Ber96b, Boo93, Bru94, Car93, Fam95, Lor99, Pos98, Zha92, Zie90h, Zie92g, Zie94-45, Zie94-44, Zie98e, Zie98z, Zie99c, Cha95, PGT95, Wil90b, Zie90e, Zie94-45, Zie98k, vT91]. Joiner [Zie95q]. Joint [Gan95, NLB^{+}98, USM99, WT98]. Jolivet [Bor97b]. Jones [Aba98, Kim91, Mor92, Zie90e, Zie91q, Yan96, Zie94z]. Jordan [Oli98]. Jöreskog [Lyn97b]. Jorgensen [Bor95b, Sen99a]. Jorner [Kel97]. Joseph [Ada92, Pro92b, Rie97, Zie92j, Zie98b]. Journel [Zie95s, Zie98t]. Journey [Zie93a]. Judd [Vin92]. Julius [Nan94, Taa91]. Jump [QY98]. Jump-Detection [QY98]. Jun [Hes97]. Juran [Zie90l, Zie90l]. Jurand [Mor93b]. Jure [Wol91]. Jurgen [Sim97, Tor99]. Just [Kal93, Zie99a]. Just-In-Time [Kal93].

K.-T [Str96]. Kadane [Zie98b]. Kahle [Tor99]. Kai [Loh92]. Kai-Tai [Loh92]. Kai-Wang [Loh92]. Kaiser [Sen94a]. Kalivas [Hag96]. Kalman [Ang92]. Kalsbeek [Zie93t]. Kalton [Zie90-28]. Kaluzny [Zie98-38]. Kalvelagen [Spr91]. Kam [Lor99]. Kam-Wah [Lor99]. Kaminsky [Zie94-43]. Kane [Bru91]. Kanji [Zie94a]. Kao [Ked98]. Kaoru [Bos92]. Kaplan [Zie98-45]. Kapur [Zie90n]. Karen [Zie99a]. Karian [Law92]. Kariya [Roh91]. Karl [Zie91r]. Karlin [Fre94]. Karprzyk [Zie90-28]. Kass [Zie96h, Zie98l]. Kateman [BR95]. Kathleen [Rab92, Wel93]. Kaufman [Gut90, Had92, Mac90a]. Kay [Sen95]. Kaye [Raz91, Zie99-32]. Kazense [Zie94-44]. Kazmierski [Fri97]. Keats [Run90, Zie92r, Zie96-37].

Kececioglu [Sch92c, Sch93a, Sch94a]. Kedem [Fre95]. Keenan [Dov93]. Keiding [Kle94]. Keith [Boo92, Zie90r]. Keki [Hoc94]. Keller [Huz95]. Keller-McNulty [Huz95]. Kelly [Zie94g, Zie96-30]. Kelton [Dea94]. Kemp [Bru94]. Kempthorne [Tom95]. Kendall [Boo92, Gov92, Zie94-28]. Kenett [Fre99, Mas99]. Kenneth [Bru96, Kan95, Sto96, Tob99, Zie99i]. Kenward [Zie91q]. Keppel [Zie92f]. Kerim [Bre97]. Kernel [Yan96]. Kessler [Zie97s]. Kewley [Zie99q]. Key [HH98a, HH98b, MTV98, Nem98, Voe94]. Kharin [Boo99]. Khattree [Zie96d]. Khosrow [Pie90b]. Khuri [Zie97y, Zie94d]. Khurshid [Zie96-43]. Kidd [Zie93u]. Kiemele [Nem92, Zie94h]. Kim [Zel94, Sam99a]. Kimber [Bar92a, Zie97r]. Kinetics [ML95]. King [Zie95-32, Zie97-32]. Kingman [Boo95]. Kit [Zie96-47]. Kitagawa [Zie99-36]. Kitsos [Zie98x]. Klaas [Bar92b]. Klaassen [Bar92b]. Klar [Zie94c]. Klaus [Tom95]. Kleijnen [Lor93]. Klein [Rao98, Arm93]. Kleinbaum [Car97, Kia95a]. Klinke [Hin97]. Kloeden [Hal95]. Klugman [Zie98-29]. Knezevic [Lyn95]. Know [Car90b, Car92, Geo94]. Knowledge [Zie90g, Zie98c]. Koch [Cam90, Zie96i]. Kocherlakota [Wel93, Wel93]. Koenig [Zie93f]. Kokoska [Zie90u]. Kolmogorov [Rag93]. Kolosova [Zie99-31]. Koopmans [Per98]. Koronacki [Gho94]. Kosecoff [Zie99-52]. Kott [Zie96g]. Kotz [Ber96b, Bru94, Car93, Cha95, Fam95, Loh92, Pos98, Zie90h, Zie92g, Zie94-45, Zie98o, Zie98p, Zie98k, Zie98z]. Kovalenko [Gri99]. Kowaka [Zie95t]. Kragh [Kle94]. Kramer [Bay99]. Krause [Zie98-46]. Krewski [Zie95-41]. Kriging [Cam96, HS93, ZZ91]. Krishnaiah [Sin90b, Zie90k, Zie91k]. Krishnamoorthi [Lee93, McC90]. Krishnan [Kus98]. Krumm [Zie93-35]. Kruskal [Zie90w]. Krzanowski [Sha95, Sha97c, Lin90, Wol96]. Kshirsagar [Zie96s]. Kubacek [Zie91m]. Kubackova [Zie91m]. Kukuca [Zie91m]. Kunio [Zie99-36]. Kuo [Sam99a, Zie93w]. Kurt [Bor97c]. Kurtz [McA90]. Kutner [Zie97-44, Kia90, Zie92b]. Kuznetsoy [Gri99].

L [Ajm99, Ang91c, Arm97, Aye98, BR95, Bar92b, Bar92a, Bax95, Ber96b, Bog92, Bow91, Bri96, Bro92, Bru92b, Bru94, Bul91, Bzi97, Car90a, Car93, Cha95, Dom90, Fai91, Fam95, Glo95, Gre94, Gri96, Had92, Ize91, Jen90, Jol90, Jon91a, Kot92, Lit91, Mac90a, Mar99, McC94, Nay98, Ped98, Pos98, Pri93, Rao98, San99, Sho90, Sin90a, Ste90c, Wat90, Zha92, Zie90-28, Zie91m, Zie91r, Zie91c, Zie92l, Zie93r, Zie93-35, Zie94e, Zie94-28, Zie94-44, Zie94s, Zie94-36, Zie94i, Zie95-49, Zie95-32, Zie95b, Zie95-44, Zie96-32, Zie98n, Zie98-43, Zie98x, Zie98e, Zie98k, Zie98z, Zie98w, Zie99z, Zie99-49, Zie99-43, Zie99-52, Zie99-30]. Lab [Spr91]. Laboratories [Zie94g]. Laboratory [Mag96, Smi93, Zie93l]. Labys [Dow95, Ord96]. Lack [Ale99]. Lack-of-Fit [Ale99]. Lad [Kat98]. Lahey [Zie97v]. Lai [Not92a, Not92c]. Lake [CDKS92, Zie99-43]. Lake-Chemistry [CDKS92]. Lambert [Per98]. Lamprecht [Zie93n]. Land [Coh92]. Lane [Dri91]. Lang [Hag96]. Language [Bat99, Sco90, Zie98h]. Lao [Zie96u]. Lapin [Ajm99, Was91]. Large [CCC90, FL95, Ste87, Ste90b, Par92]. Large-Sample [CCC90]. Larry [Gru90, Hec96, Jen93, Sar94b, Sch94b, Zie99-43, Zie99-37]. Larsen [Kot92]. Latent [Gol98, Bay92, Gol98, Zie93o]. Latin [Ste90b, Mor98a, PH99, Ste87]. Lattice [Boy96]. Latzko [Zie95p]. Launsby [Aye98, Pan92, Zie96-48, Zie97v]. Lauter [Zie90o]. Lave [Zie91n]. LaVerne [Che92]. Law [Bak96, Dea94]. Lawrence [Ajm99, Bie92, Dre91, Fai91, Gra94, Law96, Rol93, Was91, Zie90r, Zie99f, Zie99o]. Lawson [Cab94]. Lawyers [Zie92t]. Layouts [Car95]. Leader [Zie93a]. Leadership [Zie98-48]. Leading [Zie98z]. Learning [Bri96, Car97, Ful95, Kia95a, Sai96, Wu99, Ale91]. Least [Bay92, Mar96, Thi91, Ode90]. Leavenworth [Zie90t]. Lechner [Zie95o]. Lecture [Har91c, Ste99]. Ledolter [Zie93c]. Lee [Bry92, Zie97r, Ber92b, Lei92, Pri93, Zie93-27, Zie96v, Zie98j]. Leemis [Law96]. Leeuw [Zie98-43, Jol90]. Lehman [Zie90w, Zie99r]. Lehmann [Zie99-49]. Lehtonen [Zie96-29]. Leitch [Bax96b]. Leitnaker [Per97]. Lemeshaw [Zie92m]. Lemeshow [Zie94c, Cuc92]. Length [AWS94, Fel94, Luc94, WMP94a, WMP94b]. Lentner [Web90]. Lenz [Zie99n]. Leon [Dob96, Lei91]. Leon-Garcia [Lei91]. Leonard [Ada93, Mor99a, Vuk98]. LePage [KR93]. Lepingle [SN95]. Leroy [Bas90]. Leslie [Sch95]. Lessi [Zie98g]. Lessler [Zie93t]. Lester [Zie91n, Zie94-33]. Letchford [Zie95-40]. Letkowski [Zie99k]. Letter [Haw94, Yas90]. Level [BS99, Bla92, CL93, DL90a, FL97, HCV98, LIV96, Sch91c, WC92]. Levels [Ngu96b, WW92, Gri90]. Levin [Zie92t]. Levine [Wat90, Zie96-52]. Levinson [Hal98]. Levy [Zie92m, Wal92b]. Lewi [Zie98w]. Lewis [LeP90b, Bre95b, SDP^{+}95]. Li [LW98]. Liang [Zie95d]. Library [Zie95s, Zie98t]. Life [Bry92, DS93, DH92, EM95, EM99, Hul92, Kim91, KS93, ML95, MM94, Sam90, Sch91c, Sch94a, Spe91, VB94b, Zie95t, Zie96-34]. Life-Data [DS93]. Life-Testing [Bry92]. Lifetime [HLS98, Zie97r]. Likelihood [Glo99, Gri93, Har91c, Sin83, Sin92, Lei99, Bor98]. Limit [Ann99, Ano99f, BL94, JFM99, KK99, PM99]. Limited [Cal98]. Limits [AKO94, BT89, CM99, DM99a, DM99b, Gib99, Gil99, KR93, MT91, Nel95, Van92b, Van99, BT98]. Lin [WKWL95]. Lind [Zie99-42]. Linda [Zie90x]. Lindgren [Har91b, Zie97-35]. Lindley [Fox91, Sam90]. Lindman [Kia93a]. Lindsey [Bul96, Che98, Kia94, Owe96, Sha98]. Line [KD97, NW93]. Linear [Ano90f, Bor95b, Cal94, Cam96, Che98, Coo96a, Cry90, Dea95, Dea97, Emp92, Emp93, EH96, Gil92, Glo95, Gra98, Hir96, Kia90, Kia92a, LPY97, Mar96, ME99, PF91, Por93, Qiu95, RN91, SD90, Sch92b, Sch92d, Sch94b, SS89, TI99, TC90, Zie92b, Zie92n, Zie97-44, Zie99w, Zie99-51, Bor97c, Bro90a, Dea96, Hab92, Nip92, Ste99, Thi91, Zie91j, Zie98-27, Zie98-28, Zie99s]. Linear-Calibration [SS89]. Linfield [Gat96]. Link [Wol95, Zie90w]. Linking [OUH93]. Lisp [Amm92]. Lisp-Stat [Amm92]. LISREL [Ale98]. Littell [Hei90, Zie92n, Zie97-27, Zie92o, Zie99-35]. Liu [Zie94p, Zie96-39]. Lloyd [Cha92, Fox94]. Loading [SW95]. Loads [USM99]. Local [QY98, RWHH97, Sch91d]. Local-Influence [Sch91d]. Location [HC91, Pan99, Sin83, WT98, Sin92]. Lochnar [Pro92b]. Lock [Wha95].

Locks [Pla96]. Loehlin [Zie93o]. Log [RN91, Hab92, Zie98-27, Zie98-28]. Log-Linear [RN91, Hab92, Zie98-27, Zie98-28]. Logic [Alm95, Bon95, Zad95]. Logistic [Cuc92, Glo93, Zie96c, Zie98-27, Kia95a, Kia95b, Zie96x]. Logothetis [Hwa92]. Lohmöller [Bay92]. Long [Glo97, Zie98-44, Cal98]. Long-Memory [Glo97]. Long-Term [Zie98-44]. Longford [Gal95]. Longitudinal [Eub90, Zie95d, Zie97t, Zie99t, Zie94z]. Look [DP90, Hab90, HW90b, HW90a, Hir90, Nai90b, YDB90, Lin93a, Lin95a]. Looking [SC93]. Lorenzen [Gre94, MLV92]. Lori [Mag96]. Loss [Har92b, MT92, Zie98-29]. Lot [Bru96, MB98]. Lothar [LC98]. Lotus [Kus93, Sar93]. Louis [Kal93, Mor92, Ruk97a]. Louise [Boo94b]. Lowry [Ara92]. Luanga [Zie94c]. Luceño [Fal99]. Luftig [Oli98]. Luis [Nel98]. Luke [Amm92]. Lunn [Zie95a, Zin93]. Lunneborg [All96]. Lütkepohl [Jol93, Zie94x]. Lwin [Ang91a]. Lyberg [Zie90-28, Zie93r, Zie98-43]. Lyday [Bri91]. Lyman [Arm92, Koo90, Zie94f, Zie99-41]. Lynch [Zie94l]. Lynn [Gil92, KR93].

M [Alt96, Bab98, Bak98, Bar91, Bar92a, Bat99, Ber92a, Ber92b, Bil94, Bru93b, Bry99, Bzi97, Cly94, Coh92, Cox93, Dea94, Elt90, Emp93, Fam97, Fox91, Fre94, Geo94, Got94a, Gov92, Gut90, Hag96, Han97, Hes93, Hof95, Jan94, Kat91, Kia90, Kus93, Law96, LC94, Len93, LH97, Lin92, Lor92, Mac90a, Mad98, McL95, Mil90, O’D97, Per97, Pit92, Rag93, Sac93, Sar93, Sch91a, Sco90, Sea93, Sen95, Soy95, Ste90a, Sys99, Tam96, Tim96, Tor95, Ver90, Vin92, War90, Wat90, Weh93, Wes92, Yan96, Zie90l, Zie90-28, Zie90-27, Zie90i, Zie90g, Zie91k, Zie92t, Zie92h, Zie92p, Zie93w, Zie93z, Zie94h, Zie94-30, Zie94-38, Zie94m, Zie94-35, Zie95t, Zie95-47, Zie95-39, Zie96b, Zie96i, Zie96-52, Zie96g, Zie96s, Zie96-46, Zie96-49, Zie97r, Zie97-28, Zie97e, Zie97t, Zie97-37]. M [Zie97-44, Zie97n, Zie98-41, Zie98j, Zie98v, Zie98c, Zie98f, Zie98-43, Zie98g, Zie99k, Zie99-43, Zie99v, Zie99p, Zie99y]. Mac [Gat96]. MacDonald [Ped98]. Machin [Zie97-37]. Machine [Ful95, Sto92]. Machine/Process [Sto92]. Machines [DGTZ97a, DGTZ97b, Hul97, PE97, WH97]. Macros [Wit94]. Madansky [Eas90]. Maddala [Zie94v]. Maddox [Zie90-27]. Made [Bre97, Zie95-35]. Madhav [Men91]. Magdalena [Cha98]. Mager [Bur91, Yeh93]. Main [Jac92, Jac93]. Main-Effect [Jac92, Jac93]. Maintainability [Hey98, JD92, Lyn95, Pla96]. Maintenance [Zie96-35]. Make [Hoc94]. Making [Gru95, Liu92a, Liu92b]. Mal [Woo92]. Malay [Ruk98]. Malcolm [Zie96-46]. Malinowski [Bla94]. Maller [Zie98-44]. Mallios [Pig91]. Mallows [Zie91f, Zie91e, vdVtBM97]. Malone [Zie90x]. Mamzic [Zie96-41]. Management [Gut90, HH96, Kal98, Mar94, Nar93, Zie91n, Zie95q, Zie95-31, Zie97k, Zie99z, Beg90, Dog99, Dun97, Liu99, Wyc96, Zie93-37, Zie93-36, Zie97i, Zie99-28]. Manager [Bre97, Car92, Coo96b, Lit91]. Managerial [Zie97f]. Managers [Zie96-52, Ada92, Koo90, Rea93, Zie90-27, Zie93h, Zie99-41]. Managing [Voe94, Zie97s, Rea93, Zie93p, Zie93q, Zie96z, Zie96y]. Mancuso [Zie92j]. Mandel [Car95, Roz93]. Manfred [Cry90, Lin99]. Manly [Nem97, Nik93, Zie98-36]. Mann [Zie93m, Zie95-50]. Manton [Kan95]. Mantz [Ang91a]. Manual [Mag96]. Manufacturing [Fot96, Kir93, Lam92, Sam99a, Sto92, TH93, Bru92a, Che91, Kal93, McL95, Mor99a, Pla90, Pol93, Run90, Won98, Zie92r, O’D99]. Many [Oma91, vdVtBM97]. Map [HA93, HNF97]. Mapping [HLD98]. Marazzi [IHMT95]. Marchal [Zie99-42]. Marcus [Spa99]. Marden [Gro96]. Marie [Bor97b, Tim96, Zie99q]. Marie-Anne [Bor97b]. Marilyn [Pro90, Zie99k]. Mario [Far90, Pap95, Sto92]. Mark [Goo92b, Nem92]. Market [Zie96w, Zie97-36, Zie98v]. Markov [Dea91, Ped98, JR91, Yas92, Bax95, Gas97, Hof95]. Markov-Type [Yas92]. Marr [Ano93f, FMQ92]. Marriott [Sha95, Sha97c, Zie91a]. Marsal [Mye91]. Martens [Hag91]. Martin [Bax95, Mee91a, Ruk97b]. Marubini [Zie96b]. Marvin [Ang91b, Fam99, Van96b, Web90]. Marx [Kot92]. Mary [Gil92, Tob99, Wil98]. Mary-Lynn [Gil92]. Marylees [Wyc96]. Marylyn [Buc90]. Masanao [Zie98f]. Masking [FRY96]. Mason [Bow91, Zie99-42, Bru92b]. Massart [Mac90a, Zie92c, Zie94k, Zie98w]. Massey [Zie90-28]. Mastering [Hec90, Zie92j, Zie95u]. Matar [Pro92b]. Material [Beg90, Bil94]. Materials [USM99, Zie95t]. Mathematical [Gut90, RM95, Zie91f, Zie91e, Zie94w, Zie99b, Zie90-27, Gri99, Hag96, Hig98, Mat98, War90, Zie90m, Zie90n, Zie95v]. Mathematics [Smi93, Zie91c, Rig92a]. Matheron [Ste90a, Zie90i]. Mathew [BZM96]. Mathews [San99]. Mathiowetz [Zie93r]. Matrix [Zie98-30]. Matter [Ano90a, Ano90b, Ano90c, Ano90d, Ano90h, Ano90i, Ano90j, Ano90k, Ano91a, Ano91b, Ano91c, Ano91d, Ano91g, Ano91h, Ano91i, Ano91j, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93a, Ano93b, Ano93c, Ano93d, Ano93g, Ano93h, Ano93i, Ano93j, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97g, Ano97h, Ano97i, Ano97j, Ano98a, Ano98b, Ano98c, Ano98d, Ano98f, Ano98g, Ano98h, Ano98i, Ano99a, Ano99b, Ano99c, Ano99d, Ano99g, Ano99h, Ano99i, Ano99j]. Matusita [Zie90v]. Mauch [Mil94]. Maura [Zie98v]. Maurice [Boo92]. Max [Bry92, Kan95]. Maximal [Zha94]. Maximum [Gri93, KY91]. Maxwell [Pat91]. McBride [Mor92]. McCabe [Arn90, Zie94y]. McClave [Zie95-28]. McClelland [Vin92]. McConway [Zie95a]. McCuen [Bry99]. McCullagh [Sch92d]. McCulloch [Zie98l, Gho93]. McElrath [Bod97]. McGill [Buc90]. McKean [Mi99]. McLachan [Raz91]. McLachlan [Kus98, Ray93]. MCMC [Zie99j]. McNeese [Arm93]. McNeil [Zin93]. McNulty [Huz95]. McPherson [Gib91]. McReynolds [Zie95-35]. MDF [DPL^{+}94]. Mead [Der90]. Mean [BBT90, Bor97c, Gan95, NLB^{+}98, VM91, Yas95]. Means [CC99, Nel93, Tsa91, WN97]. Mears [Bre96]. Measure [JD92]. Measured [SW95]. Measurement [Bri91, BHS90, CGG92, RL95, San98, Sch94b, TI99, Tim96, WI94, Zie97i, Zie98-43, LeP90a, Arm97, Gle96, Kah94, Sea91, Sea94, Zie93r]. Measurement-Error-Model [CGG92]. Measurement-Methods [TI99]. Measurements [Oma91, TW96, WL97, Kia94, Roz93, Sha97c]. Measures [Li97, LN96, LM93, Lin92, Zie93b]. Measuring [CB93, DGTZ97a, DGTZ97b, Hul97, Peñ91, PE97, WH97, McC98, Zie95w, Zie97s, Zie98-31]. Mechanics [Zie94-44]. Mechanized [Van91]. Medhi [Ked96]. Medical [Zie93-28, Zie95x, Zie98r]. Meeden [Ruk98]. Meeker [Nel98, Jon93]. Meier [Hil95]. Meloun [LaL95]. Melvin [Rao98, Zie98-46]. Memory [Glo97]. Menard [Zie96c]. Mendenhall [Zie94-40, Arm92, Lee94]. Menzefricke [Zie96-52]. Messy [vT91]. Meta [Len93]. Meta-Analysis [Len93]. Metcalfe [Lar95]. Meteorology [CC95, Zie98g]. Method [Bre91b, Fri91, GD91, Gu91, HT91, Len93, STC99, WC92, Zie98q, Bre91a, Goo95, Pie90b]. Methodology [Gun96a, LeP90b, SE93, Sto92, Zie91r, Zie94-45, Zie90-28, Pro92a, Sea93]. Methods [Alt96, Arm93, Bar91, Beg90, Ber96a, Boo94a, Bre90, Bro90b, Bru92a, Bru93b, Bzi97, CHB90, Cha93, Che91, Chi91, Coh92, CR96, Coo96b, CDKS92, Edg98, Fai94, Got94a, Gri97a, Gri98, Gro99, Hab90, HW90b, HW90a, HL95, Han97, Hes91, Hil95, Hir90, HS97, Jac95, Jai95, Kah94, KLR91, KMP95, Kat91, KY91, KM99, Lam90, LSBW95b, LSBW95a, LN95, LeP90a, Lee93, Len90, Len95, Mil99, Mil98, Mor93b, Nai90b, Nel98, Nem97, Nik93, Nik95, OM96a, OM96b, Pol93, Pro90, Pro92b, Ram96, Rie97, Ruk97a, Ruk98, Sch94b, Sha93, STW91, Sin97, Spa99, SN95, Spe91, Spe96b, Str96, Stu94, Swe96, TI99, Tho90b, Tuc96, Van92b, Wal92b, War90, Wyc96, YDB90, Yan97, Zie90z, Zie91p, Zie92d, Zie92m, Zie93-28, Zie93-29, Zie93-27, Zie94f, Zie95-31, Zie95-38]. Methods [Zie95-37, Zie96e, Zie96-27, Zie96-29, Zie98-36, Zie99p, Zie99g, dC99, Acr97, Ang91a, Bla92, Boo93, Cal94, Che92, Chr97, Chu91, Gas94, Gov92, Gra98, Har91b, Jon91b, Kat98, Kia93b, Law96, Lor99, Mi99, Min93, Osb97, Rig92a, Rig93, Van96b, Won94, Zie91l, Zie92v, Zie95-27, Zie96j, Zie96g, Zie97-35, Zie97z, Zie98-35, Zie99u, Dea97, Mat96, McC94]. Metric [CC95]. Metterick [Pan96]. Meyers [Zie99-30, Zie99-30]. Michael [Beg90, Bel95, Jol96, Ord96, Rag91, Wha95, Zie90j, Zie91q, Zie92b, Zie96m, Zie98-48]. Michel [Dow95, Not92b, Ord96]. Michele [LaL95, Rie94a]. Michie [Ful95]. Michotte [Mac90a]. Mick [Zie99v]. Micro [Spr91]. Microcomputer [Sar94b]. Microelectronics [Sam99a]. Microsoft [Zie96m, Zie97j]. Microsoft(R) [Zie99i]. Middleton [Zie96m]. Mieczkowski [Zie99-51]. Mikel [Cab94]. Mikhail [Kin97]. Mikulak [Bel95]. Milan [LaL95]. Milík [Gri97a]. Militky [LaL95]. Millard [Zie99m]. Miller [Wyc96, Zie90-27, Zie93-35, Zie95-50, Zie96-52, Bro90b, Bru97, Sch92a, Wyc96]. Milliken [Zie97-27, vT91]. Million [Gri90]. Milton [Wol92, Kia92a]. Mine [BH93]. Mine-Product [BH93]. Mineral [Ott90]. Minimal [Jac92, Jac93]. Minimizing [Cro92]. Minimum [BS99, HCV98, RvD99, SCF97]. Minimum-Aberration [BS99, HCV98]. Mining [Van91, Zie98c].

MINITAB [Wit94, Zie90-27, Tho90b, Zie98i, Gun90]. Minja [Zel94]. Mirkin [Hig98]. Miron [Zie96-45]. Misclassified [HR92, HR94]. Missing [HLS98, Kia95b]. Mitchell [Pla96]. Mitra [Sen94a, Bem96]. Mittag [Jai95]. Mixed [BNC87, BL97, CPJ92, KM94, Ngu96b, Ste99, WW92, WL97, Zie97-27]. Mixed-Effects [WL97]. Mixed-Model [BNC87]. Mixture [CR98, CDKS92, DPL^{+}93, DPL^{+}94, Hes95, SC93, WL91, Zie92i, Raz91]. Mixtures [Czi92]. Mockus [Mar99]. Modal [Cox93]. Modarres [Geo94]. Mode [Sch96b]. Model [Ann99, Ano99f, BNC87, Bor97c, CGG92, CR98, CDKS92, Cro92, DJ94, JFM99, KK99, LT96, Len95, ML95, PM99, PJ90, Pie90a, RN91, RL95, STC99, Sen98, SS97, TW92, Van92b, Vin92, WL97, WJ92, Zie90p, Sen99b, Sch91b, Iye90, Rig99, Wha95, Zie90o]. Model-Comparison [Vin92]. Model-Free [Wha95]. Model-Independent [STC99]. Model-Oriented [Rig99, Zie90o]. Modeling [Alt99, Bax96a, Bay92, CH96b, CDKS92, Dea94, EH96, Fot96, Gro96, Kir93, Liu99, MEL98, NLB^{+}98, Pig91, Zie92p, Zie95-29, Zie96-44, Zie97-29, Zie98-39, Fre94, Gru97, Hec96, Huz95, Zie97g, Zie98f, All96]. Modelling [Bul96, Glo95, Moo91, Por97a, Zie90-27, Zie96a, Zie90n, Zie92s, Zie93-30, Zie94e, Zie95-50, Zie96-40, Zie97b, Owe96, Tho93, Zie95x, Zie99t]. Models [Ano98e, Atk96, BT89, BA96, Boo90, Bor97c, Cal94, Cal98, CPJ92, Dea95, Dob96, DH92, DP97, EMS98, Gal92, Gil92, Gla98, Gol98, Goo91, Gra98, Gue90, Gun94a, Gut90, Har91c, Hes93, HL97, Hof95, Iye98, Jol97, JR91, KM94, Khu92, Kle94, Law96, LLP^{+}99, LN96, Mad96, MSH94, Mee91b, Nip92, Ott90, Ped98, Peñ91, Qiu95, Rig93, Roe91, Sch92b, Ste99, TC90, Tim96, Tor99, Van96a, Van96b, Zel94, Zie91j, Zie94-39, Zie95-33, Zie95-39, Zie97r, Zie98-27, Zie98n, Zie98-28, Zie99-46, Zie99e, vT95, BT98, Zie97-44, Arm97, Bab98, Bor95b, Bry92, Cab98, Che98, Dea91, Dea96, Dea97, Dow95, Emp92, Emp93, Ens95, Fox97, Gal95, Glo95, Hab92, Iye90, KB94, Kia90, Kia92a, Kim91, Ord96, Raz91, Sch92d, Sch94b, Sen99a, Wel90, Wes92, Win93, Zie90a]. Models [Zie91i, Zie92b, Zie92n, Zie93o, Zie96u, Zie97-27, Zie97e, Zie97-42, Zie97-44, Zie98-29, Zie99s, Zie99-51, dV92, Kia94]. Modern [Bal93, Coa94, Dun97, Rig92a, Was91, Zie91d, Zie95-50, Ada94, Ajm99, Dre91, Fre99, Jam96, Joh91, Law92, Osb97]. Modification [DJ94]. Moen [Cha92]. Moeschberger [Rao98]. Mohan [Bil91]. Molchanov [Zie99-46]. Molenberghs [Ste99]. Money [Spo90]. Monfort [Cab98]. Monitoring [BK92a, BK92b, DP91, Fal99, Gan95, GSH98, HNF97, Mac92, NW93, NM95, Tuc92, VV92, Van96a, WG92, Yas94, Zie95-37]. Monotone [Rag95]. Monson [Gru97]. Monte [Gas97, Nik93, Zie96-27, Zie98-36, Was97, Zie96-27, Zie98-32]. Montgomery [Ale95, Por93, Zha92, Zie99l, GR97, Gun96a, Nar93, Sau92, Sri92, Zie92r, Zie96-37]. Mooney [Zie94j, Zie98-32]. Moore [Arn90, Zie94y, Moo96]. Moran [Pit92]. Morgan [Zie95-47, Cra94, Zie92d, Zie93d, Zie94s, Zie97-33]. Morgenthaler [Zie94-29]. Morris [Aba98, Kot92, Sea94]. Morton [Dav99]. Moser [Bor97c]. Moskowitz [YWMP95]. Mosteller [Sch93c, Sch93b, Zie90w]. Motiska [Zie93-34]. Motor [MSF95]. Mottley [Zie94-33]. Moubray [Zie96-35]. Moura [Hul92]. Moving [Bax90, DP91, FHT90, Gan95, Hun90b, LWCR92, LS90b, LS90c, MH90, Van96a, WM90]. Mowery [Zie96k]. MR0702171 [Sin92]. MR0887702 [Ste90b]. MR1007294 [BT98]. Mueller [Zie97b]. Muller [Pri98, Eub90, Zie99-27]. Multi [Pap95, Yin91]. Multi-Armed [Yin91]. Multi-Vari [Pap95]. Multidimensional [Jol96]. Multifactor [Fed96, JD96, KTTS90, MSB96a, MSB96b, Tob96]. Multinomial [HR92, HR94]. Multiple [Ale98, BB95, Cha93, CB93, Hir96, Jol93, MME98, PF91, RH92, SE93, SC93, Tam96, Tür94, USM99, Zie94x, Zie95-43, Chr97, Zie94-27]. Multiple-Group [Hir96]. Multiple-Outlier-Detection [PF91]. Multistage [MB98]. Multitaper [Cox96]. Multivariable [Zie92e]. Multivariate [Ano91f, Boy96, Bre91b, Chu91, Fri91, FB94, Goo91, Goo92a, Gu91, HL95, HT91, Haw91b, Hec99, HS97, Ize91, Kia92b, Kia93b, KM99, Lal99, LLP^{+}99, Lin90, Loh92, LWCR92, Mar97, Mat96, Mat98, Nai94, Nip92, NM95, Not92b, Pfa99, Pos98, Tho91, Tho92a, Tim95, Ver90, WL91, Wol96, Zie95b, Zie96d, Zie97h, Zie97l, Zie99c, Zie99-51, Bre91a, Boo98, Bur91, Glo95, Gra90, Hag91, Haw90, Iye98, Ize90, Mas99, Mat97, Mye96, Nem95, Nem97, Ord96, Ram93, Sha95, Sha97c, Tim99, Zie91k]. Multiway [Zie92k]. Muni [Chi91]. Murray [Moo91]. Murtha [Zie96n]. Myers [Gun96a, Kia92a, Rob90, Zie91d, Zie94-31]. Myles [Zie99u].

N [Ber96b, Fam95, Far93, Glo93, Gri99, Hwa92, Jam96, Kat91, Kir92, Kir93, Kus93, Mac90a, Pos98, Rie94b, Run90, Sai96, Sar93, Sin97, Weh95, Whe93, Wu99, Zie90m, Zie90h, Zie90d, Zie90n, Zie92g, Zie92s, Zie93-30, Zie93r, Zie94w, Zie94-45, Zie94s, Zie94-38, Zie94-33, Zie95n, Zie96-34, Zie96h, Zie96k, Zie97r, Zie98-43, Zie98d, Zie98e, Zie98u, Zie99i]. Nachtsheim [Zie97-44, LC90]. Naes [Zie95y, Hag91]. Nagaraja [Far93]. Naik [Zie96d]. Name [Gre94]. Nancy [Coo96b, Zie95-45]. Narahari [Kir93]. Narayan [Mat97]. Narrative [Baj98]. National [Sri93, Zie96-46]. Natural [Lyn97b]. Nature [Sai96]. Naumann [Zie97i]. Nazem [Elt90]. Neapolitan [Jon92]. Near [Ngu96b, Zie95y]. Near-Orthogonal [Ngu96b]. Nearly [WW92]. Neave [Ste90c, Pfa92]. Needed [Sne90]. Neely [Zie98y]. Neil [Zie99-51]. Nelder [Sch92d]. Nelson [Mee91b]. Nenadic [Zie94g]. Nested [LBB96, SS97]. Nesting [SS97]. Neter [Kia90, Zie92b, Zie97-44]. Nets [Ter92]. Network [Bax96a, BCL98a, BCL98b, Rob98, Sch98a]. Networks [NO96, Ste96a, Ste96b, Zie95-49, dVU96, DSSU98, Zie95-49, Ful97, Yak97, Zie96-49, Zie94-28]. Neufelder [Soy95]. Neural [Ful95, Ful97, NO96, Sim97, Ste96a, Ste96b, Zie95-49, Zie96-49, dVU96, DSSU98, Zie95-49]. Neutrinos [MJ91]. Neville [Pan96]. Nevison [Zie90u]. New-Better-Than-Used [CR93]. Newbury [Zie93x]. News [Woo91]. Next [HH96, Zie96r]. Ng [Loh92]. Nicholas [Ada94, Che92, Gal95]. Nicholls [Zie90-28]. Nick [Dri91]. Nickolaj [Gri99]. Nickulin [Kin97]. Nicolas [SN95]. Niederreiter [Zie96-27]. Niels [Kle94]. Nielsen [Har91c, Zie97-42]. Nielson [Zie94-28]. Niewiadomska [Cha98]. Niewiadomska-Bugaj [Cha98]. Nigel [Sch96c]. Nikiforov [Rie94a]. NIR [Zie95y]. No [Gre94, BT98, DPL^{+}94, HR94, Lin95a, Sen99b, Sin92, Ste90b, Bog92]. No-Name [Gre94]. Noel [Zie94-41, dV93]. Noether [Far92]. Noisy [Bre91b, Fri91, Gu91, HT91, Zie93f, Bre91a]. Nolan [Cha92, Car96]. Non [Cam96, Tsa91]. Non-Linear [Cam96]. Non-Standard [Tsa91]. Nonconformance [GS92]. Nonconstant [MM94]. Nondestructive [Ber96a, CR96, OM96a, OM96b, Spe96b, Swe96, Tuc96]. None [VV94]. Nonlinear [Arm97, Bor97b, Iye90, Kaf94, Zie91i, DSSU98, Ber91, Goo92a, KB94, Kus92, Mil90, San92, Spo90, Taa91, Tim96, Zie91l]. Nonlinearity [CS92]. Nonnegative [Bre95a, KM94]. Nonnormal [Bas97]. Nonparametric [CM99, DM99a, DM99b, ET96, Eub93, Far92, Gas94, Gib99, Gil99, HC91, HLS98, Mi99, QY98, Sch91c, Van99, Wol95, Woz91, Zie90a, Zie94j, Zie99-27, GJMS^{+}95, Ale99, Eub90, Gro99, Qiu95, Zie93s, Zie99u]. Nonprobabilistic [Per92]. Nonquadratic [Har92b]. Nonreplicated [vT91]. Nonresponse [Zie99v]. Nonsampling [Zie93t]. Nonsense [Das93]. Nonstandard [BC95, HD95, PD95, RD95a, RD95b, SWG95, SP95]. Nonstationary [Luc98]. Nonsymmetric [Har92b]. Noreen [Len90]. Norm [Spo90]. Norma [Zie99l]. Normal [Ber97, Ize91, KM99, Mee90b, WN97, Ode90]. Normality [Col92]. Normally [Mee90a, WJ92]. Norman [Ber96b, Bru94, Car93, Cha95, Fam95, Gun99, Pos98, Zie91p, Zie98q, Zie98e, Zie98k, Zie98z]. Normand [San99]. Norton [Zie98-45]. Note [Bis94, Jac92, KSJ96, Ngu96b, Rag98]. Notes [Har91c, Ste99]. Nozer [Zie98l]. NU [MQ91]. Nuclear [Beg90, Dow94, GD94, Got94c, Got94b, Jou94, MKA96, MPR99, Rut94, Tor94, War90, WB94]. Nuclear-Power-Plant [MKA96]. Nuclear-Waste-Site [Dow94, GD94, Got94c, Got94b, Jou94, Rut94, Tor94, WB94]. Number [Nel95, Str96]. Number-Theoretic [Str96]. Numbers [Woo91]. Numerical [Hal95, Rie97, SN95, Wat96, Zie99w]. Nychka [Zie99f].

O [Bax92, Pla96, Rag93, Rie97, Sin90a, Zie91l, Zie94-28, Zie97-42, Zie98g, Zie99d, vN90]. Object [Amm92, Zie99-31]. Object-Based [Zie99-31]. Object-Oriented [Amm92]. O’Brien [Sch94b]. Observational [All96, Gal92, Zie96b, Zie96-28]. Observations [Yas93, Ode90]. Observed [Sti97]. Observing [Ber92a]. Ocean [Zie96-38]. Ochi [Rag91]. O’Connell [Zie91j, Zie94t]. O’Connor [Hey92]. Odds [Zie99x]. Odeh [Emp93, Gri90]. Off [KD97]. Off-Line [KD97]. O’Hagan [Cam98, LeP90a]. Oldford [Zie95-33]. Ole [Har91c]. Oliver [Ter92, Coh92]. Olkin [Dob96]. Olson [Bel95, Zie98-46]. Omachonu [Zie99-29]. Omnibus [DP91]. On-Line [NW93]. One [CWWR97, MQ91, Oeh94, RH95, Van92b, Wan94]. One- [CWWR97]. One-Cell [Oeh94]. One-Sided [Van92b]. One-Way [Van92b]. O’Neill [Zie94q]. Open [Jon90, Jon91b]. Opening [Bel95]. Operating [EMSD91]. Operation [Van91, Zie98q]. Operational [Kat98]. Operations [HA93, LeP90b, Spr91]. Operators [McC90, Owe91]. Opinions [Sam90]. Oppenheim [Bre90, Zie95-31, Bre90, Zie95-31, Zie97-43]. Opportunities [Sri93]. Optical [WVYB97]. Optical-Fiber [WVYB97]. Optics [MA92]. Optimal [Bak98, DJ94, Har92b, JV93, MN95, Ngu94, Rag98, RJ97, Spu92, SWC97, Tür94, Bor94, Zie90q]. Optimality [Ney94]. Optimally [AD96]. Optimization [Ano97f, Ano97e, Emp97, Goo95, Gun96a, Hof95, Mar99, O’D99, Zie96-30, Fot95, Lew93, Zie99y]. Optimizing [Ott90, Sto92]. Optimum [AB97, MM94, Whe93]. Orav [LeP90b]. Ord [Boo92]. Order [Ang90b, AB97, BA96, Far93, Fre95, Lin93a, Lin95a, Sho90]. Ordered [CH96a, KTTS90]. Ordinal [vT95]. Ordinary [ZZ91]. Ordination [Sha95]. Organic [Ano97f, Ano97e, Emp97]. Organization [Zie98-47]. Organizational [Zie99q]. Oriented [Amm92, Ste99, Rig99, Zie90o]. Ørnulf [Kle94]. Orr [Sau97]. Orthogonal [CL93, DPL^{+}93, HS99, Hol98, Jac92, Jac93, LIV96, Mor98a, Ngu96b, WW92, DPL^{+}94]. Orthogonality [Hag96]. Oscar [Tom95]. Ostle [Bod97, Zie90x]. Ostrowski [Zie95a]. Other [Col92, Ped98, RJ97, Yas92]. Ott [Arm92, Ram91, Ham96, Koo90, Zie94f, Zie99-41]. Outgoing [BHS90, MT91]. Outlier [CWWR97, PF91]. Outliers [Cro94, Bre95b]. Outlying [Ano90e, Sim88]. Output [STC99]. Overdetermined [Bro92]. Owen [Gri90, Smi91, Woo92]. Ozone [GK93].

P [Arn90, Bas97, Bur91, Cha92, Dun90, Gal93, Had92, Hwa92, Ked98, Lam90, LeP90b, LuV90, Mi99, Not92a, Not92c, Osb97, Pit92, Pro93a, Qiu95, Rao98, Ruk97a, Rus94, Sac93, Sch92d, Sin90b, Sin97, Ste90c, Tho90a, Tim99, Wal92b, War90, Yan96, Yeh93, Zie90-28, Zie90a, Zie90k, Zie91k, Zie92d, Zie92c, Zie92v, Zie92m, Zie92n, Zie93e, Zie93r, Zie93g, Zie93-34, Zie94u, Zie94k, Zie94-36, Zie94-34, Zie95-33, Zie95g, Zie95-42, Zie95d, Zie95n, Zie95-50, Zie96e, Zie96-27, Zie96k, Zie96-52, Zie96g, Zie97j, Zie97n, Zie98h, Zie98-43, Zie98w, Zie99-27, Zie99v, Zie99j, Zie99n, Zie99m, dC99]. Padhraic [Zie98c]. Pagano [Zie94-30]. Pahkinen [Zie96-29]. Pairwise [LW97]. Pamela [Zie99-52]. Panchapakesan [Zie98d]. Pandit [Cox93]. Panel [NAM^{+}92, Zie90-28]. Panizzon [God95]. Panju [Zie98-29]. Pankrantz [Win93]. Pannatier [Zie96-50]. Pantula [Zie99d]. Papers [Zie99-36]. Papoulis [Gro92]. Paradigm [Glo99]. Parameter [Gru90, HL97, KGK90, MM94, MW98, NAM^{+}92, Won92, Kim91]. Parameters [AKO94, Gri97b, Kun93, Sin83, VB94b, Sin92]. Parametric [Har91c, Sha98]. Pareto [Gri93]. Park [Tan98, Zie93x]. Parker [Zie91g]. Parmar [Zie97-37]. Parsowith [Zie96q]. Part [AHT91, Sha97c, Zie98w]. Partial [Bay92, Coo93, Mar96, Zie96-47]. Partially [Jac93]. Partnership [Sne90]. Parts [Bru97, Zie93k, Gri90]. Parzen [Zie99-36]. Pass [Zie96-36]. Paszek [Mor93b]. Patel [Ber97]. Path [Bay92, Hal94]. Patil [Zie96p]. Patrick [Hag96, Hey92, Zie98-40, Zie99-43, Zie99i]. Pattern [Boo99, Nem95, Ray93, Ful97, Sim97]. Patterns [FL97]. Paul [Gat96, Mat98, Rag95, Tho92b, Van92a, Zie96-28, Zie97-38, Zie99q, HSPF92]. Pavlovic [Van91]. Pazman [KB94]. PC [LaL95, Cam90, Dun90]. PC-Aided [LaL95]. Peace [Stu94, Zie91r]. Peacock [Zie94-38]. Peak [HT92]. Peak-Preserving [HT92]. Peck [Zie99-37, Por93]. Pegg [Gri99]. Pell [Tob99]. Pellets [Tho94]. Pelosi [Zie99k]. Penalty [Qiu95]. Graph [Zie96-45, Zie96-33]. Japan [Zie95-29]. or [CR91]. Parallel [MW90]. Physical [Box90, God90, Hah90, HK90, Hoo90, Hun90a, Mac90b, Moo90, Tri90, Wan90]. Process [Sto92]. Peppen [Bar92b]. Percival [Cox96]. Perez [Pap95, Sto92]. Perez-Wilson [Pap95, Sto92]. Perform [Bru96, Sto96]. Performance [Dow94, GD94, Got94c, Got94b, Jou94, Luc98, Rut94, Tor94, WB94, Yas93, Kir93]. Periodic [Tat96]. Permutation [LN97, Nik95]. Perry [Jen91]. Personal [Bon95, Hal94]. Personalities [Zie98z]. Perspective [Bon95, Lor93, Zie98-30]. Persson [Kel97]. Persuasive [Lit91]. Pete [Dri91]. Peter [Ber92b, Bla92, Bre96, Bru93b, Bur91, Cha93, Cro99, Dri91, Gas94, Hal95, Hil95, Koh90, Mil94, Ram94, Rig99, Sea90, Sha97b, Tso92, Wit94, Yeh93, Zie94b, Zie95-35, Zie97-30, Zie98j, Zie98l, Zie99-38, Zie99-47, Zie99t]. Petroleum [Zie90j, Zie99-43]. Pfeiffer [Van92a]. Phadke [Men91]. Pharmaceutical [Zie91r, Zie96-39]. Pharmacochemistry [Yeh93]. Pharmacology [Tho90a]. Phase [HH96]. Phenomena [Lin99]. Phil [Ko95b]. Philip [Gle96, Gri99, Nik95]. Phillip [Zie97-39]. Phillips [Rag93, Bul91, Dom90, Han97, Job99]. Philosophy [Bru97]. Photographic [VD94]. Photoresist [CR98]. Photoresist-Coating [CR98]. Physical [Cox96, Hes91, Jen90, Rag91, Sha97a, Wal92a, Zie93c, Zie95-38]. Physicists [McC94]. Physics [Gal93]. Piatetski [Zie98c]. Piatetski-Shapiro [Zie98c]. Piegorsch [Zie98-42, Zie99f]. Piepel [Ano90g]. Pierre [Bil94, Zie95z, Bil92, Zie95z]. Piersol [Nan94]. Pierson [Zie93w]. Pieters [Zie90w]. Pinyuen [Joh91]. Pisani [Loh94]. Pitard [Zie95z, Bil92]. Pitman [Zie94-32]. Pitt [Hal94]. Plackett [LD92]. Plan [BHS90]. Planned [Cha92]. Planning [CM93b, CM93a, EM95, Gun93, HO93, Hah93, LST93, Pri98, WC92, Zie97w]. Plans [EMSD91, God95, Jac92, Jac93, MT91, Mee91b, Mor91, MT92, Rag98, RM95]. Plant [MKA96, Zie95t]. Plante [YWMP95]. Plants [MPR99]. Platen [Hal95]. Pleasures [Lei97]. Plot [BS99, HCV98, Mil97]. Plots [Ber98, Coo93, Coo96a, GK90]. PLS [Wol93]. PLUS [Han98, Jam96, Zie98-46, Ko95b, Ko95a, Zie99m]. Pocket [Aba98]. Point [CCC90, LT96, Zie99-49, Boo90]. Point-Process [LT96]. Poisson [LuV90, CS93, Lam92, Li97, LLP^{+}99, Boo95]. Policies [VV94]. Poling [Bel99, Zie96k]. Polyak [Mi97]. Polymerization [CFRH99]. Polynomial [MSH94, QY98, RWHH97]. Pooling [GS94, RH92]. Population [KM99, Ruk98, Zie93j]. Populations [Mee90b, Sin83, WN97, Zie92m, Sin92]. Posterior [Atk96]. Power [Ada96, Bak96, MKA96, MPR99, Per97, Zie95c]. PQ [Zie99-50]. Practical [Aba98, Bel95, Der90, Dun96, Gri99, Haw90, Lar95, Nik95, Tob99, Bay95, Bli91, Goi96, Gre98, Gun96b, Han97, Hey92, Jen90, Jen92, McL95, Mor99b, San99, Zie95-27, Zie96-29, Zie97t, Zie98-33, Zie99-27, Zie99z, Zie99-28]. Practice [Arn90, Bru97, Dog99, Moo96, Ram93, Sam99b, Ste99, Wol93, Zie94y, Zie98e, Bil92, Col93, Gas97, Lei92, Lyn94, Sha93, Zie95z]. Practices [Sha92]. Practitioners [Kra92, Jon93]. Prat [Zie98-34]. Pravin [Pet99]. Pre [Pap95]. Pre-Experimentation [Pap95]. Precision [CB93]. Predicting [GJMS^{+}95, WBS^{+}92]. Prediction [CR91, CM99, DM99a, DM99b, EM99, Gib99, Gil99, KLR91, LN96, MMY93, Roe91, Sch92b, Van99, VM91, Zie95t, DSSU98, vdVtBM97, Bor95a, Ode90]. prediction-variance [Bor95a]. Predictions [TC90]. Predictive [Law95]. Predictors [ZZ91]. Preliminary [Mor91]. Prem [Zie93m]. Prescott [SDP^{+}95]. Prescriptions [Eas90]. Presence [Bri96, Cre97, FRY96, GS95, JV93, WI94]. Presenting [Lit91]. Preserving [HT92, JS99]. Press [Gue90]. Prevention [Bru91]. Price [Ada91, Maz93]. Priestly [Zie98g]. Primary [Wal92b]. Primer [Ada92, Car90a, Rab92, Ram99, Rea93, Mil98, Nem97, Zie94b]. Principal [Lei93]. Principle [Wol95]. Principles [Boo93, Der90, Gue90, Han93, McC94, Min93, Sha92, Zie97b, Zie99-47, Zie93-33, Gol90, Job99, Lin90, Mil91, Oli97, Zie90r, Zie94-30, Zie99-29]. Prior [ST90a]. Priscilla [Kin97]. Probabilistic [LSBW95b, LSBW95a, Law96, Lyn95, Zie94-28, Zie98b, Jon92, Pap96]. Probabilities [Atk96]. Probability [Ale95, Bab98, GK90, Gun94a, Gut92, Hag90, Huz95, Joh91, Jon91a, MT91, Rag91, Rig93, SGW96, Taa99, Tho99, USM99, Wol92, Zad95, Zie91k, Zie92a, Zie96u, Jon90, Lei97, Rab92, Zie95-42, Zie98a, Bax92, Bru93a, Bry99, Cha98, Dob96, Gal93, Gro92, LeP90a, Lei91, MA92, Nar93, Rob90, Van92a, Wal92a, Was91, Zie90s, Zie91m, Zie92l, Zie94-31, Zie94-32, Zie95-28, Zie96-30, Zie99-30]. Problem [Bog92, Bru95, Fri97, HP97, Hey93, MA92, Mor92, MS95, Pit92, Pro93a, ST90a, Zie97a, Don90, Zie96-31]. Problem-Solving [Pit92]. Problems [CR98]. Procedural [Bla92]. Procedure [Mee90a, PF91]. Procedures [NW93, RH92, ST90a]. Proceedings [Bro90b, Zie95-29, Zie98-34]. Process [Acr97, Ada93, Ada96, Arm93, Bak96, Boo90, BL97, BK92a, BK92b, Boy96, Cab94, CFRH99, CS93, Col93, DP91, Edg98, FB94, Gan95, Gan97, Gho94, God95, Gri98, GSH98, Gun96a, Hal94, JV93, Kal93, Kra90, LT96, Lew92, Mac92, McC98, McG94, McL95, MS95, O’D99, OUH93, Pan96, Run90, SG99, Tat97, Tuc92, TFV93, VTFD92, VV92, Vuk98, WG92, Yas95, Zha98, Zie92r, Zie93-38, Zie94r, Zie94-34, Zie96-37, Zie96-41, Zie96f, Zie97-31, Zie98q, Zie98-43, Zie98-40, Zie99-33, Bri91, Kle94, Zie99a, Cha95, Dog93, Han93, Pla90, Ram91, Zie97u, Zie97v]. Processes [AWS94, Bax95, DH92, Fel94, HNF97, Kir92, Lei91, Luc94, MB98, NM95, Rag91, Rag93, Sen97, Van96a, WMP94a, WMP94b, Wes90, ZP94, Boo95, Fre93, Fre97, Glo97, Ked96, Ked98, Koe92, Kra92, Pit92, Sea90, SN95, Sto92, Weh91, Wen93, Zie93f, Zie94-44, Zie97k]. Processing [Gru97, Rie94b, Sen95, Rie97, Sen94a]. Producibility [Cab94]. Product [AAK99, Alt96, BH93, Gun96a, Nel95, Zie92i]. Production [Cro92, Sch95, She90b]. Productivity [Hal98, Pro90, Zie96-52]. Products [Bis97, Zie95k]. Profession [Sne98]. Professional [Gun96b]. Profile [FB94, Len93]. Profit [Zie94-33]. Profitability [HH96]. Programming [Bro90a, LaL92, RM95, Zie96-36, Zie98h, Bax95, Bat99, Zie99-31]. Progressively [VB94b]. Project [Zie99-28]. Projection [LD92]. Proof [SW95, USM99]. Proof-Loading [SW95]. Propagation [HA93]. Proper [Wil90a]. Properties [Bax90, FHT90, Hun90b, LD92, LuV90, LS90b, LS90c, LG99, MH90, Ste87, WM90, Bor95a, Ste90b, Sin90a]. Proportional [TWN98]. Proportional-Integral [TWN98]. Proschan [Cly94]. Prototype [BS97]. Prove [Zie99-32]. Provost [Cha92]. Proximity [Lee99]. Pruett [Zie94r]. Pukelsheim [Bor94]. Purves [Loh94]. Puterman [Bax95]. Puzey [Aba98]. Pyzdek [Zie93v, Car90b, Car92, Hea94, Spe96a, Wan94, Hea94, Wan94].

Q [Deh90, Jon93, Nel98, Sam99b]. QFD [Coo96b]. QSAR [Bur91]. Quadratic [CCC90, MT92, Spu93]. Qualimetrics [Zie98w]. Quality [Aba98, Ada92, Ada94, Ano91f, AHT91, Bel95, Bem96, Bos92, Bru93b, BHS90, Car90b, Car93, Cha92, Die92, Dun96, Dun97, Edg92, Fre99, Fri97, Gan97, GR97, Gho94, Gri90, HH96, Haw91b, Hoc94, Jai95, KD97, Kal98, Kin93, Lam90, LC94, Lew92, Mar94, Mas99, McG94, McL95, MT91, Mil91, Mil94, NW93, O’D97, Oli98, Pan96, Pla90, Pro90, Pro92b, Pro92a, Ram91, Rea93, Rol93, Ros96, Sac93, Sar94b, Sau92, Sea94, Spa99, Tan98, Tor99, WVYB97, Won98, Wyc96, Zei94, Zie90l, Zie90t, Zie91p, Zie92w, Zie93l, Zie93-37, Zie93q, Zie93-36, Zie93-35, Zie94-44, Zie94-43, Zie95-45, Zie96o, Zie96q, Zie96-52, Zie96-46, Zie97-39, Zie97-32, Zie97o, Zie97a, Zie99-50, Zie99o, Zie99y, Zie99n, dC99, Zie99-33, Bre90, Car92, Lew93, Neu99, Smi91, Zie92i, Zie93p]. Quality [Zie93a, Zie93-34, Zie96k, Zie96-51, Zie96z, Zie97-34, Zie98-43, Zie99-29, BR95, Ben93, Bre96, Coo96b, Dog99, Hwa92, Lew92, Maz98, McC90, McC91, Men91, Pie90b, She90a, She90b, Sin90b, Spe96a, Wal92b, Zie93w, Zie93v, Zie93u, Zie94-33, Zie95-30, Zie95-31, Zie96-32, Zie97w, Zie98-35, Zie99-33]. Quality-Monitoring [NW93]. Quantal [Cra94]. Quantified [Dav99]. Quantiles [YE94]. Quantitative [Bay99, Gro99, STC99, Sch94b, Che92, Pro90, Zie99-34]. Quantities [Baj98]. Quark [HS97]. Quasi [Zie96-27, Bor98]. Quasi-Likelihood [Bor98]. Quasi-Monte [Zie96-27]. Quesenberry [Ano93f, FMQ92, dC99]. Questions [Zie97q]. Queueing [Bax92, Gla98, Sys99]. Quick [Zie96-33, Zie97x].

R [Aba98, Ada92, Ada94, Arm97, Baj98, Bar92a, Bel95, Bel99, Bil91, Bla94, Bor97a, Bre90, Bri91, Bzi97, Che92, Coh92, Dea91, Dea97, Der90, Dri92, Eas99, Emp93, Fre93, Gas97, Gho93, Gho94, Gol95, Got94a, Gri90, Gun99, Hab92, Ham96, Hes91, Hoc94, Hor98, Jan94, Jen90, Jen93, Kia92a, Kia93a, Kir92, KY91, Kra90, Lit91, MA92, McC98, McG94, Mil98, Moo99, Mor99b, Mor97a, Nem92, Pan92, Per97, Pfa92, Pie90a, Pit92, Pit96, Rag93, Ram99, Ram91, Rie94b, Rut93, Sco90, Sco92, She90a, Sho90, Sin90a, Sin90b, Ste90c, Ste98, Ter92, Thi91, Tob99, Wal92a, War90, Wat96, Weh93, Wei90, Wei94, Wil98, Won94, Zie90-28, Zie90t, Zie90w, Zie90k, Zie91q, Zie91k, Zie92k, Zie92c, Zie92v, Zie92n, Zie92o, Zie92u, Zie93c, Zie93-28, Zie93p]. R [Zie93v, Zie93r, Zie93-36, Zie94e, Zie94k, Zie94j, Zie94-31, Zie94l, Zie94-44, Zie94-43, Zie94-33, Zie94z, Zie94q, Zie95-33, Zie95-31, Zie95n, Zie95-28, Zie95-50, Zie96d, Zie96e, Zie96-44, Zie96h, Zie96-40, Zie96-52, Zie96-48, Zie96-29, Zie96r, Zie96-36, Zie96v, Zie97b, Zie97-27, Zie97-42, Zie97v, Zie97g, Zie97k, Zie98-45, Zie98s, Zie98q, Zie98-44, Zie99-47, Zie99s, Zie99t, Zin93]. R&R [VV99]. Rabe [Lee99]. Rabe-hesketh [Lee99]. Rade [Zie91c]. Raghavarao [Sea93]. Ragsdale [Zie98-39]. Raheja [Sha92]. Rahim [Zie99y]. Raj [Bas90]. Raja [Tim99]. Ralph [Goi96]. Ramasamy [Zie98c]. Ramaswamy [Zie97k]. Ramon [Hei90, Zie99-35]. Ramsay [Geo98]. Ramsey [Zie96-52]. Rand [Hor98]. Random [Ann99, Ano99f, BT89, Gol90, JV93, JFM99, Khu92, KK99, Lei91, Lin99, MJ91, Oma91, PM99, PH99, RH95, Sen97, Taa91, Van92b, VV99, WT98, BT98, Bax96a, Gal95, Nai96]. Random-Effects [BT89, Van92b, VV99, BT98]. Randomization [Nik93, Zie98-36, Zie98-37]. Randy [Tob99]. Range [CC99]. Rank [Gro96, Tim99, Gov92]. Rao [Zie91k, Zie93-28, Zie94v, Zie98g, Eas99, Rie94b, Sin90b, Wei94, Zie90k, Zie91k, Zie95l, Zie96p]. Raoul [KR93]. Rasmussen [Sau94]. Rate [Hir96, Hul92, MPR99]. Rates [GS92, GS94, TW96]. Rather [Zad95, Ano97f]. Ratio [BBT90, MS90, Sin83, SGW96, Sin92]. Rationale [Nay98]. Ratios [Zie99x]. Ratkowsky [Zie94m, Zie91i]. Ratliff [Zie94g]. Raudenbush [Cal94]. Raun [Zie90d]. Rausand [Van96b]. Rauwendaal [Zie95-34]. Rawlings [Zie99d, vN90]. Raymond [Bel95, Gun96a, Rob90, Zie91d, Zie99-30]. Rayner [Kin91, PBR92]. Reaction [AB97]. Read [Zie98o, Zie98p, Ber97, Zie90h, Zie92g]. Readings [Han90]. Real [GSH98, KMP95, Zie94-34]. Real-Time [GSH98, Zie94-34]. Really [Arm90]. Reanalysis [MSF95]. Reasoning [Jon92, Lor99]. Reavis [Zie96k]. Recognition [Ful97, JR91, Nem95, Boo99, Ray93]. Rectification [GS92]. Recurrence [Nel95]. Recurrent [LN95]. Recursive [Ano91e, Gri97b, Haw91a]. Red [Zie95y]. Redesign [Zie97u]. Redman [Mar94]. Reduce [DJ94]. Reduced [Tim99]. Reduced-Rank [Tim99]. Reduction [Jol90, Lew93]. Regimes [MQ91]. Region [CCC90]. Regions [AAK99]. Regression [Ano91e, Ano91f, Atk96, BB95, Ber98, Bre95a, CS92, Coo96a, Dal90, ET96, Eub90, Fai91, Far97, FF93b, FF93a, Gle96, Gol95, Gra98, Gun99, HM93, Haw91a, Haw91b, HLD98, Iye90, Kaf94, Kia90, Kia92b, Kia95a, Kia95b, Lam92, Law98, MQ91, Mar96, ME99, MS90, MER91, Mil90, Osb97, PF91, Peñ91, Por93, Pro92a, Qiu95, QY98, Sch96a, SD90, STR99, Spu93, St.94, TW90, WVYB97, Win93, Zie90b, Zie91i, Zie91l, Zie91d, Zie96x, Zie96c, Zie97-44, Zie99d, DSSU98, vN90, Ale98, Ber91, Bor97b, Cuc92, Eub93, Fai94, Gol93, Sch92a, Tim99, Wei90, Zie92o, Zie98-27, Alt99, Ang91b, Bar95, Cal98, Chi91, Fox97, Gra94, Maz93, Moo99, Pet99, Zie93x, Zie94-35].

Regression-Adjusted [Ano91f, Haw91b]. Regression-Based [MS90]. Regressions [Moo99]. Regressors [vdVtBM97]. Reid [Zie92s, Zie93-30]. Reinhard [Spe91]. Reinsel [Tim99, Tim95, Zie95-46, Zie97l]. Rejection [CWWR97]. Rejoinder [WMP94b]. Reklaitis [Mar99]. Related [Gra98, Hab90, HW90b, HW90a, Hir90, Jac95, Loh92, Nai90b, Woo91, YDB90]. Relations [Zie91l]. Relationship [TI99]. Reliability [AHT91, Bar92b, Bar92a, BCL98a, BCL98b, Bry92, Bry99, Car90a, Dog99, EMS98, Geo94, Gle91, Gri99, Han97, Hey92, JD92, Kim91, MW90, MKA96, ML95, Nel98, Pap96, Rag98, RM95, Rig93, Rob98, Sch98a, Sen98, Sha93, Sin90b, Tor95, Tor99, Van91, Van96b, Wal92b, WJ92, Zie91p, Zie97r, Sen99b, Boo90, Die92, Fre99, Gri97a, McC99, Sam90, Soy95, Zie96-34, Zie96e, Bax96b, Car96, Fox94, Gru95, Hey98, Kha95, Law96, Lee93, Lyn95, Lyn97a, Pla96, Rig92b, Sam99a, Sch92c, Sch93a, Sch94a, Zie96-35]. Reliability-Centered [Zie96-35]. Reliability-Maintainability [JD92]. Relyea [Jen92]. Rencher [Boo98, Mat96]. René [Spo90]. Renewals [LT96]. Repairable [AHT91, LT96, SP95]. Repairable-Systems [SP95]. Repairs [Nel95]. Repeatability [OUH93]. Repeated [Kia94, Lin92, LN96, Sha97c, Tim96, Zie93b]. Repeated-Measures [LN96]. Repetitive [GS95]. Replicated [Jac93, Lwi94]. Reply [DGTZ97b, Got94b, LSBW95a, MSB96b, OM96b, RD95b, Ste96b]. Report [Kaf99, Mor97b, Mor98c, Nai90a, Nai91a, Nai92, Var94a]. Reporting [Woo91]. Reproduction [VD94]. Resampling [HR92, Nik95, HR94, Cha93]. Resampling-Based [Cha93]. Research [Bor97a, Kaf94, Sch95, Spr91, Sri93, Zie91r, Zie92f, Zie92u, Zie93u, Zie95-44, Zie96w, Zie97w, Zie98v, Zie99d, Zie99-52, Zie99g, Zie99-48, vN90, Ing92, Sri93, Zie90x, Zie94n, Zie95x, Zie98r]. Residual [Coo93, PF91, Sch91d]. Residuals [Ano91e, Haw91a, MQ91]. Residue [McA90]. Resistant [CWWR97]. Resnick [Wen93]. Resolution [Bis94, BL97, LIV96, SCF97]. Resolutions [DL90a]. Resolve [Fed96, JD96, MSB96a, MSB96b, Tob96]. Resource [Coh92]. Resources [Got94a]. Response [Alt99, Ano99f, BCL98b, BA96, BK92b, Bre91b, CCC90, CM93a, Cra94, DM99b, DL90b, Far97, FF93a, HW90a, Khu92, LS90c, MME98, Pro92a, RAA90b, SE93, SC93, VM91, Gun96a, Zie97y]. Response-Surface [DL90b]. Rest [Hal94]. Restricted [Sho90]. Results [Zie96-33]. Retherford [Zel94]. Reuven [Goo95]. Reversal [GS94]. Review [Aba98, Acr97, Ada91, Ada92, Ada93, Ada94, Ada96, Ajm99, Alb93, Ale91, Ale95, Ale98, Ale99, All96, Alt96, Alt99, Amm92, And90, Ang90a, Ang90b, Ang91a, Ang91b, Ang91c, Ang92, Ara92, Arm92, Arm93, Arm97, Arn90, Aye98, Bab98, Baj98, Bak98, Bal93, BR95, Bar91, Bar92b, Bar92a, Bar95, Bas90, Bas97, Bas99, Bat99, Bax92, Bax95, Bax96b, Bax96a, Bay92, Bay95, Bay99, Beg90, Bel95, Bel99, Bem96, Ben93, Ber91, Ber92a, Ber96b, Ber97, Ber92b, Bie92, Bil91, Bil92, Bil94, Bla94, Bla92, Bli91, Bod97, Bog92, Boo90, Boo95, Boo92, Boo93, Boo94b, Boo94a, Boo97, Boo98, Boo99, Bor97a, Bor94, Bor97b, Bor95b, Bor97c, Bor98, Bos92, Bow91, Boy90, Bre95b, Bre97, Bre90, Bre96, Bri91, Bri96, Bro90a, Bro90b, Bro92, Bru91, Bru92a, Bru92b].

Review [Bru93b, Bru93a, Bru94, Bru95, Bru96, Bru97, Bry92, Bry99, Buc90, Bul91, Bul96, Bur91, Bzi97, Cab94, Cab98, Cal94, Cal98, Cam90, Cam96, Cam98, Car90a, Car90b, Car92, Car93, Car95, Car96, Car97, Cha95, Cha92, Cha98, Cha93, Che92, Che94, Che98, Che91, Chi91, Chr97, Chu91, Cly94, Coa94, Cob91, Coh92, Col92, Col93, Coo96b, Cox93, Cox96, Cra94, Cro94, Cro99, Cry90, Cuc92, Czi92, Das93, Dav99, DeV99, Dea91, Dea94, Dea95, Dea96, Dea97, Deh90, Der90, Dic91, Dic97, Die92, Dob96, Dog93, Dog99, Dom90, Don90, Dov93, Dow95, Dre91, Dri91, Dri92, Dun90, Dun96, Dun97, Eas90, Eas99, Edg92, Edg98, Elt90, Emp92, Emp93, Emp97, Emp99, Ens95, Eub90, Eub93, Fai91, Fai94, Fal99, Fam95, Fam97, Fam99, Far90]. Review [Far92, Far93, Fot95, Fot96, Fox91, Fox94, Fox97, Fox98, Fre92, Fre99, Fre93, Fre94, Fre95, Fre97, Fri97, Ful95, Ful97, Gal95, Gal92, Gal93, Gan97, Gas97, Gas94, Gat96, GR97, Geo94, Geo98, Gho93, Gho94, Gib91, Gil92, Gla98, Gle91, Gle96, Glo93, Glo95, Glo97, Glo99, God93, God95, Goi96, Gol99, Gol95, Gol93, Gol98, Gol90, Goo91, Goo92a, Goo95, Goo92b, Got92, Got94a, Gov92, Gra90, Gra94, Gra98, Gre94, Gre98, Gri90, Gri96, Gri97a, Gri99, Gri98, Gri95, Gro92, Gro96, Gro99, Gru90, Gru95, Gru97, Gue90, Gun90, Gun96a, Gun99, Gun94b, Gun94a, Gun96b, Gut90, Gut92, Hab92, Had92, Hag90, Hag91, Hag96, Hal95, Hal98, Hal94, HH96, Ham96, Han90, Han93, Han97, Han98, Hap91, Hap92, Har91a, Har91b, Har91d, Har91c]. Review [Har92a, Has91, Haw90, Hea94, Hec90, Hec95, Hec96, Hec99, Hei90, Hem99, Hes91, Hes93, Hes97, Hey92, Hey98, Hey93, Hig98, Hil95, Hin97, Hoc94, Hof95, Hor98, Hul92, Huz95, Hwa92, IHMT95, Ing92, Iye98, Iye90, Ize90, Ize91, Jac95, Jai95, Jam96, Jan94, Jen90, Jen91, Jen92, Jen93, Job99, Joh91, Jol90, Jol93, Jol96, Jol97, Jon90, Jon91b, Jon91a, Jon92, Jon93, Kah94, Kal93, Kal98, Kan95, Kat91, KR93, KB94, Kat98, Ked96, Ked98, Kel97, Kha95, Kia90, Kia92a, Kia92b, Kia93a, Kia93b, Kia94, Kia95a, Kia95b, Kim91, Kin91, Kin93, Kin97, Kir92, Kir93, Kle94, Ko95b, Ko95a, Koe92, Koh90, Koh94, Koo90, KY91, Kot92, Kra90, Kra92, Kus92, Kus98, Kus93, LaL92, LaL95, Lal99, Lam90, Lar95, Law95, Law92, Law98]. Review [Law96, LC94, LC98, LeP90a, LeP90b, Lee90, Lee93, Lee94, Lee99, Lei91, Lei92, Lei93, Lei97, Lei99, Len90, Len93, Len95, Lew92, Lew93, LH97, Lin99, Lin92, Lin90, Lit91, Liu92a, Liu92b, Liu99, Loh92, Loh94, Lor92, Lor93, Lor99, LuV90, Lyn94, Lyn95, Lyn97a, Lyn97b, Lyn99, Mac90a, Mad96, Mad98, Mag90, Mag96, Mar97, Mar94, MA92, Mar99, Mas99, Mat96, Mat97, Mat98, Maz93, Maz98, McA90, McC90, McC91, McC98, McC94, McC99, McG94, McL95, Mee91a, Mee91b, Men91, Mi97, Mi99, Mil91, Mil94, Mil99, Mil90, Mil98, Min93, Moo91, Moo96, Moo99, Mor93a, Mor99a, Mor92, Mor99b, Mor97a, Mor93b, Mul97, Mye91, Mye96, Nai94, Nai96, Nan94, Nar93, Nay98, Nel98, Nem92, Nem95, Nem97, Neu99, Nik93, Nik95, Nip92, Not92a, Not92c, Not92b, O’D97, O’D99]. Review [Oli97, Oli98, Ord96, Osb97, Ott90, Owe91, Owe96, Pan92, Pan96, Pap95, Pap96, Par92, Pat91, Ped98, Per92, Per98, Per97, Pet99, Pfa92, Pfa99, Pie90b, Pie90a, Pig91, Pit92, Pit96, Pla90, Pla96, Pol93, Por93, Por97b, Por97a, Pos98, Pot93, Pri93, Pri98, Pro90, Pro92b, Pro93a, Pro93b, Pro94, Pro96, Qiu95, Rab92, Rag91, Rag93, Rag95, Ram99, Ram93, Ram91, Ram94, Ram96, Rao98, Ray93, Raz91, Rea93, Rie94b, Rie94a, Rie97, Rig92b, Rig92a, Rig93, Rig99, Rob90, Roh91, Rol93, Roz93, Ruk97a, Ruk97b, Ruk98, Run90, Rus94, Rut93, Sac93, Sai96, Sam90, Sam99b, Sam99a, San98, San92, San99, Sar90, Sar93, Sar94b, Sar94a, Sau92, Sau94, Sau97, Sau98, Sch91a, Sch92a, Sch96a, Sch91b, Sch95, Sch92c, Sch93a, Sch94a, Sch96b]. Review [Sch92d, Sch93c, Sch93b, Sch94b, Sch96c, Sco90, Sco91, Sco92, Sea90, Sea91, Sea93, Sea94, Sen99a, Sen94b, Sen94a, Sen95, Sen97, Sha92, Sha93, Sha95, Sha97c, Sha97b, Sha97a, Sha98, She90a, She90b, Sho90, Sim97, Sim91, Sin90a, Sin90b, Sin97, Sin98, Smi91, Smi93, Soy95, Spa99, SN95, Spe96a, Spe91, Spo90, Spr91, Sri92, Sri93, St.94, Ste90a, Ste91, Ste90c, Ste98, SG99, Ste99, Sti97, Sto96, Sto92, Str96, Stu94, Sys99, Taa91, Taa99, Tam96, Tan98, Tan93, Ter92, Thi91, Tho99, Tho90a, Tho93, Tho90b, Tho92b, Tim95, Tim96, Tim99, Tob99, Tom95, Tor95, Tor99, Tro97, Tsa91, Tso92, Van92a, Van91, Van96b, Ver90, Vin92, Voe94, Vuk98, Wad99, Wal92a, Wal92b, Wan94, War90, Was91, Was97, Wat90, Wat96, Web90]. Review [Weh91, Weh93, Weh95, Wei90, Wei94, Wel90, Wel93, Wen93, Wen94, Wes92, Wha95, Whe93, Wil90a, Wil90b, Wil98, Win90, Win93, Wit94, Wol91, Wol96, Wol92, Won94, Won98, Woo91, Woo92, Woz91, Wu99, Wyc96, Yak97, Yan96, Yan97, Yeh93, Yin91, Zei94, Zel94, Zha92, Zie90l, Zie90p, Zie90x, Zie90i, Zie90m, Zie90o, Zie90v, Zie90h, Zie90q, Zie90j, Zie90z, Zie90e, Zie90y, Zie90s, Zie90d, Zie90f, Zie90n, Zie90t, Zie90r, Zie90w, Zie90u, Zie90b, Zie90c, Zie90a, Zie90k, Zie90g, Zie91g, Zie91n, Zie91i, Zie91m, Zie91p, Zie91l, Zie91k, Zie91f, Zie91e, Zie91a, Zie91c, Zie91j, Zie91d, Zie91b, Zie91h, Zie91o, Zie92b, Zie92j, Zie92f, Zie92a, Zie92t, Zie92h, Zie92g, Zie92d, Zie92q, Zie92k, Zie92e, Zie92c, Zie92x, Zie92i, Zie92r, Zie92w, Zie92v, Zie92p, Zie92m]. Review [Zie92l, Zie92n, Zie92o, Zie92u, Zie92s, Zie93-30, Zie93e, Zie93c, Zie93j, Zie93t, Zie93b, Zie93-28, Zie93-31, Zie93o, Zie93w, Zie93l, Zie93-37, Zie93p, Zie93v, Zie93h, Zie93d, Zie93x, Zie93-33, Zie93r, Zie93-29, Zie93m, Zie93z, Zie93-27, Zie93-38, Zie93a, Zie93u, Zie93q, Zie93k, Zie93g, Zie93f, Zie93i, Zie93s, Zie93-32, Zie93-36, Zie93n, Zie93-34, Zie93-35, Zie93y, Zie94o, Zie94u, Zie94e, Zie94h, Zie94f, Zie94k, Zie94-42, Zie94d, Zie94w, Zie94r, Zie94t, Zie94x, Zie94-28, Zie94-29, Zie94j, Zie94a, Zie94v, Zie94-30, Zie94-41, Zie94-39, Zie94y, Zie94-31, Zie94g, Zie94l, Zie94-44, Zie94b, Zie94s, Zie94-32, Zie94-40, Zie94-38, Zie94-43, Zie94-33, Zie94-36, Zie94-37, Zie94-27, Zie94z, Zie94i, Zie94p, Zie94c, Zie94m, Zie94n, Zie94-34, Zie94q, Zie94-35, Zie95q, Zie95-48, Zie95p, Zie95-30, Zie95u, Zie95r, Zie95-33]. Review [Zie95g, Zie95-42, Zie95-29, Zie95d, Zie95-41, Zie95i, Zie95t, Zie95-35, Zie95-45, Zie95-31, Zie95o, Zie95-38, Zie95e, Zie95-32, Zie95b, Zie95h, Zie95-40, Zie95z, Zie95s, Zie95y, Zie95v, Zie95f, Zie95l, Zie95n, Zie95j, Zie95x, Zie95w, Zie95-37, Zie95-47, Zie95-46, Zie95-28, Zie95m, Zie95-39, Zie95-36, Zie95-27, Zie95-34, Zie95k, Zie95c, Zie95a, Zie95-44, Zie95-43, Zie96m, Zie96-47, Zie96-34, Zie96-32, Zie96o, Zie96q, Zie96b, Zie96-28, Zie96t, Zie96d, Zie96-45, Zie96w, Zie96-30, Zie96n, Zie96-37, Zie96-41, Zie96e, Zie96-31, Zie96-44, Zie96-50, Zie96j, Zie96-43, Zie96i, Zie96h, Zie96-40, Zie96-27, Zie96a, Zie96x, Zie96-33, Zie96-35, Zie96k, Zie96-51, Zie96g, Zie96-48, Zie96-29, Zie96-39, Zie96-42, Zie96p, Zie96-38, Zie96r, Zie96s, Zie96c, Zie96u, Zie96l, Zie96-46, Zie96f, Zie96z, Zie96-49, Zie96-36, Zie96v, Zie96y, Zie97-39, Zie97b]. Review [Zie97h, Zie97-32, Zie97-27, Zie97-38, Zie97m, Zie97-29, Zie97s, Zie97r, Zie97-34, Zie97-28, Zie97l, Zie97e, Zie97u, Zie97f, Zie97x, Zie97p, Zie97-36, Zie97-33, Zie97-42, Zie97-35, Zie97o, Zie97v, Zie97c, Zie97d, Zie97t, Zie97-37, Zie97-43, Zie97q, Zie97j, Zie97-41, Zie97z, Zie97g, Zie97-30, Zie97a, Zie97w, Zie97-31, Zie97k, Zie97i, Zie97y, Zie97-40, Zie97n, Zie98o, Zie98-42, Zie98-45, Zie98-48, Zie98-39, Zie98h, Zie98-34, Zie98-41, Zie98-32, Zie98-27, Zie98a, Zie98-38, Zie98t, Zie98s, Zie98y, Zie98j, Zie98n, Zie98-35, Zie98q, Zie98p, Zie98-29, Zie98-47, Zie98v, Zie98-31, Zie98-33, Zie98-46, Zie98c, Zie98f, Zie98-43, Zie98l, Zie98m, Zie98-36, Zie98-44, Zie98-28, Zie98-40, Zie98g, Zie98-37, Zie98-30, Zie98i, Zie98r, Zie98x, Zie98d, Zie98e, Zie98k, Zie98z, Zie98b, Zie98w, Zie98u, Zie99-27, Zie99l, Zie99-42, Zie99k, Zie99-38]. Review [Zie99-31, Zie99-47, Zie99w, Zie99f, Zie99-45, Zie99-46, Zie99z, Zie99-49, Zie99-43, Zie99u, Zie99i, Zie99-28, Zie99-34, Zie99v, Zie99-29, Zie99-33, Zie99-37, Zie99e, Zie99-39, Zie99j, Zie99-36, Zie99s, Zie99c, Zie99-51, Zie99d, Zie99t, Zie99p, Zie99q, Zie99r, Zie99-35, Zie99-50, Zie99a, Zie99b, Zie99o, Zie99y, Zie99n, Zie99-44, Zie99m, Zie99g, Zie99-48, Zie99h, Zie99-30, Zie99-41, Zie99-32, Zie99-40, Zie99x, Zin93, dV92, dV93, dC99, vN90, vT91, vT95]. Reviews [Pro92a, Zie90-28, Zie90-27, Zie91q, Zie91r, Zie94-45, Zie95-49, Zie95-50, Zie96-52, Zie97-44, Zie99-52]. Reweighted [Mar96]. Reyment [Lyn97b]. Rhinehart [Jan94]. Rhyder [O’D99]. Rice [Zie95v]. Richard [Ara92, Bay99, Boo93, Bow91, Bro92, Bry99, Fox97, Glo99, Gru95, Gut92, Hil95, Jon91a, Jon92, Kin93, Kle94, Kot92, Kra90, Lei97, Lor99, McC99, Nem95, Pot93, Pro93a, Sar90, Sco91, Sen94b, Sha97b, Smi93, Tho90b, Zie91j, Zie94o, Zie94t, Zie99c, Zie99h]. Richardson [Gas97]. Ricky [McL95]. Ridge [Pet93]. Rieder [Zie97z]. Riedwyl [Haw90]. Right [Ada91]. Rijckevorsel [Jol90]. Rindskopf [Wat90]. Rinne [Jai95]. Ripley [Ful97, Jam96, Koe92]. Rising [Zie90w]. Risk [Geo94, Zie94q, Zie95m, Zie98y, Zie99-28, Zie99-34, Zie97-29, Liu99, Zie91n]. Risks [ACV99]. Rivoirard [Cam96]. Road [Zie93i]. Roadmaps [Zie95-32]. Robert [Arm93, Aye98, Ben93, Bow91, Bri96, Bru92b, Cha98, Gan97, Got92, Gri95, HH96, Har91a, Jan94, Liu92b, Loh94, Pan92, Pro90, Pro92b, Rig92b, Sar93, Tam96, Wil90b, Win90, Won98, Zel94, Zie90-28, Zie90-27, Zie90c, Zie95-37, Zie96f, Zie98l, Zie99-42, Zie99v, Zie99-52, dV92, Zie99j]. Roberto [Zie96l]. Roberts [Zie90-27, Zie93h, Zie94-37, Zie97w]. Robertson [Sho90]. Robin [Por97a]. Robot [OUH93]. Robust [Alt96, CV97, EH96, Hor98, IHMT95, LN95, Li97, MSH94, Men91, Pie90b, STW91, Tat97, WVYB97, WT98, Zie91l, Zie93k, Zie98-47, Fai91, Har91a, Mi99, Pri98, Tan98, Zie97z]. Robustness [BL97, TWN98, Zie94-29, Boo99, Roh91]. Rock [Zie90m]. Roe [Gal93]. Roebuck [Sch91a]. Roger [Ang91c, Bax96b, Ber92a, Loh94].

Rogerson [Zie97-30]. Roland [Dun96, Zie91h]. Role [Zie90y]. Rolf [Emp97, Kel97]. Ron [Fre99, Mas99]. Ronald [Cab94, Cha92, Dea97, Gri96, Han93, Hem99, Kah94, Mor92, Nip92, Rob90, Zie98h, Zie98-27, Zie99-30]. Ronchetti [Zie94-29]. Root [Zie94-36]. Rosenbaum [Zie96-28]. Rosenfeld [Zie99q]. Rosenkrantz [Taa99]. Ross [Len93, Zie99-29, Bab98, Fre97, Kus92, Lor92, O’D97, San92, Zie95n, Zie97-39, Zie97-28, Zie98a]. Rossi [Zie98l]. Rossman [Mul97]. Rotatability [DP90]. Roughness [Qiu95]. Rousseeuw [Had92]. Route [Kal93]. Routines [IHMT95]. Row [Tür94]. Row-Column [Tür94]. Roxy [Zie99-37]. Royall [Glo99]. Ruanaidh [Rie97]. Rubinstein [Goo95]. Rudas [Zie99x]. Rudolf [Ale91, Alt99, Dav99]. Rule [VB94a]. Rules [Wad99]. Run [AWS94, Fel94, God93, Luc94, Mor99a, Ngu96b, WMP94a, WMP94b, Zie92i]. Run-Length [AWS94, Fel94, Luc94, WMP94a, WMP94b]. Runger [Zie99l, Ale95]. Runs [Cro92, WW92]. Ruppert [Arm97, Wei90]. Russek [Zie99t]. Russek-Cohen [Zie99t]. Russell [Jan94, Zie99t]. Rustagi [Fot95]. Ryan [Coo96b, Lam90, Osb97].

S [Aba98, Alb93, Arn90, Bas90, Bat99, Bli91, Bre90, Bru96, Buc90, Bzi97, Cal94, Cam90, Cha93, Chu91, Col92, Cuc92, Fot95, Fox98, Fre99, Fre93, Fre94, Gas97, Gre98, Gue90, Gun94b, Hin97, Huz95, IHMT95, Iye90, Kal98, Kia92a, Kin97, Ko95b, Ko95a, Koh94, Kus92, Lee93, Lee99, Len95, Mac90a, Mas99, Maz98, McC90, Moo96, Nai94, Nan94, Oli98, Pat91, Pig91, Ram96, Rea93, Sam99b, San92, Sco90, Sea94, Sto96, Taa91, Tho90a, Wal92b, Was97, Wol92, Yan97, Zha92, Zie90-27, Zie90h, Zie90u, Zie92g, Zie92k, Zie92m, Zie93r, Zie93k, Zie93s, Zie94h, Zie94-28, Zie94-29, Zie94-39, Zie94-45, Zie94p, Zie94c, Zie94q, Zie95d, Zie95-38, Zie95-32, Zie95-40, Zie95l, Zie95j, Zie96k, Zie96-48, Zie96-39, Zie97-34, Zie97-28, Zie97-41, Zie97-30, Zie98o, Zie98-38, Zie98p, Zie98-29, Zie98-33]. S [Zie98-46, Zie98l, Zie98m, Zie98d, Zie98w, Zie99z, Zie99s, Zie99o, vT95, Han98, Hes93, Jam96, Ko95b, Ko95a, Zie98-46, Zie99m, Zie98-38]. S-PLUS [Han98, Jam96, Zie98-46, Ko95b, Ko95a, Zie99m]. Sacco [Zie98b]. Sachs [LC98]. Safeguards [War90]. Safety [Boo90, Tor99]. Sahai [Zie96-43]. Saichev [Sha97a]. Sall [Zie99r]. Sally [Nai96]. Salsburg [Zie93-31]. Sample [CWWR97, CCC90, HP97, HC91, Hul92, Mad98, Ste87, Wil90a, Zie94c, Ste90b, Emp93, Min93, Sea93]. Sampled [ME99]. Samples [Mor93a, Sin83, Sin92]. Sampling [Bil92, Bil94, BL94, GS92, HR92, HL95, Hea90, Hes95, KD97, LS90a, McA90, MT91, Mor99b, Mor91, MT92, PH99, RAA90a, RAA90b, Ste87, Woo90, Zie93-33, Zie95z, HR94, Ste90b, Bru96, Por97b, Ruk98, Sto96, Tan93, Zie90r, Zie90k, Zie93j, Zie93-32, Bil94, Zie92m, Zie93y]. Samprit [Maz93]. Samuel [Ber96b, Bru94, Car93, Cha95, Col92, Fam95, Loh92, Pos98, Zie98k, Zie98z, Zie99-31]. Sander [Zie96e]. Sanders [Per97, Fre92]. Sandifer [Zie99k]. Sandra [Hei90, Zie99-35]. Sanford [Gol95]. Sanjit [Sen94a]. Santner [Sim91, Tam96]. SAS [Ale91, Hec90, Ste99, Tho92b, Zie95u, Zie95b, Zie96-45, Zie96i, Zie96x, Zie96-33, Zie97-38, Zie97-36, Zie98h, Zie99-31, Zie99-47, Zie99r, Zie99g, Zie99-48, Zie96w, Zie96x, Zie97m, Zie97-36, Zie97x, Zie97p, Hei90, LaL92, Zie92n, Zie92o, Zie93z, Zie95-32, Zie96-36]. SAS-Oriented [Ste99]. SAS/Graph [Zie96-45, Zie96-33]. SAS(R) [Hec96, Zie96d, Zie96w, Zie97-32, Zie97m, Zie98r, Zie97-27, Zie99-35]. Sastry [Zie99d]. Satellite [CC95]. Satisfaction [Zie97i, San98, Zie95w, Zie97s, Zie98-31]. Saturated [HO95, Kun97, Lin93a, Lin95a]. Saunders [RN91, Zie95p]. Saville [Mil98, Zie93-29]. Sawyer [Zie98v]. Scale [HP97, MM94]. Scaled [KS93]. Scaling [Jol96]. Scatterplots [Deu96]. Schader [Zie90g]. Scheaffer [Zie95-28, Jon91a]. Scheiner [Zie95j]. Schemes [Bax90, DP91, FHT90, Hun90b, LS90b, LS90c, Luc98, LG99, MH90, SWC97, TWN98, WM90, Yas92, Yas93]. Scherkenbach [Zie93i]. Schiff [Goi96]. Schilling [Ram91]. Schimek [Zie98-41]. Schleifer [Law98]. Schlotzhauer [Hei90, Zie99-35]. Schmidt [Pan92, Zie96-48, Nem92, Zie94h]. Schneider [Zie94r]. Schofield [Zie98u]. Schucany [Zie97-34]. Schuler [Zie93p]. Schum [Zie98b]. Schürmann [Sim97]. Schurz [Hal95]. Schwartz [Zie98-43]. Science [Alm95, Bax92, Bro90b, Gla98, KY91, Nem95, Zie91g, Bow91, Gut90, Nar93, Zie95-38, Zie96-39, Zie98z, Zie99z]. Sciences [Zie90h, Zie92g, Zie98o, Zie98p, Cro99, Got92, Hes91, Jen90, Lyn97b, Rag91, Ram94, Ram96, Rus94, Sha97a, Wal92a, Wat96, Zie90z, Zie91r, Zie92d, Zie92l, Zie93-28, Zie94o, Zie94w, Zie94-40, Zie96t, Zie96-38]. Scientific [Gib91, Tho90a, Zie96-27, Bro92]. Scientists [Box90, Gut92, Hah90, HK90, Hoo90, Hun90a, Mac90b, Moo90, Taa99, Tri90, Wan90, God90, Bar91, Bzi97, Rob90, Tho99, Zie93c, Zie94-31, Zie96-42, Zie99p, Zie99-30]. Score [Goo95]. Scorecard [Zie98-45]. Scott [Cal98, Zie90-27, Zie94k, Zie96c, Ram93]. Scram [MPR99]. Screening [Mee90a, Tam96, WBS^{+}92]. SDE [Hal95]. Searle [Gho93, Zie99s]. Seasholtz [Tob99]. Seber [Ber91, Por97b]. Second [BA96, Dri91, Gra94]. Second-Order [BA96]. Seeber [Zie96-40]. Seeing [BB95, Fam97]. Seheult [Die92]. Seismic [Ano00, SBD98]. Selected [Roe91, Zie99-36]. Selecting [ES94, Zie95-33]. Selection [BL94, CHW97, Len95, Sch92a, Tam96, Zie94-27]. Self [Car97, Kal93, Kia95a]. Self-Instructional [Kal93]. Self-Learning [Car97, Kia95a]. Selvin [Zie95-36, Zie95-27]. Semiparametric [Zie99-27]. Semivariogram [ZZ91]. Sen [Chi91, Zie94u]. Senn [Zie94n, Zie99-39]. Sense [Das93]. Sensitivity [ET96, Goo95, Ney94, STC99, Tho94, YE94]. Sequences [Hol98]. Sequential [Ber92a, SGW96, Zie94u, Zie94-37]. Serial [MQ91, Zie94z]. Serially [Yas93]. Series [Ang92, BH97, Boo94b, Boo97, Cab98, Chu91, Dic91, Dow95, Elt90, Ens95, Fre95, Har92a, Jol93, Li97, MW90, Nip92, Ord96, Rag98, RM95, Sha97b, Ste91, Tim95, Tso92, Viv91, Wel90, Win90, Zha92, Zie90c, Zie92v, Zie94t, Zie94x, Zie95l, Zie95-46, Zie97l, Zie97-42, Zie98f, Zie98g, Boo92, Dic97, Ped98, Per98, Zie92p, Zie95e, Zie97-40]. Series/Parallel [MW90]. Service [Zie97k]. Services [Gun96b]. Sets [Kan95, Zie95a]. Setting [Cam98]. Settle [Zie99-52]. Several [Spu92]. Severity [BCL98a, BCL98b, Rob98, Sch98a]. Sevin [Gun90]. Seymour [Law95]. Shachter [Len93]. Shanubhogue [Bas97]. Shao [Hes97]. Shape [KGK90, Won92]. Shapiro [Zie98c, Col92, Goo95, Zie94-39]. Sharma [Mar97]. Sharon [Zie99-30]. Shayle [Gho93]. Shearer [Gun96b]. Sheather [Har91a]. Shecter [Rea93]. Sheila [Bel99, Zie97s]. Shein [Dea95]. Shein-Chung [Dea95]. Sheldon [Bab98, Fre97, Lor92, O’D97, Zie98a, Zei94]. Shelemyahu [Fre99, Rig93]. Shelley [Sau94, Zie98-38]. Shenton [Sin90a]. Sherlock [Sch91a]. Shewhart [Ada96, BC95, HD95, PD95, RD95b, SWG95, RD95a]. Shewhart-Type [BC95, HD95, PD95, RD95b, SWG95, RD95a]. Shiavi [Sen94b]. Shiffler [Zie95-50]. Shift [HC91]. Shihadeh [vT95]. Shilliff [Zie93-34]. Shipley [War90]. Shipp [Zie96-33]. Shirland [Sar94b]. Shirley [Ing92]. Shiue [Zie96-27]. Short [Cro92, Mor99a, God93]. Should [Car90b, Car92, Geo94]. Shrinkage [Fam99]. Shumway [Win90, Zie90c]. Sided [BT89, Van92b, BT98, Ode90]. Sidney [Wen93]. Siegel [Zie97-33]. Sigma [Cab94]. Signal [Gru97, Rie97, Sen94b, Sen94a, Sen95, Rie94b]. Signals [Ano00, Kun93, ME99, SBD98]. Significance [Nay98, Zie99-40]. Sikorski [Mor93b]. Silverman [Die92, Geo98, Qiu95]. Similarities [Gol98]. Similarity [Jol97, Koh90]. Simiu [Zie95o]. Simon [Han90, Har91a, Zie91b]. Simonoff [Yan97]. Simple [Bru91, DJ94, LN95]. Simpson [Dri92]. Simulated [Deu96]. Simulation [Boo94b, Dea91, Dow94, GD94, Got94c, Got94b, Jou94, Par92, Rut94, Tor94, WB94, Law92, Lor92, Zie98-32, Bre97, Dea94, LeP90b, Lor93, Zie97-29, Zie97-28]. Simulations [Ste87, Ste90b, Zie97n]. Simultaneous [CM99, Dal90, DM99a, DM99b, Gib99, Gil99, Mee90b, MER91, Spu93, Van99, Wes90]. Sincich [Zie94-40]. Singh [Zie90-28]. Single [Sch91c]. Single-Stress [Sch91c]. Singpurwalla [Zie96h, Zie98l]. Sinha [Roh91, Zie93j]. Site [Dow94, GD94, Got94c, Got94b, Jou94, Rut94, Tor94, WB94]. Situations [BC95, HD95, PD95, RD95a, RD95b, SWG95, Tsa91]. Siu [Nay98, Zie99-40]. Sivia [Fox98]. Six [Bog92, Cab94]. Size [Emp93, HC91, Hul92, Ngu96b, Sea93, Zie94c, Zie90z, Wil90a]. Sizes [Kva96]. Skiadas [Zie96a]. Skip [Bru96]. Skip-Lot [Bru96]. Small [Ano90f, DL90b, SS89, WW92, Zie95a]. Small-Disturbance [Ano90f, SS89]. Smeyers [Zie98w]. Smeyers-Verbeke [Zie98w]. Smith [Bar92a, Cro99, Deh90, Fox91, Gun99, KY91, Sam99b, Ter92, Zie95g, Zie96s, Zie97-34, Zie98h]. Smooth [Kin91]. Smoothing [Ale99, DeV99, Deu96, HT92, Yan96, Zie98-41, Yan97]. Smyth [Zie98c]. Snell [Dri92, Zie91q, Zie92s, Zie93-30]. Soares [Zie98s, Zie95r]. Social [Got92]. Society [Lew92, Wen94, Zie99-33]. Software [DM98, Sam99a, Soy95, Zie95-32, Zie95s, Zie96-50, Zie98t, Zie96d, Zie96-45, Zie96-33, Zie99-31, Zie99-47]. Soil [Coh92]. Solution [Hal95]. Solving [Don90, Fri97, MA92, Pit92, Zie96-31, Bog92, Bru95, Hey93, Mor92, Pro93a, Zie97a]. Som [Mor99b]. Some [ACV99, Bak96, CL93, DP91, FF93b, FF93a, HM93, KM94, LN95, Ros96, Tsa91, ZM94]. Song [Dea95]. Song-Gui [Dea95]. Source [AHT91]. Southern [GK93]. Space [Zie92p, Zie94z, Cox93]. Spagon [Zie98-40]. Spall [Wel90]. Span [Kim91]. Sparsity [Kun97]. Spatial [GC91, Koe92, Nip92, TH93, VD94, Zie94-41, Zie96-50, Zie98-38, Zie98-33, ZZ91, dV93, Cob91, Got92, Zie97-30]. Spatially [HNF97, Zie99t]. SPC [Boo94a, Hea94, Hem99, Jen92, Pro93b, Wan94, Zie94r, Cro92, HP92, NM95, God93, Mil94, Hal98, Hal94, Kra92, Kus93, Mor99a, Owe91, Pro94, Sar93, Woo92, Zie95-35, Zie95-34, dC99]. SPC/TQM [Pro93b]. Special [AWS94, Fel94, Hea94, Kra92, Luc94, WMP94a, WMP94b, Zie91p]. Special-Cause [AWS94, Fel94, Luc94, WMP94a, WMP94b]. Specifications [Bak98, Zie97-31]. Specified [LIV96]. Spector [Ko95b, Zie92n]. Spectral [Hag96, Kii92, Lau94, Nan94, Per92, Per98, Cox96]. Spectroscopy [Zie95y, Goo92b]. Spectrum [Cox93, Gru90]. Speech [JR91]. Speed [Ste91]. Spherical [Bor95a]. Spiegelhalter [Ful95, Gas97]. Spiegelman [Zie92c, Zie94k]. Spizzichino [Gru95]. Spline [ME99, Gal92]. Splines [Zha94, Ize90]. Split [BS99, HCV98, MB98]. Split-Lot [MB98]. Split-Plot [BS99, HCV98]. Sposito [Bro90a]. Spreadsheet [Zie98-39]. Sprent [Gas94, Zie90a]. Spurrier [Mag96]. Squared [Kin97, VM91]. Squares [BBT90, Mar96, Mor98a, Thi91, Bay92]. Srivastava [Bil91, Chi91, Zie90s]. Stability [RH92]. Staff [Zie96w, Zie96x, Zie97m, Zie97-36]. Stage [HLD98, ST90a, Sto92]. Stahel [Zie93k, Zie94-29]. Stamatis [Sch96b]. Stamping [JS99]. Standard [BA96, Spu92, Tat97, TW96, Tsa91]. Standard-Error [TW96]. Standards [Cam98]. Stanely [Cha93]. Stanford [Zie95-38]. Stangl [Zie97d]. Stanislav [Zie99-46]. Stanley [Ing92, Zie94s]. Stanton [Che92]. Stapleton [Dea96]. Start [Bor97a, Zie97x, Zie99r]. Started [Zie98v]. Stat [Amm92]. STATA [Bie92]. State [Cox93, Zie94z, Zie92p]. State-Space [Zie94z]. Stationary [CCC90, Luc98, Zha98]. Statistic [Ano00, SBD98]. Statistical [Ada94, Ada96, Alt99, Amm92, Bar91, Bas99, Bel95, Ber96a, Boo98, Boo99, BK92a, BK92b, Bre95b, Bro90a, Bru91, Bru92a, Cam98, CFRH99, CHB90, Cha98, Cly94, CR96, CC95, Cro99, Cry90, Das93, Dea96, Der90, Dic97, DP91, DGTZ97a, DGTZ97b, Dun97, EM99, FF93b, FF93a, FB94, Ful95, Gas94, GR97, Glo95, Gun94a, Hal94, Han93, Hap92, Har91c, HM93, Hei90, Hes91, Hin97, Hul97, Kal93, KMP95, KB94, Kat98, Kia92a, Ko95a, LaL95, LSBW95b, LSBW95a, Law92, Law96, Len93, Len95, Lor93, LPY97, Mac92, MA92, Mat97, McL95, Mee91b, Mi99, Nem97, OM96a, OM96b, Oli97, Per97, PE97, Pro92b, Ray93, Rig93, Roh91, Ruk97b, Sai96, Sau92, Sau98, Sch95, Sen94b, Sen95, Sha98, Sho90, Sim97, Sim91, Spe96b, Sri93, Swe96, Tam96, TI99]. Statistical [Tho90b, Tuc92, TFV93, Tuc96, VTFD92, VV92, Van96b, WG92, War90, Win90, WH97, Woo91, Zha98, Zie90p, Zie90h, Zie90z, Zie90y, Zie90a, Zie91a, Zie91j, Zie92b, Zie92g, Zie93z, Zie93-38, Zie93s, Zie94e, Zie94f, Zie94-28, Zie94-29, Zie94j, Zie94a, Zie94-34, Zie95-29, Zie96j, Zie97p, Zie97o, Zie97-44, Zie98o, Zie98-35, Zie98q, Zie98p, Zie98d, Zie98z, Zie99-31, Zie99u, Zie99c, Zie99p, Zie99-35, Zie99n, Zie99g, Zie91r, Acr97, Ada93, Ang91c, Arm93, Bar92a, Bas97, Beg90, Boo94a, Bor97b, Bow91, Boy90, Bru93b, Bry92, Bzi97, Che91, Coa94, Coh92, Col93, Dov93, Edg98, Fal99, Fri97, Gho94, Glo99, Got94a, Gru97, Has91, Hec99, Hes93, Hey93, Hil95, Jac95, Jai95, Jon91b, Jon93, Kan95, Kat91, Kin93, Kle94, Koe92, Koo90, Kra90, Lam90, Lei99, Lor99, McG94]. Statistical [Mil99, Mil98, Moo91, Mor93b, Nay98, Nel98, O’D97, Ott90, Per92, Pfa99, Pig91, Pol93, Por97a, Pro93a, Ram96, Run90, Sar94b, Sha93, Spa99, Spe91, Vuk98, Weh95, Won94, Wu99, Wyc96, Zel94, Zie90v, Zie90t, Zie90u, Zie91r, Zie91o, Zie92q, Zie92r, Zie92s, Zie93-30, Zie93-28, Zie93-29, Zie93-27, Zie94-39, Zie94-38, Zie95-38, Zie95-37, Zie95-39, Zie95-36, Zie96-37, Zie96-41, Zie96-40, Zie96-39, Zie96-38, Zie97-32, Zie98-41, Zie98-40, Zie99-38, Zie99-37, Zie99-39, Zie99-41, Zie99-40, vT95]. Statistician [Aba98, Don90, Not92c, Zie96-31, Zie98-30]. Statisticians [Arm90, Box90, God90, Hah90, HH98a, HH98b, HK90, Hoo90, Hun90a, LeP90b, Mac90b, MTV98, Moo90, Nem98, Tri90, Wan90, Eas90]. Statistics [Ale91, Ale95, Bal93, Bax92, Bax96a, Ber92b, Bod97, Bro90b, Bry99, Bul96, Cly94, Cro99, Dri92, Dun90, Far92, Fox91, Fre99, Gal93, Gue90, Ham96, Har91c, Haw90, Jam96, Jen90, Kel97, LC98, Lee94, Lin99, Lit91, Liu92a, Mag96, Mi97, Mul97, Nar93, Nem92, NO96, Pan96, Rie94b, Rob90, Sar94a, Sau94, Sau97, Sch96a, Sin90b, Smi93, Sne98, Spr91, Ste98, Ste96a, Ste96b, Ste99, Taa99, Wal92a, Was91, Wat90, Wei94, Yeh93, Zie90k, Zie91m, Zie91k, Zie92a, Zie92l, Zie93c, Zie93k, Zie94h, Zie94v, Zie94-31, Zie94-45, Zie94g, Zie95v, Zie95n, Zie95-28, Zie95-50, Zie96d, Zie96-30, Zie96h, Zie96-52, Zie97o, Zie97c, Zie97z, Zie98h, Zie98j, Zie98n, Zie98l, Zie98k, Zie99k, Zie99b, Zie99-30, dVU96, Zie90-27, Zie95-50, Zie96-52, Ajm99, Ang90b, Arm92, Arn90, Ben93]. Statistics [Bli91, Bro90b, Buc90, Bul91, Dom90, Dre91, Fam97, Far90, Far93, Fot95, Goi96, Goo91, Gra94, Gri96, Gro92, Gun90, IHMT95, Jan94, Joh91, Lal99, Lit91, Mee91a, Moo96, Rig92b, Sch91a, Str96, Wei94, Wil90b, Wol92, Woz91, Yan97, Zie90s, Zie91b, Zie93e, Zie93m, Zie94d, Zie94y, Zie95b, Zie95f, Zie95c, Zie96p, Zie97-43, Zie98-34, Zie98-33, Zie98x, Zie98e, Zie99-27, Zie99l, Zie99w, Zie99f, Zie99r, Bie92, Boo93, Bri96, Bru95, Eas99, Fre92, Gat96, Gib91, Glo97, Goo92b, Har91b, Hes91, Ing92, Kot92, Lar95, Lei92, Loh94, Lyn94, Mye91, Rus94, Wen94, Wil98, Zie90x, Zie90w, Zie92t, Zie92u, Zie93-31, Zie94-42, Zie94-41, Zie94-40, Zie94-43, Zie95-41, Zie95-40, Zie95-50, Zie96-43, Zie96-52, Zie96-42, Zie97-34, Zie97-33, Zie97-35, Zie98-42, Zie99-42, Zie99-45, Zie99-43, Zie99-44]. Statistics [dV93]. Stats [Zie99m, Zie98-38]. Staudte [Har91a]. Stechel [Zie96-40]. Stechel-Berger [Zie96-40]. Stefanski [Arm97]. Stein [HH96, Oma91, Won98]. Stenger [Zie93-32]. Step [Hec96, MT92, Sar94a, Zie95b, Zie99-52]. Step-by-Step [Hec96, Zie95b, Zie99-52]. Step-Loss [MT92]. Stepanski [Zie95b]. Stephanie [Zie99q]. Stephen [Bay95, Bru95, Cal94, Nem92, Spa99, Zie90y, Zie94s, Zie94n, Zie99-28, Zie99-39]. Stephens [Bru96, Sto96]. Stepping [Ano90e, Sim88]. Steps [Bog92]. Stepwise [VS98]. Steve [Zie95-36, Zie95-27]. Steven [Kus93, Pan92, Por97b, Sar93, Sen95, Zie93y, Zie99m]. Stevens [Bal93].

Stinson [Zie98-47]. Stochastic [Ano97e, Bax95, Fre94, Goo95, Huz95, Ked98, LH97, Luc98, Rag91, Rag93, SN95, Tor99, Weh91, Weh93, Wen93, Zie96a, Fot96, Fre93, Fre97, Ked96, Kir92, Mad96, Zie96-44, Zie99-46]. Stock [Zie97-36]. Stokes [Zie96i, Zie98v]. Straight [Aye98]. Straker [Zie97a]. Strata [Zie90h]. Strategic [Zie99-47]. Strategies [KTTS90]. Strategy [STR99]. Stratton [Zie99a]. Strength [WJ92]. Stress [DH92, Sam99a, Sch91c, WJ92]. Stress-Strength [WJ92]. Strip [Mil97]. Strip-Plot [Mil97]. Stroup [Zie97-27]. Structural [Ang92, Gri97a, Hec96, Zie97b]. Structure [LBB96, Rag95]. Structures [OUH93, Sha97c]. Stuart [Stu94]. Studentized [PF91]. Studies [BL97, KTTS90, Nem95, RH92, VV99, Zie96h, Zie98l, Zie98-40, Zie99f, Gle91, Jen93, Zie90i, Zie94p, Zie94c, Zie96b, Zie96-28, Zie99-37]. Study [Kaf94, RH95, Sto92, YE94, Ang91b]. Studying [Moo99]. Subba [Zie98g]. Subhash [Mar97]. Subir [Zie91o]. Subject [USM99, Yas95]. Subjective [Kat98, Zie95-42]. Subrahmaniam [Wel93, Rab92]. Subset [Bre95a, Roe91, Sch92a]. Subset-Selected [Roe91]. Subsets [STR99]. Subspaces [CC95]. Subsystems [MW90]. Success [Zie97v, Neu99]. Sudhakar [Cox93]. Sudman [Zie93r]. Sufi [Elt90]. Suhir [Tho99]. Sum [SGW96]. Summary [HH98b]. Summed [BBT90]. Summers [Maz98]. Sung [Tan98]. Supersaturated [ACV99, LW97, Lin93b, Lin95b, Ngu96a]. Supplement [Gun90, Zie92g]. Supportability [Lyn95]. Surface [BA96, CCC90, DL90b, Gun96a, Khu92, MME98, MMY93, Pro92a, SE93, VM91]. Surfaces [And90, Zie90e, Zie97y]. Survey [Min93, Zie90-28, Zie96-47, Zie96g, Zie97q, Zie99-52, Coh92, Zie90-28, Zie93-33, Zie93-32, Zie98-43, Zie99-48]. Surveys [Zie99-52, Zie90-28, Mad98, Sau97, Zie93t, Zie93r, Zie96-29, Zie99v, Zie99q]. Survival [CR93, Gol99, WL91, Zie93-27, Zie95x, Zie96b, Zie97r, Zie98r, Car97, Rao98, Zie97-38, Zie97-37, Zie98-44]. Survivors [Zie98-44]. Susan [LC94]. Suter [Zie95m]. Sutherland [Kin93]. Sweeny [Pan96]. Sweeting [Bar92a]. Swift [Zie99-29, Boo94b, Dun97]. Sybil [Zie91g]. Sylvie [Bor97b]. Symmetric [Loh92]. Symposium [Bro90b]. Synthesis [Emp97, Len93, Ano97f, Ano97f, Ano97e, Emp97, Zie95-44]. System [Ano98e, DP97, EMSD91, Gri97b, Hec96, Hei90, Kva96, Rag98, RM95, Taa91, Viv91, Zie92n, Zie92o, Zie93z, Zie95b, Zie97-27, Zie97-36, Zie99-35, Ale91, Hec90, Zie95u, Zie96w, Zie96i, Zie96x, Zie97-32, Zie97-38, Zie97m, Zie99-48, Bar92b, Van96b]. System-Based [EMSD91]. Systematic [CM93b, CM93a, FL97, Gun93, HO93, Hah93, LST93, Lin95b]. Systems [Ano97e, AHT91, Bil94, Cox93, Dog99, Goo95, Gri99, Jon92, LT96, Law92, LH97, MW90, Sam99a, SDP^{+}95, SP95, Tor95, Zie99-50, Bro92, Cry90, Fot96, Kir93, Koh90, Rag95, She90a, She90b, Smi93, Van91, Weh93, Dea91].

T [Aba98, Ang91a, Bar92a, Bzi97, Cox96, Cra94, Gal95, Gol99, Hag91, Hey92, Job99, Liu92b, Mi99, Not92a, Not92c, Oli98, Ott90, Owe91, Por97a, Pro94, Rus94, Sac93, Sam99a, She90b, Sho90, Str96, Wen94, Zie90e, Zie93v, Zie93-27, Zie94g, Zie94l, Zie94-40, Zie95y, Zie95l, Zie96u, Zie98-38, Zie98g, Zie99n, Zie99x]. Tabak [Weh91]. Taber [Zie97g]. Tables [Ano90e, Nel93, Sim88, Zie90u, Zie97o, Alb93, Zie99x]. Taguchi [Coo96b, Edg92, Pie90b, Pro92b, She90b, GD91, NAM^{+}92, NW93, Stu94, Zie91p, Zie97-39]. Tague [Zie95-45]. Tai [Loh92]. Takeaki [Roh91]. Talbot [Pit92]. Talk [Aye98]. Tammy [Zie99-51]. Tanabe [Zie99-36]. Tandberg [Zie95y]. Taneja [Joh91]. Tanner [Mee91a, Ruk97b]. Tanur [Zie90w]. Tanya [Zie99-31]. Tapiero [Kal98]. Tarter [Wha95]. Taylor [Fre94, Dov93, Ful95, Lew93, She90a, Tan93]. Team [Mor92, Zie93-34, Jon90, Jon91b]. Teams [Zie94l]. Teboul [Zie93q]. Technical [Dog99, Zie92j, Zie97-36]. Technique [KD97, Pap95]. Techniques [Bay99, DeV99, Dov93, Fot95, Kva96, Par92, Rao98, Tor95, Zie97-39, Zie98r, Zie99-31, GJMS^{+}95, Bre96, Cox96, LC98, Mar97, Mor99b]. Technologies [Sha92]. Technologists [Ram99]. Technology [Ada92, Bon95, LPR92, Nem95, Mar94, Zie91g]. Technometrics [BT98, DPL^{+}94, HR94, Lin95a, Sen99b, Sin92, Ste90b]. Telephone [Zie90-28]. Tenner [Zie93-37, Zie97u]. Term [Zie98-44]. Terms [VM91, Zie96o, Zie91a]. Terraza [Dow95, Ord96]. Terry [San98]. Test [AKO94, AAK99, CWWR97, CC99, Col92, EMSD91, ES94, LN97, MQ91, Mee91b, ML95, Ney94, Rag98, RM95, Sin83, WN97, Wol95, Sin92]. Test-Based [CWWR97]. Testing [Bry92, Cha93, DM98, DH92, GS95, KGK90, KS93, LBB96, Len90, MA92, Mee91b, Nik95, Pol93, Sam90, ST90a, Sch94a, VS96, VS98, WJ92, Zie96-34, Har91a, Hor98, Hul92, Kin97, Spe91]. Tests [Bak96, BE94, CB93, CS93, Dow95, EM95, GK90, HC91, Kin91, MEL98, MM94, Nik95, Ord96, Sch91c, SGW96, YE94, ZM94, Ale99, Roh91, Ste90c, Zie94a, Zie98-37]. Text [Car97, Kia95a]. Textbook [Mac90a]. Their [LLP^{+}99, Zie92a]. Theo [Zie90p]. Theoretic [Str96]. Theories [Bil94]. Theory [Ang90a, Ang90b, Ano90f, Bax92, Bil92, Bor95b, Chi91, Chr97, Col93, Cry90, Dea95, Dog99, Gri99, Har91b, Jac95, Jon92, Kia92a, Lei92, Mas99, Mor93a, Ram93, Rie94a, Sch96b, Sea90, Sen99a, Sha93, Sin97, SS89, Van96b, Won98, Zad95, Zie90v, Zie92v, Zie92s, Zie93-30, Zie95o, Zie95z, Zie97-35, Zie98-41, Zie98d, Zie98e, Car90a, Hag90, Per92, Sai96, Sen95, Sys99, Wu99, Zie94-45, Mad98, Zie99-49]. There [Arm90]. Theresa [Zie99k]. Thermal [Tho94]. Thermal-Battery [Tho94]. Therrien [Mag90]. Thiebaux [Zie96-38]. Think [Bul91, Dom90]. Thinking [Che95, Koo90, Zie99-41, Per97]. Thomas [Car90b, Car92, Cha92, Edg92, Fri97, Gre94, Hea94, Lam90, Osb97, Ruk97a, Sim91, Spe96a, Tam96, Wan94, Web90, Zie95-30, Zie96-45, Zie99q]. Thombs [Mag96]. Thompson [Gho94, Por97b, Boo90, Mad98, Pie90a, Zie93y]. Three [HL97]. Three-Parameter [HL97]. Thriyambakam [Kus98]. Tibshirani [Gri95, dV92]. Tichý [Gri97a]. Tierney [Amm92]. Tijms [Mad96, Spr91]. Tim [Sho90]. Time [Ada91, Ang92, Dic91, Dic97, Dow95, Elt90, Gri99, GSH98, Joh90, Jol93, Kal93, LT96, Li97, LM93, LPY97, Nip92, Ord96, Ped98, Per98, Sha97b, Ste91, Tim95, Wel90, Win90, Zha92, Zie90c, Zie92p, Zie94t, Zie94x, Zie95e, Zie95l, Zie97l, Zie97-40, Zie98f, Zie98g, Boo92, Boo94b, Boo97, Cab98, Chu91, Ens95, Fre95, Har92a, Tso92, Zie92v, Zie94-34, Zie95-46, Zie97-42]. Time-to-Failure [LM93, LPY97]. Times [HLS98, Rag98]. Timm [Zie99-51]. Timothy [Neu99, Zie99t]. Timpone [Zie97g]. Titterington [LH97, Weh93]. Toby [Bre95b]. Todeschini [Zie96l]. Tolerance [BT89, CR91, KM99, Mee90b, MER91, Van92b, WI94, BT98, Bak98]. Tolerancing [Bis97]. Tolley [Kan95]. Ton [Gol98]. Tong [Ize91]. Tonnage [JS99]. Tool [McL95, Zie99d, vN90]. Toolbox [Zie97a, Bel95, SG99, Zie95-45]. Tools [Bor97b, Bre96, FF93b, FF93a, HM93, Mat98, Nem92, Owe91, Pro94, Zie94h, Zie95-31, Bru91, Han93, Zie94l, Zie99-50, Bre90, Ruk97b]. Top [HS97]. Topics [Ada96, Hea94, Ver90]. Total [HH96, Pan96, Zie99-29, Zie92w, Zie93-37, Zie93-36, Zie99-50]. Toxicologists [Zie93-31]. Toxicology [Zie95-41, Zie98-42]. TQM [HH96, Pro93b, Boo94a, Zie95k, Zie94-44]. Trace [McA90]. Trait [Gol98]. Transformation [CR91, KSJ96, Roc93, TW90, TW92, VB94a]. Transformation-Model [TW92]. Transformations [Wei90]. Treatment [LBB96, Spu92]. Treatments [Spu92]. Tree [Jol97]. Trend [AD96]. Trends [Joh90, LT96, MPR99]. Trevor [Hes93, Jol96, dV92]. Trewin [Zie98-43]. Trials [Zie94n, Zie96b, Zie91q]. Trimble [Zie96-51]. Triola [Far90]. Trivedi [Pet99]. Troia [Zie95r]. Troubleshooting [Hem99, Ram91]. True [WI94]. Truncated [Rao98, Mor93a]. Truncation [MJ91]. Truth [Eas99]. Tsong [Kin93]. Tsong-how [Kin93]. Tsui [Lor99]. Tu [Hes97]. Tuck [SDP^{+}95]. Tuckwell [Hag90]. Tufte [Baj98, Sco92]. Tukey [Zie91f, Zie91e, Zie95-43, Sch93c, Sch93b, Zie91f, Zie91e, Zie93e, Zie95-43]. Tumay [Bre97]. Tumbelty [Hal98]. Tunner [Ada92]. Turbulence [Rag93]. Turkman [Zie94-42, Zie99-45, Zie99-44]. Turlach [Hin97]. Turner [Bod97]. Tutorial [Fox98]. Tutorials [Zie94k, Zie92c]. Tutz [Glo95, Zie94e]. Two [AAK99, Ano90e, BT89, BT98, BS99, CWWR97, Car95, DL90a, EM95, FL97, HP97, Hea94, HLD98, HCV98, LIV96, Ode90, RL95, Sch91c, ST90a, Sim88, SS97, VV99, Wes90, WC92]. Two-Component [RL95]. Two-Level [BS99, DL90a, FL97, HCV98, LIV96, Sch91c, WC92]. Two-Sample [CWWR97, HP97]. Two-Sided [BT89, BT98, Ode90]. Two-Stage [HLD98, ST90a]. Two-Way [Ano90e, Car95, Sim88, SS97, VV99]. Type [BC95, HD95, PD95, RD95a, RD95b, Sin92, SWG95, WN97, Yas92, Sin83].

U [Zie96-52, Lei92]. U-Statistics [Lei92]. U.S [Zie95-29]. U.S./Japan [Zie95-29]. Ulf [Kel97]. Unbalanced [KM94]. Uncertainty [Cam98, Kah94, Liu92a, ME96, Zie90p, Zie95g, Zie96n, Zie95-47]. Uncontrollable [RH95]. Undamped [Kun93]. Understanding [Cro99, SDP^{+}95, Arm92, Pan92, Ram99, St.94, Voe94, Zie92x, Zie93-38, Zie96-48, Zie96-49]. Unequal [CC99, Dal90]. Unidentified [Pan99]. Unified [Sam99a, Sim97]. Univariate [Ber96b, Chu91, Cox96, Fam95, HL95, Zie95b, Bru94, Dow95, Zie99-51]. University [Jon90, Jon91b]. Unknown [Dal90, Zie90w]. Unreplicated [BH97, LN97, Pan99]. Update [Zie98o, Zie98p]. Upton [Cob91]. Urban [Len95]. Usama [Zie98c]. Use [Ano91e, Bru91, Dow94, GD94, Got94c, Got94b, Haw91a, Hes91, Jou94, Kun97, Lau94, Lit91, MMY93, Pro92a, Rut94, Tor94, WB94, Zie90z]. Used [CR93, DS93]. Useful [CCC90, Ros96]. User [Lei93, LIV96, Zie95s, Zie98t]. User-Specified [LIV96]. Uses [Nel93]. Ushakov [Pap96]. Using [Ada91, AAK99, Bre95a, BHS90, Dea91, DGTZ97a, DGTZ97b, Gan95, GC91, Gun96a, Hec96, Hoc94, Hul97, JS99, Kan95, Ko95a, Kus93, LM93, Lwi94, Men91, PE97, Sar93, STR99, Ste87, SW95, Tho90b, Van96a, WH97, Yas92, Zie95b, Zie96m, Zie96w, Zie96i, Zie96x, Zie97-32, Zie97-38, Zie97m, Zie97-29, Zie97p, Zie97-36, Zie98i, Zie99-47, Zie99-48, GJMS^{+}95, Ste90b, ZM94, Pan96]. Uthurusamy [Zie98c]. Utility [Nay98]. Utts [Fam97]. Uwe [Tor99].

V [Bas97, Cam98, Fre93, Lar95, Oli98, Pap96, Rie94a, Rig99, Sai96, Sam90, Sen97, Yak97, Zie90o, Zie90q, Zie92e, Zie94-42, Zie97h, Zie97-42, Zie98t, Zie99-45, DL90a]. Vach [Kia95b]. Vajda [Tho90a]. Valerii [Rig99]. Valid [Sch92b]. Validation [Len95, Zie97v]. Validity [Nay98]. Valkó [Tho90a]. Valsecchi [Zie96b]. Value [Ang90a, Cha93, Sne98, Zie95o, Zie96y, Zie98-48]. Valued [Ped98]. Values [Gri90, Kia95b, Nel93, WI94, ZM94]. Vance [MLV92]. Vandeginote [Zie98w]. Vandeginste [Mac90a]. Vanzetti [Zie98b]. Vapnik [Sai96, Wu99]. Vardeman [Spa99, Bru95, Zie95-38]. Vari [Pap95]. Variability [BH93]. Variable [Bay92, CHW97, Coo96a, DH92, Hea90, LS90a, Mee90a, RH95, RAA90a, RAA90b, Woo90, Zie93o, Alt99]. Variable-Selection [CHW97]. Variable-Stress [DH92]. Variables [Ano91f, CS92, ES94, HL95, Haw91b, MT92, OUH93, SD90, SW95, Tho91, Tho92a, Cal98, God95, Zie94-35, vT95]. Variably [KS93]. Variance [Ano99e, BNC87, Gan95, KM94, Kia93a, Kun97, LBB96, Mee90b, Pot93, RWHH97, SS97, Yas94, Bor95a, ZM94, Emp99, Sch93c, Sch93b, Gho93]. Variance-Function [RWHH97]. Variances [CC99, Dal90, Spu92, WN97]. Variate [Gol90, Kii92]. Variation [Lew93, Voe94, Bri96, Cam98]. Variowin [Zie96-50]. Vavra [San98]. Vecchio [Ram99]. Vega [Zie98-38]. Velu [Tim99]. Venables [Jam96]. Verbeke [Ste99, Zie98w]. Verification [Dav99]. Vermunt [Zie98-28]. Veronica [Zie98-40]. versus [Luc98]. Very [KMP95]. Via [Ano90e, Sim88, RM95, DSSU98]. Vic [Bre95b, Len93, Min93, Zie99-44]. Victor [Bru91, Woo91]. Viertl [Spe91]. View [FF93b, FF93a, HM93, Lam98, LSBW95b, LSBW95a, Sim97]. VIII [Zie95-43]. Vincent [Bro90a, Zie99-29]. Vining [Mil99]. Vinod [Zie94v]. Virgil [Gre94]. Vision [Pro93b]. Visual [Baj98]. Visualization [Ram93]. Visualizing [Hec99, Pfa99, Gun94b]. Viswanadham [Kir93]. Vladimir [Van91, Wu99]. Vol [Cob91, KY91, LeP90b, Nem95, Sch92c, Sin90b]. Voltage [HL97]. Volume [Ano90l, Ano91k, Ano92i, Ano93k, Ano94i, Ano95i, Ano96i, Ano97k, Ano98j, Ano99k, Ber96b, Fam95, Hea94, Kia92b, Kia93b, LaL95, Sch93a, Tom95, Wan94, Zie90h, Zie90k, Zie91f, Zie91e, Zie91h, Zie94-45, Zie95-43, Zie96h, Zie98o, Zie98p, Zie98l, Zie98k, vT91, Zie92g]. Volumes [Zie96-47]. Vose [Zie99-34].

W [Beg90, Boo90, Bru93b, Bru94, Bru97, Cal94, Cha92, Che92, Chu91, Col93, DeV99, Dea94, Fox94, Gas97, Gut92, Har91b, Hin97, Jam96, Kia90, Kia95b, Kin93, LeP90b, Len90, Lin90, Lyn99, Mag90, Mar99, Mi99, Nar93, Pit92, Pol93, Pro96, Qiu95, Ram93, Sam99a, Sch93c, Sch93b, Sha95, Sha97c, Thi91, Wit94, Won94, Woz91, Zie90-28, Zie90-27, Zie90w, Zie90g, Zie91f, Zie91e, Zie92c, Zie93t, Zie93w, Zie93k, Zie93i, Zie94k, Zie94-28, Zie94-29, Zie94g, Zie94-40, Zie95p, Zie95e, Zie96s, Zie97-27, Zie97-34, Zie97f, Zie97-44, Zie98-42, Zie98-41, Zie98y, Zie98m, Zie99f, Zie99-43, Zie99c]. Wackernagel [Mye96]. Wadsworth [Bar91, Zie99p]. Wafer [HNF97]. Wah [Lor99]. Wahba [Gal92]. Waksberg [Zie90-28]. Walden [Rus94, Cox96]. Walford [Sch96c]. Walker [Zie99g]. Wall [Has91]. Wallgren [Kel97]. Walpole [Rob90, Zie94-31, Zie99-30]. Walter [Dow95, Ord96, Ped98, Taa99, Wen94, Zie98-42, Zie99f]. Waltraud [Tor99]. Wan [Ale98]. Wand [Yan96]. Wander [Van96a]. Wang [Dea95, Loh92, Das93, Str96]. Wardell [YWMP95]. Wardrop [Bri96]. Warehousing [Zie99-47]. Warranty [KLR91]. Warren [Gun90, Zie99t]. Wasserman [Kia90, Zie92b, Zie97-44, Zie97-44]. Waste [Dow94, GD94, Got94c, Got94b, Jou94, Rut94, Tor94, WB94]. Water [Got94a]. Watson [McL95]. Watts [Mil90]. Wavelet [Han98]. Wavelets [JS99, Zie97-43]. Way [Ano90e, Car95, Sim88, SS97, Van92b, VV99, Far92]. Waymire [Kir92]. Wayne [Dic97, Ham96, Lew93, Mee91b, Tan93, Woz91, Zie95h, Zie97-29, Zie99z]. Weak [Gol93]. Weakest [Wol95]. Weakest-Link [Wol95]. Wearden [Ing92]. Webster [Coh92]. Weerahandi [PGT95]. Weese [Aye98]. Wegman [Bro90b]. Wegscheider [Zie92c, Zie94k]. Wei [Chu91]. Weibull [HL97, MSF95, Zie95-48]. Weight [Lwi94]. Weighted [Bax90, DP91, FHT90, Gan95, Hes95, Hun90b, LWCR92, LS90b, LS90c, MH90, TW90, WM90]. Weighting [CR91, Wei90]. Weir [Zie90-27]. Weisberg [Gol95]. Welbrock [Zie99-47]. Welsh [Bas99]. Wendell [Hey93]. Werner [Zie94l]. Wersberg [Zie93k]. Wesley [Boy90]. Wesner [Zie96-51]. West [Wes92, Zie97e]. Westergren [Zie91c]. Western [Pro92b]. Westfall [Cha93]. Wetherill [Col93]. Whats [Zie93-35]. Wheeler [Bel99, Bri91, Zie93-38, Ada96, God93, Voe94, Zie92x]. Where [LIV96]. Whitmore [Zie97r]. Whittaker [Goo91]. Whitten [Kim91]. Whys [Zie93-35]. Wichern [Zie99c]. Wiener [DH92]. Wignall [Ott90]. Wilcox [Hor98]. Wild [Ber91]. Wilks [Ram96, Sco90]. Willem [Lor93]. Willems [Sch91b]. Willet [Koh90]. William [Alt96, Alt99, Arm92, Arm93, Buc90, Chu91, Coo96b, Gun94b, Hal98, Hec99, Jon93, Lee94, Nar93, Nel98, Per92, Pfa99, Pig91, Pla90, Rie97, St.94, Tor95, Zie90z, Zie92b, Zie97-41, Zie98-47, Zie99-42, Zie99t]. Williams [Mor97a, Rag93]. Willmot [Zie98-29]. Wilrich [Zie99n]. Wilson [Zie94-36, Alt99, Pan96, Pap95, Sto92, Won94]. Wind [Ste91]. Windows [Zie97j]. Winning [Zie96-51]. Winston [Zie99z, Zie97-29]. Wire [Jen92]. Witmer [Che94]. Wittenmark [Har91d]. Witzling [Lit91]. Wojbor [Lin99, Sha97a, Zie99-46]. Wojtek [Wol96]. Wolfe [Zie99u]. Wolfgang [Eub93]. Wolfinger [Zie97-27, Zie99t]. Wollongong [Zie98u]. Wood [Mil98, Zie93-29]. Woodbury [Kan95]. Woods [Zie96o]. Woodward [Gil92]. Work [Zie93u]. Working [Eas90]. Workman [Goo92b]. Workplace [Zie94-43]. Works [Zie91f, Zie91e, Zie95-43, Tho92b]. Workshop [Mul97]. World [Rea93, Hoc94]. World-Class [Rea93]. Worthington [Ste90c]. Woyczynski [Lin99, Sha97a, Zie99-46]. Wright [Sho90, Zie95-42]. Writing [Zie94w, Zie92j]. Wu [Car93, LW98]. Wulich [Weh91]. Wynn [Hwa92, Zie90q].

X [Zie98-44]. Xavier [Bax96a]. Xishi [Car93]. XploRe [Hin97].

Yacov [Liu99]. Yadolah [Fai94, Zie92q]. Yamaguchi [And90]. Yandell [Gre98]. Yao [Fot96, Rag95]. Yarus [Zie96-44]. Yates [Gra90]. Ye [Han98]. Years [Rag93]. Yefim [Dog93]. Yield [Sam99a]. Young [Cha93]. Yu [Gri99]. Yurij [Boo99].

Z [Lee94]. Zacks [Fre99, Rig93]. Zaven [Law92]. Zayadezny [Weh91]. Zbigniew [Mor93b]. Zedick [Zie92f]. Zeger [Zie95d]. Zehna [Wit94]. Zeisel [Zie99-32]. Zero [GS92, Lam92, LLP^{+}99]. Zero-Defect [GS92]. Zero-Inflated [Lam92, LLP^{+}99]. Zhang [Sch98b]. Zhou [BZM96, Zie98-44]. Zimmerman [Kus93, Sar93, Kus93, Sar93]. Zopounidis [Zie96a]. Zucchini [Ped98]. Zünd [Hil95]. Zupan [Wol91]. Zwanziger [Zie98m].

References

Albers:1999:TRU

[AAK99] Willem Albers, Gerda R. J. Arts, and Wilbert C. M. Kallenberg. Test regions using two or more correlated product characteristics. Technometrics, 41(2):153–165, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

           jstor.org/stable/1270731.

Atkinson:1997:COD

[AB97] Anthony C. Atkinson and Barbara Bogacka. Compound D- and D_{s}-optimum designs for determining the order of a chemical reaction. Technometrics, 39(4):347–356, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

           http://www.jstor.org/stable/1271499.

Abate:1998:BRB

[Aba98] Marcey Abate. Book review: The Pocket Statistician: A Practical Guide to Quality Improvement by S. Coleman; T. Greenfield; R. Jones; C. Morris; I. Puzey. Technometrics, 40(1):75–76, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271399.

Acree:1997:BRB

[Acr97] David A. Acree. Book review: Statistical Process Control Methods by Gary K. Griffith. Technometrics, 39(1):107–108, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270789.

Abraham:1999:SRC

[ACV99] B. Abraham, H. Chipman, and K. Vijayan. Some risks in the construction and analysis of supersaturated designs. Technometrics, 41(2):135–141, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270729.

Atkinson:1996:EDO

[AD96] A. C. Atkinson and A. N. Donev. Experimental designs optimally balanced for trend. Technometrics, 38(4):333–341, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271304.

Adams:1991:BRB

[Ada91] Benjamin M. Adams. Book review: Right Every Time: Using the Deming Approach by Frank Price. Technometrics, 33(3):357– 358, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268789.

Adams:1992:BRB

[Ada92] James A. Adams. Book review: A Quality Technology Primer for Managers by Joseph R. Tunner. Technometrics, 34(3):362– 363, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270055.

Adams:1993:BRB

[Ada93] Benjamin M. Adams. Book review: Statistical Process Control by Leonard A. Doty. Technometrics, 35(1):90–91, February 1993.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269299.

Adams:1994:BRB

[Ada94] Benjamin M. Adams. Book review: Modern Statistical Quality Control and Improvement by Nicholas R. Farnum. Technometrics, 36(2):223–224, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270244.

Adams:1996:BRB

[Ada96] Benjamin M. Adams. Book review: Advanced Topics in Statistical Process Control: The Power of Shewhart’s Charts by Donald J. Wheeler. Technometrics, 38(3):286, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270614.

Arjas:1991:HPQ

[AHT91] Elja Arjas, Christian Kornerup Hansen, and Poul Thyregod. Heterogeneous part quality as a source of reliability improvement in repairable systems. Technometrics, 33(1):1–12, February 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269003.

Ajmani:1999:BRB

[Ajm99] Vivek B. Ajmani. Book review: Modern Engineering Statistics by Lawrence L. Lapin. Technometrics, 41(4):373, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271361.

Albers:1994:ATL

[AKO94] Willem Albers, Wilbert C. M. Kallenberg, and Gertjan D. Otten. Accurate test limits with estimated parameters. Technometrics, 36(1):92–101, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269202.

Albert:1993:BRB

[Alb93] Jim Albert. Book review: The Analysis of Contingency Tables by B. S. Everitt. Technometrics, 35(3):327–328, August 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269530.

Alexander:1991:BRB

[Ale91] Melvin T. Alexander. Book review: Learning Statistics with the SAS System by Rudolf Freund. Technometrics, 33(1):110–112, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269017.

Alexander:1995:BRB

[Ale95] Melvin T. Alexander. Book review: Applied Statistics and Probability for Engineers by Douglas C. Montgomery; George Runger. Technometrics, 37(4):455–457, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269738.

Alexander:1998:BRB

[Ale98] William P. Alexander. Book review: LISREL Approaches to Interaction Effects in Multiple Regression by James Jaccard; Choi K. Wan. Technometrics, 40(3):259–260, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271188.

Alexander:1999:BRB

[Ale99] William Alexander. Book review: Nonparametric Smoothing and Lack-of-Fit Tests by Jeffrey D. Hart. Technometrics, 41(2):175– 176, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270744.

Allen:1996:BRB

[All96] David Allen. Book review: Modeling Experimental and Observational Data by Clifford E. Lunneborg. Technometrics, 38(3): 291, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270620.

Almond:1995:DFL

[Alm95] Russell G. Almond. Discussion: Fuzzy logic: Better science? or better engineering? Technometrics, 37(3):267–270, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269907.

Altland:1996:BRB

[Alt96] Henry W. Altland. Book review: Engineering Methods for Robust Product Design by William Y. Fowlkes; Clyde M. Creveling. Technometrics, 38(3):286–287, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270615.

Altland:1999:BRB

[Alt99] Henry W. Altland. Book review: Regression Analysis: Statistical Modeling of a Response Variable by Rudolf J. Freund; William J. Wilson. Technometrics, 41(4):368, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271353.

Ammann:1992:BRB

[Amm92] Larry P. Ammann. Book review: Lisp-Stat, an Object-Oriented Environment for Statistical Computing and Dynamic Graphics by Luke Tierney. Technometrics, 34(2):238–239, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269258.

Anderson:1990:BRB

[And90] Charles Anderson. Book review: Curves and Surfaces in Computer Aided Geometric Design by F. Yamaguchi. Technometrics, 32(2):229–230, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268878.

Angus:1990:BRBa

[Ang90a] John E. Angus. Book review: Extreme Value Theory in Engineering by Enrique Castillo. Technometrics, 32(1):107–110, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269867.

Angus:1990:BRBb

[Ang90b] John E. Angus. Book review: The Asymptotic Theory of Extreme Order Statistics by Janos Galambos. Technometrics, 32(1): 110–111, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269868.

Angus:1991:BRBa

[Ang91a] John E. Angus. Book review: Empirical Bayes Methods by J. Mantz; T. Lwin. Technometrics, 33(2):243–245, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269060.

Angus:1991:BRBb

[Ang91b] John E. Angus. Book review: Regression Estimators: A Comparative Study by Marvin Gruber. Technometrics, 33(4):480–482, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269428.

Angus:1991:BRBc

[Ang91c] John E. Angus. Book review: Statistical Inference by George Casella; Roger L. Berger. Technometrics, 33(4):493, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269442.

Angus:1992:BRB

[Ang92] John E. Angus. Book review: Forecasting, Structural Time Series and the Kalman Filter by Andrew C. Harvey. Technometrics, 34 (4):496–497, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268960.

Annis:1999:EFC

[Ann99] Charles Annis. [Estimating Fatigue Curves with the Random Fatigue-Limit Model]: Discussion. Technometrics, 41(4):290–291, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271343. See [PM99].

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. Technometrics, 32(1):116–118, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269876.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. Technometrics, 32(2):??, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268896.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. Technometrics, 32(3):368, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269143.

Anonymous:1990:BMd

[Ano90d] Anonymous. Back matter. Technometrics, 32(4):470–471, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270156.

Anonymous:1990:CDO

[Ano90e] Anonymous. Correction: “Detecting Outlying Cells in Two-Way Contingency Tables Via Backwards-Stepping”. Technometrics, 32(1):115, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269875. See [Sim88].

Anonymous:1990:CSD

[Ano90f] Anonymous. Correction: “Small-Disturbance Asymptotic Theory for Linear Calibration Estimators”. Technometrics, 32(4):468, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270154. See [SS89].

Anonymous:1990:CPB

[Ano90g] Anonymous. Corrections to Piepel, Beckman and Carroll. Technometrics, 32(2):241, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268895. See [Pie90b, BNC87, Wei90].

Anonymous:1990:FMa

[Ano90h] Anonymous. Front matter. Technometrics, 32(1):93, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269834.

Anonymous:1990:FMb

[Ano90i] Anonymous. Front matter. Technometrics, 32(2):217, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268852.

Anonymous:1990:FMc

[Ano90j] Anonymous. Front matter. Technometrics, 32(3):339, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269091.

Anonymous:1990:FMd

[Ano90k] Anonymous. Front matter. Technometrics, 32(4):445, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270112.

Anonymous:1990:VI

[Ano90l] Anonymous. Volume information. Technometrics, 32(4):472–477, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270111.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Technometrics, 33(1):123–124, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269036.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Technometrics, 33(2):250, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269072.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Technometrics, 33(3):??, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268811.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Technometrics, 33(4):495–496, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269446.

Anonymous:1991:CDU

[Ano91e] Anonymous. Correction: “Diagnostics for Use with Regression Recursive Residuals”. Technometrics, 33(4):494, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269445. See

[Haw91a].

Anonymous:1991:CMQ

[Ano91f] Anonymous. Correction: “Multivariate Quality Control Based on Regression-Adjusted Variables”. Technometrics, 33(3):376, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268810. See

[Haw91b].

Anonymous:1991:FMa

[Ano91g] Anonymous. Front matter. Technometrics, 33(1):105, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269002.

Anonymous:1991:FMb

[Ano91h] Anonymous. Front matter. Technometrics, 33(2):235, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269037.

Anonymous:1991:FMc

[Ano91i] Anonymous. Front matter. Technometrics, 33(3):355, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268778.

Anonymous:1991:FMd

[Ano91j] Anonymous. Front matter. Technometrics, 33(4):469, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269408.

Anonymous:1991:VI

[Ano91k] Anonymous. Volume information. Technometrics, 33(4):497–501, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269407.

Anonymous:1992:BMa

[Ano92a] Anonymous. Back matter. Technometrics, 34(1):124–126, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269595.

Anonymous:1992:BMb

[Ano92b] Anonymous. Back matter. Technometrics, 34(2):??, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269278.

Anonymous:1992:BMc

[Ano92c] Anonymous. Back matter. Technometrics, 34(3):??, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270077.

Anonymous:1992:BMd

[Ano92d] Anonymous. Back matter. Technometrics, 34(4):503–504, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268970.

Anonymous:1992:FMa

[Ano92e] Anonymous. Front matter. Technometrics, 34(1):97, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269546.

Anonymous:1992:FMb

[Ano92f] Anonymous. Front matter. Technometrics, 34(2):223, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269230.

Anonymous:1992:FMc

[Ano92g] Anonymous. Front matter. Technometrics, 34(3):353, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270027.

Anonymous:1992:FMd

[Ano92h] Anonymous. Front matter. Technometrics, 34(4):??, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268936.

Anonymous:1992:VI

[Ano92i] Anonymous. Volume information. Technometrics, 34(4):505–510, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268935.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Technometrics, 35(1):107–108, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269328. See

correction [Ano93e].

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Technometrics, 35(2):242, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269700.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Technometrics, 35(3):??, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269545.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Technometrics, 35(4):467–468, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270309.

Anonymous:1993:C

[Ano93e] Anonymous. Correction. Technometrics, 35(2):241, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269699. See

[Ano93a].

Anonymous:1993:CCM

[Ano93f] Anonymous. Correction: Comments on Marr and Quesenberry (1991). Technometrics, 35(1):105, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269327. See [FMQ92].

Anonymous:1993:FMa

[Ano93g] Anonymous. Front matter. Technometrics, 35(1):83, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269279.

Anonymous:1993:FMb

[Ano93h] Anonymous. Front matter. Technometrics, 35(2):219, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269655.

Anonymous:1993:FMc

[Ano93i] Anonymous. Front matter. Technometrics, 35(3):321, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269516.

Anonymous:1993:FMd

[Ano93j] Anonymous. Front matter. Technometrics, 35(4):449, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270267.

Anonymous:1993:VI

[Ano93k] Anonymous. Volume information. Technometrics, 35(4):469–473, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270266.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Technometrics, 36(1):??, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269229.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Technometrics, 36(2):235–236, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270265.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Technometrics, 36(3):??, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269406.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Technometrics, 36(4):439–440, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269988.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Technometrics, 36(1):110, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269189.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Technometrics, 36(2):213, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270219.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Technometrics, 36(3):313, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269367.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Technometrics, 36(4):419, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269947.

Anonymous:1994:VI

[Ano94i] Anonymous. Volume information. Technometrics, 36(4):441–445, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269946.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Technometrics, 37(1):??, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269188.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Technometrics, 37(2):246–248, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269654.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Technometrics, 37(3):360, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269945.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter. Technometrics, 37(4):474–475, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269766.

Anonymous:1995:FMa

[Ano95e] Anonymous. Front matter. Technometrics, 37(1):112, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269144.

Anonymous:1995:FMb

[Ano95f] Anonymous. Front matter. Technometrics, 37(2):226, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269614.

Anonymous:1995:FMc

[Ano95g] Anonymous. Front matter. Technometrics, 37(3):340, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269904.

Anonymous:1995:FMd

[Ano95h] Anonymous. Front matter. Technometrics, 37(4):361–472, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269728.

Anonymous:1995:VI

[Ano95i] Anonymous. Volume information. Technometrics, 37(4):476–480, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269727.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. Technometrics, 38(1):94, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268934.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. Technometrics, 38(2):204, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270443.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. Technometrics, 38(3):302, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270641.

Anonymous:1996:BMd

[Ano96d] Anonymous. Back matter. Technometrics, 38(4):415–417, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271339.

Anonymous:1996:FMa

[Ano96e] Anonymous. Front matter. Technometrics, 38(1):74, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268897.

Anonymous:1996:FMb

[Ano96f] Anonymous. Front matter. Technometrics, 38(2):181–196, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270398.

Anonymous:1996:FMc

[Ano96g] Anonymous. Front matter. Technometrics, 38(3):284, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270600.

Anonymous:1996:FMd

[Ano96h] Anonymous. Front matter. Technometrics, 38(4):400, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271296.

Anonymous:1996:VI

[Ano96i] Anonymous. Volume information. Technometrics, 38(4):418–422, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271295.

Anonymous:1997:BMa

[Ano97a] Anonymous. Back matter. Technometrics, 39(1):117–118, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270806.

Anonymous:1997:BMb

[Ano97b] Anonymous. Back matter. Technometrics, 39(2):??, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270938.

Anonymous:1997:BMc

[Ano97c] Anonymous. Back matter. Technometrics, 39(3):??, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271165.

Anonymous:1997:BMd

[Ano97d] Anonymous. Back matter. Technometrics, 39(4):436, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271534.

Anonymous:1997:CBOb

[Ano97e] Anonymous. Correction: Design and Optimization in Organic Synthesis and Complex Stochastic Systems and Engineering. Technometrics, 39(4):435, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271533. See [LH97].

Anonymous:1997:CBOa

[Ano97f] Anonymous. Correction: Design and Optimization in Organic Synthesis [rather than Experimental Design and Chemical Synthesis]. Technometrics, 39(4):435, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271532. See [Emp97].

Anonymous:1997:FMa

[Ano97g] Anonymous. Front matter. Technometrics, 39(1):100–114, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270762.

Anonymous:1997:FMb

[Ano97h] Anonymous. Front matter. Technometrics, 39(2):225, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270899.

Anonymous:1997:FMc

[Ano97i] Anonymous. Front matter. Technometrics, 39(3):331, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271128.

Anonymous:1997:FMd

[Ano97j] Anonymous. Front matter. Technometrics, 39(4):423, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271498.

Anonymous:1997:VI

[Ano97k] Anonymous. Volume information. Technometrics, 39(4):437–440, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271497.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. Technometrics, 40(1):86–87, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271422.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. Technometrics, 40(2):??, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270677.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. Technometrics, 40(3):??, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271209.

Anonymous:1998:BMd

[Ano98d] Anonymous. Back matter. Technometrics, 40(4):363, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270561.

Anonymous:1998:CCD

[Ano98e] Anonymous. Correction: “Cumulative Damage Models for System Failure with Application to Carbon Fibers and Composites”. Technometrics, 40(2):167, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270675. See [DP97].

Anonymous:1998:FMa

[Ano98f] Anonymous. Front matter. Technometrics, 40(1):72, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271386.

Anonymous:1998:FMb

[Ano98g] Anonymous. Front matter. Technometrics, 40(2):153, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270642.

Anonymous:1998:FMc

[Ano98h] Anonymous. Front matter. Technometrics, 40(3):254, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271166.

Anonymous:1998:FMd

[Ano98i] Anonymous. Front matter. Technometrics, 40(4):348–361, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270526.

Anonymous:1998:VI

[Ano98j] Anonymous. Volume information. Technometrics, 40(4):364–368, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270525.

Anonymous:1999:BMa

[Ano99a] Anonymous. Back matter. Technometrics, 41(1):86–87, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271023.

Anonymous:1999:BMb

[Ano99b] Anonymous. Back matter. Technometrics, 41(2):??, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270761.

Anonymous:1999:BMc

[Ano99c] Anonymous. Back matter. Technometrics, 41(3):??, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270599.

Anonymous:1999:BMd

[Ano99d] Anonymous. Back matter. Technometrics, 41(4):386–387, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271385.

Anonymous:1999:CIE

[Ano99e] Anonymous. Correction: “Interaction Effects in Factorial Analysis of Variance”. Technometrics, 41(3):275, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270598. See [Emp99].

Anonymous:1999:EFC

[Ano99f] Anonymous. [Estimating Fatigue Curves with the Random Fatigue-Limit Model]: Response. Technometrics, 41(4):297–302, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271346. See [PM99].

Anonymous:1999:FMa

[Ano99g] Anonymous. Front matter. Technometrics, 41(1):71, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270989.

Anonymous:1999:FMb

[Ano99h] Anonymous. Front matter. Technometrics, 41(2):166, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270720.

Anonymous:1999:FMc

[Ano99i] Anonymous. Front matter. Technometrics, 41(3):261, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270562.

Anonymous:1999:FMd

[Ano99j] Anonymous. Front matter. Technometrics, 41(4):367, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271341.

Anonymous:1999:VI

[Ano99k] Anonymous. Volume information. Technometrics, 41(4):388–392, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271340.

Anonymous:2000:CCS

[Ano00] Anonymous. Correction: “A Cepstral F Statistic for Detecting Delay-Fired Seismic Signals”. Technometrics, 42(2):226, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271495. See

[SBD98].

Aragon:1992:BRB

[Ara92] Jorge Aragón. Book review: The Architecture of Chance by Richard Lowry. Technometrics, 34(2):242–243, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269265.

Armstrong:1990:CTR

[Arm90] Robert P. Armstrong. Commentary: Is there really a communications gap between statisticians and engineers? Technometrics, 32(3):249–250, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269093. See [HK90].

Armitage:1992:BRB

[Arm92] Stan J. Armitage. Book review: Understanding Statistics by Lyman Ott; William Mendenhall. Technometrics, 34(3):368–369, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270063.

Armitage:1993:BRB

[Arm93] Stanley J. Armitage. Book review: Statistical Methods for the Process Industries by William H. McNeese; Robert A. Klein. Technometrics, 35(2):229–230, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269679.

Armstrong:1997:BRB

[Arm97] Ben Armstrong. Book review: Measurement Error in Nonlinear Models by R. J. Carroll; D. Ruppert; L. A. Stefanski. Technometrics, 39(2):231–232, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270919.

Arnold:1990:BRB

[Arn90] Harvey J. Arnold. Book review: Introduction to the Practice of Statistics by David S. Moore; George P. McCabe. Technometrics, 32(3):347–348, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269120.

Atkinson:1996:DPP

[Atk96] A. C. Atkinson. Discussion: Posterior probabilities and experimental designs for discriminating between regression models. Technometrics, 38(4):314–317, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271298.

Adams:1994:RLD

[AWS94] Benjamin M. Adams, William H. Woodall, and Claude R. Superville. [Run-Length Distributions of Special-Cause Control Charts for Correlated Processes]: Discussion. Technometrics, 36 (1):19–22, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269193. See [WMP94a].

Ayeni:1998:BRB

[Aye98] Babatunde J. Ayeni. Book review: Straight Talk on Designing Experiments by Robert G. Launsby; Daniel L. Weese. Technometrics, 40(1):75, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271398.

Bisgaard:1996:SEE

[BA96] Søren Bisgaard and Bruce Ankenman. Standard errors for the eigenvalues in second-order response surface models. Technometrics, 38(3):238–246, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270607.

Babu:1998:BRB

[Bab98] Gutti J. Babu. Book review: Introduction to Probability Models by Sheldon M. Ross. Technometrics, 40(1):78, February 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271404.

Bajorski:1998:BRB

[Baj98] Peter Bajorski. Book review: Visual Explanations: Images and Quantities, Evidence and Narrative by Edward R. Tufte. Technometrics, 40(1):81, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271409.

Baker:1996:SNT

[Bak96] R. D. Baker. Some new tests of the power law process. Technometrics, 38(3):256–265, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270609.

Bakir:1998:BRB

[Bak98] Saad T. Bakir. Book review: Tolerance Design: A Handbook for Developing Optimal Specifications by C. M. Creveling. Technometrics, 40(1):74–75, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271397.

Balestracci:1993:BRB

[Bal93] Davis Balestracci. Book review: Intermediate Statistics, A Modern Approach by James Stevens. Technometrics, 35(1):98, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269309.

Barker:1991:BRB

[Bar91] Anne M. Barker. Book review: Handbook of Statistical Methods for Engineers and Scientists by Harrison M. Wadsworth. Technometrics, 33(2):238–239, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269051.

Barker:1992:BRBb

[Bar92a] Lawrence Barker. Book review: Statistical Analyses of Reliability Data by M. J. Crowder; A. C. Kimber; R. L. Smith; T. J. Sweeting. Technometrics, 34(4):485–486, November 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268947.

Barker:1992:BRBa

[Bar92b] Lawrence Barker. Book review: System Reliability: Concepts and Applications by Klaas B. Klaassen; Jack C. L. van Peppen. Technometrics, 34(2):234–235, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269254.

Barrett:1995:BRB

[Bar95] Bruce Barrett. Book review: Regression Analysis: Concepts and Applications by Franklin A. Graybill; Hariharan K. Iyer. Technometrics, 37(2):229, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269625.

Basu:1990:BRB

[Bas90] Asit P. Basu. Book review: The Inverse Gaussian Distribution by Raj S. Chhikara; J. Leroy Folks. Technometrics, 32(2):235, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268886.

Basu:1997:BRB

[Bas97] Asit P. Basu. Book review: Statistical Analysis of Nonnormal Data by J. V. Deshpande; A. P. Gore; A. Shanubhogue. Technometrics, 39(3):341, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271150.

Basu:1999:BRB

[Bas99] Asit P. Basu. Book review: Aspects of Statistical Inference by A. H. Welsh. Technometrics, 41(3):270–271, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270585.

Bates:1999:BRB

[Bat99] Douglas M. Bates. Book review: Programming with Data: A Guide to the S Language by John M. Chambers. Technometrics, 41(3):266–267, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270578.

Baxley:1990:EWM

[Bax90] Robert V. Baxley, Jr. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]: Discussion. Technometrics, 32(1):13–16, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269836. See [LS90b].

Baxter:1992:BRB

[Bax92] Laurence A. Baxter. Book review: Probability, Statistics, and Queueing Theory with Computer Science Applications by Arnold O. Allen. Technometrics, 34(2):240–241, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269262.

Baxter:1995:BRB

[Bax95] Laurence A. Baxter. Book review: Markov Decision Processes: Discrete Stochastic Dynamic Programming by Martin L. Puterman. Technometrics, 37(3):353, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269932.

Baxter:1996:BRBb

[Bax96a] Laurence A. Baxter. Book review: Random Fields on a Network: Modeling, Statistics, and Applications by Xavier Guyon. Technometrics, 38(4):409–410, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271325.

Baxter:1996:BRBa

[Bax96b] Laurence A. Baxter. Book review: Reliability Analysis for Engineers: An Introduction by Roger D. Leitch. Technometrics, 38 (2):185, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270417.

Bayne:1992:BRB

[Bay92] Charles K. Bayne. Book review: Latent Variable Path Modeling with Partial Least Squares by Jan-Bernd Lohmöller. Technometrics, 34(1):110–111, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269575.

Bayne:1995:BRB

[Bay95] Charles K. Bayne. Book review: Practical Guide to Chemometrics by Stephen John Haswell. Technometrics, 37(2):230–231, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269627.

Bayne:1999:BRB

[Bay99] Charles K. Bayne. Book review: Chemometric Techniques for Quantitative Analysis by Richard Kramer. Technometrics, 41(2): 173–174, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270741.

Berk:1995:SCM

[BB95] Kenneth N. Berk and David E. Booth. Seeing a curve in multiple regression. Technometrics, 37(4):385–398, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269731.

Birch:1990:CIB

[BBT90] Nancy J. Birch, Richard K. Burdick, and Naitee Ting. Confidence intervals and bounds for a ratio of summed expected mean squares. Technometrics, 32(4):437–444, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270123.

Blazek:1995:STC

[BC95] Linda W. Blazek and David E. Coleman. [Shewhart-Type Charts in Nonstandard Situations]: Discussion. Technometrics, 37(1):24– 26, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269147. See [RD95a].

Becker:1998:EDDa

[BCL98a] Richard A. Becker, Linda A. Clark, and Diane Lambert. Events defined by duration and severity, with an application to network reliability. Technometrics, 40(3):177–189, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271167. See discussion [Rob98,

Sch98a] and response [BCL98b].

Becker:1998:EDDb

[BCL98b] Richard A. Becker, Linda A. Clark, and Diane Lambert. [Events Defined by Duration and Severity, with an Application to Network Reliability]: Response. Technometrics, 40(3):192–194, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271170. See

[BCL98a].

Boulanger:1994:EDC

[BE94] Michèle Boulanger and Luis A. Escobar. Experimental design for a class of accelerated degradation tests. Technometrics, 36 (3):260–272, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269370.

Beggs:1990:BRB

[Beg90] William J. Beggs. Book review: Statistical Methods for Nuclear Material Management by W. Michael Bowen; Carl A. Bennett. Technometrics, 32(4):462–463, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270145.

Belunis:1995:BRB

[Bel95] Nancy Belunis. Book review: A Practical Guide to Statistical Quality Improvement: Opening up the Statistical Toolbox by Michael R. Beauregard; Raymond J. Mikulak; Barbara A. Olson. Technometrics, 37(2):233–234, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269632.

Belunis:1999:BRB

[Bel99] Nancy Belunis. Book review: Building Continual Improvement by Donald J. Wheeler; Sheila R. Poling. Technometrics, 41(2): 169–170, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270736.

Bement:1996:BRB

[Bem96] Thomas R. Bement. Book review: Fundamentals of Quality Control and Improvement by Amitava Mitra. Technometrics, 38(1):80– 81, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268912.

Benbow:1993:BRB

[Ben93] Don Benbow. Book review: Quality Engineering Statistics by Robert A. Dovich. Technometrics, 35(3):332–333, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269539.

Berk:1991:BRB

[Ber91] Kenneth Berk. Book review: Nonlinear Regression by G. A. F. Seber; C. J. Wild. Technometrics, 33(3):362–363, August 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268793.

Berk:1992:BRB

[Ber92a] Robert H. Berk. Book review: Observing Interaction: An Introduction to Sequential Analysis by Roger Bakeman; John M. Gottman. Technometrics, 34(1):112–113, February 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269577.

Berliner:1992:BRB

[Ber92b] L. Mark Berliner. Book review: Bayesian Statistics: An Introduction by Peter M. Lee. Technometrics, 34(1):115, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269580.

Berens:1996:ASM

[Ber96a] Alan P. Berens. [Applications of Statistical Methods to Nondestructive Evaluation]: Discussion. Technometrics, 38(2):113–116, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270400. See

[OM96a].

Berk:1996:BRB

[Ber96b] Robert H. Berk. Book review: Continuous Univariate Distributions, Volume 2 by Norman L. Johnson; Samuel Kotz; N. Balakrishnan. Technometrics, 38(2):189, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270425.

Berk:1997:BRB

[Ber97] Robert Berk. Book review: Handbook of the Normal Distribution by Jagdish K. Patel; Campbell B. Read. Technometrics, 39(3): 340, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271149.

Berk:1998:RDP

[Ber98] Kenneth N. Berk. Regression diagnostic plots in 3-D. Technometrics, 40(1):39–47, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271391.

Bancroft:1993:MPV

[BH93] Bruce A. Bancroft and Gerald R. Hobbs. Mine-product variability estimators with emphasis on the case of coal. Technometrics, 35 (4):390–402, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270272.

Bergman:1997:DEU

[BH97] Bo Bergman and Anders Hynén. Dispersion effects from unreplicated designs in the 2^{k−p} series. Technometrics, 39(2):191–198, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270907.

Brush:1990:EOQ

[BHS90] Gary G. Brush, Bruce Hoadley, and Bernard Saperstein. Estimating outgoing quality using the quality measurement plan. Technometrics, 32(1):31–41, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269842.

Bieger:1992:BRB

[Bie92] Charles D. Bieger. Book review: Statistics with STATA by Lawrence C. Hamilton. Technometrics, 34(1):117–118, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269583.

Bilonick:1991:BRB

[Bil91] Richard A. Bilonick. Book review: An Introduction to Applied Geostatistics by Edward H. Isaaks; R. Mohan Srivastava. Technometrics, 33(4):483–485, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269430.

Bilonick:1992:BRB

[Bil92] Richard A. Bilonick. Book review: Pierre Gy’s Sampling Theory and Sampling Practice by Francis F. Pitard. Technometrics, 34(3):363–364, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270056.

Bilonick:1994:BRB

[Bil94] Richard A. Bilonick. Book review: Sampling of Heterogeneous and Dynamic Material Systems: Theories of Heterogeneity, Sampling and Homogenizing by Pierre M. Gy. Technometrics, 36(2): 218–219, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270238.

Bisgaard:1994:NDR

[Bis94] Søren Bisgaard. A note on the definition of resolution for blocked 2^{k−p} designs. Technometrics, 36(3):308–311, August 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269375.

Bisgaard:1997:DET

[Bis97] Søren Bisgaard. Designing experiments for tolerancing assembled products. Technometrics, 39(2):142–152, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270902.

Box:1992:SPMa

[BK92a] George Box and Tim Kramer. Statistical process monitoring and feedback adjustment — a discussion. Technometrics, 34(3):251– 267, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270028. See discussion [Mac92, Tuc92, VV92, WG92] and response [BK92b].

Box:1992:SPMb

[BK92b] George Box and Tim Kramer. [Statistical Process Monitoring and Feedback Adjustment]: Response. Technometrics, 34(3):282– 285, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270034. See [BK92a].

Box:1994:SSI

[BL94] George Box and Alberto Luceño. Selection of sampling interval and action limit for discrete feedback adjustment. Technometrics, 36(4):369–378, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269952.

Borkowski:1997:DMR

[BL97] John J. Borkowski and James M. Lucas. Designs of mixed resolution for process robustness studies. Technometrics, 39(1):63– 70, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270773.

Blazek:1992:BRB

[Bla92] Linda W. Blazek. Book review: Interactive Computer Graphics: Functional, Procedural and Device-Level Methods by Peter Burger; Duncan Gillies. Technometrics, 34(1):105–106, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269567.

Blackwood:1994:BRB

[Bla94] Larry G. Blackwood. Book review: Factor Analysis in Chemistry by Edmund R. Malinowski. Technometrics, 36(1):115, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269209.

Blischke:1991:BRB

[Bli91] Wallace R. Blischke. Book review: Practical Statistics by S. S. Cohen. Technometrics, 33(4):489–490, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269437.

Beckman:1987:DMM

[BNC87] Richard J. Beckman, Christopher J. Nachtsheim, and R. Dennis Cook. Diagnostics for mixed-model analysis of variance. Technometrics, 29(4):413–426, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269452. See corrections[Ano90g].

Bodt:1997:BRB

[Bod97] Barry A. Bodt. Book review: Engineering Statistics: The Industrial Experience by B. Ostle; K. Turner; C. Hicks; G. McElrath. Technometrics, 39(1):107, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270788.

Boggs:1992:BRB

[Bog92] Brady Boggs. Book review: No Downtime: Six Steps to Industrial Problem Solving by L. E. Cook. Technometrics, 34(2):234, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269253.

Bonissone:1995:DFL

[Bon95] Piero P. Bonissone. Discussion: Fuzzy logic control technology: A personal perspective. Technometrics, 37(3):262–266, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269906.

Booker:1990:BRB

[Boo90] Jane M. Booker. Book review: Point Process Models with Applications to Safety and Reliability by W. A. Thompson. Technometrics, 32(3):344–345, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269117.

Booth:1992:BRB

[Boo92] David E. Booth. Book review: Time Series by Maurice Kendall; J. Keith Ord. Technometrics, 34(1):118–119, February 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269585.

Booth:1993:BRB

[Boo93] David E. Booth. Book review: Statistics, Principles and Methods by Richard A. Johnson; Gouri K. Bhattacharyya. Technometrics, 35(2):237–238, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269691.

Booth:1994:BRBb

[Boo94a] David E. Booth. Book review: Statistical Methods for SPC and TQM by Derek Bissell. Technometrics, 36(4):420–421, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269959.

Booth:1994:BRBa

[Boo94b] David E. Booth. Book review: Time Series, Forecasting, Simulation, Applications by Gareth Janacek; Louise Swift. Technometrics, 36(2):217–218, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270237.

Booker:1995:BRB

[Boo95] Jane Booker. Book review: Poisson Processes by J. F. C. Kingman. Technometrics, 37(1):124, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269173.

Booth:1997:BRB

[Boo97] David E. Booth. Book review: Time Series Analysis by James D. Hamilton. Technometrics, 39(1):102–103, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270781.

Booth:1998:BRB

[Boo98] David E. Booth. Book review: Multivariate Statistical Inference and Applications by Alvin C. Rencher. Technometrics, 40 (4):353–354, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270542.

Booth:1999:BRB

[Boo99] David E. Booth. Book review: Robustness in Statistical Pattern Recognition by Yurij Kharin. Technometrics, 41(4):370–371, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271357.

Borkowski:1994:BRB

[Bor94] John J. Borkowski. Book review: Optimal Design of Experiments by Friedrich Pukelsheim. Technometrics, 36(2):214–216, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270234.

Borkowski:1995:SPV

[Bor95a] John J. Borkowski. Spherical prediction-variance properties of central composite and Box-Behnken designs. Technometrics, 37 (4):399–410, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269732.

Borowiak:1995:BRB

[Bor95b] Dale Borowiak. Book review: The Theory of Linear Models by Bent Jorgensen. Technometrics, 37(2):228–229, May 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269624.

Bordley:1997:BRB

[Bor97a] Robert F. Bordley. Book review: Evaluating Research Articles: From Start to Finish by Ellen R. Girden. Technometrics, 39(4): 431–432, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271521.

Borkowski:1997:BRB

[Bor97b] John J. Borkowski. Book review: Statistical Tools for Nonlinear Regression by Sylvie Huet; Annie Bouvier; Marie-Anne Gruet; Emmanuel Jolivet. Technometrics, 39(3):333–334, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271139.

Borowiak:1997:BRB

[Bor97c] Dale Borowiak. Book review: Linear Models: A Mean Model Approach by Barry Kurt Moser. Technometrics, 39(4):425–426, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271508.

Borowiak:1998:BRB

[Bor98] Dale Borowiak. Book review: Quasi-Likelihood and Its Application by Christopher C. Heyde. Technometrics, 40(4):356, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270545.

Bossert:1992:BRB

[Bos92] James L. Bossert. Book review: Introduction to Quality Control by Kaoru Ishikawa. Technometrics, 34(4):494–495, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268958.

Bowerman:1991:BRB

[Bow91] Bruce L. Bowerman. Book review: Statistical Design and Analysis of Experiments with Applications to Engineering and Science by Robert L. Mason; Richard F. Gunst; James L. Hess. Technometrics, 33(1):106, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269012.

Box:1990:CCB

[Box90] George Box. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 251–252, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269094. See [HK90].

Boyer:1990:BRB

[Boy90] John E. Boyer, Jr. Book review: Statistical Analysis for Engineers: A Computer-Based Approach by J. Wesley Barnes. Technometrics, 32(2):222–223, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268870.

Boyles:1996:MPA

[Boy96] Russell A. Boyles. Multivariate process analysis with lattice data. Technometrics, 38(1):37–49, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268901.

Bandeen-Roche:1995:BRB

[BR95] Karen J. Bandeen-Roche. Book review: Quality Control in Analytic Chemistry by G. Kateman; L. Buydens. Technometrics, 37(1):115–116, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269161.

Bresky:1990:BRB

[Bre90] Norman P. Bresky. Book review: Tools and Methods for the Improvement of Quality by H. Gitlow; S. Gitlow; A. Oppenheim; R. Oppenheim. Technometrics, 32(3):340–341, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269112.

Breiman:1991:MEMa

[Bre91a] Leo Breiman. The ∏ method for estimating multivariate functions from noisy data. Technometrics, 33(2):125–143, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269038. See discussion

[Bre91a, Gu91, HT91] and response [Bre91b].

Breiman:1991:MEMb

[Bre91b] Leo Breiman. [The ∏ Method for Estimating Multivariate Functions from Noisy Data]: Response. Technometrics, 33(2):156–160, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269042. See

[Bre91a].

Breiman:1995:BSR

[Bre95a] Leo Breiman. Better subset regression using the nonnegative garrote. Technometrics, 37(4):373–384, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269730.

Bremer:1995:BRB

[Bre95b] Ronald Bremer. Book review: Outliers in Statistical Data by Vic Barnett; Toby Lewis. Technometrics, 37(1):117–118, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269164.

Bresky:1996:BRB

[Bre96] Norman Bresky. Book review: Quality Improvement Tools and Techniques by Peter Mears. Technometrics, 38(2):185, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270418.

Breneman:1997:BRB

[Bre97] James E. Breneman. Book review: Simulation Made Easy, a Manager’s Guide by Charles Harrell; Kerim Tumay. Technometrics, 39(4):431, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271518.

Brick:1991:BRB

[Bri91] J. Michael Brick. Book review: Evaluating the Measurement Process by D. Wheeler; R. Lyday. Technometrics, 33(4):474, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269422.

Brill:1996:BRB

[Bri96] Bob Brill. Book review: Statistics: Learning in the Presence of Variation by Robert L. Wardrop. Technometrics, 38(4):405–406, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271319.

Brooks:1990:BRBa

[Bro90a] Daniel G. Brooks. Book review: Linear Programming with Statistical Applications by Vincent A. Sposito. Technometrics, 32 (3):355–356, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269131.

Brooks:1990:BRBb

[Bro90b] Daniel G. Brooks. Book review: Proceedings of the 20th Symposium on the Interface: Computing Science and Statistics: Computationally Intensive Methods in Statistics by Edward J. Wegman; Donald Gantz; John Miller. Technometrics, 32(4):465, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270150.

Brooks:1992:BRB

[Bro92] Daniel G. Brooks. Book review: Scientific Data Analysis: An Introduction to Overdetermined Systems by Richard L. Branham. Technometrics, 34(2):239, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269259.

Brugger:1991:BRB

[Bru91] Richard M. Brugger. Book review: Defect Prevention: Use of Simple Statistical Tools by Victor E. Kane. Technometrics, 33(1): 117, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269027.

Brugger:1992:BRBa

[Bru92a] Richard M. Brugger. Book review: Handbook of Statistical Methods in Manufacturing by Richard Barrett Clements. Technometrics, 34(3):360–361, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270052.

Brugger:1992:BRBb

[Bru92b] Richard M. Brugger. Book review: How to Construct Fractional Factorial Experiments by Richard F. Gunst; Robert L. Mason. Technometrics, 34(4):492–493, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268956.

Brugger:1993:BRBb

[Bru93a] Richard M. Brugger. Book review: Probability and Its Applications for Engineers by David H. Evans. Technometrics, 35(4): 459–460, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270291.

Brugger:1993:BRBa

[Bru93b] Richard M. Brugger. Book review: Statistical Methods in Engineering and Quality Assurance by Peter W. M. John. Technometrics, 35(1):91–92, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269300.

Brugger:1994:BRB

[Bru94] Richard M. Brugger. Book review: Univariate Discrete Distributions by Norman L. Johnson; Samuel Kotz; Adrienne W. Kemp. Technometrics, 36(1):118, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269213.

Brugger:1995:BRB

[Bru95] Richard M. Brugger. Book review: Statistics for Engineering Problem Solving by Stephen B. Vardeman. Technometrics, 37(3): 348, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269926.

Brugger:1996:BRB

[Bru96] Richard M. Brugger. Book review: How to Perform Skip-Lot and Chain Sampling by Kenneth S. Stephens. Technometrics, 38(1): 82, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268915.

Brugger:1997:BRB

[Bru97] Richard M. Brugger. Book review: Fitting Frequency Distributions: Philosophy and Practice, Parts I and II by David W. Miller. Technometrics, 39(2):234–235, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270923.

Bryson:1992:BRB

[Bry92] Maurice C. Bryson. Book review: Statistical Analysis of Reliability and Life-Testing Models by Lee J. Bain; Max Engelhardt. Technometrics, 34(4):486–487, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268948.

Bryson:1999:BRB

[Bry99] Maurice Bryson. Book review: Probability, Statistics, and Reliability for Engineers by Bilal M. Ayyub; Richard H. McCuen. Technometrics, 41(2):172–173, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270740.

Bisgaard:1997:DAP

[BS97] Søren Bisgaard and David M. Steinberg. The design and analysis of 2^{k−p} × s prototype experiments. Technometrics, 39(1):52– 62, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270772.

Bingham:1999:MAT

[BS99] Derek Bingham and Randy R. Sitter. Minimum-aberration twolevel fractional factorial split-plot designs. Technometrics, 41 (1):62–70, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270995.

Beckman:1989:TST

[BT89] Richard J. Beckman and Gary L. Tietjen. Two-sided tolerance limits for balanced random-effects ANOVA models. Technometrics, 31(2):185–197, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268816. See correction [BT98].

Beckman:1998:CTS

[BT98] Richard J. Beckman and Gary L. Tietjen. Correction: “Twosided tolerance limits for balanced random-effects ANOVA models” [Technometrics 31 (1989), no. 2, 185–197; MR1007294 (90f:62230)]. Technometrics, 40(3):269, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271208. See [BT89].

Buckingham:1990:BRB

[Buc90] Janet Buckingham. Book review: Dynamic Graphics for Statistics by William S. Cleveland; Marylyn E. McGill. Technometrics, 32 (4):446–447, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270124.

Bulmahn:1991:BRB

[Bul91] Barbara J. Bulmahn. Book review: How to Think about Statistics by John L. Phillips. Technometrics, 33(1):112, February 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269018.

Bulmahn:1996:BRB

[Bul96] Barbara J. Bulmahn. Book review: Introductory Statistics: A Modelling Approach by J. K. Lindsey. Technometrics, 38(2):190, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270427.

Burdick:1991:BRB

[Bur91] Donald Burdick. Book review: Multivariate Chemometrics in QSAR: A Dialogue by Peter P. Mager. Technometrics, 33(1): 119–120, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269031.

Bzik:1997:BRB

[Bzi97] Thomas J. Bzik. Book review: Statistical Methods for Engineers and Scientists by R. M. Bethea; B. S. Duran; T. L. Boullion. Technometrics, 39(2):229, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270915.

Burdick:1996:CZM

[BZM96] Richard K. Burdick, Leping Zhou, and Thomas Mathew. Comment on Zhou and Mathew (1994). Technometrics, 38(1):92–93, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1268933. See [ZM94].

Caby:1994:BRB

[Cab94] Errol Caby. Book review: Six Sigma Producibility Analysis and Process Characterization by Mikel J. Harry; J. Ronald Lawson. Technometrics, 36(1):122, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269219.

Caby:1998:BRB

[Cab98] Errol Caby. Book review: Time Series and Dynamic Models by Christian Gourieroux; Alain Monfort. Technometrics, 40(2):158, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270656.

Calvin:1994:BRB

[Cal94] James A. Calvin. Book review: Hierarchical Linear Models: Applications and Data Analysis Methods by Anthony S. Bryk; Stephen W. Raudenbush. Technometrics, 36(1):116–117, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269211.

Calvin:1998:BRB

[Cal98] James A. Calvin. Book review: Regression Models for Categorical and Limited Dependent Variables by J. Scott Long. Technometrics, 40(1):80–81, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271407.

Campbell:1990:BRB

[Cam90] Katherine Campbell. Book review: Exploration-Geochemical Data Analysis with the IBM PC by George S. Koch. Technometrics, 32(3):358, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269135.

Campbell:1996:BRB

[Cam96] Katherine Campbell. Book review: Introduction to Disjunctive Kriging and Non-Linear Geostatistics by Jacques Rivoirard. Technometrics, 38(2):187–188, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270423.

Campbell:1998:BRB

[Cam98] Katherine Campbell. Book review: Setting Environmental Standards: The Statistical Approach to Handling Uncertainty and Variation by V. Barnett; A. O’Hagan. Technometrics, 40(3): 262, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271192.

Carey:1990:BRB

[Car90a] Michele Boulanger Carey. Book review: A Primer of Reliability Theory by D. L. Grosh. Technometrics, 32(3):345–346, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269118.

Carlson:1990:BRB

[Car90b] Richard R. Carlson. Book review: What Every Engineer Should Know about Quality Control by Thomas Pyzdek. Technometrics, 32(3):342–343, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269114.

Carlson:1992:BRB

[Car92] Richard R. Carlson. Book review: What Every Manager Should Know about Quality by Thomas Pyzdek. Technometrics, 34(4): 495–496, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268959.

Carlson:1993:BRB

[Car93] Richard R. Carlson. Book review: Inspection Errors for Attributes in Quality Control by Norman L. Johnson; Samuel Kotz; Xishi Wu. Technometrics, 35(2):232, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269682.

Carniello:1995:BRB

[Car95] Michael L. Carniello. Book review: Analysis of Two-Way Layouts by John Mandel. Technometrics, 37(4):461, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269745.

Carpenter:1996:BRB

[Car96] Mark Carpenter. Book review: Reliability Engineering Bible by Bryan Dodson; Dennis Nolan. Technometrics, 38(4):406–407, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271320.

Carpenter:1997:BRB

[Car97] Mark Carpenter. Book review: Survival Analysis: A Self-Learning Text by David G. Kleinbaum. Technometrics, 39(2):228–229, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270914.

Christensen:1993:TPA

[CB93] Ronald Christensen and Larry G. Blackwood. Tests for precision and accuracy of multiple measuring devices. Technometrics, 35 (4):411–420, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270274.

Cutler:1994:AA

[CB94] Adele Cutler and Leo Breiman. Archetypal analysis. Technometrics, 36(4):338–347, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269949.

Crone:1995:SAM

[CC95] L. J. Crone and D. S. Crosby. Statistical applications of a metric on subspaces to satellite meteorology. Technometrics, 37(3):324– 328, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269916.

Chen:1999:RTE

[CC99] Shun-Yi Chen and Hubert J. Chen. A range test for the equivalency of means under unequal variances. Technometrics, 41(3): 250–260, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270569.

Carter:1990:LSC

[CCC90] Walter H. Carter, Jr., Vernon M. Chinchilli, and Eleanor D. Campbell. A large-sample confidence region useful in characterizing the stationary point of a quadratic response surface. Technometrics, 32(4):425–435, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270122.

Crawford:1992:MLC

[CDKS92] Sybil L. Crawford, Morris H. DeGroot, Joseph B. Kadane, and Mitchell J. Small. Modeling lake-chemistry distributions: Approximate Bayesian methods for estimating a finite-mixture model. Technometrics, 34(4):441–453, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268943.

Capilla:1999:ISE

[CFRH99] Carmen Capilla, Alberto Ferrer, Rafael Romero, and Angel Hualda. Integration of statistical and engineering process control in a continuous polymerization process. Technometrics, 41 (1):14–28, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270991.

Carrillo-Gamboa:1992:MEM

[CGG92] Olivia Carrillo-Gamboa and Richard F. Gunst. Measurementerror-model collinearities. Technometrics, 34(4):454–464, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268944.

Chipman:1996:BAO

[CH96a] Hugh Chipman and Michael Hamada. Bayesian analysis of ordered categorical data from industrial experiments. Technometrics, 38(1):1–10, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268898.

Chipman:1996:DFB

[CH96b] Hugh A. Chipman and Michael S. Hamada. Discussion: Factorbased or effect-based modeling? implications for design. Technometrics, 38(4):317–320, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271299.

Chang:1992:BRB

[Cha92] Phil T. Chang. Book review: Improving Quality Through Planned Experimentation by Ronald D. Moen; Thomas W. Nolan; Lloyd P. Provost. Technometrics, 34(4):491–492, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268954.

Chaubey:1993:BRB

[Cha93] Yogendra P. Chaubey. Book review: Resampling-Based Multiple Testing: Examples and Methods for p-Value Adjustment by Peter H. Westfall; S. Stanely Young. Technometrics, 35(4):450– 451, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270279.

Champ:1995:BRB

[Cha95] Charles W. Champ. Book review: Process Capability Indices by Samuel Kotz; Norman L. Johnson. Technometrics, 37(2):233, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269631.

Chang:1998:BRB

[Cha98] Joseph T. Chang. Book review: Probability and Statistical Inference by Robert Bartoszynski; Magdalena Niewiadomska-Bugaj. Technometrics, 40(4):354–356, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270544.

Caspary:1990:DAS

[CHB90] W. F. Caspary, W. Haen, and H. Borutta. Deformation analysis by statistical methods. Technometrics, 32(1):49–57, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269844.

Chernick:1991:BRB

[Che91] Michael R. Chernick. Book review: Statistical Methods in Engineering and Manufacturing by John E. Brown. Technometrics, 33(3):356, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268787.

Chen:1992:BRB

[Che92] Chung Chen. Book review: Quantitative Forecasting Methods by Nicholas R. Farnum; LaVerne W. Stanton. Technometrics, 34(3): 358, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270047.

Chen:1994:BRB

[Che94] Hanfeng Chen. Book review: Data Analysis: An Introduction by Jeffrey A. Witmer. Technometrics, 36(4):427, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269967.

Cheeseman:1995:DFT

[Che95] Peter Cheeseman. Discussion: Fuzzy thinking. Technometrics, 37(3):282–283, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269910.

Chen:1998:BRB

[Che98] Hanfeng Chen. Book review: Applying Generalized Linear Models by James K. Lindsey. Technometrics, 40(2):156–157, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270654.

Chiu:1991:BRB

[Chi91] Shean-Tsong Chiu. Book review: Regression Analysis: Theory, Methods, and Applications by Ashish Sen; Muni Srivastava. Technometrics, 33(4):479–480, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269427.

Christensen:1997:BRB

[Chr97] Ronald Christensen. Book review: Multiple Comparisons: Theory and Methods by Jason C. Hsu. Technometrics, 39(2):232–233, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270920.

Chuang:1991:BRB

[Chu91] Alice Chuang. Book review: Time Series Analysis: Univariate and Multivariate Methods by William W. S. Wei. Technometrics, 33(1):108–109, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269015.

Chipman:1997:BVS

[CHW97] H. Chipman, M. Hamada, and C. F. J. Wu. A Bayesian variableselection approach for analyzing designed experiments with complex aliasing. Technometrics, 39(4):372–381, November 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271501.

Cheng:1993:COF

[CL93] Ching-Shui Cheng and Chang Chung Li. Constructing orthogonal fractional factorial designs when some factor-level combinations are debarred. Technometrics, 35(3):277–283, August 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269520.

Clyde:1994:BRB

[Cly94] Merlise Clyde. Book review: The Statistical Exorcist: Dispelling Statistics Anxiety by M. Hollander; F. Proschan. Technometrics, 36(3):330, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269394.

Coleman:1993:SAPb

[CM93a] David E. Coleman and Douglas C. Montgomery. [A Systematic Approach to Planning for a Designed Industrial Experiment]: Response. Technometrics, 35(1):25–27, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269285. See [CM93b].

Coleman:1993:SAPa

[CM93b] David E. Coleman and Douglas C. Montgomery. A systematic approach to planning for a designed industrial experiment. Technometrics, 35(1):1–12, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1269280. See discussion [Gun93, Hah93,

HO93, LST93] and response [CM93a].

Chou:1999:SNP

[CM99] Youn-Min Chou and Robert L. Mason. [Simultaneous Nonparametric Prediction Limits]: Discussion. Technometrics, 41(2):102– 103, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270722. See [DM99a].

Coakley:1994:BRB

[Coa94] Kevin J. Coakley. Book review: Statistical Challenges in Modern Astronomy by Eric D. Feigelson; G. Jogesh Babu. Technometrics, 36(3):328, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269390.

Coburn:1991:BRB

[Cob91] Timothy C. Coburn. Book review: Spatial Data Analysis by Example (Vol. 2: Categorical and Directional Data) by Graham J. G. Upton; Bernard Fingleton. Technometrics, 33(1):119, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269030.

Cohen:1992:BRB

[Coh92] Jonathan Cohen. Book review: Statistical Methods in Soil and Land Resource Survey by R. Webster; M. A. Oliver. Technometrics, 34(4):497–498, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268961.

Coleman:1992:BRB

[Col92] David Coleman. Book review: How to Test Normality and Other Distributional Assumptions by Samuel S. Shapiro. Technometrics, 34(3):359–360, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270050.

Coleman:1993:BRB

[Col93] David E. Coleman. Book review: Statistical Process Control: Theory and Practice by G. Barrie Wetherill; Don W. Brown. Technometrics, 35(1):89–90, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269298.

Cook:1993:EPR

[Coo93] R. Dennis Cook. Exploring partial residual plots. Technometrics, 35(4):351–362, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270269.

Cook:1996:AVP

[Coo96a] R. Dennis Cook. Added-variable plots and curvature in linear regression. Technometrics, 38(3):275–278, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270611.

Coons:1996:BRB

[Coo96b] Lori Coons. Book review: Quality up, Costs down: A Manager’s Guide to Taguchi Methods and QFD by William E. Eureka; Nancy E. Ryan. Technometrics, 38(2):183–184, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270415.

Cox:1993:BRB

[Cox93] Dennis D. Cox. Book review: Modal and Spectrum Analysis: Data Dependent Systems in State Space by Sudhakar M. Pandit. Technometrics, 35(2):228, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269677.

Cox:1996:BRB

[Cox96] Dennis D. Cox. Book review: Spectral Analysis for Physical Applications: Multitaper and Conventional Univariate Techniques by Donald B. Percival; Andrew T. Walden. Technometrics, 38(3): 294, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270624.

Christensen:1992:CDD

[CPJ92] Ronald Christensen, Larry M. Pearson, and Wesley Johnson. Case-deletion diagnostics for mixed models. Technometrics, 34 (1):38–45, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269550.

Carroll:1991:PTI

[CR91] Raymond J. Carroll and David Ruppert. Prediction and tolerance intervals with transformation and/or weighting. Technometrics, 33(2):197–210, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269046.

Chang:1993:IES

[CR93] Myron N. Chang and P. V. Rao. Improved estimation of survival functions in the new-better-than-used class. Technometrics, 35(2):192–203, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269664.

Coleman:1996:ASM

[CR96] David Coleman and Philip Ramsey. [Applications of Statistical Methods to Nondestructive Evaluation]: Discussion. Technometrics, 38(2):117–121, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270401. See [OM96a].

Cornell:1998:GMM

[CR98] John A. Cornell and Phillip J. Ramsey. A generalized mixture model for categorized-components problems with an application to a photoresist-coating experiment. Technometrics, 40(1):48– 61, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271392.

Crawford:1994:BRB

[Cra94] Sybil L. Crawford. Book review: Analysis of Quantal Response Data by B. J. T. Morgan. Technometrics, 36(1):117–118, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269212.

Cressie:1997:JPI

[Cre97] Noel Cressie. Jackknifing in the presence of inhomogeneity. Technometrics, 39(1):45–51, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270771.

Crowder:1992:SMS

[Cro92] Stephen V. Crowder. An SPC model for short production runs: Minimizing expected cost. Technometrics, 34(1):64–73, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269553.

Crosby:1994:BRB

[Cro94] Teri Crosby. Book review: How to Detect and Handle Outliers by Boris Iglewicz; David C. Hoaglin. Technometrics, 36(3):315– 316, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269377.

Crosby:1999:BRB

[Cro99] Teri A. Crosby. Book review: Into Statistics: A Guide to Understanding Statistical Concepts in Engineering and the Sciences by Peter J. Smith. Technometrics, 41(4):373–374, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271363.

Cryer:1990:BRB

[Cry90] Jonathan D. Cryer. Book review: The Statistical Theory of Linear Systems by E. J. Hannan; Manfred Deistler. Technometrics, 32(1):111, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269869.

Carroll:1992:DNH

[CS92] R. J. Carroll and C. H. Spiegelman. Diagnostics for nonlinearity and heteroscedasticity in errors-in-variables regression. Technometrics, 34(2):186–196, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269234.

Cohen:1993:ETI

[CS93] Arthur Cohen and H. B. Sackrowitz. Evaluating tests for increasing intensity of a Poisson process. Technometrics, 35(4):446–448, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270277.

Cucchiara:1992:BRB

[Cuc92] Andrew Cucchiara. Book review: Applied Logistic Regression by D. Hosmer; S. Lemeshow. Technometrics, 34(3):358–359, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270048.

Cheng:1997:RC

[CV97] Chi-Lun Cheng and John W. Van Ness. Robust calibration. Technometrics, 39(4):401–411, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271504.

Carey:1997:RTB

[CWWR97] V. J. Carey, E. E. Walters, C. G. Wager, and B. A. Rosner. Resistant and test-based outlier rejection: Effects on Gaussian oneand two-sample inference. Technometrics, 39(3):320–330, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271136.

Czitrom:1992:BRB

[Czi92] Veronica Czitrom. Book review: Experiments with Mixtures by John A. Cornell. Technometrics, 34(1):104–105, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269566.

Dalal:1990:SCB

[Dal90] S. R. Dalal. Simultaneous confidence bands for regression with unknown unequal variances. Technometrics, 32(2):173–186, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268861.

DasGupta:1993:BRB

[Das93] A. DasGupta. Book review: Sense and Nonsense of Statistical Inference by Chamont Wang. Technometrics, 35(4):462, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270295.

Davis:1999:BRB

[Dav99] Charles B. Davis. Book review: Compliance Quantified: An Introduction to Data Verification by Rudolf Avenhaus; Morton John Canty. Technometrics, 41(2):174, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270742.

delCastillo:1999:BRB

[dC99] Enrique del Castillo. Book review: SPC Methods for Quality Improvement by Charles P. Quesenberry. Technometrics, 41(2): 167–168, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270733.

Draper:1994:IDF

[DDPG94] Norman R. Draper, Timothy P. Davis, Lourdes Pozueta, and Daniel M. Grove. Isolation of degrees of freedom for Box–Behnken designs. Technometrics, 36(3):283–291, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269372.

Deal:1991:BRB

[Dea91] R. B. Deal, Jr. Book review: Systems Analysis Using Simulation and Markov Models by John R. Clymer. Technometrics, 33 (3):371–372, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268804.

Deal:1994:BRB

[Dea94] R. B. Deal. Book review: Simulation Modeling and Analysis by Averill M. Law; W. David Kelton. Technometrics, 36(4):429–430, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269971.

Dean:1995:BRB

[Dea95] Angela Dean. Book review: Advanced Linear Models: Theory and Applications by Song-Gui Wang; Shein-Chung Chow. Technometrics, 37(1):119–120, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269167.

Dean:1996:BRB

[Dea96] Angela M. Dean. Book review: Linear Statistical Models by James Stapleton. Technometrics, 38(4):402–403, November 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271313.

Dean:1997:BRB

[Dea97] Angela Dean. Book review: Methods and Applications of Linear Models by Ronald R. Hocking. Technometrics, 39(3):332–333, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271138.

Dehnad:1990:BRB

[Deh90] Khosrow Dehnad. Book review: Decision Analysis: A Bayesian Approach by J. Q. Smith. Technometrics, 32(2):232–233, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268883.

Derringer:1990:BRB

[Der90] George C. Derringer. Book review: The Design of Experiments: Statistical Principles for Practical Application by R. Mead. Technometrics, 32(4):448, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270125.

Deutsch:1996:CSH

[Deu96] Clayton V. Deutsch. Constrained smoothing of histograms and scatterplots with simulated annealing. Technometrics, 38(3):266– 274, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270610.

DeVeaux:1999:BRB

[DeV99] Richard D. DeVeaux. Book review: Applied Smoothing Techniques for Data Analysis by Adrian W. Bowman; Adelchi Azzalini. Technometrics, 41(3):263, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270572.

Dowling:1997:SIGa

[DGTZ97a] M. M. Dowling, P. M. Griffin, K.-L. Tsui, and C. Zhou. Statistical issues in geometric feature inspection using coordinate measuring machines. Technometrics, 39(1):3–17, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270764. See discussion [Hul97,

WH97, PE97] and reply [DGTZ97b].

Dowling:1997:SIGb

[DGTZ97b] M. M. Dowling, P. M. Griffin, K.-L. Tsui, and C. Zhou. [Statistical Issues in Geometric Feature Inspection Using Coordinate Measuring Machines]: Reply. Technometrics, 39(1):24, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270768. See

[DGTZ97a].

Doksum:1992:MVS

[DH92] Kjell A. Doksum and Arnljot Hóyland. Models for variable-stress accelerated life testing experiments based on Wiener processes and the inverse Gaussian distribution. Technometrics, 34(1):74– 82, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269554.

Dickey:1991:BRB

[Dic91] D. A. Dickey. Book review: The Analysis of Time Series: An Introduction by C. Chatfield. Technometrics, 33(3):363–364, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268794.

Dickey:1997:BRB

[Dic97] David A. Dickey. Book review: Introduction to Statistical Time Series by Wayne A. Fuller. Technometrics, 39(1):103–104, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270782.

Diegert:1992:BRB

[Die92] Kathleen Diegert. Book review: Industrial Quality and Reliability by G. Barnard; G. Box; D. Cox; A. Seheult; B. Silverman. Technometrics, 34(3):360, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270051.

DuMouchel:1994:SBM

[DJ94] William DuMouchel and Bradley Jones. A simple Bayesian modification of D-optimal designs to reduce dependence on an assumed model. Technometrics, 36(1):37–47, February 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269197.

Draper:1990:CBT

[DL90a] Norman R. Draper and Dennis K. J. Lin. Connections between two-level designs of resolutions III^{∗} and V. Technometrics, 32 (3):283–288, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269105.

Draper:1990:SRS

[DL90b] Norman R. Draper and Dennis K. J. Lin. Small response-surface designs. Technometrics, 32(2):187–194, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268862.

Dalal:1998:FCD

[DM98] Siddhartha R. Dalal and Colin L. Mallows. Factor-covering designs for testing software. Technometrics, 40(3):234–243, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271179.

Davis:1999:SNPa

[DM99a] Charles B. Davis and Roger J. McNichols. Simultaneous nonparametric prediction limits. Technometrics, 41(2):89–101, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270721. See

discussion [CM99, Gib99, Gil99, Van99] and response [DM99b].

Davis:1999:SNPb

[DM99b] Charles B. Davis and Roger J. McNichols. [Simultaneous Nonparametric Prediction Limits]: Response. Technometrics, 41(2): 108–112, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270726. See [DM99a].

Dobrow:1996:BRB

[Dob96] Robert P. Dobrow. Book review: Probability Models and Applications by Ingram Olkin; Leon J. Gleser; Cyrus Derman. Technometrics, 38(4):408–409, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271323.

Doganaksoy:1993:BRB

[Dog93] Necip Doganaksoy. Book review: Process Improvement in the Electronics Industry by Yefim Fasser; Donald Brettner. Technometrics, 35(3):334–335, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269542.

Doganaksoy:1999:BRB

[Dog99] Necip Doganaksoy. Book review: Quality and Reliability of Technical Systems: Theory, Practice, Management by Alessandro Birolini. Technometrics, 41(2):171–172, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270739.

Domangue:1990:BRB

[Dom90] Rickie J. Domangue. Book review: How to Think about Statistics by John L. Phillips. Technometrics, 32(1):94, February 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269848.

Donnelly:1990:BRB

[Don90] Patrick P. Donnelly. Book review: Problem Solving: A Statistician’s Guide by Christopher Chatfield. Technometrics, 32(3):349– 350, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269123.

Dovich:1993:BRB

[Dov93] Robert A. Dovich. Book review: Statistical Techniques for Data Analysis by John Keenan Taylor. Technometrics, 35(3):332, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269538.

Downing:1994:UCS

[Dow94] Darryl J. Downing. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Discussion. Technometrics, 36(2):147–149, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270223. See [Got94c].

Downing:1995:BRB

[Dow95] Darryl J. Downing. Book review: Univariate Tests for Time Series Models by Jeff B. Cromwell; Walter C. Labys; Michel Terraza. Technometrics, 37(3):346, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269923.

Draper:1990:ALR

[DP90] Norman R. Draper and Friedrich Pukelsheim. Another look at rotatability. Technometrics, 32(2):195–202, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268863.

Domangue:1991:SOE

[DP91] Rickie Domangue and Steven C. Patch. Some omnibus exponentially weighted moving average statistical process monitoring schemes. Technometrics, 33(3):299–313, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268782.

Durham:1997:CDM

[DP97] S. D. Durham and W. J. Padgett. Cumulative damage models for system failure with application to carbon fibers and composites. Technometrics, 39(1):34–44, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270770. See correction [Ano98e].

Draper:1993:MDF

[DPL^{+}93] N. R. Draper, P. Prescott, S. M. Lewis, A. M. Dean, P. W. M. John, and M. G. Tuck. Mixture designs for four components in orthogonal blocks. Technometrics, 35(3):268–276, August 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269519. See correction

[DPL^{+}94] and comments [SDP^{+}95]..

Draper:1994:CMD

[DPL^{+}94] N. R. Draper, P. Prescott, S. M. Lewis, A. M. Dean, P. W. M. John, and M. G. Tuck. Correction: “Mixture designs for four components in orthogonal blocks” [Technometrics 35 (1993), no. 3, 268–276; Draper:1993:MDF (94c:62110)]. Technometrics, 36 (2):234, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270264. See [DPL^{+}93].

Drew:1991:BRB

[Dre91] James H. Drew. Book review: Modern Data Analysis: A First Course in Applied Statistics by Lawrence C. Hamilton. Technometrics, 33(4):487–488, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269434.

Driscoll:1991:BRB

[Dri91] Michael F. Driscoll. Book review: Genstat 5: A Second Course by Pete Digby; Nick Galwey; Peter Lane. Technometrics, 33(3): 371, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268803.

Driscoll:1992:BRB

[Dri92] Michael F. Driscoll. Book review: Applied Statistics: A Handbook of GENSTAT Analyses by E. J. Snell; H. R. Simpson. Technometrics, 34(4):499, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268963.

Doganaksoy:1993:CAC

[DS93] Necip Doganaksoy and Josef Schmee. Comparisons of approximate confidence intervals for distributions used in life-data analysis. Technometrics, 35(2):175–184, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269662.

deVeaux:1998:PIN

[DSSU98] Richard D. De Veaux, Jennifer Schumi, Jason Schweinsberg, and Lyle H. Ungar. Prediction intervals for neural networks via nonlinear regression. Technometrics, 40(4):273–282, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270528.

Dunn:1990:BRB

[Dun90] Mark R. Dunn. Book review: Exploring Statistics with the IBM PC by David P. Doane. Technometrics, 32(3):350, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269124.

Dunn:1996:BRB

[Dun96] Mark R. Dunn. Book review: Achieving Quality Improvement: A Practical Guide by Roland Caulcutt. Technometrics, 38(2):183, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270413.

Dunn:1997:BRB

[Dun97] Mark R. Dunn. Book review: Introduction to Modern Statistical Quality Control and Management by J. A. Swift. Technometrics, 39(2):228, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270913.

deVeaux:1992:BRB

[dV92] Richard D. de Veaux. Book review: Generalized Additive Models by Trevor J. Hastie; Robert J. Tibshirani. Technometrics, 34(2):225–226, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269240.

deVeaux:1993:BRB

[dV93] Richard D. de Veaux. Book review: Statistics for Spatial Data by Noel A. C. Cressie. Technometrics, 35(3):322–323, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269525.

deVeaux:1996:NNA

[dVU96] Richard D. de Veaux and Lyle H. Ungar. [Neural Networks in Applied Statistics]: Discussion. Technometrics, 38(3):215–218, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270602. See

[Ste96a].

Easterling:1990:BRB

[Eas90] Robert G. Easterling. Book review: Prescriptions for Working Statisticians by Albert Madansky. Technometrics, 32(2):218–219, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268865.

Easterling:1999:BRB

[Eas99] Robert G. Easterling. Book review: Statistics and Truth by C. R. Rao. Technometrics, 41(1):72–73, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270996.

Edgeman:1992:BRB

[Edg92] Rick L. Edgeman. Book review: Engineering Quality by Design: Interpreting the Taguchi Approach by Thomas B. Barker. Technometrics, 34(2):230, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269247.

Edgeman:1998:BRB

[Edg98] Rick L. Edgeman. Book review: Statistical Methods for Industrial Process Control by David Drain. Technometrics, 40(2):154, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270649.

Engel:1996:GLM

[EH96] Jan Engel and A. Freek Huele. A generalized linear modeling approach to robust design. Technometrics, 38(4):365–373, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271307.

Eltinge:1990:BRB

[Elt90] John L. Eltinge. Book review: Applied Time Series Analysis for Business and Economic Forecasting by Sufi M. Nazem. Technometrics, 32(1):101–102, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269859.

Escobar:1995:PAL

[EM95] Luis A. Escobar and William Q. Meeker, Jr. Planning accelerated life tests with two or more experimental factors. Technometrics, 37(4):411–427, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269733.

Escobar:1999:SPB

[EM99] Luis A. Escobar and William Q. Meeker. Statistical prediction based on censored life data. Technometrics, 41(2):113–124, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270727.

Emptage:1992:BRB

[Emp92] Michael Emptage. Book review: An Introduction to Generalized Linear Models by A. J. Dobson. Technometrics, 34(2):224–225, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269239.

Emptage:1993:BRB

[Emp93] Michael R. Emptage. Book review: Sample Size Choice: Charts for Experiments with Linear Models by R. E. Odeh; M. Fox. Technometrics, 35(2):234–235, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269687.

Emptage:1997:BRB

[Emp97] Michael R. Emptage. Book review: Design and Optimization in Organic Synthesis; Experimental Design and Chemical Synthesis by Rolf Carlson. Technometrics, 39(2):231, May 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270918. See correction

[Ano97f].

Emptage:1999:BRB

[Emp99] Michael Emptage. Book review: Interaction Effects in Factorial Analysis of Variance by James Jaccard. Technometrics, 41(1): 80, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271008. See correction [Ano99e].

Erkanli:1998:BCC

[EMS98] Alaattin Erkanli, Thomas A. Mazzuchi, and Refik Soyer. Bayesian computations for a class of reliability growth models. Technometrics, 40(1):14–23, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271389.

Easterling:1991:SBC

[EMSD91] Robert G. Easterling, Mainak Mazumdar, Floyd W. Spencer, and Kathleen V. Diegert. System-based component-test plans and operating characteristics: Binomial data. Technometrics, 33(3):287– 298, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268781.

Ensor:1995:BRB

[Ens95] Katherine B. Ensor. Book review: Time Series Models by Andrew C. Harvey. Technometrics, 37(4):464–465, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269749.

Evans:1994:TAA

[ES94] J. C. Evans and S. J. Schwager. A test of additional accuracy for selecting groups of allocation variables. Technometrics, 36(2): 202–211, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270232.

Efromovich:1996:ANB

[ET96] Sam Efromovich and Edward V. Thomas. Application of nonparametric binary regression to evaluate the sensitivity of explosives. Technometrics, 38(1):50–58, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268902.

Eubank:1990:BRB

[Eub90] Randall Eubank. Book review: Nonparametric Regression Analysis of Longitudinal Data by Hans-Georg Müller. Technometrics, 32(2):220–221, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268867.

Eubank:1993:BRB

[Eub93] R. L. Eubank. Book review: Applied Nonparametric Regression by Wolfgang Härdle. Technometrics, 35(2):225–226, May 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269674.

Fairley:1991:BRB

[Fai91] David Fairley. Book review: Robust Regression: Analysis and Applications by K. D. Lawrence; J. L. Arthur. Technometrics, 33(2):239–240, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269052.

Fairley:1994:BRB

[Fai94] David Fairley. Book review: Alternative Methods of Regression by David Birkes; Yadolah Dodge. Technometrics, 36(3):319–321, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269381.

Faltin:1999:BRB

[Fal99] Frederick W. Faltin. Book review: Statistical Control by Monitoring and Freedback Adjustment by George Box; Alberto Luceño. Technometrics, 41(2):167, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270732.

Famoye:1995:BRB

[Fam95] Felix Famoye. Book review: Continuous Univariate Distributions, Volume 1 by Norman L. Johnson; Samuel Kotz; N. Balakrishnan. Technometrics, 37(4):466, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269751.

Famoye:1997:BRB

[Fam97] Felix Famoye. Book review: Seeing through Statistics by Jessica M. Utts. Technometrics, 39(1):108–109, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270791.

Famoye:1999:BRB

[Fam99] Felix Famoye. Book review: Improving Efficiency by Shrinkage by Marvin H. J. Gruber. Technometrics, 41(2):176, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270745.

Farnsworth:1990:BRB

[Far90] David L. Farnsworth. Book review: Elementary Statistics by Mario F. Triola. Technometrics, 32(4):456–457, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270136.

Farnsworth:1992:BRB

[Far92] David L. Farnsworth. Book review: Introduction to Statistics: The Nonparametric Way by Gottfried E. Noether. Technometrics, 34(2):240, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269261.

Farnsworth:1993:BRB

[Far93] David L. Farnsworth. Book review: A First Course in Order Statistics by Barry C. Arnold; N. Balakrishnan; H. N. Nagaraja. Technometrics, 35(4):452, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270281.

Faraway:1997:RAF

[Far97] Julian J. Faraway. Regression analysis for a functional response. Technometrics, 39(3):254–261, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271130.

Fuchs:1994:MPC

[FB94] Camil Fuchs and Yoav Benjamini. Multivariate profile charts for statistical process control. Technometrics, 36(2):182–195, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270230.

Fedorov:1996:FDR

[Fed96] Valerii Fedorov. [Follow-Up Designs to Resolve Confounding in Multifactor Experiments]: Discussion. Technometrics, 38(4):321– 322, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271300. See [MSB96a].

Fellner:1994:RLD

[Fel94] William H. Fellner. [Run-Length Distributions of Special-Cause Control Charts for Correlated Processes]: Discussion. Technometrics, 36(1):22–23, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269194. See [WMP94a].

Frank:1993:SVSb

[FF93a] Ildiko E. Frank and Jerome H. Friedman. [A Statistical View of Some Chemometrics Regression Tools]: Response. Technometrics, 35(2):143–148, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269659. See [FF93b].

Frank:1993:SVSa

[FF93b] Ildiko E. Frank and Jerome H. Friedman. A statistical view of some chemometrics regression tools. Technometrics, 35(2):109– 135, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269656. See discussion [HM93].

Faltin:1990:EWM

[FHT90] Frederick W. Faltin, Gerald J. Hahn, and William T. Tucker. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]: Discussion. Technometrics, 32(1):19– 20, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269838. See [LS90b].

Freeny:1995:GAL

[FL95] Anne E. Freeny and James M. Landwehr. Graphical analysis for a large designed experiment. Technometrics, 37(1):1–14, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269145.

Filliben:1997:SAA

[FL97] James J. Filliben and Ker-Chau Li. A systematic approach to the analysis of complex interaction patterns in two-level factorial designs. Technometrics, 39(3):286–297, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271133.

Farebrother:1992:CMQ

[FMQ92] R. W. Farebrother, Ray L. Marr, and Charles P. Quesenberry. Comments on Marr and Quesenberry (1991). Technometrics, 34 (3):375–376, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1270076. See [MQ91] and correction [Ano93f].

Fotopoulos:1995:BRB

[Fot95] Stergios B. Fotopoulos. Book review: Optimization Techniques in Statistics by Jagdish S. Rustagi. Technometrics, 37(4):466–467, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269752.

Fotopoulos:1996:BRB

[Fot96] Stergios B. Fotopoulos. Book review: Stochastic Modeling and Analysis of Manufacturing Systems by David D. Yao. Technometrics, 38(1):84–85, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268919.

Fox:1991:BRB

[Fox91] Eric P. Fox. Book review: Bayesian Statistics 3 by J. Bernardo; M. DeGroot; D. Lindley; A. Smith. Technometrics, 33(2):245, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269061.

Fox:1994:BRB

[Fox94] Eric P. Fox. Book review: Reliability Improvement with Design of Experiments by Lloyd W. Condra. Technometrics, 36(3):316– 317, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269378.

Fox:1997:BRB

[Fox97] Eric P. Fox. Book review: Regression Models by Richard Breen. Technometrics, 39(2):233, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270921.

Fox:1998:BRB

[Fox98] Eric P. Fox. Book review: Data Analysis: A Bayesian Tutorial by D. S. Sivia. Technometrics, 40(2):155–156, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270652.

Freitas:1992:BRB

[Fre92] Stephen A. Freitas. Book review: Statistics: A Fresh Approach by Donald H. Sanders. Technometrics, 34(2):244–245, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269269.

Frey:1993:BRB

[Fre93] Michael Frey. Book review: Stochastic Processes by D. R. Cox; V. S. Isham. Technometrics, 35(3):329–330, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269533.

Frey:1994:BRB

[Fre94] Michael Frey. Book review: An Introduction to Stochastic Modeling by H. M. Taylor; S. Karlin. Technometrics, 36(4):428–429, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269970.

Frey:1995:BRB

[Fre95] Michael Frey. Book review: Time Series Analysis by Higher Order Crossings by Benjamin Kedem. Technometrics, 37(2):235, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269635.

Frey:1997:BRB

[Fre97] Michael Frey. Book review: Stochastic Processes by Sheldon Ross. Technometrics, 39(2):230–231, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270917.

Freitas:1999:BRB

[Fre99] Stephen A. Freitas. Book review: Modern Industrial Statistics: Design and Control of Quality and Reliability by Ron S. Kenett; Shelemyahu Zacks. Technometrics, 41(3):263–264, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270573.

Friedman:1991:MEM

[Fri91] Jerome H. Friedman. [The ∏ Method for Estimating Multivariate Functions from Noisy Data]: Discussion. Technometrics, 33 (2):145–148, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269039. See [Bre91a].

Fricker:1997:BRB

[Fri97] Ronald D. Fricker, Jr. Book review: Statistical Problem Solving in Quality Engineering by Thomas J. Kazmierski. Technometrics, 39(2):227–228, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270912.

Flehinger:1996:IAD

[FRY96] Betty J. Flehinger, Benjamin Reiser, and Emmanuel Yashchin. Inference about defects in the presence of masking. Technometrics, 38(3):247–255, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270608.

Fulkerson:1995:BRB

[Ful95] Bill Fulkerson. Book review: Machine Learning, Neural and Statistical Classification by D. Michie; D. J. Spiegelhalter; C. C. Taylor. Technometrics, 37(4):459, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269742.

Fulkerson:1997:BRB

[Ful97] Bill Fulkerson. Book review: Pattern Recognition and Neural Networks by B. D. Ripley. Technometrics, 39(2):233–234, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270922.

Galway:1992:BRB

[Gal92] Lionel Galway. Book review: Spline Models for Observational Data by Grace Wahba. Technometrics, 34(1):113–114, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269578.

Galway:1993:BRB

[Gal93] Lionel Galway. Book review: Probability and Statistics in Experimental Physics by Byron P. Roe. Technometrics, 35(4):460–461, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270292.

Gallo:1995:BRB

[Gal95] Paul P. Gallo. Book review: Random Coefficient Models by Nicholas T. Longford. Technometrics, 37(1):120, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269168.

Gan:1995:JMP

[Gan95] F. F. Gan. Joint monitoring of process mean and variance using exponentially weighted moving average control charts. Technometrics, 37(4):446–453, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269736.

Ganocy:1997:BRB

[Gan97] Stephen J. Ganocy. Book review: Design of Experiments for Process Improvement and Quality Assurance by Robert F. Brewer. Technometrics, 39(4):425, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271507.

Gastwirth:1994:BRB

[Gas94] J. L. Gastwirth. Book review: Applied Nonparametric Statistical Methods by Peter Sprent. Technometrics, 36(2):216–217, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270236.

Gasparini:1997:BRB

[Gas97] Mauro Gasparini. Book review: Markov Chain Monte Carlo in Practice by W. R. Gilks; S. Richardson; D. J. Spiegelhalter. Technometrics, 39(3):338, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271145.

Gatliffe:1996:BRB

[Gat96] Thomas R. Gatliffe. Book review: Statistics for Environmental Engineers by Paul Mac Berthouex; Linfield C. Brown. Technometrics, 38(2):182–183, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270412.

Grondona:1991:USC

[GC91] Martin O. Grondona and Noel Cressie. Using spatial considerations in the analysis of experiments. Technometrics, 33(4):381– 392, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269411.

Grove:1991:TIC

[GD91] D. M. Grove and T. P. Davis. Taguchi’s idle column method. Technometrics, 33(3):349–353, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268786.

Gilbert:1994:UCS

[GD94] Richard O. Gilbert and Pamela G. Doctor. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]:

Discussion. Technometrics, 36(2):141–144, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270221. See [Got94c].

George:1994:BRB

[Geo94] Larry George. Book review: What Every Engineer Should Know about Reliability and Risk Analysis by M. Modarres. Technometrics, 36(2):226–227, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270248.

Georgiev:1998:BRB

[Geo98] Alexander A. Georgiev. Book review: Functional Data Analysis by Jim Ramsay; Bernard Silverman. Technometrics, 40(3):260– 261, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271190.

Ghosh:1993:BRB

[Gho93] Subir Ghosh. Book review: Variance Components by Shayle R. Searle; George Casella; Charles E. McCulloch. Technometrics, 35(2):221–222, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269669.

Ghosh:1994:BRB

[Gho94] Subir Ghosh. Book review: Statistical Process Control for Quality Improvement by James R. Thompson; Jacek Koronacki. Technometrics, 36(1):120, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269216.

Goldberg:1992:BEB

[GI92] Kenneth M. Goldberg and Boris Iglewicz. Bivariate extensions of the boxplot. Technometrics, 34(3):307–320, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270037.

Gibbons:1991:BRB

[Gib91] Ned M. Gibbons. Book review: Statistics in Scientific Investigation by Glen McPherson. Technometrics, 33(3):356–357, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268788.

Gibbons:1999:SNP

[Gib99] Robert D. Gibbons. [Simultaneous Nonparametric Prediction Limits]: Discussion. Technometrics, 41(2):104–106, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270723. See [DM99a].

Gill:1992:BRB

[Gil92] P. S. Gill. Book review: Introduction to Linear Models and Experimental Design by J. Arthur Woodward; Douglas G. Bonett; Mary-Lynn Brecht. Technometrics, 34(3):354, August 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270042.

Gilbert:1999:SNP

[Gil99] Richard O. Gilbert. [Simultaneous Nonparametric Prediction Limits]: Discussion. Technometrics, 41(2):106–107, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270724. See [DM99a].

Garcia-Jurado:1995:PUB

[GJMS^{+}95] I. García-Jurado, W. González Manteiga, J. M. Prada Sánchez, M. Febrero-Bande, and R. Cao. Predicting using Box-Jenkins, nonparametric, and bootstrap techniques. Technometrics, 37(3): 303–310, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269914.

Gan:1990:GFT

[GK90] F. F. Gan and K. J. Koehler. Goodness-of-fit tests based on P – P probability plots. Technometrics, 32(3):289–303, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269106.

Gunst:1993:CAI

[GK93] Richard F. Gunst and Nelson A. Kelly. Captive-air irradiation experiments on ozone formation in Southern California. Technometrics, 35(3):256–267, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269518.

Glaz:1998:BRB

[Gla98] Joseph Glaz. Book review: Frontiers in Queueing: Models and Applications in Science and Engineering by Jewgeni H. Dshalalow.

Technometrics, 40(2):162–163, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270662.

Gleser:1991:BRB

[Gle91] Leon Jay Gleser. Book review: Design and Analysis of Reliability Studies by Graham Dunn. Technometrics, 33(4):472–473, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269420.

Gleser:1996:BRB

[Gle96] Leon Jay Gleser. Book review: Measurement, Regression, and Calibration by Philip J. Brown. Technometrics, 38(1):75, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268905.

Glosup:1993:BRB

[Glo93] Jeffrey Glosup. Book review: Handbook of the Logistic Distribution by N. Balakrishnan. Technometrics, 35(2):236–237, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269689.

Glosup:1995:BRB

[Glo95] Jeffrey Glosup. Book review: Multivariate Statistical Modelling Based on Generalized Linear Models by L. Fahrmeir; G. Tutz. Technometrics, 37(3):350–351, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269929.

Glosup:1997:BRB

[Glo97] Jeffrey Glosup. Book review: Statistics for Long-Memory Processes by Jan Beran. Technometrics, 39(1):105–106, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270785.

Glosup:1999:BRB

[Glo99] Jeffrey Glosup. Book review: Statistical Evidence: A Likelihood Paradigm by Richard Royall. Technometrics, 41(1):80–81, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271009.

Godfrey:1990:CCB

[God90] A. Blanton Godfrey. Commentary: Communications between Statisticians and Engineers/Physical Scientists: Does anyone care? Technometrics, 32(3):253–255, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269095.

Godsey:1993:BRB

[God93] E. B. Godsey. Book review: Short Run SPC by Donald J. Wheeler. Technometrics, 35(2):233, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269684.

Godsey:1995:BRB

[God95] E. B. Godsey. Book review: Bilateral Plans for Inspection and Process Control by Variables by Gaetano Panizzon. Technometrics, 37(4):463–464, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269748.

Goik:1996:BRB

[Goi96] Patrick P. Goik. Book review: Practical Engineering Statistics by Daniel Schiff; Ralph D’Agostino. Technometrics, 38(4):404, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271316.

Golland:1990:BRB

[Gol90] Ronald W. Golland. Book review: Principles of Random Variate Generation by John Dagpunar. Technometrics, 32(2):226–227, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268875.

Goldstein:1993:BRB

[Gol93] Richard Goldstein. Book review: Conditioning Diagnostics: Collinearity and Weak Data in Regression by David A. Belsley. Technometrics, 35(1):85–86, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269293.

Goldman:1995:BRB

[Gol95] Aaron Goldman. Book review: An Introduction to Regression Graphics by R. Dennis Cook; Sanford Weisberg. Technometrics, 37(3):343–344, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269921.

Goldstein:1998:BRB

[Gol98] Richard Goldstein. Book review: Latent Class and Discrete Latent Trait Models: Similarities and Differences by Ton Heinen. Technometrics, 40(1):81, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271408.

Goldberg:1999:BRB

[Gol99] Judith D. Goldberg. Book review: Applied Survival Analysis by Chap T. Le. Technometrics, 41(3):267–268, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270580.

Goodall:1991:BRB

[Goo91] Colin R. Goodall. Book review: Graphical Models in Applied Multivariate Statistics by J. Whittaker. Technometrics, 33(4): 476–478, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269424.

Goodall:1992:BRB

[Goo92a] Colin Goodall. Book review: Nonlinear Multivariate Analysis by Albert Gifi. Technometrics, 34(3):357–358, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270046.

Goodman:1992:BRB

[Goo92b] Jay H. Goodman. Book review: Statistics in Spectroscopy by Howard Mark; Jerry Workman. Technometrics, 34(4):498–499, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268962.

Goodfriend:1995:BRB

[Goo95] Jason H. Goodfriend. Book review: Discrete Event Systems: Sensitivity Analysis and Stochastic Optimization by the Score Function Method by Reuven Y. Rubinstein; Alexander Shapiro. Technometrics, 37(1):124–125, February 1995. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269174.

Gotway:1992:BRB

[Got92] Carol A. Gotway. Book review: Spatial Data Analysis in the Social and Environmental Sciences by Robert Haining. Technometrics, 34(4):499–500, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268964.

Gotway:1994:BRB

[Got94a] Carol A. Gotway. Book review: Statistical Methods in Water Resources by D. R. Helsel; R. M. Hirsch. Technometrics, 36(3):323– 324, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269385.

Gotway:1994:UCSb

[Got94b] Carol A. Gotway. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Reply. Technometrics, 36 (2):156–161, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270227. See [Got94c].

Gotway:1994:UCSa

[Got94c] Carol A. Gotway. The use of conditional simulation in nuclearwaste-site performance assessment. Technometrics, 36(2):129–141, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270220. See

discussion [GD94, WB94, Dow94, Jou94, Tor94, Rut94] and reply [Got94b].

Govindarajulu:1992:BRB

[Gov92] Z. Govindarajulu. Book review: Rank Correlation Methods by M. Kendall; J. D. Gibbons. Technometrics, 34(1):108, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269571.

Gaudard:1997:BRB

[GR97] Marie Gaudard and Philip J. Ramsey. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 39(3):332, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271137.

Gray:1990:BRB

[Gra90] J. Brian Gray. Book review: Multivariate Exploratory Data Analysis by Allen Yates. Technometrics, 32(3):356–357, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269132.

Gray:1994:BRB

[Gra94] J. Brian Gray. Book review: Regression with Graphics: A Second Course in Applied Statistics by Lawrence C. Hamilton. Technometrics, 36(1):112–113, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269206.

Gray:1998:BRB

[Gra98] J. Brian Gray. Book review: Applied Regression Analysis, Linear Models, and Related Methods by John Fox. Technometrics, 40 (2):156, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270653.

Greer:1994:BRB

[Gre94] Daniel A. Greer. Book review: Design of Experiments: A No-Name Approach by Thomas J. Lorenzen; Virgil L. Anderson. Technometrics, 36(3):317–318, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269379.

Grego:1998:BRB

[Gre98] John M. Grego. Book review: Practical Data Analysis for Designed Experiments by Brian S. Yandell. Technometrics, 40(2): 154–155, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270650.

Grice:1990:BRB

[Gri90] John V. Grice. Book review: Parts Per Million Values for Estimating Quality Levels by R. E. Odeh; D. B. Owen. Technometrics, 32(1):98, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269855.

Grimshaw:1993:CML

[Gri93] Scott D. Grimshaw. Computing maximum likelihood estimates for the generalized Pareto distribution. Technometrics, 35(2):185–191, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269663.

Grimshaw:1995:BRB

[Gri95] Scott D. Grimshaw. Book review: An Introduction to the Bootstrap by Bradley Efron; Robert J. Tibshirani. Technometrics, 37 (3):341, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269918.

Grice:1996:BRB

[Gri96] John V. Grice. Book review: A Data-Based Approach to Statistics by Ronald L. Iman. Technometrics, 38(4):405, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271318.

Griffith:1997:BRB

[Gri97a] William S. Griffith. Book review: Applied Methods of Structural Reliability by Milík Tichý. Technometrics, 39(1):104–105, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270784.

Grillenzoni:1997:RGE

[Gri97b] Carlo Grillenzoni. Recursive generalized M-estimators of system parameters. Technometrics, 39(2):211–224, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270909.

Griffiths:1998:BRB

[Gri98] Kristi L. Griffiths. Book review: Handbook of Experimental Methods for Process Improvement by David Drain. Technometrics, 40(1):74, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271396.

Griffith:1999:BRB

[Gri99] William S. Griffith. Book review: Mathematical Theory of Reliability of Time Dependent Systems with Practical Applications by Igor N. Kovalenko; Nickolaj Yu. Kuznetsoy; Philip A. Pegg. Technometrics, 41(3):269, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270582.

Groggel:1992:BRB

[Gro92] David J. Groggel. Book review: Probability and Statistics by Athanasios Papoulis. Technometrics, 34(2):243, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269266.

Groggel:1996:BRB

[Gro96] David J. Groggel. Book review: Analyzing and Modeling Rank Data by John I. Marden. Technometrics, 38(4):403, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271314.

Groggel:1999:BRB

[Gro99] David J. Groggel. Book review: Nonparametric Methods for Quantitative Analysis by Jean Dickinson Gibbons. Technometrics, 41(1):75, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271000.

Gruber:1990:BRB

[Gru90] Marvin H. J. Gruber. Book review: Bayesian Spectrum Analysis and Parameter Estimation by G. Larry Bretthorst. Technometrics, 32(2):226, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268874.

Gruber:1995:BRB

[Gru95] Marvin H. J. Gruber. Book review: Reliability and Decision Making by Richard E. Barlow; Carlo A. Clarotti; Fabio Spizzichino. Technometrics, 37(3):346–347, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269924.

Gruber:1997:BRB

[Gru97] Marvin H. J. Gruber. Book review: Statistical Digital Signal Processing and Modeling by Monson H. Hayes. Technometrics, 39 (3):335–336, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271141.

Greenberg:1992:ENR

[GS92] Betsy S. Greenberg and S. Lynne Stokes. Estimating nonconformance rates after zero-defect sampling with rectification. Technometrics, 34(2):203–213, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269236.

Gurland:1994:RIF

[GS94] John Gurland and Jayaram Sethuraman. Reversal of increasing failure rates when pooling failure data. Technometrics, 36 (4):416–418, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269957.

Greenberg:1995:RTP

[GS95] Betsy S. Greenberg and S. Lynne Stokes. Repetitive testing in the presence of inspection errors. Technometrics, 37(1):102–111, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269157.

Grimshaw:1998:RTP

[GSH98] Scott D. Grimshaw, Scott D. Shellman, and Arnon M. Hurwitz. Real-time process monitoring for changing inputs. Technometrics, 40(4):283–296, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270529.

Gu:1991:MEM

[Gu91] Chong Gu. [The ∏ Method for Estimating Multivariate Functions from Noisy Data]: Discussion. Technometrics, 33(2):149–154, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269040. See

[Bre91a].

Guess:1990:BRB

[Gue90] Frank M. Guess. Book review: Bayesian Statistics: Principles, Models, and Applications by S. James Press. Technometrics, 32 (4):453–454, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270132.

Gunst:1990:BRB

[Gun90] Richard F. Gunst. Book review: MINITAB Supplement to Accompany General Statistics, by Warren Chase and Fred Brown by Anne Sevin. Technometrics, 32(4):456, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270135.

Gunter:1993:SAP

[Gun93] Berton H. Gunter. [A Systematic Approach to Planning for a Designed Industrial Experiment]: Discussion. Technometrics, 35 (1):13–14, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269281. See [CM93b].

Gunter:1994:BRBb

[Gun94a] Bert Gunter. Book review: A First Course in Probability Models and Statistical Inference by J. H. C. Creighton. Technometrics, 36(4):428, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269969.

Gunter:1994:BRBa

[Gun94b] Bert Gunter. Book review: Visualizing Data by William S. Cleveland. Technometrics, 36(3):314–315, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269376.

Gunst:1996:BRB

[Gun96a] Richard F. Gunst. Book review: Response Surface Methodology: Process and Product Optimization Using Designed Experiments by Raymond H. Myers; Douglas C. Montgomery. Technometrics, 38 (3):285–286, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270613.

Gunter:1996:BRB

[Gun96b] Berton H. Gunter. Book review: Practical Continuous Improvement for Professional Services by Clive Shearer. Technometrics, 38(1):84, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268917.

Gunst:1999:BRB

[Gun99] Richard F. Gunst. Book review: Applied Regression Analysis by Norman R. Draper; Harry Smith. Technometrics, 41(3):265– 266, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270576.

Guttorp:1990:BRB

[Gut90] Peter Guttorp. Book review: Fuzzy Mathematical Models in Engineering and Management Science by A. Kaufman; M. Gupta. Technometrics, 32(2):238, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268889.

Guttorp:1992:BRB

[Gut92] Peter Guttorp. Book review: The Art of Probability for Scientists and Engineers by Richard W. Hamming. Technometrics, 34 (2):242, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269264.

Haining:1993:EPT

[HA93] Robert Haining and Giuseppe Arbia. Error propagation through map operations. Technometrics, 35(3):293–305, August 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269522.

Haberman:1990:CLA

[Hab90] Shelby J. Haberman. [A Critical Look at Accumulation Analysis and Related Methods]: Discussion. Technometrics, 32(2):131–132, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268854. See

[HW90b].

Habibullah:1992:BRB

[Hab92] Mohamed Habibullah. Book review: Log-Linear Models by R. Christensen. Technometrics, 34(2):227, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269241.

Hadi:1992:BRB

[Had92] Ali S. Hadi. Book review: Finding Groups in Data: An Introduction to Cluster Analysis by L. Kaufman; P. J. Rousseeuw.

Technometrics, 34(1):111–112, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269576.

Hagwood:1990:BRB

[Hag90] Charles Hagwood. Book review: Elementary Applications of Probability Theory by H. Tuckwell. Technometrics, 32(1):112–113, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269872.

Hagwood:1991:BRB

[Hag91] Charles Hagwood. Book review: Multivariate Calibration by H. Martens; T. Naes. Technometrics, 33(3):366–367, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268798.

Hagwood:1996:BRB

[Hag96] Charles Hagwood. Book review: Mathematical Analysis of Spectral Orthogonality by John H. Kalivas; Patrick M. Lang. Technometrics, 38(1):86, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268921.

Hahn:1990:CCB

[Hah90] Gerald J. Hahn. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 257–258, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269096. See [HK90].

Hahn:1993:SAP

[Hah93] Gerald J. Hahn. [A Systematic Approach to Planning for a Designed Industrial Experiment]: Discussion. Technometrics, 35 (1):15–17, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269282. See [CM93b].

Halteman:1994:BRB

[Hal94] Edward J. Halteman. Book review: SPC for the Rest of Us: A Personal Path to Statistical Process Control by Hy Pitt. Technometrics, 36(4):424, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269963.

Hall:1995:BRB

[Hal95] Timothy Hall. Book review: Numerical Solution of SDE Through Computer Experiments by Peter E. Kloeden; Eckhard Platen; Henri Schurz. Technometrics, 37(1):125, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269175.

Hall:1998:BRB

[Hal98] Kathryn Hall. Book review: SPC Essentials and Productivity Improvement by William A. Levinson; Frank Tumbelty. Technometrics, 40(1):73, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271394.

Hamilton:1996:BRB

[Ham96] Martin A. Hamilton. Book review: Environmental Statistics and Data Analysis by Wayne R. Ott. Technometrics, 38(3):292– 293, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270622.

Hand:1990:BRB

[Han90] David J. Hand. Book review: Readings in Decision Analysis by Simon French. Technometrics, 32(4):463–464, November 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270147.

Hansen:1993:BRB

[Han93] James L. Hansen. Book review: Process Control: Statistical Principles and Tools by Ronald Gulezian. Technometrics, 35(2):230– 231, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269680.

Hansen:1997:BRB

[Han97] Christian K. Hansen. Book review: Practical Methods for Reliability Data Analysis by J. I. Ansell; M. J. Phillips. Technometrics, 39(2):226, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270910.

Hansen:1998:BRB

[Han98] Mark Hansen. Book review: Applied Wavelet Analysis with S-PLUS by Andrew Bruce; Hong-Ye Gao. Technometrics, 40(2):159, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270657.

Hapeman:1991:BRB

[Hap91] David E. Hapeman. Book review: Categorical Data Analysis by Alan Agresti. Technometrics, 33(2):241, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269056.

Hapeman:1992:BRB

[Hap92] David E. Hapeman. Book review: The Statistical Analysis of Categorical Data by Erling B. Anderson. Technometrics, 34(3):364, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270057.

Hardin:1991:BRB

[Har91a] J. Michael Hardin. Book review: Robust Estimation and Testing by Robert G. Staudte; Simon J. Sheather. Technometrics, 33 (4):478–479, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269426.

Harper:1991:BRB

[Har91b] William V. Harper. Book review: Statistics: Theory and Methods by Donald A. Berry; Bernard W. Lindgren. Technometrics, 33(3):369–370, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268801.

Harris:1991:BRBb

[Har91c] Bernard Harris. Book review: Parametric Statistical Models and Likelihood (Lecture Notes in Statistics 50) by Ole E. Barndorff-Nielsen. Technometrics, 33(4):493–494, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269443.

Harris:1991:BRBa

[Har91d] Thomas J. Harris. Book review: Adaptive Control by J. Astrom; B. Wittenmark. Technometrics, 33(4):487, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269433.

Harris:1992:BRB

[Har92a] Thomas Harris. Book review: Time Series Analysis, Identification and Adaptive Filtering by Daniel Graupe. Technometrics, 34(3):355–356, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270044.

Harris:1992:OCN

[Har92b] Thomas J. Harris. Optimal controllers for nonsymmetric and nonquadratic loss functions. Technometrics, 34(3):298–306, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270036.

Hassett:1991:BRB

[Has91] Nancy C. Hassett. Book review: Statistical Data Analysis Handbook by Francis Wall. Technometrics, 33(1):113–114, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269020.

Hawkins:1990:BRB

[Haw90] Douglas M. Hawkins. Book review: Multivariate Statistics: A Practical Approach by Bernhard Flury; Hans Riedwyl. Technometrics, 32(1):98–99, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269856.

Hawkins:1991:DUR

[Haw91a] Douglas M. Hawkins. Diagnostics for use with regression recursive residuals. Technometrics, 33(2):221–234, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269048. See correction

[Ano91e].

Hawkins:1991:MQC

[Haw91b] Douglas M. Hawkins. Multivariate quality control based on regression-adjusted variables. Technometrics, 33(1):61–75, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269008. See

correction [Ano91f].

Hawkins:1994:LE

[Haw94] Daniel B. Hawkins. Letter to the Editor. Technometrics, 36(4): 438, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269987.

Hamilton:1991:DAS

[HC91] Martin A. Hamilton and Bruce Jay Collings. Determining the appropriate sample size for nonparametric tests for location shift. Technometrics, 33(3):327–337, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268784.

Huang:1998:MAT

[HCV98] Peng Huang, Dechang Chen, and Joseph O. Voelkel. Minimumaberration two-level split-plot designs. Technometrics, 40(4):314– 326, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270532.

Hahn:1995:STC

[HD95] Gerald J. Hahn and Necip Doganaksoy. [Shewhart-Type Charts in Nonstandard Situations]: Discussion. Technometrics, 37(1):29– 31, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269149. See [RD95a].

Healy:1990:CCV

[Hea90] John Healy. [CUSUM Charts with Variable Sampling Intervals]: Discussion. Technometrics, 32(4):385–386, November 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270115. See [RAA90a].

Heavlin:1994:BRB

[Hea94] William D. Heavlin. Book review: Pyzdek’s Guide to SPC: Volume Two, Applications and Special Topics by Thomas Pyzdek. Technometrics, 36(4):422–424, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269962.

Heckler:1990:BRB

[Hec90] Charles E. Heckler. Book review: Mastering the SAS System by Jay A. Jaffe. Technometrics, 32(3):350–351, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269125.

Heckler:1995:BRB

[Hec95] Charles E. Heckler. Book review: Applied Discriminant Analysis by Carl J. Huberty. Technometrics, 37(3):349–350, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269928.

Heckler:1996:BRB

[Hec96] Charles E. Heckler. Book review: A Step-by-Step Approach to Using the SAS(R) System for Factor Analysis and Structural Equation Modeling by Larry Hatcher. Technometrics, 38(3):296– 297, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270628.

Heckler:1999:BRB

[Hec99] Charles E. Heckler. Book review: Statistical Graphics for Visualizing Multivariate Data by William G. Jacoby. Technometrics, 41(4):369, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271355.

Heitkamp:1990:BRB

[Hei90] Norman Heitkamp. Book review: SAS System for Elementary Statistical Analysis by Sandra D. Schlotzhauer; Ramon C. Littell; Cary. Technometrics, 32(2):230–231, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268879.

Hembree:1999:BRB

[Hem99] Barry Hembree. Book review: The SPC Troubleshooting Guide by Ronald Blank. Technometrics, 41(3):266, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270577.

Hess:1991:BRB

[Hes91] James L. Hess. Book review: Statistics: A Guide to the Use of Statistical Methods in the Physical Sciences by R. J. Barlow. Technometrics, 33(1):108, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269014.

Hesterberg:1993:BRB

[Hes93] Tim Hesterberg. Book review: Statistical Models in S by John M. Chambers; Trevor J. Hastie. Technometrics, 35(2):227–228, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269676.

Hesterberg:1995:WAI

[Hes95] Tim Hesterberg. Weighted average importance sampling and defensive mixture distributions. Technometrics, 37(2):185–194, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269620.

Hesterberg:1997:BRB

[Hes97] Tim Hesterberg. Book review: The Jackknife and Bootstrap by Jun Shao; Dongsheng Tu. Technometrics, 39(4):429, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271515.

Heydorn:1992:BRB

[Hey92] Richard P. Heydorn. Book review: Practical Reliability Engineering by Patrick D. T. O’Connor. Technometrics, 34(4):487–488, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268949.

Heyes:1993:BRB

[Hey93] Gerald B. Heyes. Book review: Statistical Problem Solving by Wendell E. Carr. Technometrics, 35(3):333–334, August 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269541.

Heydorn:1998:BRB

[Hey98] Richard P. Heydorn. Book review: Reliability and Maintainability Engineering by Charles E. Ebeling. Technometrics, 40(3):256– 257, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271182.

Halverson:1996:BRB

[HH96] Tim Halverson and Ian Hau. Book review: The Next Phase of Total Quality Management: TQM II and Focus on Profitability by Robert Stein. Technometrics, 38(3):295, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270626.

Hahn:1998:KCSa

[HH98a] Gerry Hahn and Roger Hoerl. Key challenges for statisticians in business and industry. Technometrics, 40(3):195–200, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271171. See dis-

cussion [MTV98, Nem98] and summary [HH98b].

Hahn:1998:KCSb

[HH98b] Gerry Hahn and Roger Hoerl. [Key Challenges for Statisticians in Business and Industry]: Summary. Technometrics, 40(3):211– 213, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271176. See [HH98a].

Higbee:1998:BRB

[Hig98] Kenneth Higbee. Book review: Mathematical Classification and Clustering by Boris Mirkin. Technometrics, 40(1):80, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271406.

Hillyer:1995:BRB

[Hil95] Martin Hillyer. Book review: Statistical Methods in Analytical Chemistry by Peter C. Meier; Richard E. Zünd. Technometrics, 37(1):113–114, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269159.

Hinde:1997:BRB

[Hin97] John Hinde. Book review: XploRe: An Interactive Statistical Computing Environment by W. Härdle; S. Klinke; B. A. Turlach. Technometrics, 39(4):430–431, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271517.

Hirotsu:1990:CLA

[Hir90] Chihiro Hirotsu. [A Critical Look at Accumulation Analysis and Related Methods]: Discussion. Technometrics, 32(2):133–136, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268855. See

[HW90b].

Hirst:1996:ERE

[Hir96] David Hirst. Error-rate estimation in multiple-group linear discriminant analysis. Technometrics, 38(4):389–399, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271310.

Hoadley:1990:CBS

[HK90] A. Bruce Hoadley and J. R. Kettenring. Communications between statisticians and engineers/physical scientists. Technometrics, 32(3):243–247, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1269092. See commentary [Arm90, Box90, God90, Hah90, Hoo90, Hun90a, Mac90b, Moo90, Sne90, Tri90, Wan90].

Hamilton:1995:CMU

[HL95] David C. Hamilton and Mary L. Lesperance. A comparison of methods for univariate and multivariate acceptance sampling by variables. Technometrics, 37(3):329–339, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269917.

Hirose:1997:IGD

[HL97] Hideo Hirose and Tze Leung Lai. Inference from grouped data in three-parameter Weibull models with applications to breakdownvoltage experiments. Technometrics, 39(2):199–210, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270908.

Holliday:1998:EME

[HLD98] Timothy Holliday, Anthony J. Lawrance, and Timothy P. Davis. Engine-mapping experiments: A two-stage regression approach. Technometrics, 40(2):120–126, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270646.

Hu:1998:NEL

[HLS98] X. Joan Hu, Jerald F. Lawless, and Kazuyuki Suzuki. Nonparametric estimation of a lifetime distribution when censoring times are missing. Technometrics, 40(1):3–13, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271388.

Hastie:1993:SVS

[HM93] Trevor Hastie and Colin Mallows. [A Statistical View of Some Chemometrics Regression Tools]: Discussion. Technometrics, 35 (2):140–143, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1269658. See [FF93b] and response [FF93a].

Hansen:1997:MWM

[HNF97] Mark H. Hansen, Vijayan N. Nair, and David J. Friedman. Monitoring wafer map data from integrated circuit fabrication processes for spatially clustered defects. Technometrics, 39(3):241–253, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271129.

Haaland:1993:SAP

[HO93] Perry D. Haaland and M. A. O’Connell. [A Systematic Approach to Planning for a Designed Industrial Experiment]: Discussion. Technometrics, 35(1):17–20, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269283. See [CM93b].

Haaland:1995:IES

[HO95] Perry D. Haaland and Michael A. O’Connell. Inference for effectsaturated fractional factorials. Technometrics, 37(1):82–93, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269155.

Hockman:1994:BRB

[Hoc94] Kymm K. Hockman. Book review: World Class Quality: Using Design of Experiments to Make It Happen by Keki R. Bhote. Technometrics, 36(4):425–426, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269965.

Hoffman:1995:BRB

[Hof95] Lorrie L. Hoffman. Book review: Markov Models and Optimization by M. H. A. Davis. Technometrics, 37(2):236, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269636.

Holms:1998:DEE

[Hol98] Arthur G. Holms. Design of experiments as expansible sequences of orthogonal blocks with crossed-classification block effects. Technometrics, 40(3):244–253, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271180.

Hooper:1990:CCB

[Hoo90] Jeffrey H. Hooper. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3):259–260, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269097. See [HK90].

Horn:1998:BRB

[Hor98] Paul S. Horn. Book review: Introduction to Robust Estimation and Hypothesis Testing by Rand R. Wilcox. Technometrics, 40 (1):77–78, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271402.

Hoerl:1992:DIS

[HP92] Roger W. Hoerl and Andrew C. Palm. Discussion: Integrating SPC and APC. Technometrics, 34(3):268–272, August 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270029.

Hall:1997:AIT

[HP97] Peter Hall and A. R. Padmanabhan. Adaptive inference for the two-sample scale problem. Technometrics, 39(4):412–422, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271505.

Haitovsky:1992:CRM

[HR92] Yoel Haitovsky and Joel Rapp. Conditional resampling for misclassified multinomial data with applications to sampling inspection. Technometrics, 34(4):473–483, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268946. See correction [HR94].

Haitovsky:1994:CCR

[HR94] Yoel Haitovsky and Joel Rapp. Corrections: “Conditional resampling for misclassified multinomial data with applications to sampling inspection” [Technometrics 34 (1992), no. 4, 473–482]. Technometrics, 36(3):334, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269404. See [HR92].

Handcock:1993:BAK

[HS93] Mark S. Handcock and Michael L. Stein. A Bayesian analysis of kriging. Technometrics, 35(4):403–410, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270273.

Holmstrom:1997:MDM

[HS97] Lasse Holmström and Stephan R. Sain. Multivariate discrimination methods for top quark analysis. Technometrics, 39(1):91– 99, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270776.

Hedayat:1999:COA

[HS99] A. S. Hedayat and John Stufken. Compound orthogonal arrays. Technometrics, 41(1):57–61, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270994.

Hadi:1992:CPF

[HSPF92] Ali S. Hadi, Jeffrey S. Simonoff, S. R. Paul, and K. Y. Fung. Comments on Paul and Fung (1991). Technometrics, 34(3):373– 375, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270075. See [PF91].

Hastie:1991:MEM

[HT91] Trevor Hastie and Rob Tibshirani. [The ∏ Method for Estimating Multivariate Functions from Noisy Data]: Discussion. Technometrics, 33(2):155, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269041. See [Bre91a].

Hall:1992:EPP

[HT92] Peter Hall and D. M. Titterington. Edge-preserving and peakpreserving smoothing. Technometrics, 34(4):429–440, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268942.

Hulting:1992:BRB

[Hul92] Fred L. Hulting. Book review: How to Determine Sample Size and Estimate Failure Rate in Life Testing by Eduardo C. Moura. Technometrics, 34(4):488–489, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268950.

Hulting:1997:SIG

[Hul97] Fred L. Hulting. [Statistical Issues in Geometric Feature Inspection Using Coordinate Measuring Machines]: Discussion. Technometrics, 39(1):18–20, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270765. See [DGTZ97a].

Hunter:1990:CCB

[Hun90a] J. Stuart Hunter. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 261, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269098. See [HK90].

Hunter:1990:EWM

[Hun90b] J. Stuart Hunter. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]: Discussion. Technometrics, 32(1):21–22, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269839. See [LS90b].

Huzurbazar:1995:BRB

[Huz95] Aparna V. Huzurbazar. Book review: Concepts in Probability and Stochastic Modeling by J. Higgins; S. Keller-McNulty. Technometrics, 37(4):467, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269753.

Hamada:1990:CLAb

[HW90a] M. Hamada and C. F. J. Wu. [A Critical Look at Accumulation Analysis and Related Methods]: Response. Technometrics, 32(2):159–162, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268859.

Hamada:1990:CLAa

[HW90b] M. Hamada and C. F. J. Wu. A critical look at accumulation analysis and related methods. Technometrics, 32(2):119–130, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268853. See dis-

cussion [Hab90, Hir90, Nai90b, YDB90] and response [HW90a].

Hamada:1991:ACD

[HW91] M. Hamada and C. F. J. Wu. Analysis of censored data from highly fractionated experiments. Technometrics, 33(1):25–38, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269005.

Hwan:1992:BRB

[Hwa92] Marilyn Hwan. Book review: Quality Through Design by N. Logothetis; H. P. Wynn. Technometrics, 34(2):230–231, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269248.

Iglewicz:1995:BRB

[IHMT95] Boris Iglewicz, Richard M. Heiberger, Dirk Moore, and Yun Tan. Book review: Algorithms, Routines, and S Functions, for Robust Statistics by A. Marazzi. Technometrics, 37(3):351–353, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269931.

Inglis:1992:BRB

[Ing92] James Inglis. Book review: Statistics for Research by Shirley Dowdy; Stanley Wearden. Technometrics, 34(3):367–368, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270062.

Iyer:1990:BRB

[Iye90] Hari Iyer. Book review: Model Discrimination for Nonlinear Regression Models by Dale S. Borowiak. Technometrics, 32(4):448– 450, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270126.

Iyengar:1998:BRB

[Iye98] Satish Iyengar. Book review: Multivariate Models and Dependence Concepts by H. Joe. Technometrics, 40(4):353, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270541.

Izenman:1990:BRB

[Ize90] Alan Julian Izenman. Book review: Multivariate Splines by Charles K. Chui. Technometrics, 32(2):235–237, May 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268887.

Izenman:1991:BRB

[Ize91] Alan Julian Izenman. Book review: The Multivariate Normal Distribution by Y. L. Tong. Technometrics, 33(4):492–493, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269441.

Jacroux:1992:NDC

[Jac92] Mike Jacroux. A note on the determination and construction of minimal orthogonal main-effect plans. Technometrics, 34(1):92– 96, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269556.

Jacroux:1993:CMP

[Jac93] Mike Jacroux. On the construction of minimal partially replicated orthogonal main-effect plans. Technometrics, 35(1):32–36, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269287.

Jackson:1995:BRB

[Jac95] J. Edward Jackson. Book review: Statistical Factor Analysis and Related Methods: Theory and Applications by Alexander Basilevsky. Technometrics, 37(2):227–228, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269623.

Jain:1995:BRB

[Jai95] Kamlesh Jain. Book review: Statistical Methods of Quality Assurance by H.-J. Mittag; H. Rinne. Technometrics, 37(2):234, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269633.

James:1996:BRB

[Jam96] David A. James. Book review: Modern Applied Statistics with S-PLUS by W. N. Venables; B. D. Ripley. Technometrics, 38 (1):77–78, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268908.

Janky:1994:BRB

[Jan94] Denis G. Janky. Book review: Applied Engineering Statistics by Robert M. Bethea; R. Russell Rhinehart. Technometrics, 36(2): 221–222, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270241.

Jobe:1992:BCB

[JD92] J. Marcus Jobe and H. T. David. Buehler confidence bounds for a reliability-maintainability measure. Technometrics, 34(2):214– 222, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269237.

Jones:1996:FDR

[JD96] Bradley Jones and William Dumouchel. [Follow-Up Designs to Resolve Confounding in Multifactor Experiments]: Discussion. Technometrics, 38(4):323–326, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271301. See [MSB96a].

Jensen:1990:BRB

[Jen90] Louis Jensen. Book review: Practical Statistics for the Physical Sciences by L. L. Havlicek; R. C. Crain. Technometrics, 32 (3):357–358, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269134.

Jensen:1991:BRB

[Jen91] Karen L. Jensen. Book review: Experimental Design in Biotechnology by Perry D. Haaland. Technometrics, 33(1):106–108, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269013.

Jensen:1992:BRB

[Jen92] Karen L. Jensen. Book review: Practical Application of SPC in the Wire and Cable Industry by Douglas B. Relyea. Technometrics, 34(1):101–102, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269562.

Jensen:1993:BRB

[Jen93] Karen L. Jensen. Book review: Concepts for R & R Studies by Larry B. Barrentine. Technometrics, 35(1):93–94, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269303.

Johnson:1999:EFC

[JFM99] Valen E. Johnson, Mark Fitzgerald, and Harry F. Martz. [Estimating Fatigue Curves with the Random Fatigue-Limit Model]: Discussion. Technometrics, 41(4):294–296, November 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271345. See [PM99].

Jobe:1999:BRB

[Job99] J. Marcus Jobe. Book review: Principles of Experimental Design and Analysis by Alberto Garcia-Diaz; Don T. Phillips. Technometrics, 41(1):77, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271004.

John:1990:TTF

[Joh90] Peter W. M. John. Time trends and factorial experiments. Technometrics, 32(3):275–282, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269104.

Johnson:1991:BRB

[Joh91] Douglas H. Johnson. Book review: Modern Elementary Probability and Statistics by Edward J. Dudewicz; Pinyuen Chen; Baldeo K. Taneja. Technometrics, 33(3):370–371, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268802.

Jolliffe:1990:BRB

[Jol90] Ian T. Jolliffe. Book review: Component and Correspondence Analysis: Dimension Reduction by Functional Approximation by Jan L. A. van Rijckevorsel; Jan de Leeuw. Technometrics, 32 (4):463, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270146.

Jolliffe:1993:BRB

[Jol93] Ian T. Jolliffe. Book review: Introduction to Multiple Time Series Analysis by Helmut Lütkepohl. Technometrics, 35(1):88–89, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269296.

Jolliffe:1996:BRB

[Jol96] Ian Jolliffe. Book review: Multidimensional Scaling by Trevor F. Cox; Michael A. A. Cox. Technometrics, 38(4):403–404, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271315.

Jolliffe:1997:BRB

[Jol97] Ian T. Jolliffe. Book review: Tree Models of Similarity and Association by James E. Corter. Technometrics, 39(2):237, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270927.

Jones:1990:BRB

[Jon90] Peter W. Jones. Book review: Applications of Probability by British Open University Course Team. Technometrics, 32(4):457, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270137.

Jones:1991:BRBb

[Jon91a] Jason Jones. Book review: Introduction to Probability and Its Applications by Richard L. Scheaffer. Technometrics, 33(3):374– 375, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268807.

Jones:1991:BRBa

[Jon91b] Peter W. Jones. Book review: Statistical Methods by The Open University Course Team. Technometrics, 33(1):114, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269021.

Jones:1992:BRB

[Jon92] P. W. Jones. Book review: Probabilistic Reasoning in Expert Systems, Theory and Algorithms by Richard E. Neapolitan. Technometrics, 34(1):99–100, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269559.

Jones:1993:BRB

[Jon93] Jason R. Jones. Book review: Statistical Intervals: A Guide for Practitioners by Gerald J. Hahn; William Q. Meeker. Technometrics, 35(3):323, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269526.

Journel:1994:UCS

[Jou94] André G. Journel. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Discussion. Technometrics, 36(2):149–150, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270224. See [Got94c].

Juang:1991:HMM

[JR91] B. H. Juang and L. R. Rabiner. Hidden Markov models for speech recognition. Technometrics, 33(3):251–272, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268779.

Jones:1999:BCC

[JR99] Geoffrey Jones and David M. Rocke. Bootstrapping in controlled calibration experiments. Technometrics, 41(3):224–233, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270567.

Jin:1999:FPD

[JS99] Jionghua Jin and Jianjun Shi. Feature-preserving data compression of stamping tonnage information using wavelets. Technometrics, 41(4):327–339, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271349.

Jensen:1993:OAP

[JV93] Karen L. Jensen and Stephen B. Vardeman. Optimal adjustment in the presence of deterministic process drift and random adjustment error. Technometrics, 35(4):376–389, November 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270271.

Kafadar:1994:ANR

[Kaf94] Karen Kafadar. An application of nonlinear regression in research and development: A case study from the electronics industry. Technometrics, 36(3):237–248, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269368.

Kafadar:1999:ER

[Kaf99] Karen Kafadar. Editor’s report. Technometrics, 41(4):365–366, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271352.

Kahn:1994:BRB

[Kah94] William D. Kahn. Book review: Measurement Uncertainty: Methods and Applications by Ronald H. Dieck. Technometrics, 36(4): 432–433, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269975.

Kalemkarian:1993:BRB

[Kal93] George S. Kalemkarian. Book review: The Self-Instructional Route to Statistical Process Control and Just-In-Time Manufacturing by I. Louis Bare; Bebe Bare. Technometrics, 35(1):92, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269301.

Kalemkarian:1998:BRB

[Kal98] George S. Kalemkarian. Book review: The Management of Quality and Its Control by Charles S. Tapiero. Technometrics, 40(4): 351–352, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270539.

Kandel:1995:BRB

[Kan95] Abraham Kandel. Book review: Statistical Applications Using Fuzzy Sets by Kenneth G. Manton; Max A. Woodbury; H. Dennis Tolley. Technometrics, 37(2):237, May 1995. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269638.

Katti:1991:BRB

[Kat91] S. K. Katti. Book review: Statistical Methods and Concepts by M. N. Das. Technometrics, 33(1):114, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269022.

Katti:1998:BRB

[Kat98] S. K. Katti. Book review: Operational Subjective Statistical Methods by Frank Lad. Technometrics, 40(2):157–158, May 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270655.

Katti:1994:BRB

[KB94] S. K. Katti and Prabha Betne. Book review: Nonlinear Statistical Models by Andrej Pazman. Technometrics, 36(4):433–434, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269976.

Kalagnanam:1997:EST

[KD97] Jayant R. Kalagnanam and Urmila M. Diwekar. An efficient sampling technique for off-line quality control. Technometrics, 39 (3):308–319, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271135.

Kedem:1996:BRB

[Ked96] Benjamin Kedem. Book review: Stochastic Processes by J. Medhi. Technometrics, 38(1):85–86, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268920.

Kedem:1998:BRB

[Ked98] Benjamin Kedem. Book review: An Introduction to Stochastic Processes by Edward P. C. Kao. Technometrics, 40(1):78, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271403.

Kelley:1997:BRB

[Kel97] D. Lynn Kelley. Book review: Graphing Statistics and Data: Creating Better Charts by Anders Wallgren; Britt Wallgren; Rolf Persson; Ulf Jorner; Jan-Aage Haaland. Technometrics, 39(4): 429–430, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271516.

Keating:1990:THA

[KGK90] Jerome P. Keating, Ronald E. Glaser, and Norma S. Ketchum. Testing hypotheses about the shape parameter of a gamma distribution. Technometrics, 32(1):67–82, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269846.

Khamis:1995:BRB

[Kha95] Imad H. Khamis. Book review: Reliability Engineering by K. K. Aggarwal. Technometrics, 37(2):234–235, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269634.

Khuri:1992:RSM

[Khu92] A. I. Khuri. Response surface models with random block effects. Technometrics, 34(1):26–37, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269549.

Kianifard:1990:BRB

[Kia90] Farid Kianifard. Book review: Applied Linear Regression Models by J. Neter; W. Wasserman; M. H. Kutner. Technometrics, 32(3):352–353, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269127.

Kianifard:1992:BRBa

[Kia92a] Farid Kianifard. Book review: A First Course in the Theory of Linear Statistical Models by R. H. Myers; J. S. Milton. Technometrics, 34(2):228–229, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269243.

Kianifard:1992:BRBb

[Kia92b] Farid Kianifard. Book review: Applied Multivariate Data Analysis, Volume I: Regression and Experimental Design by J. D. Jobson. Technometrics, 34(4):490–491, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268953.

Kianifard:1993:BRBa

[Kia93a] Farid Kianifard. Book review: Analysis of Variance in Experimental Design by Harold R. Lindman. Technometrics, 35(2):222–223, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269670.

Kianifard:1993:BRBb

[Kia93b] Farid Kianifard. Book review: Applied Multivariate Data Analysis, Volume II: Categorical and Multivariate Methods by J. D. Jobson. Technometrics, 35(3):326–327, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269529.

Kianifard:1994:BRB

[Kia94] Farid Kianifard. Book review: Models for Repeated Measurements by J. K. Lindsey. Technometrics, 36(4):426–427, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269966.

Kianifard:1995:BRBa

[Kia95a] Farid Kianifard. Book review: Logistic Regression: A Self-Learning Text by D. G. Kleinbaum. Technometrics, 37(1):116– 117, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269163.

Kianifard:1995:BRBb

[Kia95b] Farid Kianifard. Book review: Logistic Regression with Missing Values in the Covariates by W. Vach. Technometrics, 37 (4):460–461, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269744.

Kiiveri:1992:CVA

[Kii92] H. T. Kiiveri. Canonical variate analysis of high-dimensional spectral data. Technometrics, 34(3):321–331, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270038.

Kim:1991:BRB

[Kim91] Jee Soo Kim. Book review: Parameter Estimation in Reliability and Life Span Models by Clifford Cohen; Betty Jones Whitten. Technometrics, 33(1):117–118, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269028.

King:1991:BRB

[Kin91] Terry King. Book review: Smooth Tests of Goodness of Fit by J. Rayner; D. Best. Technometrics, 33(4):491, November 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269440.

King:1993:BRB

[Kin93] Dennis W. King. Book review: Statistical Quality Design and Control by Richard E. DeVor; Tsong-how Chang; John W. Sutherland. Technometrics, 35(3):331–332, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269537.

King:1997:BRB

[Kin97] Terry King. Book review: A Guide to Chi-Squared Testing by Priscilla E. Greenwood; Mikhail S. Nickulin. Technometrics, 39 (4):431, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271520.

Kirmani:1992:BRB

[Kir92] S. N. U. A. Kirmani. Book review: Stochastic Processes with Applications by R. N. Bhattacharya; E. C. Waymire. Technometrics, 34(1):99, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269558.

Kirmani:1993:BRB

[Kir93] Syed N. U. A. Kirmani. Book review: Performance Modeling of Automated Manufacturing Systems by N. Viswanadham; Y. Narahari. Technometrics, 35(4):456, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270286.

Kvam:1999:EFC

[KK99] Paul H. Kvam and Eric P. Kvam. [Estimating Fatigue Curves with the Random Fatigue-Limit Model]: Discussion. Technometrics, 41(4):292–294, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271344. See [PM99].

Klein:1994:BRB

[Kle94] John Klein. Book review: Statistical Models Based on Counting Process by Per Kragh Andersen; Ørnulf Borgan; Richard D. Gill; Niels Keiding. Technometrics, 36(1):111–112, February 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269205.

Kalbfleisch:1991:MAP

[KLR91] J. D. Kalbfleisch, J. F. Lawless, and J. A. Robinson. Methods for the analysis and prediction of warranty claims. Technometrics, 33(3):273–285, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268780.

Kelly:1994:INE

[KM94] Robert J. Kelly and Thomas Mathew. Improved nonnegative estimation of variance components in some mixed models with unbalanced data. Technometrics, 36(2):171–181, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270229.

Krishnamoorthy:1999:CAM

[KM99] K. Krishnamoorthy and Thomas Mathew. Comparison of approximation methods for computing tolerance factors for a multivariate normal population. Technometrics, 41(3):234–249, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270568.

Kandel:1995:DVR

[KMP95] Abraham Kandel, Alejandro Martins, and Roberto Pacheco. Discussion: On the very real distinction between fuzzy and statistical methods. Technometrics, 37(3):276–281, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269909.

Ko:1995:BRBb

[Ko95a] Daijin Ko. Book review: A Handbook of Statistical Analyses Using S-Plus by Brian S. Everitt. Technometrics, 37(2):232–233, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269630.

Ko:1995:BRBa

[Ko95b] Daijin Ko. Book review: An Introduction to S and S-Plus by Phil Spector. Technometrics, 37(2):232, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269629.

Koehler:1992:BRB

[Koe92] James R. Koehler. Book review: Statistical Inference for Spatial Processes by B. D. Ripley. Technometrics, 34(1):109, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269573.

Kohberger:1990:BRB

[Koh90] Robert Kohberger. Book review: Similarity and Clustering in Chemical Information Systems by Peter Willet. Technometrics, 32(3):359–360, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269137.

Kohberger:1994:BRB

[Koh94] Robert C. Kohberger. Book review: Cluster Analysis by Brian S. Everitt. Technometrics, 36(2):216, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270235.

Koons:1990:BRB

[Koo90] George F. Koons. Book review: Statistical Thinking for Managers by David K. Hildebrand; Lyman Ott. Technometrics, 32 (2):234–235, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268885.

Kotlovker:1992:BRB

[Kot92] Debbi L. Kotlovker. Book review: Statistics by Richard J. Larsen; Morris L. Marx. Technometrics, 34(2):245, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269270.

Katti:1993:BRB

[KR93] S. K. Katti and Larry Ries. Book review: Exploring the Limits of Bootstrap by Raoul LePage; Lynn Billard. Technometrics, 35(2):226–227, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269675.

Kramer:1990:BRB

[Kra90] Tim Kramer. Book review: Statistical Process Control and beyond by Richard R. Clements. Technometrics, 32(1):97, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269853.

Kramer:1992:BRB

[Kra92] Tim Kramer. Book review: SPC for Practitioners: Special Cases and Continuous Processes by Gary Fellers. Technometrics, 34(4): 493–494, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268957.

Kvam:1993:LTV

[KS93] Paul H. Kvam and Francisco J. Samaniego. Life testing in variably scaled environments. Technometrics, 35(3):306–314, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269523.

Kim:1996:NBC

[KSJ96] Choongrak Kim, Barry E. Storer, and Meeseon Jeong. A note on Box–Cox transformation diagnostics. Technometrics, 38(2): 178–180, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270411.

Koch:1990:DSI

[KTTS90] Gary G. Koch, Catherine Tangen, Gail Tudor, and Maura E. Stokes. Discussion: Strategies and issues for the analysis of ordered categorical data from multifactor studies in industry. Technometrics, 32(2):137–149, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268856.

Kundu:1993:EPU

[Kun93] Debasis Kundu. Estimating the parameters of undamped exponential signals. Technometrics, 35(2):215–218, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269666.

Kunert:1997:UFS

[Kun97] Joachim Kunert. On the use of the factor-sparsity assumption to get an estimate of the variance in saturated designs. Technometrics, 39(1):81–90, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270775.

Kushary:1992:BRB

[Kus92] Debashis Kushary. Book review: Nonlinear Estimation by Gavin J. S. Ross. Technometrics, 34(1):108, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269572.

Kushler:1993:BRB

[Kus93] Robert Kushler. Book review: SPC Using Lotus 1–2–3 by Steven M. Zimmerman; Robert N. Zimmerman. Technometrics, 35(4): 457–458, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270288.

Kushary:1998:BRB

[Kus98] Debashis Kushary. Book review: The EM Algorithm and Extensions by Geoffrey J. McLachlan; Thriyambakam Krishnan. Technometrics, 40(3):260, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271189.

Kvam:1996:ETC

[Kva96] Paul H. Kvam. Estimation techniques for common cause failure data with different system sizes. Technometrics, 38(4):382–388, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271309.

Koschat:1991:BRB

[KY91] Martin A. Koschat and Richard A. Young. Book review: Maximum Entropy Methods in Science and Engineering (Vol. 2) by G.

J. Ericson; C. R. Smith. Technometrics, 33(4):486–487, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269432.

LaLonde:1992:BRB

[LaL92] Steven M. LaLonde. Book review: SAS Applications Programming, a Gentle Introduction by Frank C. DiIorio. Technometrics, 34(3):366–367, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270061.

LaLonde:1995:BRB

[LaL95] Steven M. LaLonde. Book review: Chemometrics for Analytical Chemistry Volume 1: PC-Aided Statistical Data Analysis by Milan Meloun; Jiri Militky; Michele Forina. Technometrics, 37(1): 114–115, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269160.

Lalonde:1999:BRB

[Lal99] Steven M. Lalonde. Book review: A First Course in Multivariate Statistics by Bernard Flury. Technometrics, 41(3):262–263, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270571.

Lambert:1990:BRB

[Lam90] Diane Lambert. Book review: Statistical Methods for Quality Improvement by Thomas P. Ryan. Technometrics, 32(3):340, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269111.

Lambert:1992:ZIP

[Lam92] Diane Lambert. Zero-inflated Poisson regression, with an application to defects in manufacturing. Technometrics, 34(1):1–14, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269547.

Lambert:1998:DAV

[Lam98] Diane Lambert. Discussion: Another view. Technometrics, 40 (3):201–203, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271172.

Larsen:1995:BRB

[Lar95] Greg Larsen. Book review: Statistics in Engineering: A Practical Approach by Andrew V. Metcalfe. Technometrics, 37(4):455, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269737.

Laubscher:1994:CUS

[Lau94] Nico F. Laubscher. Comments on the use of spectral control charts. Technometrics, 36(3):335–336, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269405.

Lawrence:1992:BRB

[Law92] Kenneth D. Lawrence. Book review: Modern Statistical Systems and GPSS Simulation by Zaven A. Karian; Edward J. Dudewicz. Technometrics, 34(3):366, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270060.

Lawera:1995:BRB

[Law95] Martin Lawera. Book review: Predictive Inference: An Introduction by Seymour Geisser. Technometrics, 37(1):121, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269169.

Lawson:1996:BRB

[Law96] John Lawson. Book review: Reliability: Probabilistic Models and Statistical Methods by Lawrence M. Leemis. Technometrics, 38 (3):287–288, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270616.

Lawrence:1998:BRB

[Law98] Kenneth D. Lawrence. Book review: Data Analysis, Regression and Forecasting by Arthur Schleifer,; David E. Bell. Technometrics, 40(3):258–259, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271186.

Lamb:1996:TEV

[LBB96] R. H. Lamb, Dennis D. Boos, and Cavell Brownie. Testing for effects on variance in experiments with factorial treatment structure and nested errors. Technometrics, 38(2):170–177, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270410.

Lucas:1990:CCN

[LC90] James M. Lucas and R. Dennis Cook. Comments on Cook and Nachtsheim (1989). Technometrics, 32(3):363–365, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269141.

Le-Cointe:1994:BRB

[LC94] Lillian Le-Cointe. Book review: Basic Quality Improvement by Susan M. Garrity. Technometrics, 36(1):124, February 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269221.

Le-Cointe:1998:BRB

[LC98] Lillian Le-Cointe. Book review: Applied Statistics: A Handbook of Techniques by Lothar Sachs. Technometrics, 40(2):155, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270651.

Lin:1992:PPP

[LD92] Dennis K. J. Lin and Norman R. Draper. Projection properties of Plackett and Burman designs. Technometrics, 34(4):423–428, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268941.

Lee:1990:BRB

[Lee90] Haelie Lee. Book review: Interpreting Data by Alan Anderson. Technometrics, 32(4):452–453, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270130.

Lee:1993:BRB

[Lee93] Kwan R. Lee. Book review: Reliability Methods for Engineers by K. S. Krishnamoorthi. Technometrics, 35(3):333, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269540.

Lee:1994:BRB

[Lee94] Carl Lee. Book review: Beginning Statistics A to Z by William Mendenhall. Technometrics, 36(3):329, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269392.

Lee:1999:BRB

[Lee99] Kwan Lee. Book review: The Analysis of Proximity Data by B. S. Everitt; S. Rabe-hesketh. Technometrics, 41(1):73–74, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270998.

Leigh:1991:BRB

[Lei91] Stefan Leigh. Book review: Probability and Random Processes for Electrical Engineering by Alberto Leon-Garcia. Technometrics, 33(3):372–373, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268805.

Leigh:1992:BRB

[Lei92] Stefan D. Leigh. Book review: U-Statistics Theory and Practice by A. J. Lee. Technometrics, 34(3):364–365, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270058.

Leigh:1993:BRB

[Lei93] Stefan Leigh. Book review: A User’s Guide to Principal Components by J. Edward Jackson. Technometrics, 35(1):84–85, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269292.

Leigh:1997:BRB

[Lei97] Stefan Leigh. Book review: The Pleasures of Probability by Richard Isaac. Technometrics, 39(1):109, February 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270792.

Leigh:1999:BRB

[Lei99] Stefan Leigh. Book review: Statistical Inference Based on the Likelihood by Adelchi Azzalini. Technometrics, 41(2):176–177, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270746.

Lenth:1990:BRB

[Len90] Russell V. Lenth. Book review: Computer Intensive Methods for Testing Hypotheses: An Introduction by Eric W. Noreen. Technometrics, 32(4):453, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270131.

Lenth:1993:BRB

[Len93] Russell V. Lenth. Book review: Meta-Analysis by the Confidence Profile Method: The Statistical Synthesis of Evidence by David M. Eddy; Vic Hasselblad; Ross Shachter. Technometrics, 35(2):237, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269690.

Lenth:1995:BRB

[Len95] Russell V. Lenth. Book review: Computer Intensive Statistical Methods: Validation, Model Selection, and Bootstrap by J. S. Urban Hjorth. Technometrics, 37(4):458–459, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269741.

LePage:1990:BRBa

[LeP90a] Raoul LePage. Book review: Probability: Methods and Measurement by Anthony O’Hagan. Technometrics, 32(1):112, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269871.

LePage:1990:BRBb

[LeP90b] Raoul LePage. Book review: Simulation Methodology for Statisticians, Operations Analysts, and Engineers (Vol. 1) by P. A. W. Lewis; E. J. Orav. Technometrics, 32(3):351–352, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269126.

Lewis:1992:BRB

[Lew92] Richard A. Lewis. Book review: Quality Assurance for the Chemical and Process Industries by American Society for Quality Control. Technometrics, 34(1):102–103, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269563.

Lewis:1993:BRB

[Lew93] Richard A. Lewis. Book review: Optimization and Variation Reduction in Quality by Wayne A. Taylor. Technometrics, 35(1):92– 93, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269302.

Luceno:1999:EDP

[LG99] Alberto Luceño and Francisco J. González. Effects of dynamics on the properties of feedback adjustment schemes with dead band. Technometrics, 41(2):142–152, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270730.

Li:1997:BRB

[LH97] Ta-Hsin Li and Tailen Hsing. Book review: Complex Stochastic Systems and Engineering by D. M. Titterington. Technometrics, 39(3):336, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1271142. See correction [Ano97e].

Li:1997:TSC

[Li97] Ta-Hsin Li. Time series characterization, Poisson integral, and robust divergence measures. Technometrics, 39(4):357–371, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271500.

Ling:1990:BRB

[Lin90] Robert F. Ling. Book review: Principles of Multivariate Analysis by W. Krzanowski. Technometrics, 32(2):225, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268873.

Lindstrom:1992:BRB

[Lin92] Mary J. Lindstrom. Book review: The Analysis of Repeated Measures by M. J. Crowder; D. J. Hand. Technometrics, 34(2):227– 228, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269242.

Lin:1993:ALF

[Lin93a] Dennis K. J. Lin. Another look at first-order saturated designs: the p-efficient designs. Technometrics, 35(3):284–292, August 1993.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269521. See cor-

rection [Lin95a].

Lin:1993:NCS

[Lin93b] Dennis K. J. Lin. A new class of supersaturated designs. Technometrics, 35(1):28–31, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269286. See comments [WKWL95].

Lin:1995:CAL

[Lin95a] Dennis K. J. Lin. Correction: “Another look at first-order saturated designs: the p-efficient designs” [Technometrics 35 (1993), no. 3, 284–292]. Technometrics, 37(4):473, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269765. See [Lin93a].

Lin:1995:GSS

[Lin95b] Dennis K. J. Lin. Generating systematic supersaturated designs. Technometrics, 37(2):213–225, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269622.

Lin:1999:BRB

[Lin99] Yachen Lin. Book review: Introductory Statistics and Random Phenomena by Manfred Denker; Wojbor A. Woyczynski. Technometrics, 41(3):270, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270584.

Little:1991:BRB

[Lit91] Kevin Little. Book review: Presenting Statistics: A Manager’s Guide to the Persuasive Use of Statistics by L. Witzling; R. Greenstreet. Technometrics, 33(4):478, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269425.

Liu:1992:BRBa

[Liu92a] Shin Ta Liu. Book review: Decision-Making under Uncertainty: An Applied Statistics Approach by George K. Chacko. Technometrics, 34(2):243–244, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269267.

Liu:1992:BRBb

[Liu92b] Shin Ta Liu. Book review: Making Hard Decisions, an Introduction to Decision Analysis by Robert T. Clemen. Technometrics, 34(3):365–366, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270059.

Liu:1999:BRB

[Liu99] Shin Ta Liu. Book review: Risk Modeling, Assessment, and Management by Yacov Y. Haimes. Technometrics, 41(4):376–377, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271367.

Liao:1996:COT

[LIV96] C. T. Liao, H. K. Iyer, and D. F. Vecchia. Construction of orthogonal two-level designs of user-specified resolution where N ≠ 2^{k}. Technometrics, 38(4):342–353, November 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271305.

Li:1999:MZI

[LLP^{+}99] Chin-Shang Li, Jye-Chyi Lu, Jinho Park, Kyungmoo Kim, Paul A. Brinkley, and John P. Peterson. Multivariate zero-inflated Poisson models and their applications. Technometrics, 41(1):29–38, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270992.

Lucas:1996:CCE

[LLW^{+}96] James M. Lucas, J. M. Lucas, William J. Welch, Robert J. Buck, Jerome Sacks, Henry P. Wynn, Max D. Morris, and Matthias Schonlau. Comments on computer experiments. Technometrics, 38(2):197–203, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270442. See [WBS^{+}92].

Lu:1993:UDM

[LM93] C. Joseph Lu and William Q. Meeker. Using degradation measures to estimate a time-to-failure distribution. Technometrics, 35(2):161–174, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269661.

Lawless:1995:SSR

[LN95] J. F. Lawless and C. Nadeau. Some simple robust methods for the analysis of recurrent events. Technometrics, 37(2):158–168, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269617.

Liski:1996:PRM

[LN96] Erkki P. Liski and Tapio Nummi. Prediction in repeated-measures models with engineering applications. Technometrics, 38(1):25– 36, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268900.

Loughin:1997:PTE

[LN97] Thomas M. Loughin and William Noble. A permutation test for effects in an unreplicated factorial design. Technometrics, 39(2): 180–190, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270906.

Loh:1992:BRB

[Loh92] Wei-Yin Loh. Book review: Symmetric Multivariate and Related Distributions by Kai-Tai Fang; Samuel Kotz; Kai-Wang Ng. Technometrics, 34(2):235–236, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269255.

Lohr:1994:BRB

[Loh94] Sharon Lohr. Book review: Statistics by David Freeman; Robert Pisani; Roger Purves; Ani Adhikari. Technometrics, 36(1):119– 120, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269215.

Lordo:1992:BRB

[Lor92] Robert A. Lordo. Book review: A Course in Simulation by Sheldon M. Ross. Technometrics, 34(2):236–238, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269257.

Lordo:1993:BRB

[Lor93] Robert A. Lordo. Book review: Simulation: A Statistical Perspective by Jack Kleijnen; Willem van Groenendaal. Technometrics, 35(4):453–454, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270283.

Lordo:1999:BRB

[Lor99] Robert A. Lordo. Book review: Statistical Reasoning and Methods by Richard A. Johnson; Kam-Wah Tsui. Technometrics, 41 (3):269–270, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270583.

Leger:1992:BTA

[LPR92] Christian Léger, Dimitris N. Politis, and Joseph P. Romano. Bootstrap technology and applications. Technometrics, 34(4):378–398, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268938.

Lu:1997:SIT

[LPY97] Jye-Chyi Lu, Jinho Park, and Qing Yang. Statistical inference of a time-to-failure distribution derived from linear degradation data. Technometrics, 39(4):391–400, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271503.

Lucas:1990:CCV

[LS90a] James M. Lucas and Michael S. Saccucci. [CUSUM Charts with Variable Sampling Intervals]: Discussion. Technometrics, 32(4): 387–388, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270116. See [RAA90a].

Lucas:1990:EWMa

[LS90b] James M. Lucas and Michael S. Saccucci. Exponentially weighted moving average control schemes: Properties and enhancements. Technometrics, 32(1):1–12, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1269835. See discussion [Bax90, WM90,

FHT90, Hun90b, MH90] and response [LS90c].

Lucas:1990:EWMb

[LS90c] James M. Lucas and Michael S. Saccucci. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]:

Response. Technometrics, 32(1):27–29, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269841.

Laviolette:1995:PSVb

[LSBW95a] Michael Laviolette, John W. Seaman, Jr., J. Douglas Barrett, and William H. Woodall. [A Probabilistic and Statistical View of Fuzzy Methods]: Reply. Technometrics, 37(3):287–292, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269912. See

[LSBW95b].

Laviolette:1995:PSVa

[LSBW95b] Michael Laviolette, John W. Seaman, Jr., J. Douglas Barrett, and William H. Woodall. A probabilistic and statistical view of fuzzy methods. Technometrics, 37(3):249–261, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269905. See discussion and re-

ply [LSBW95a].

Leon:1993:SAP

[LST93] Ramon V. Leon, Anne C. Shoemaker, and Kwok-Leung Tsui. [A Systematic Approach to Planning for a Designed Industrial Experiment]: Discussion. Technometrics, 35(1):21–24, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269284. See

[CM93b].

Lawless:1996:PPM

[LT96] J. F. Lawless and K. Thiagarajah. A point-process model incorporating renewals and time trends, with application to repairable systems. Technometrics, 38(2):131–138, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270406.

Lucas:1994:RLD

[Luc94] James M. Lucas. [Run-Length Distributions of Special-Cause Control Charts for Correlated Processes]: Discussion. Technometrics, 36(1):17–19, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269192. See [WMP94a].

Luceno:1998:PDF

[Luc98] Alberto Luceño. Performance of discrete feedback adjustment schemes with dead band, under stationary versus nonstationary stochastic disturbance. Technometrics, 40(3):223–233, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271178.

LuValle:1990:BRB

[LuV90] M. J. LuValle. Book review: Generalized Poisson Distributions: Properties and Applications by P. C. Consul. Technometrics, 32 (3):346–347, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269119.

Li:1997:CPA

[LW97] William W. Li and C. F. Jeff Wu. Columnwise-pairwise algorithms with applications to the construction of supersaturated designs. Technometrics, 39(2):171–179, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270905. See comments [LW98].

Li:1998:CLW

[LW98] William W. Li and C. F. Jeff Wu. Comments on Li and Wu (1997). Technometrics, 40(4):362, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270560. See [LW97].

Lowry:1992:MEW

[LWCR92] Cynthia A. Lowry, William H. Woodall, Charles W. Champ, and Steven E. Rigdon. A multivariate exponentially weighted moving average control chart. Technometrics, 34(1):46–53, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269551.

Lwin:1994:AWF

[Lwi94] Thaung Lwin. Analysis of weight frequency distributions using replicated data. Technometrics, 36(1):28–36, February 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269196.

Lynch:1994:BRB

[Lyn94] Richard O. Lynch. Book review: Statistics in Practice by Ernest A. Blaisdell. Technometrics, 36(2):222–223, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270243.

Lynch:1995:BRB

[Lyn95] James Lynch. Book review: Reliability, Maintainability and Supportability: A Probabilistic Approach by J. Knezevic. Technometrics, 37(1):122, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269171.

Lynch:1997:BRBa

[Lyn97a] James Lynch. Book review: Reliability Engineering by E. A. Elsayed. Technometrics, 39(2):226–227, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270911.

Lynch:1997:BRBb

[Lyn97b] Richard Otis Lynch. Book review: Applied Factor Analysis in the Natural Sciences by Richard Reyment; K. G. Jöreskog. Technometrics, 39(4):427, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271512.

Lynch:1999:BRB

[Lyn99] Richard O. Lynch. Book review: Introduction to Design and Analysis of Experiments by George W. Cobb. Technometrics, 41(2): 170, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270737.

Maryak:1992:BRB

[MA92] John L. Maryak and David J. Abeshouse. Book review: Probability, Statistical Optics, and Data Testing: A Problem Solving Approach by B. R. Frieden. Technometrics, 34(3):369–370, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270065.

Mack:1990:BRB

[Mac90a] Gregory Mack. Book review: Chemometrics: A Textbook by D. L. Massart; B. G. M. Vandeginste; S. N. Deming; Y. Michotte; L. Kaufman. Technometrics, 32(1):104–105, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269863.

MacKay:1990:CCB

[Mac90b] Jock MacKay. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 263–264, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269099. See [HK90].

MacGregor:1992:SPM

[Mac92] John F. MacGregor. [Statistical Process Monitoring and Feedback Adjustment]: Discussion. Technometrics, 34(3):273–275, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270030. See

[BK92a].

Madrigal:1996:BRB

[Mad96] J. L. Madrigal. Book review: Stochastic Models: An Algorithmic Approach by Henk C. Tijms. Technometrics, 38(1):84, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268918.

Madrigal:1998:BRB

[Mad98] J. L. Madrigal. Book review: Theory of Sample Surveys by M. E. Thompson. Technometrics, 40(4):356–357, November 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270546.

Magoun:1990:BRB

[Mag90] A. Dale Magoun. Book review: Decision Estimation and Classification by Charles W. Therrien. Technometrics, 32(4):464, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270148.

Magoun:1996:BRB

[Mag96] A. Dale Magoun. Book review: Elementary Statistics Laboratory Manual by John D. Spurrier; Don Edwards; Lori A. Thombs. Technometrics, 38(2):190, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270428.

Mallows:1995:MC

[Mal95] Colin L. Mallows. More comments on C_{P}. Technometrics, 37 (4):362–372, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269729.

Margavio:1994:BRB

[Mar94] Thomas M. Margavio. Book review: Data Quality: Management and Technology by Thomas C. Redman. Technometrics, 36(1): 122–123, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269220.

Marx:1996:IRP

[Mar96] Brian D. Marx. Iteratively reweighted partial least squares estimation for generalized linear regression. Technometrics, 38(4): 374–381, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271308.

Marcucci:1997:BRB

[Mar97] Mark Marcucci. Book review: Applied Multivariate Techniques by Subhash Sharma. Technometrics, 39(1):101, February 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270777.

Maryak:1999:BRB

[Mar99] John Maryak. Book review: Bayesian Heuristic Approach to Discrete and Global Optimization by J. Mockus; W. Eddy; A. Mockus; L. Mockus; G. Reklaitis. Technometrics, 41(4):378–379, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271369.

Mason:1999:BRB

[Mas99] Robert L. Mason. Book review: Multivariate Quality Control: Theory and Applications by Camil Fuchs; Ron S. Kenett. Technometrics, 41(2):168–169, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270734.

Mathew:1996:BRB

[Mat96] Thomas Mathew. Book review: Methods of Multivariate Analysis by Alvin C. Rencher. Technometrics, 38(1):76–77, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268907.

Mathew:1997:BRB

[Mat97] Thomas Mathew. Book review: Multivariate Statistical Analysis by Narayan C. Giri. Technometrics, 39(1):101, February 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270778.

Mathew:1998:BRB

[Mat98] Thomas Mathew. Book review: Mathematical Tools for Applied Multivariate Analysis by J. Douglas Carroll; Paul E. Green; Anil Chaturvedi. Technometrics, 40(3):259, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271187.

Mazu:1993:BRB

[Maz93] Michael J. Mazu. Book review: Regression Analysis by Example by Samprit Chatterjee; Bertram Price. Technometrics, 35 (1):86–87, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269294.

Mazu:1998:BRB

[Maz98] Michael J. Mazu. Book review: Quality by Donna C. S. Summers. Technometrics, 40(3):257–258, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271184.

Mee:1998:SLD

[MB98] Robert W. Mee and Rodney L. Bates. Split-lot designs: Experiments for multistage batch processes. Technometrics, 40(2):127– 140, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270647.

McArthur:1990:BRB

[McA90] Richard D. McArthur. Book review: Trace Residue Analysis: Chemometric Estimations of Sampling, Amount, and Error by D. A. Kurtz. Technometrics, 32(4):461–462, November 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270144.

McComb:1990:BRB

[McC90] Mark Anthony McComb. Book review: Quality Control for Operators and Foremen by K. S. Krishnamoorthi. Technometrics, 32 (4):459, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270141.

McComb:1991:BRB

[McC91] Mark Anthony McComb. Book review: Quality Control by Dale H. Besterfield. Technometrics, 33(4):473–474, November 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269421.

McCool:1994:BRB

[McC94] John I. McCool. Book review: Methods for Experimental Design: Principles and Applications for Physicists and Chemists by Jacques L. Goupy. Technometrics, 36(3):318–319, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269380.

McComb:1998:BRB

[McC98] Mark A. McComb. Book review: Measuring Process Capability by Davis R. Bothe. Technometrics, 40(4):350–351, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270538.

McCool:1999:BRB

[McC99] John I. McCool. Book review: Engineering Reliability by Richard E. Barlow. Technometrics, 41(1):75–76, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271001.

McGuire:1994:BRB

[McG94] Stephen A. McGuire. Book review: Statistical Process Control for Quality Improvement by James R. Evans. Technometrics, 36 (4):420, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269958.

McLaughlin:1995:BRB

[McL95] Gregory C. McLaughlin. Book review: Practical Statistical Process Control: A Tool for Quality Manufacturing by Fred Alsup; Ricky M. Watson. Technometrics, 37(4):461–462, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269746.

Mee:1996:CUC

[ME96] Robert W. Mee and Keith R. Eberhardt. A comparison of uncertainty criteria for calibration. Technometrics, 38(3):221–229, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270605.

Marx:1999:GLR

[ME99] Brian D. Marx and Paul H. C. Eilers. Generalized linear regression on sampled signals and curves: A P -spline approach. Technometrics, 41(1):1–13, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270990.

Mee:1990:IPS

[Mee90a] Robert W. Mee. An improved procedure for screening based on a correlated, normally distributed variable. Technometrics, 32 (3):331–337, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269110.

Mee:1990:STI

[Mee90b] Robert W. Mee. Simultaneous tolerance intervals for normal populations with common variance. Technometrics, 32(1):83–92, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269847.

Meek:1991:BRB

[Mee91a] Gary E. Meek. Book review: Investigations for a Course in Statistics by Martin A. Tanner. Technometrics, 33(4):494, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269444.

Meeker:1991:BRB

[Mee91b] William Q. Meeker, Jr. Book review: Accelerated Testing: Statistical Models, Test Plans, and Data Analyses by Wayne Nelson. Technometrics, 33(2):236–238, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269050.

Meeker:1998:ADT

[MEL98] William Q. Meeker, Luis A. Escobar, and C. Joseph Lu. Accelerated degradation tests: Modeling and analysis. Technometrics, 40(2):89–99, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270643.

Menten:1991:BRB

[Men91] Tom Menten. Book review: Quality Engineering Using Robust Design by Madhav Phadke. Technometrics, 33(2):236, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269049.

Mee:1991:CST

[MER91] Robert W. Mee, Keith R. Eberhardt, and Charles P. Reeve. Calibration and simultaneous tolerance intervals for regression. Technometrics, 33(2):211–219, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269047.

MacGregor:1990:EWM

[MH90] John F. MacGregor and Thomas J. Harris. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]: Discussion. Technometrics, 32(1):23–26, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269840. See

[LS90b].

Mi:1997:BRB

[Mi97] Jie Mi. Book review: Computational Statistics in Climatology by Ilya Polyak. Technometrics, 39(4):428–429, November 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271514.

Mi:1999:BRB

[Mi99] Jie Mi. Book review: Robust Nonparametric Statistical Methods by T. P. Hettmansperger; J. W. McKean. Technometrics, 41(1): 74, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270999.

Milliken:1990:BRB

[Mil90] George A. Milliken. Book review: Nonlinear Regression Analysis and Its Applications by Douglas M. Bates; Donald G. Watts. Technometrics, 32(2):219–220, May 1990. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268866.

Miller:1991:BRB

[Mil91] Bill Miller. Book review: Principles of Quality Control by Jerry Banks. Technometrics, 33(3):358–359, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268790.

Miller:1994:BRB

[Mil94] John E. Miller. Book review: A Basic Approach to Quality Control and SPC by Peter D. Mauch. Technometrics, 36(4):424–425, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269964.

Miller:1997:SPC

[Mil97] Arden Miller. Strip-plot configurations of fractional factorials. Technometrics, 39(2):153–161, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270903.

Milliken:1998:BRB

[Mil98] George A. Milliken. Book review: Statistical Methods: A Geometric Primer by David J. Saville; Graham R. Wood. Technometrics, 40(1):78–80, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271405.

Miller:1999:BRB

[Mil99] Melinda A. Miller. Book review: Statistical Methods for Engineers by G. Geoffrey Vining. Technometrics, 41(3):264–265, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270574.

Minton:1993:BRB

[Min93] Paul D. Minton. Book review: Sample Survey Principles and Methods by Vic Barnett. Technometrics, 35(2):235, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269688.

Morrell:1991:RTN

[MJ91] Christopher H. Morrell and Richard A. Johnson. Random truncation and neutrinos. Technometrics, 33(4):429–440, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269415.

Martz:1996:EBE

[MKA96] Harry F. Martz, Paul H. Kvam, and Lee R. Abramson. Empirical Bayes estimation of the reliability of nuclear-power-plant emergency Diesel generators. Technometrics, 38(1):11–24, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268899.

Meeker:1995:ALT

[ML95] William Q. Meeker and Michael J. LuValle. An accelerated life test model based on reliability kinetics. Technometrics, 37(2):133–146, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269615.

McWilliams:1992:CLV

[MLV92] Thomas P. McWilliams, Thomas J. Lorenzen, and Lonnie C. Vance. Comments on Lorenzen and Vance (1986). Technometrics, 34(2):248–249, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269276.

Meeter:1994:OAL

[MM94] Carol A. Meeter and William Q. Meeker, Jr. Optimum accelerated life tests with a nonconstant scale parameter. Technometrics, 36(1):71–83, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269200.

Merchant:1998:IMC

[MME98] Aparna Merchant, Melinda McCann, and Don Edwards. Improved multiple comparisons with a control in response surface analysis. Technometrics, 40(4):297–303, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270530.

Morris:1993:BDA

[MMY93] Max D. Morris, Toby J. Mitchell, and Donald Ylvisaker. Bayesian design and analysis of computer experiments: Use of derivatives in surface prediction. Technometrics, 35(3):243–255, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269517.

Meyer:1995:CEA

[MN95] Ruth K. Meyer and Christopher J. Nachtsheim. The coordinateexchange algorithm for constructing exact optimal experimental designs. Technometrics, 37(1):60–69, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269153.

Moore:1990:CCB

[Moo90] David S. Moore. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 265–266, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269100. See [HK90].

Moore:1991:BRB

[Moo91] Leslie M. Moore. Book review: Statistical Modelling in GLIM by Murray Aitkin; Dorothy Anderson; Brian Francis; John Hinde. Technometrics, 33(1):109–110, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269016.

Moore:1996:BRB

[Moo96] Leslie M. Moore. Book review: The Basic Practice of Statistics by David S. Moore. Technometrics, 38(4):404–405, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271317.

Moore:1999:BRB

[Moo99] Roger H. Moore. Book review: Regression Graphics: Ideas for Studying Regressions through Graphics by R. Dennis Cook. Technometrics, 41(4):368–369, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271354.

Morris:1991:FSP

[Mor91] Max D. Morris. Factorial sampling plans for preliminary computational experiments. Technometrics, 33(2):161–174, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269043.

Morganstein:1992:BRB

[Mor92] David Morganstein. Book review: An Introduction to Team Approach to Problem Solving by Louis Jones; Ronald McBride. Technometrics, 34(1):103–104, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269564.

Morgan:1993:BRB

[Mor93a] Carolyn Morgan. Book review: Truncated and Censored Samples, Theory and Applications by A. Clifford Cohen. Technometrics, 35(2):220–221, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269667.

Morton:1993:BRB

[Mor93b] Michael J. Morton. Book review: Statistical Methods in Applied Chemistry by Jurand Czerminski; Andrzej Iwasiewicz; Zbigniew Paszek; Andrzej Sikorski. Technometrics, 35(2):228–229, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269678.

Morris:1997:BRB

[Mor97a] Max D. Morris. Book review: Cyclic and Computer Generated Designs by J. A. John; E. R. Williams. Technometrics, 39(1): 104, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270783.

Morris:1997:ER

[Mor97b] Max D. Morris. Editor’s report. Technometrics, 39(1):1–2, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270763.

Morgan:1998:OCL

[Mor98a] J. P. Morgan. Orthogonal collections of Latin squares. Technometrics, 40(4):327–333, November 1998. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270533.

Morris:1998:EA

[Mor98b] Max Morris. Editorial announcement. Technometrics, 40(1):1– 2, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271387.

Morris:1998:ER

[Mor98c] Max D. Morris. Editor’s report. Technometrics, 40(4):271–272, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270527.

Morgan:1999:BRB

[Mor99a] Walter T. Morgan. Book review: SPC for Short Run Manufacturing by Leonard A. Doty. Technometrics, 41(2):169, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270735.

Morganstein:1999:BRB

[Mor99b] David R. Morganstein. Book review: Practical Sampling Techniques by R. K. Som. Technometrics, 41(2):175, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270743.

Martz:1999:ETS

[MPR99] Harry F. Martz, Robert L. Parker, and Dale M. Rasmuson. Estimation of trends in the scram rate at nuclear power plants. Technometrics, 41(4):352–364, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271351.

Marr:1991:NTS

[MQ91] Ray L. Marr and C. P. Quesenberry. A NU test for serial correlation of residuals from one or more regression regimes. Technometrics, 33(4):441–457, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269416.

McDonald:1990:DAR

[MS90] Gary C. McDonald and William J. Studden. Design aspects of regression-based ratio estimation. Technometrics, 32(4):417–424, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270121.

Morris:1995:DAI

[MS95] M. D. Morris and A. D. Solomon. Design and analysis for an inverse problem arising from an advection-dispersion process. Technometrics, 37(3):293–302, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269913.

Meyer:1996:FDRa

[MSB96a] R. Daniel Meyer, David M. Steinberg, and George Box. Follow-up designs to resolve confounding in multifactor experiments. Technometrics, 38(4):303–313, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1271297. See discussion [Fed96, JD96, Tob96] and reply [MSB96b].

Meyer:1996:FDRb

[MSB96b] R. Daniel Meyer, David S. Steinberg, and George Box. [Follow-Up Designs to Resolve Confounding in Multifactor Experiments]: Reply. Technometrics, 38(4):327–332, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271303. See [MSB96a].

Mudholkar:1995:EWF

[MSF95] Govind S. Mudholkar, Deo Kumar Srivastava, and Marshall Freimer. The exponentiated Weibull family: A reanalysis of the bus-motor-failure data. Technometrics, 37(4):436–445, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269735.

McKean:1994:RHB

[MSH94] Joseph W. McKean, Simon J. Sheather, and Thomas P. Hettmansperger. Robust and high-breakdown fits of polynomial models. Technometrics, 36(4):409–415, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269956.

McShane:1991:PLO

[MT91] Lisa M. McShane and Bruce W. Turnbull. Probability limits on outgoing quality for continuous sampling plans. Technometrics, 33(4):393–404, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269412.

Moskowitz:1992:BVA

[MT92] Herbert Moskowitz and Kwei Tang. Bayesian variables acceptancesampling plans: Quadratic loss function and step-loss function. Technometrics, 34(3):340–347, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1270040.

Meyer:1998:KCS

[MTV98] R. Daniel Meyer, Donald C. Trost, and Kimberly E. Vukovinsky. [Key Challenges for Statisticians in Business and Industry]: Discussion. Technometrics, 40(3):203–205, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271173. See [HH98a].

Mulekar:1997:BRB

[Mul97] Madhuri S. Mulekar. Book review: Workshop Statistics, Discovery with Data by Allan J. Rossman. Technometrics, 39(2):235–236, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270925.

Martz:1990:BRA

[MW90] H. F. Martz and R. A. Waller. Bayesian reliability analysis of complex series/parallel systems of binomial subsystems and components. Technometrics, 32(4):407–416, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270120.

Moorhead:1998:CDP

[MW98] P. R. Moorhead and C. F. J. Wu. Cost-driven parameter design. Technometrics, 40(2):111–119, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270645.

Myers:1991:BRB

[Mye91] Donald E. Myers. Book review: Statistics for Geoscientists by D. Marsal. Technometrics, 33(1):118, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269029.

Myers:1996:BRB

[Mye96] Donald E. Myers. Book review: Multivariate Geostatistics by Hans Wackernagel. Technometrics, 38(4):401–402, November 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271311.

Nair:1990:ER

[Nai90a] Vijay Nair. Editor’s report. Technometrics, 32(4):369–370, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270113.

Nair:1990:CLA

[Nai90b] Vijayan N. Nair. [A Critical Look at Accumulation Analysis and Related Methods]: Discussion. Technometrics, 32(2):151–152, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268857. See

[HW90b].

Nair:1991:ER

[Nai91a] Vijay Nair. Editor’s report. Technometrics, 33(4):378–380, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269410.

Nair:1991:NE

[Nai91b] Vijay Nair. New Editor. Technometrics, 33(4):377, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269409.

Nair:1992:ER

[Nai92] Vijay Nair. Editor’s report. Technometrics, 34(4):377, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268937.

Naik:1994:BRB

[Nai94] Dayanand N. Naik. Book review: Applied Multivariate Data Analysis by Brian S. Everitt; Graham Dunn. Technometrics, 36(2):214, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270233.

Naik:1996:BRB

[Nai96] Dayanand N. Naik. Book review: Random Factors in ANOVA by Sally Jackson; Dale E. Brashers. Technometrics, 38(2):187, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270421.

Nair:1992:TPD

[NAM^{+}92] Vijayan N. Nair, Bovas Abraham, Jock MacKay, John A. Nelder, George Box, Madhav S. Phadke, Raghu N. Kacker, Jerome Sacks, William J. Welch, Thomas J. Lorenzen, Anne C. Shoemaker, Kwok L. Tsui, James M. Lucas, Shin Taguchi, Raymond H. Myers, G. Geoffrey Vining, and C. F. Jeff Wu. Taguchi’s parameter design: A panel discussion. Technometrics, 34(2):127–161, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269231. Edited

by Vijayan N. Nair.

Nandagopalan:1994:BRB

[Nan94] S. Nandagopalan. Book review: Engineering Applications of Correlation and Spectral Analysis by Julius S. Bendat; Allan G. Piersol. Technometrics, 36(2):220–221, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270240.

Narayanan:1993:BRB

[Nar93] A. Narayanan. Book review: Probability and Statistics in Engineering and Management Science by William W. Hines; Douglas C. Montgomery. Technometrics, 35(2):238–239, May 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269692.

Nayak:1998:BRB

[Nay98] Tapan K. Nayak. Book review: Statistical Significance: Rationale, Validity and Utility by Siu L. Chow. Technometrics, 40(2): 161–162, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270661.

Nelson:1993:AUA

[Nel93] Peter R. Nelson. Additional uses for the analysis of means and extended tables of critical values. Technometrics, 35(1):61–71, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269290.

Nelson:1995:CLR

[Nel95] Wayne Nelson. Confidence limits for recurrence data: Applied to cost or number of product repairs. Technometrics, 37(2):147–157, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269616.

Nelson:1998:BRB

[Nel98] Wayne Nelson. Book review: Statistical Methods for Reliability Data by William Q. Meeker; Luis A. Escobar. Technometrics, 40(3):255–256, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271181.

Nemeth:1992:BRB

[Nem92] Margaret A. Nemeth. Book review: Basic Statistics: Tools for Continuous Improvement by Mark J. Kiemele; Stephen R. Schmidt. Technometrics, 34(2):231–232, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269249.

Nemeth:1995:BRB

[Nem95] Margaret A. Nemeth. Book review: Multivariate Pattern Recognition in Chemometrics, Illustrated by Case Studies (Data Handling in Science and Technology, Vol. 9) by Richard G. Brereton. Technometrics, 37(1):116, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269162.

Nemeth:1997:BRB

[Nem97] Margaret A. Nemeth. Book review: Multivariate Statistical Methods: A Primer by Bryan F. J. Manly. Technometrics, 39(4): 426–427, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271511.

Nemeth:1998:KCS

[Nem98] Margaret A. Nemeth. [Key Challenges for Statisticians in Business and Industry]: Discussion. Technometrics, 40(3):206–207, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271174. See

[HH98a].

Neubauer:1999:BRB

[Neu99] Dean V. Neubauer. Book review: Success through Quality by Timothy J. Clark. Technometrics, 41(4):375, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271365.

Neyer:1994:OBS

[Ney94] Barry T. Neyer. A D-optimality-based sensitivity test. Technometrics, 36(1):61–70, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269199.

Nguyen:1994:COB

[Ngu94] Nam-Ky Nguyen. Construction of optimal block designs by computer. Technometrics, 36(3):300–307, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269374.

Nguyen:1996:AAC

[Ngu96a] Nam-Ky Nguyen. An algorithmic approach to constructing supersaturated designs. Technometrics, 38(1):69–73, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268904.

Nguyen:1996:NCN

[Ngu96b] Nam-Ky Nguyen. A note on the construction of near-orthogonal arrays with mixed levels and economic run size. Technometrics, 38(3):279–283, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270612.

Niknian:1993:BRB

[Nik93] Minoo Niknian. Book review: Randomization and Monte Carlo Methods in Biology by Bryan F. J. Manly. Technometrics, 35 (1):96–97, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269308.

Niknian:1995:BRB

[Nik95] Minoo Niknian. Book review: Permutation Tests: A Practical Guide to Resampling Methods for Testing Hypotheses by Philip Good. Technometrics, 37(3):341–342, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269919.

Niple:1992:BRB

[Nip92] Edward R. Niple. Book review: Linear Models for Multivariate, Time Series, and Spatial Data by Ronald Christensen. Technometrics, 34(3):354–355, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270043.

Nelder:1998:JMM

[NLB^{+}98] John A. Nelder, Youngjo Lee, Bo Bergman, Anders Hynén, A. Freek Huele, and Jan Engel. Joint modeling of mean and dispersion. Technometrics, 40(2):168–175, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270676.

Nomikos:1995:MSC

[NM95] Paul Nomikos and John F. MacGregor. Multivariate SPC charts for monitoring batch processes. Technometrics, 37(1):41–59, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269152.

Nychka:1996:NNA

[NO96] Douglas Nychka and Michael O’Connell. [Neural Networks in Applied Statistics]: Discussion. Technometrics, 38(3):218–220, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270603. See

[Ste96a].

Notz:1992:BRBa

[Not92a] William I. Notz. Book review: Continuous Bivariate Distributions Emphasising Applications by T. P. Hutchinson; C. D. Lai. Technometrics, 34(2):236, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269256.

Notz:1992:BRBc

[Not92b] William I. Notz. Book review: Exploratory and Multivariate Data Analysis by Michel Jambu. Technometrics, 34(4):489–490, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268952.

Notz:1992:BRBb

[Not92c] William I. Notz. Book review: The Engineering Statistician’s Guide to Continuous Bivariate Distributions by T. P. Hutchinson; C. D. Lai. Technometrics, 34(4):489, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268951.

Nayebpour:1993:ATL

[NW93] Mohamad R. Nayebpour and William H. Woodall. An analysis of Taguchi’s on-line quality-monitoring procedures for attributes. Technometrics, 35(1):53–60, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269289.

ODonnell:1997:BRB

[O’D97] Robert O’Donnell. Book review: Statistical Aspects of Quality Control by Cyrus Derman; Sheldon M. Ross. Technometrics, 39 (4):424, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271506.

ODonnell:1999:BRB

[O’D99] Robert O’Donnell. Book review: Manufacturing Process Design and Optimization by Robert F. Rhyder. Technometrics, 41(1):78– 79, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271006.

Odeh:1990:TSP

[Ode90] Robert E. Odeh. Two-sided prediction intervals to contain at least k out of m future observations from a normal distribution. Technometrics, 32(2):203–216, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268864.

Oehlert:1994:IOC

[Oeh94] Gary W. Oehlert. Isolating one-cell interactions. Technometrics, 36(4):403–408, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269955.

Olin:1997:BRB

[Oli97] Bryan Olin. Book review: Principles of Statistical Data Handling by Fred Davidson. Technometrics, 39(2):235, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270924.

Olin:1998:BRB

[Oli98] Bryan Olin. Book review: Design of Experiments in Quality Engineering by J. T. Luftig; V. S. Jordan. Technometrics, 40(4): 350, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270537.

Olin:1996:ASMa

[OM96a] Bryan D. Olin and William Q. Meeker. Applications of statistical methods to nondestructive evaluation. Technometrics, 38(2): 95–112, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270399. See discussion [Ber96a, CR96, Spe96b, Swe96, Swe96, Tuc96] and reply [OM96b].

Olin:1996:ASMb

[OM96b] Bryan D. Olin and William Q. Meeker. [Applications of Statistical Methods to Nondestructive Evaluation]: Reply. Technometrics, 38(2):128–130, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270405. See [OM96a].

Oman:1991:RCM

[Oma91] Samuel D. Oman. Random calibration with many measurements: An application of Stein estimation. Technometrics, 33(2):187–195, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269045.

Ord:1996:BRB

[Ord96] Keith Ord. Book review: Multivariate Tests for Time Series Models by Jeff B. Cromwell; Michael J. Hannan; Walter C. Labys; Michel Terraza. Technometrics, 38(2):187, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270422.

Osborne:1997:BRB

[Osb97] Deborah M. Osborne. Book review: Modern Regression Methods by Thomas P. Ryan. Technometrics, 39(4):426, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271509.

Otto:1990:BRB

[Ott90] Gordon H. Otto. Book review: Statistical Models for Optimizing Mineral Exploration by T. K. Wignall; J. de Geoffroy. Technometrics, 32(4):460–461, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270143.

Offodile:1993:ACS

[OUH93] O. Felix Offodile, Kingsley O. Ugwu, and Leslie Hayduk. Analysis of the causal structures linking process variables to robot repeatability and accuracy. Technometrics, 35(4):421–435, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270275.

Owen:1991:BRB

[Owe91] Christopher R. Owen. Book review: SPC Tools for Operators by John T. Burr. Technometrics, 33(1):116, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269025.

Owen:1996:BRB

[Owe96] Willis L. Owen. Book review: Modelling Frequency and Count Data by J. K. Lindsey. Technometrics, 38(3):290–291, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270619.

Pankratz:1992:BRB

[Pan92] Peter C. Pankratz. Book review: Understanding Industrial Designed Experiments by Steven R. Schmidt; Robert G. Launsby. Technometrics, 34(4):492, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268955.

Pankratz:1996:BRB

[Pan96] Peter C. Pankratz. Book review: Using Statistics in Industry: Quality Improvement Through Total Process Control by Gordon Betteley; Neville Metterick; Edward Sweeny; David Wilson. Technometrics, 38(1):81, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268913.

Pan:1999:IUL

[Pan99] Guohua Pan. The impact of unidentified location effects on dispersion-effects identification from unreplicated factorial designs. Technometrics, 41(4):313–326, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271348.

Papp:1995:BRB

[Pap95] Zoltan Papp. Book review: Multi-Vari Chart and Analysis: A Pre-Experimentation Technique by Mario Perez-Wilson. Technometrics, 37(1):121, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269170.

Papp:1996:BRB

[Pap96] Zoltan Papp. Book review: Probabilistic Reliability Engineering by Boris V. Gnedenko; Igor A. Ushakov; James A. Falk. Technometrics, 38(2):184–185, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270416.

Parmigiani:1992:BRB

[Par92] Giovanni Parmigiani. Book review: Large Deviation Techniques in Decision, Simulation and Estimation by James A. Bucklew. Technometrics, 34(1):120–121, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269587.

Patch:1991:BRB

[Pat91] Steven Patch. Book review: Designing Experiments and Analyzing Data by S. Maxwell; H. Delaney. Technometrics, 33(3):375– 376, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268809.

Paul:1992:CBR

[PBR92] S. R. Paul, D. J. Best, and J. C. W. Rayner. Comment on Best and Rayner (1987). Technometrics, 34(2):249–250, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269277.

Palm:1995:STC

[PD95] Andrew C. Palm and Robert L. DeAmico. [Shewhart-Type Charts in Nonstandard Situations]: Discussion. Technometrics, 37(1):26– 29, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269148. See [RD95a].

Phillips:1997:SIG

[PE97] Steven D. Phillips and Keith R. Eberhardt. [Statistical Issues in Geometric Feature Inspection Using Coordinate Measuring Machines]: Discussion. Technometrics, 39(1):22–24, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270767. See

[DGTZ97a].

Pederson:1998:BRB

[Ped98] Shane P. Pederson. Book review: Hidden Markov and Other Models for Discrete-Valued Time Series by Iain L. MacDonald; Walter Zucchini. Technometrics, 40(3):263, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271194.

Pena:1991:MID

[Peñ91] Daniel Peña. Measuring influence in dynamic regression models. Technometrics, 33(1):93–101, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269010.

Percival:1992:BRB

[Per92] Donald B. Percival. Book review: Statistical Spectral Analysis: A Nonprobabilistic Theory by William A. Gardner. Technometrics, 34(1):109–110, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269574.

Perry:1997:BRB

[Per97] Robert L. Perry. Book review: The Power of Statistical Thinking by M. G. Leitnaker; R. D. Sanders; C. Hild. Technometrics, 39(1):108, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270790.

Percival:1998:BRB

[Per98] Donald B. Percival. Book review: The Spectral Analysis of Time Series by Lambert H. Koopmans. Technometrics, 40(4): 354, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270543.

Peterson:1993:GAR

[Pet93] John J. Peterson. A general approach to ridge analysis with confidence intervals. Technometrics, 35(2):204–214, May 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269665.

Peterson:1999:BRB

[Pet99] John J. Peterson. Book review: Regression Analysis of Count Data by A. Colin Cameron; Pravin K. Trivedi. Technometrics, 41 (4):371, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271358.

Paul:1991:GES

[PF91] S. R. Paul and Karen Y. Fung. A generalized extreme Studentized residual multiple-outlier-detection procedure in linear regression. Technometrics, 33(3):339–348, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268785. See comments [HSPF92].

Pfahler:1992:BRB

[Pfa92] Lori Pfahler. Book review: The Deming Dimension by H. R. Neave. Technometrics, 34(1):104, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269565.

Pfahler:1999:BRB

[Pfa99] Lori B. Pfahler. Book review: Statistical Graphics for Visualizing Multivariate Data by William G. Jacoby. Technometrics, 41(3):262, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270570.

Pham-Gia:1995:CWJ

[PGT95] T. Pham-Gia and N. Turkkan. Comment on Weerahandi and Johnson (1992). Technometrics, 37(1):130–131, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269186. See

[WJ92].

Pebesma:1999:LHS

[PH99] Edzer J. Pebesma and Gerard B. M. Heuvelink. Latin hypercube sampling of Gaussian random fields. Technometrics, 41(4): 303–312, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271347.

Piepel:1990:BRBb

[Pie90a] Gregory F. Piepel. Book review: Empirical Model Building by James R. Thompson. Technometrics, 32(3):353, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269128.

Piepel:1990:BRBa

[Pie90b] Gregory F. Piepel. Book review: Quality Control, Robust Design, and the Taguchi Method by Khosrow Dehnad. Technometrics, 32(2):221, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268868. See corrections [Ano90g].

Pigeon:1991:BRB

[Pig91] Joseph G. Pigeon. Book review: Statistical Modeling: Applications in Contemporary Issues by William S. Mallios. Technometrics, 33(3):375, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268808.

Pitts:1992:BRB

[Pit92] Karen A. Pitts. Book review: A Guide to Graphical Problem-Solving Processes by J. W. Moran; R. P. Talbot; R. M. Benson. Technometrics, 34(1):106, February 1992. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269568.

Pitts:1996:BRB

[Pit96] Karen A. Pitts. Book review: Fundamentals Concepts in the Design of Experiments by Charles R. Hicks. Technometrics, 38 (1):82–84, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268916.

Pederson:1990:EMD

[PJ90] Shane P. Pederson and Mark E. Johnson. Estimating model discrepancy. Technometrics, 32(3):305–314, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269107.

Placzkowski:1990:BRB

[Pla90] Gene Placzkowski. Book review: In-Process Quality Control for Manufacturing by William E. Barkman. Technometrics, 32(4): 458, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270139.

Placzkowski:1996:BRB

[Pla96] Gene Placzkowski. Book review: Reliability, Maintainability and Availability Assessment by Mitchell O. Locks. Technometrics, 38 (3):288–289, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270617.

Pascual:1999:EFC

[PM99] Francis G. Pascual and William Q. Meeker. Estimating fatigue curves with the random fatigue-limit model. Technometrics, 41 (4):277–290, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1271342. See discussion [Ann99, KK99, JFM99] and response [Ano99f].

Polsky:1993:BRB

[Pol93] Boris R. Polsky. Book review: Statistical Methods for Testing, Development, and Manufacturing by Forrest W. Breyfogle. Technometrics, 35(4):455–456, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270285.

Porter:1993:BRB

[Por93] Donald R. Porter. Book review: Introduction to Linear Regression Analysis by Douglas C. Montgomery; Elizabeth A. Peck. Technometrics, 35(2):224–225, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269673.

Portier:1997:BRBb

[Por97a] Kenneth Portier. Book review: Statistical Modelling in Hydrology by Robin T. Clarke. Technometrics, 39(4):427–428, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271513.

Portier:1997:BRBa

[Por97b] Kenneth M. Portier. Book review: Adaptive Sampling by Steven K. Thompson; George A. F. Seber. Technometrics, 39(2):236, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270926.

Poston:1998:BRB

[Pos98] Wendy L. Poston. Book review: Discrete Multivariate Distributions by Norman L. Johnson; Samuel Kotz; N. Balakrishnan. Technometrics, 40(2):160–161, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270659.

Potter:1993:BRB

[Pot93] Randall W. Potter. Book review: Confidence Intervals on Variance Components by Richard K. Burdick; Franklin A. Graybill. Technometrics, 35(3):328, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269531.

Price:1993:BRB

[Pri93] Barbara Price. Book review: A First Course in Factor Analysis by Andrew L. Comrey; Howard B. Lee. Technometrics, 35 (4):453, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270282.

Price:1998:BRB

[Pri98] Barbara Price. Book review: Robust Planning and Analysis of Experiments by Christine H. Muller. Technometrics, 40(3):258, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271185.

Propst:1990:BRB

[Pro90] Annabeth Propst. Book review: Quantitative Methods for Quality and Productivity Improvement by Marilyn K. Hart; Robert F. Hart. Technometrics, 32(4):458–459, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270140.

Propst:1992:BRBb

[Pro92a] Annabeth Propst. Book reviews: How to Apply Response Surface Methodology by John A. Cornell; How to Use Regression Analysis in Quality Control by Douglas C. Crocker. Technometrics, 34(3):359, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270049.

Propst:1992:BRBa

[Pro92b] Annabeth L. Propst. Book review: Designing for Quality, an Introduction to the Best of Taguchi and Western Methods of Statistical Experimental Design by Robert H. Lochnar; Joseph E. Matar. Technometrics, 34(1):100, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269560.

Propst:1993:BRBa

[Pro93a] Annabeth L. Propst. Book review: Statistical Problem Solving by Hans J. Bajaria; Richard P. Copp. Technometrics, 35(2):234, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269686.

Propst:1993:BRBb

[Pro93b] Annabeth L. Propst. Book review: The Deming Vision: SPC/TQM for Administrators by Gary Fellers. Technometrics, 35(4):456–457, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270287.

Propst:1994:BRB

[Pro94] Annabeth L. Propst. Book review: SPC Tools for Everyone by John T. Burr. Technometrics, 36(3):322–323, August 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269384.

Propst:1996:BRB

[Pro96] Annabeth L. Propst. Book review: The New Economics: For Industry, Government, Education by W. Edwards Deming. Technometrics, 38(3):294–295, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270625.

Qiu:1995:BRB

[Qiu95] Chunfu Qiu. Book review: Nonparametric Regression and Generalized Linear Models: A Roughness Penalty Approach by P. J. Green; B. W. Silverman. Technometrics, 37(3):342–343, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269920.

Qiu:1998:LPJ

[QY98] Peihua Qiu and Brian Yandell. A local polynomial jump-detection algorithm in nonparametric regression. Technometrics, 40(2):141– 152, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270648.

Reynolds:1990:CCVa

[RAA90a] Marion R. Reynolds, Jr., Raid W. Amin, and Jesse C. Arnold. CUSUM charts with variable sampling intervals. Technometrics, 32(4):371–384, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1270114. See discussion [Hea90, LS90a, Woo90] and response [RAA90b].

Reynolds:1990:CCVb

[RAA90b] Marion R. Reynolds, Jr., Raid W. Amin, and Jesse C. Arnold. [CUSUM Charts with Variable Sampling Intervals]: Response. Technometrics, 32(4):393–396, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270118.

Rabinowitz:1992:BRB

[Rab92] Larry Rabinowitz. Book review: A Primer in Probability by Kathleen Subrahmaniam. Technometrics, 34(3):369, August 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270064.

Raghavachari:1991:BRB

[Rag91] Madabhushi Raghavachari. Book review: Applied Probability and Stochastic Processes in Engineering and Physical Sciences by Michael K. Ochi. Technometrics, 33(4):491, November 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269439.

Raghavachari:1993:BRB

[Rag93] M. Raghavachari. Book review: Turbulence and Stochastic Processes: Kolmogorov’s Ideas 50 Years on by J. C. R. Hunt; O. M. Phillips; D. Williams. Technometrics, 35(3):330, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269534.

Raghavachari:1995:BRB

[Rag95] Madabhushi Raghavachari. Book review: Monotone Structure in Discrete-Event Systems by Paul Glasserman; David D. Yao. Technometrics, 37(4):467–468, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269754.

Raghavachari:1998:NOC

[Rag98] M. Raghavachari. A note on optimal component test plans for series system reliability with exponential failure times. Technometrics, 40(4):345–347, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270535.

Ramsey:1991:BRB

[Ram91] Philip Ramsey. Book review: Process Quality Control: Troubleshooting and Interpretation of Data by Ellis R. Ott; Edward G. Schilling. Technometrics, 33(3):359–360, August 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268791.

Ramsay:1993:BRB

[Ram93] Philip H. Ramsay. Book review: Multivariate Density Estimation, Theory, Practice, and Visualization by David W. Scott. Technometrics, 35(4):451–452, November 1993. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270280.

Ramsey:1994:BRB

[Ram94] Philip H. Ramsey. Book review: Data Fitting in the Chemical Sciences by Peter Gans. Technometrics, 36(4):430–431, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269973.

Ramsey:1996:BRB

[Ram96] Philip H. Ramsey. Book review: Statistical Methods in the Atmospheric Sciences by Daniel S. Wilks. Technometrics, 38(4):402, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271312.

Ramirez:1999:BRB

[Ram99] José G. Ramírez. Book review: Understanding Design of Experiments: A Primer for Technologists by R. J. del Vecchio. Technometrics, 41(3):265, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270575.

Rao:1998:BRB

[Rao98] M. B. Rao. Book review: Survival Analysis Techniques for Censored and Truncated Data by John P. Klein; Melvin L. Moeschberger. Technometrics, 40(2):159–160, May 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270658.

Rayens:1993:BRB

[Ray93] William S. Rayens. Book review: Discriminant Analysis and Statistical Pattern Recognition by Geoffrey J. McLachlan. Technometrics, 35(3):324–326, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269528.

Razzaghi:1991:BRB

[Raz91] Mehdi Razzaghi. Book review: Mixture Models by Geoffrey J. McLachan; Kaye E. Basford. Technometrics, 33(3):365–366, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268796.

Roes:1995:STCa

[RD95a] Kit C. B. Roes and Ronald J. M. M. Does. Shewhart-type charts in nonstandard situations. Technometrics, 37(1):15–24, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269146. See dis-

cussion [BC95, PD95, HD95, SWG95] and reply [RD95b].

Roes:1995:STCb

[RD95b] Kit C. B. Roes and Ronald J. M. M. Does. [Shewhart-Type Charts in Nonstandard Situations]: Reply. Technometrics, 37 (1):36–40, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269151. See [RD95a].

Read:1993:BRB

[Rea93] Sherry L. Read. Book review: Managing for World-Class Quality: A Primer for Executives and Managers by E. S. Shecter. Technometrics, 35(3):335–336, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269543.

Ruberg:1992:MCP

[RH92] Stephen J. Ruberg and Jason C. Hsu. Multiple comparison procedures for pooling batches in stability studies. Technometrics, 34(4):465–472, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268945.

Radson:1995:AFE

[RH95] Darrell Radson and Gary D. Herrin. Augmenting a factorial experiment when one factor is an uncontrollable random variable: A case study. Technometrics, 37(1):70–81, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269154.

Riedel:1994:BRBb

[Rie94a] Kurt S. Riedel. Book review: Detection of Abrupt Changes: Theory and Application by Michèle Basseville; Igor V. Nikiforov. Technometrics, 36(3):326–327, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269388.

Riedel:1994:BRBa

[Rie94b] Kurt S. Riedel. Book review: Signal Processing and Its Applications (Handbook of Statistics 10) by N. K. Bose; C. R. Rao. Technometrics, 36(3):325–326, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269387.

Riedel:1997:BRB

[Rie97] Kurt S. Riedel. Book review: Numerical Bayesian Methods Applied to Signal Processing by Joseph J. K. O. Ruanaidh; William J. Fitzgerald. Technometrics, 39(1):106, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270786.

Rigdon:1992:BRBb

[Rig92a] Steven E. Rigdon. Book review: Mathematics in Chemistry: An Introduction to Modern Methods by Harry G. Hecht. Technometrics, 34(2):239–240, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269260.

Rigdon:1992:BRBa

[Rig92b] Steven E. Rigdon. Book review: Reliability Statistics by Robert A. Dovich. Technometrics, 34(1):118, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269584.

Rigdon:1993:BRB

[Rig93] Steven E. Rigdon. Book review: Introduction to Reliability Analysis: Probability Models and Statistical Methods by Shelemyahu Zacks. Technometrics, 35(2):221, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269668.

Rigdon:1999:BRB

[Rig99] Steven E. Rigdon. Book review: Model-Oriented Design of Experiments by Valerii V. Federov; Peter Hackl. Technometrics, 41 (1):77–78, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271005.

Rocke:1997:ODE

[RJ97] David M. Rocke and Geoffrey Jones. Optimal design for ELISA and other forms of immunoassay. Technometrics, 39(2):162–170, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270904.

Rocke:1995:TCM

[RL95] David M. Rocke and Stefan Lorenzato. A two-component model for measurement error in analytical chemistry. Technometrics, 37(2):176–184, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269619.

Rajgopal:1995:DCT

[RM95] Jayant Rajgopal and Mainak Mazumdar. Designing component test plans for series system reliability via mathematical programming. Technometrics, 37(2):195–212, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269621.

Rieck:1991:LLM

[RN91] James R. Rieck and Jerry R. Nedelman. A log-linear model for the Birnbaum–Saunders distribution. Technometrics, 33(1):51– 60, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269007.

Robinson:1990:BRB

[Rob90] Jeffrey A. Robinson. Book review: Probability and Statistics for Engineers and Scientists by Ronald E. Walpole; Raymond H. Myers. Technometrics, 32(3):348–349, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269122.

Robinson:1998:EDD

[Rob98] Jeff Robinson. [Events Defined by Duration and Severity, with an Application to Network Reliability]: Discussion. Technometrics, 40(3):190, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271168. See [BCL98a].

Rocke:1993:BTF

[Roc93] David M. Rocke. On the beta transformation family. Technometrics, 35(1):72–81, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269291.

Roecker:1991:PEE

[Roe91] Ellen B. Roecker. Prediction error and its estimation for subsetselected models. Technometrics, 33(4):459–468, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269417.

Rohatgi:1991:BRB

[Roh91] Vijay K. Rohatgi. Book review: Robustness of Statistical Tests by Takeaki Kariya; Bimal K. Sinha. Technometrics, 33(2):246, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269062.

Rollins:1993:BRB

[Rol93] Derrick K. Rollins. Book review: Fundamentals of Industrial Quality Control by Lawrence Aft. Technometrics, 35(2):232–233, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269683.

Rosenbaum:1996:SUC

[Ros96] Paul R. Rosenbaum. Some useful compound dispersion experiments in quality design. Technometrics, 38(4):354–364, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271306.

Rosenbaum:1999:BCD

[Ros99] Paul R. Rosenbaum. Blocking in compound dispersion experiments. Technometrics, 41(2):125–134, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270728.

Rousseeuw:1995:DFC

[Rou95] Peter J. Rousseeuw. Discussion: Fuzzy clustering at the intersection. Technometrics, 37(3):283–286, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269911.

Rozum:1993:BRB

[Roz93] Michael A. Rozum. Book review: Evaluation and Control of Measurements by John Mandel. Technometrics, 35(2):223, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269671.

Rukhin:1997:BRBa

[Ruk97a] Andrew L. Rukhin. Book review: Bayes and Empirical Bayes Methods for Data Analysis by Bradley P. Carlin; Thomas A. Louis. Technometrics, 39(3):337, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271144.

Rukhin:1997:BRBb

[Ruk97b] Andrew L. Rukhin. Book review: Tools for Statistical Inference by Martin A. Tanner. Technometrics, 39(3):341, August 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271152.

Rukhin:1998:BRB

[Ruk98] Andrew L. Rukhin. Book review: Bayesian Methods for Finite Population Sampling by Malay Ghosh; Glen Meeden. Technometrics, 40(3):262–263, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271193.

Runger:1990:BRB

[Run90] George C. Runger. Book review: Statistical Process Control in Automated Manufacturing by J. B. Keats; N. F. Hubele. Technometrics, 32(3):341–342, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269113.

Russell:1994:BRB

[Rus94] Edmund L. Russell III. Book review: Statistics in the Environmental and Earth Sciences by A. T. Walden; P. Guttorp. Technometrics, 36(4):431–432, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269974.

Rutherford:1993:BRB

[Rut93] James A. Rutherford. Book review: The Experimenter’s Companion by R. B. Clements. Technometrics, 35(1):94, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269304.

Rutherford:1994:UCS

[Rut94] Brian M. Rutherford. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Discussion. Technometrics, 36(2):152–156, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270226. See [Got94c].

Rousseeuw:1999:FAM

[RvD99] Peter J. Rousseeuw and Katrien van Driessen. A fast algorithm for the minimum covariance determinant estimator. Technometrics, 41(3):212–223, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270566.

Ruppert:1997:LPV

[RWHH97] David Ruppert, M. P. Wand, Ulla Holst, and Ola Hössjer. Local polynomial variance-function estimation. Technometrics, 39 (3):262–273, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271131.

Saccucci:1993:BRB

[Sac93] Michael S. Saccucci. Book review: Engineering Quality and Experimental Design by D. M. Groves; T. P. Davis. Technometrics, 35(4):454–455, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270284.

Sain:1996:BRB

[Sai96] Stephan R. Sain. Book review: The Nature of Statistical Learning Theory by V. N. Vapnik. Technometrics, 38(4):409, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271324.

Samaniego:1990:BRB

[Sam90] Francisco J. Samaniego. Book review: Accelerated Life Testing and Experts’ Opinions in Reliability by C. A. Clarotti; D. V. Lindley. Technometrics, 32(2):233–234, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268884.

Samaniego:1999:BRBb

[Sam99a] Francisco J. Samaniego. Book review: Reliability, Yield and Stress Burn-In: A Unified Approach for Microelectronics Systems Manufacturing and Software Development by W. Kuo; W. T. K. Chien; T. Kim. Technometrics, 41(4):371–372, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271359.

Samaniego:1999:BRBa

[Sam99b] Francisco J. Samaniego. Book review: The Practice of Bayesian Analysis by S. French; J. Q. Smith. Technometrics, 41(1):73, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270997.

Sandry:1992:BRB

[San92] Thomas D. Sandry. Book review: Nonlinear Estimation by Gavin J. S. Ross. Technometrics, 34(3):356–357, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270045.

Sande:1998:BRB

[San98] Innis G. Sande. Book review: Improving Your Measurement of Customer Satisfaction by Terry G. Vavra. Technometrics, 40(3): 264, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271195.

Sandry:1999:BRB

[San99] Thomas D. Sandry. Book review: Practical Guide to Experimental Design by Normand L. Frigon; David Mathews. Technometrics, 41(2):170–171, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270738.

Sarle:1990:BRB

[Sar90] Warren S. Sarle. Book review: Algorithms for Clustering Data by Anil K. Jain; Richard C. Dubes. Technometrics, 32(2):227–229, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268876.

Sartori:1993:BRB

[Sar93] David E. Sartori. Book review: SPC Using Lotus 1–2–3 by Steven M. Zimmerman; Robert N. Zimmerman. Technometrics, 35(3): 336, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269544.

Sartori:1994:BRBb

[Sar94a] Dave Sartori. Book review: Elementary Statistics: A Step by Step Approach by Allan G. Bluman. Technometrics, 36(4):427– 428, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269968.

Sartori:1994:BRBa

[Sar94b] Dave Sartori. Book review: Statistical Quality Control with Microcomputer Applications by Larry E. Shirland. Technometrics, 36(2):224–225, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270245.

Sauter:1992:BRB

[Sau92] Roger M. Sauter. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 34(2): 232–233, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269251.

Sauter:1994:BRB

[Sau94] Roger M. Sauter. Book review: An Introduction to Statistics with Data Analysis by Shelley Rasmussen. Technometrics, 36(1):118– 119, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269214.

Sauter:1997:BRB

[Sau97] Roger M. Sauter. Book review: Fundamentals of Applied Statistics and Surveys by David B. Orr. Technometrics, 39(1):106– 107, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270787.

Sauter:1998:BRB

[Sau98] Roger M. Sauter. Book review: Introduction to the Statistical Analysis of Categorical Data by Erling B. Andersen. Technometrics, 40(3):261–262, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271191.

Shumway:1998:CSD

[SBD98] R. H. Shumway, D. R. Baumgardt, and Z. A. Der. A cepstral F statistic for detecting delay-fired seismic signals. Technometrics, 40(2):100–110, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1270644. See correction [Ano00].

Smith:1993:BDL

[SC93] Wendell F. Smith, Jr. and John A. Cornell. Biplot displays for looking at multiple response data in mixture experiments. Technometrics, 35(4):337–350, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270268.

Sitter:1997:FRM

[SCF97] Randy R. Sitter, Jiahua Chen, and Moshe Feder. Fractional resolution and minimum aberration in blocked 2^{n−k} designs. Technometrics, 39(4):382–390, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271502.

Schaefer:1991:BRB

[Sch91a] R. L. Schaefer. Book review: Concise Statistics by M. Godfrey; E. Roebuck; A. Sherlock. Technometrics, 33(1):112–113, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269019.

Schiess:1991:BRB

[Sch91b] James R. Schiess. Book review: From Data to Model by Jan C. Willems. Technometrics, 33(4):489, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269436.

Schmoyer:1991:NAT

[Sch91c] Richard L. Schmoyer. Nonparametric analyses for two-level singlestress accelerated life tests. Technometrics, 33(2):175–186, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269044.

Schwarzmann:1991:CBL

[Sch91d] Bernd Schwarzmann. A connection between local-influence analysis and residual diagnostics. Technometrics, 33(1):103–104, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269011.

Schaefer:1992:BRB

[Sch92a] Robert L. Schaefer. Book review: Subset Selection in Regression by A. J. Miller. Technometrics, 34(2):229, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269244.

Schmoyer:1992:AVP

[Sch92b] Richard L. Schmoyer. Asymptotically valid prediction intervals for linear models. Technometrics, 34(4):399–408, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268939.

Schneider:1992:BRB

[Sch92c] Helmut Schneider. Book review: Reliability Engineering Handbook (Vol. 1) by Dimitri Kececioglu. Technometrics, 34(3):361–362, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270054.

Schuenemeyer:1992:BRB

[Sch92d] John H. Schuenemeyer. Book review: Generalized Linear Models by P. McCullagh; J. A. Nelder. Technometrics, 34(2):224, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269238.

Schneider:1993:BRB

[Sch93a] Helmut Schneider. Book review: Reliability Engineering Handbook, Volume 2 by Dimitri Kececioglu. Technometrics, 35(1):94– 95, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269305.

Schuenemeyer:1993:BRBb

[Sch93b] John H. Schuenemeyer. Book review: Fundamentals of Exploratory Analysis of Variance by David C. Hoaglin, Frederick Mosteller, and John W. Tukey. Technometrics, 35(1):89, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269297.

Schuenemeyer:1993:BRBa

[Sch93c] John H. Schuenemeyer. Book review: Fundamentals of Exploratory Analysis of Variance by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 35(1):89, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269297.

Schneider:1994:BRB

[Sch94a] Helmut Schneider. Book review: Reliability Life Testing Handbook by Dimitri Kececioglu. Technometrics, 36(1):111, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269204.

Schuenemeyer:1994:BRB

[Sch94b] John H. Schuenemeyer. Book review: Introducing Quantitative Geography: Measurement, Methods and Generalised Linear Models by Larry O’Brien. Technometrics, 36(1):114–115, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269208.

Schiferl:1995:BRB

[Sch95] Elizabeth A. Schiferl. Book review: Efficiency in Research, Development and Production: The Statistical Design and Analysis of Chemical Experiments by Leslie Davies. Technometrics, 37 (4):457–458, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269740.

Schaefer:1996:BRB

[Sch96a] Robert L. Schaefer. Book review: An Introduction to Computational Statistics: Regression Analysis by Robert I. Jennrich. Technometrics, 38(2):185–186, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270419.

Schneider:1996:BRB

[Sch96b] Helmut Schneider. Book review: Failure Mode and Effect Analysis: FMEA from Theory to Execution by D. H. Stamatis. Technometrics, 38(1):80, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268911.

Schuenemeyer:1996:BRB

[Sch96c] John H. Schuenemeyer. Book review: Geographical Data Analysis by Nigel Walford. Technometrics, 38(2):188, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270424.

Schmee:1998:EDD

[Sch98a] Josef Schmee. [Events Defined by Duration and Severity, with an Application to Network Reliability]: Discussion. Technometrics, 40(3):191–192, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271169. See [BCL98a].

Schmid:1998:CZ

[Sch98b] Wolfgang Schmid. Comments on Zhang (1998). Technometrics, 40(4):362, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270559. See [Zha98].

Scott:1990:BRB

[Sco90] David W. Scott. Book review: The New S Language by R. A. Becker; J. M. Chambers; A. R. Wilks. Technometrics, 32(1): 103–104, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269862.

Scott:1991:BRB

[Sco91] Del Scott. Book review: Computations for the Analysis of Designed Experiments by Richard Heiberger. Technometrics, 33(4): 471–472, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269419.

Scott:1992:BRB

[Sco92] Douglas M. Scott. Book review: Envisioning Information by Edward R. Tufte. Technometrics, 34(1):106–107, February 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269569.

Schall:1990:IVL

[SD90] Robert Schall and Timothy T. Dunne. Influential variables in linear regression. Technometrics, 32(3):323–330, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269109.

Snee:1995:CDP

[SDP^{+}95] Ronald D. Snee, N. R. Draper, P. Prescott, S. M. Lewis, A. M. Dean, P. W. M. John, and M. G. Tuck. Comments on Draper, Prescott, Lewis, Dean, John, and Tuck (1993): Understanding formulation systems. Technometrics, 37(1):131–132, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269187. See

[DPL^{+}93].

Sa:1993:MCC

[SE93] Ping Sa and Don Edwards. Multiple comparisons with a control in response surface methodology. Technometrics, 35(4):436–445, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270276.

Seaman:1990:BRB

[Sea90] John W. Seaman, Jr. Book review: Introduction to the Theory of Coverage Processes by Peter Hall. Technometrics, 32(2):237–238, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268888.

Seaman:1991:BRB

[Sea91] John W. Seaman, Jr. Book review: Measurement Evaluation by Andor Boros. Technometrics, 33(4):474–476, November 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269423.

Seaman:1993:BRB

[Sea93] John W. Seaman, Jr. Book review: Sample Size Methodology by M. M. Desu; D. Raghavarao. Technometrics, 35(1):95–96, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269307.

Seaman:1994:BRB

[Sea94] John W. Seaman, Jr. Book review: Measurement and Calibration for Quality Assurance by Alan S. Morris. Technometrics, 36(2):225–226, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270247.

Sengupta:1994:BRBb

[Sen94a] Sailes K. Sengupta. Book review: Handbook for Digital Signal Processing by Sanjit K. Mitra; James F. Kaiser. Technometrics, 36 (4):430, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269972.

Sengupta:1994:BRBa

[Sen94b] Sailes K. Sengupta. Book review: Introduction to Applied Statistical Signal Analysis by Richard Shiavi. Technometrics, 36(1): 113–114, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269207.

Sengupta:1995:BRB

[Sen95] Sailes K. Sengupta. Book review: Fundamentals of Statistical Signal Processing: Estimation Theory by Steven M. Kay. Technometrics, 37(4):465–466, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269750.

Sengupta:1997:BRB

[Sen97] Sailes K. Sengupta. Book review: Introduction to Random Processes in Engineering by A. V. Balakrishnan. Technometrics, 39 (2):230, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270916.

Sen:1998:ECR

[Sen98] Ananda Sen. Estimation of current reliability in a Duane-based reliability growth model. Technometrics, 40(4):334–344, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270534. See

correction [Sen99b].

Sen:1999:BRB

[Sen99a] Ananda Sen. Book review: The Theory of Dispersion Models by Bent Jørgensen. Technometrics, 41(2):177–178, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270747.

Sen:1999:CEC

[Sen99b] Ananda Sen. Correction: “Estimation of current reliability in a Duane-based reliability growth model” [Technometrics 40 (1998), no. 4, 334–344; Sen:1998:ECR (99h:62126)]. Technometrics, 41 (4):385, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271384. See [Sen98].

Stephenson:1999:BRB

[SG99] Hal W. Stephenson and Ned M. Gibbons. Book review: Business Process Improvement Toolbox by Bjørn Andersen. Technometrics, 41(4):374–375, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271364.

Steiner:1996:GDS

[SGW96] Stefan H. Steiner, P. Lee Geyer, and George O. Wesolowsky. Grouped data-sequential probability ratio tests and cumulative sum control charts. Technometrics, 38(3):230–237, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270606.

Shanmugam:1992:BRB

[Sha92] Ramalingam Shanmugam. Book review: Assurance Technologies: Principles and Practices by Dev G. Raheja. Technometrics, 34 (2):233–234, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269252.

Shanmugam:1993:BRB

[Sha93] Ramalingam Shanmugam. Book review: Statistical Methods in Reliability Theory and Practice by Brian D. Bunday. Technometrics, 35(1):95, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269306.

Shanmugam:1995:BRB

[Sha95] Ramalingam Shanmugam. Book review: Multivariate AnalysisPart 1: Distributions, Ordination and Inference by W. J. Krzanowski; F. H. C. Marriott. Technometrics, 37(3):351, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269930.

Shanmugam:1997:BRBc

[Sha97a] Ramalingam Shanmugam. Book review: Distributions in the Physical and Engineering Sciences by Alexander I. Saichev; Wojbor A. Woyczynski. Technometrics, 39(4):431, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271519.

Shanmugam:1997:BRBb

[Sha97b] Ramalingam Shanmugam. Book review: Introduction to Time Series and Forecasting by Peter J. Brockwell; Richard A. Davis. Technometrics, 39(4):426, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271510.

Shanmugam:1997:BRBa

[Sha97c] Ramalingam Shanmugam. Book review: Multivariate Analysis,Part 2: Classification, Covariance Structures, and Repeated Measurements by W. J. Krzanowski; F. H. C. Marriott. Technometrics, 39(1):101, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270779.

Shanmugam:1998:BRB

[Sha98] Ram Shanmugam. Book review: Parametric Statistical Inference by J. K. Lindsey. Technometrics, 40(2):161, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270660.

Shea:1990:BRB

[She90a] Gerald Shea. Book review: Quality Control Systems by James R. Taylor. Technometrics, 32(2):221–222, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268869.

Sheesley:1990:BRB

[She90b] John H. Sheesley. Book review: Quality Engineering in Production Systems by G. Taguchi; E. Elsayed; T. Hsiang. Technometrics, 32 (4):457–458, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270138.

Shoemaker:1990:BRB

[Sho90] Lewis H. Shoemaker. Book review: Order Restricted Statistical Inference by Tim Robertson; F. T. Wright; R. L. Dykstra. Technometrics, 32(1):111–112, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269870.

Simonoff:1988:DOC

[Sim88] Jeffrey S. Simonoff. Detecting outlying cells in two-way contingency tables via backwards-stepping. Technometrics, 30(3):339– 345, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1270088. See correction [Ano90e].

Simpson:1991:BRB

[Sim91] Douglas G. Simpson. Book review: The Statistical Analysis of Discrete Data by Thomas J. Santner; Diane E. Duffy. Technometrics, 33(2):242, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269057.

Simonson:1997:BRB

[Sim97] Katherine Hansen Simonson. Book review: Pattern Classification: A Unified View of Statistical and Neural Approaches by Jurgen Schürmann. Technometrics, 39(1):102, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270780.

Singh:1983:LRT

[Sin83] N. Singh. The likelihood ratio test for the equality of location parameters of k(≥ 2) exponential populations based on Type II censored samples. Technometrics, 25(2):193–195, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268553. See correction [Sin92].

Singh:1990:BRBa

[Sin90a] Ashok K. Singh. Book review: Properties of Estimators for the Gamma Distribution by K. O. Bowman; L. R. Shenton. Technometrics, 32(1):102–103, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269861.

Singh:1990:BRBb

[Sin90b] Ashok K. Singh. Book review: Quality Control and Reliability (Handbook of Statistics, Vol. 7) by P. R. Krishnaiah; C. R. Rao. Technometrics, 32(3):344, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269116.

Singh:1992:CLR

[Sin92] N. Singh. Correction to: “The likelihood ratio test for the equality of location parameters of k (≥ 2) exponential populations based on Type II censored samples” [Technometrics 25 (1983), no. 2, 193– 195; MR0702171 (84g:62031)]. Technometrics, 34(4):502, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268969. See

[Sin83].

Singh:1997:BRB

[Sin97] A. K. Singh. Book review: The Exponential Distribution: Theory, Methods and Applications by N. Balakrishnan; A. P. Basu. Technometrics, 39(3):341, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271151.

Singh:1998:BRB

[Sin98] Anita Singh. Book review: Astrostatistics by G. J. Babu; E. D. Feigelson. Technometrics, 40(1):76–77, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271400.

Smiley:1991:BRB

[Smi91] Richard A. Smiley. Book review: Beating Your Competition Through Quality by D. B. Owen. Technometrics, 33(1):116–117, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269026.

Smith:1993:BRB

[Smi93] Patricia L. Smith. Book review: Chemometrics: Applications of Mathematics and Statistics to Laboratory Systems by Richard G. Brereton. Technometrics, 35(3):330–331, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269535.

Spall:1995:BRB

[SN95] James C. Spall and J. Barry Nelson. Book review: Numerical Methods for Stochastic Processes by Nicolas Bouleau; Dominique Lepingle. Technometrics, 37(2):236–237, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269637.

Snee:1990:CPN

[Sne90] Ronald D. Snee. Commentary: A partnership is needed. Technometrics, 32(3):267–269, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269101. See [HK90].

Snee:1998:DIV

[Sne98] Ronald D. Snee. Discussion: Increasing the value of the statistics profession. Technometrics, 40(3):207–210, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271175.

Soyer:1995:BRB

[Soy95] Refik Soyer. Book review: Ensuring Software Reliability by Ann M. Neufelder. Technometrics, 37(1):122–123, February 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269172.

Sweeting:1995:ANA

[SP95] Trevor J. Sweeting and Michael J. Phillips. An application of nonstandard asymptotics to the analysis of repairable-systems data. Technometrics, 37(4):428–435, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269734.

Spagon:1999:BRB

[Spa99] Patrick D. Spagon. Book review: Statistical Quality Assurance Methods for Engineers by Stephen B. Vardeman; J. Marcus Jobe. Technometrics, 41(4):372–373, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271360.

Spencer:1991:BRB

[Spe91] Floyd W. Spencer. Book review: Statistical Methods in Accelerated Life Testing by Reinhard Viertl. Technometrics, 33(3):360– 362, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268792.

Speck:1996:BRB

[Spe96a] Daniel Speck. Book review: Quality Engineering Bible by Thomas Pyzdek. Technometrics, 38(4):407–408, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271321.

Spencer:1996:ASM

[Spe96b] Floyd W. Spencer. [Applications of Statistical Methods to Nondestructive Evaluation]: Discussion. Technometrics, 38(2):122–124, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270402. See

[OM96a].

Sposito:1990:BRB

[Spo90] Vince A. Sposito. Book review: Nonlinear L_{p}-Norm Estimation by René Gonin; Arthur H. Money. Technometrics, 32(4): 450–451, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270127.

Spring:1991:BRB

[Spr91] Raymond V. Spring. Book review: Exploring Operations Research and Statistics in the Micro Lab by Erwin Kalvelagen; Henk C. Tijms. Technometrics, 33(1):120, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269033.

Spurrier:1992:ODC

[Spu92] John D. Spurrier. Optimal designs for comparing the variances of several treatments with that of a standard treatment. Technometrics, 34(3):332–339, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270039.

Spurrier:1993:CSC

[Spu93] John D. Spurrier. Comparison of simultaneous confidence bands for quadratic regression over a finite interval. Technometrics, 35 (3):315–320, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269524.

Srinivasan:1992:BRB

[Sri92] Syamala Srinivasan. Book review: Design and Analysis of Experiments by Douglas C. Montgomery. Technometrics, 34(2):230, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269246.

Srinivasan:1993:BRB

[Sri93] Cidambi Srinivasan. Book review: Combining Information: Statistical Issues and Opportunities for Research by National Research Council. Technometrics, 35(4):462–463, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270296.

Srivastava:1989:SDA

[SS89] V. K. Srivastava and N. Singh. Small-disturbance asymptotic theory for linear-calibration estimators. Technometrics, 31(3):373– 378, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

3556147. See correction [Ano90f].

Song:1997:VCT

[SS97] Ruiguang Song and Stanley A. Shulman. Variance components in the two-way nested model with incomplete nesting information. Technometrics, 39(1):71–80, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270774.

Schneider:1990:APT

[ST90a] Helmut Schneider and Kwei Tang. Adaptive procedures for the two-stage group-testing problem based on prior distributions and costs. Technometrics, 32(4):397–405, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270119.

Spurrier:1990:CCD

[ST90b] John D. Spurrier and L. A. Thombs. Control charts for detecting cyclical behavior. Technometrics, 32(2):163–171, May 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268860.

StLaurent:1994:BRU

[St.94] Roy T. St. Laurent. Book review: Understanding Regression Assumptions by William D. Berry. Technometrics, 36(3):321, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269382.

Saltelli:1999:QMI

[STC99] A. Saltelli, S. Tarantola, and K. P.-S. Chan. A quantitative model-independent method for global sensitivity analysis of model output. Technometrics, 41(1):39–56, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270993.

Stein:1987:LSP

[Ste87] Michael Stein. Large sample properties of simulations using Latin hypercube sampling. Technometrics, 29(2):143–151, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269769. See cor-

rection [Ste90b].

Stein:1990:BRB

[Ste90a] Michael Stein. Book review: Estimating and Choosing by G. Matheron; A. M. Hasofer. Technometrics, 32(3):358–359, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269136.

Stein:1990:CLS

[Ste90b] Michael Stein. Correction: “Large sample properties of simulations using Latin hypercube sampling” [Technometrics 29 (1987), no. 2, 143–151; MR0887702 (88e:62055)]. Technometrics, 32(3): 367, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269142. See [Ste87].

Stephenson:1990:BRB

[Ste90c] W. R. Stephenson. Book review: Distribution Free Tests by H. R. Neave; P. L. Worthington. Technometrics, 32(4):454–455, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270133.

Stein:1991:BRB

[Ste91] Michael Stein. Book review: The Analysis of Directional Time Series: Applications to Wind Speed and Direction by J. Breckling. Technometrics, 33(4):485–486, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269431.

Stern:1996:NNAa

[Ste96a] Hal S. Stern. Neural networks in applied statistics. Technometrics, 38(3):205–214, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1270601. See discussion [dVU96, NO96] and reply [Ste96b].

Stern:1996:NNAb

[Ste96b] Hal S. Stern. [Neural Networks in Applied Statistics]: Reply. Technometrics, 38(3):220, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270604. See [Ste96a].

Stephenson:1998:BRB

[Ste98] W. Robert Stephenson. Book review: Applied Statistics: A First Course in Inference by F. A. Graybill; H. K. Iyer; R. K. Burdick. Technometrics, 40(4):352–353, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270540.

Stewart:1999:BRB

[Ste99] Paul Stewart. Book review: Linear Mixed Models in Practice: A SAS-Oriented Approach (Lecture Notes in Statistics 126) by Geert Verbeke; Geert Molenberghs. Technometrics, 41(1):79–80, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271007.

Stine:1997:BRB

[Sti97] Robert A. Stine. Book review: Analysis of Observed Chaotic Data by Henry D. I. Abarbanel. Technometrics, 39(3):334–335, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271140.

Stone:1992:BRB

[Sto92] Ricardo R. Stone. Book review: Machine/Process Capability Study: A Five Stage Methodology for Optimizing Manufacturing Processes by Mario Perez-Wilson. Technometrics, 34(2):232, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269250.

Stokes:1996:BRB

[Sto96] S. Lynne Stokes. Book review: How to Perform Continuous Sampling by Kenneth S. Stephens. Technometrics, 38(1):82, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268914.

Stromberg:1996:BRB

[Str96] Arnold J. Stromberg. Book review: Number-Theoretic Methods in Statistics by K.-T. Fang; Y. Wang. Technometrics, 38(2):189–190, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270426.

Seaver:1999:IIS

[STR99] Bill Seaver, Konstantinos Triantis, and Chip Reeves. The identification of influential subsets in regression using a fuzzy clustering strategy. Technometrics, 41(4):340–351, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271350.

Stufken:1994:BRB

[Stu94] John Stufken. Book review: Taguchi Methods: A Hands-On Approach by Glen Stuart Peace. Technometrics, 36(1):121–122, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269218.

Shoemaker:1991:EEM

[STW91] Anne C. Shoemaker, Kwok-Leung Tsui, and C. F. Jeff Wu. Economical experimentation methods for robust design. Technometrics, 33(4):415–427, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269414.

Steiner:1995:ECB

[SW95] S. H. Steiner and G. O. Wesolowsky. Estimating the correlation between destructively measured variables using proof-loading.

Technometrics, 37(1):94–101, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269156.

Sun:1997:OBS

[SWC97] Don X. Sun, C. F. Jeff Wu, and Youyi Chen. Optimal blocking schemes for 2^{n} and 2^{n−p} designs. Technometrics, 39(3):298– 307, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271134.

Sweeting:1996:ASM

[Swe96] Trevor J. Sweeting. [Applications of Statistical Methods to Nondestructive Evaluation]: Discussion. Technometrics, 38(2):124–126, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270403. See

[OM96a].

Sullivan:1995:STC

[SWG95] Joe H. Sullivan, William H. Woodall, and Martha M. Gardner. [Shewhart-Type Charts in Nonstandard Situations]: Discussion. Technometrics, 37(1):31–35, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269150. See [RD95a].

Syski:1999:BRB

[Sys99] R. Syski. Book review: Fundamentals of Queueing Theory by D. Gross; C. M. Harris. Technometrics, 41(1):76–77, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271003.

Taam:1991:BRB

[Taa91] Winson Taam. Book review: Nonlinear System Analysis and Identification from Random Data by Julius S. Bendat. Technometrics, 33(4):482–483, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269429.

Taam:1999:BRB

[Taa99] Winson Taam. Book review: Introduction to Probability and Statistics for Scientists and Engineers by Walter A. Rosenkrantz. Technometrics, 41(4):373, November 1999. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271362.

Tamhane:1996:BRB

[Tam96] Ajit C. Tamhane. Book review: Design and Analysis of Experiments for Statistical Selection, Screening and Multiple Comparisons by Robert E. Bechhofer; Thomas J. Santner; David M. Goldsman. Technometrics, 38(3):289–290, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270618.

Tang:1993:BRB

[Tan93] Kwei Tang. Book review: Guide to Acceptance Sampling by Wayne A. Taylor. Technometrics, 35(4):458, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270289.

Tanaydin:1998:BRB

[Tan98] Suat Tanaydin. Book review: Robust Design and Analysis for Quality Engineering by Sung H. Park. Technometrics, 40(4):349, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270536.

Tatum:1996:CCD

[Tat96] Lawrence G. Tatum. Control charts for the detection of a periodic component. Technometrics, 38(2):152–160, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270408.

Tatum:1997:REP

[Tat97] Lawrence G. Tatum. Robust estimation of the process standard deviation for control charts. Technometrics, 39(2):127–141, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270901.

Thomas:1990:AIP

[TC90] William Thomas and R. Dennis Cook. Assessing influence on predictions from generalized linear models. Technometrics, 32 (1):59–65, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269845.

Terrell:1992:BRB

[Ter92] George R. Terrell. Book review: Influence Diagrams, Belief Nets, and Decision Analysis by R. Oliver; J. Smith. Technometrics, 34(1):114–115, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269579.

Tucker:1993:ASP

[TFV93] William T. Tucker, Frederick W. Faltin, and Scott A. Vander Wiel. Algorithmic statistical process control: An elaboration. Technometrics, 35(4):363–375, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270270.

Taam:1993:DSE

[TH93] Winson Taam and Michael Hamada. Detecting spatial effects from factorial experiments: An application from integrated-circuit manufacturing. Technometrics, 35(2):149–160, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269660.

Thisted:1991:BRB

[Thi91] Ronald A. Thisted. Book review: Linear Least Squares Computations by R. W. Farebrother. Technometrics, 33(3):368–369, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268800.

Thomas:1990:BRB

[Tho90a] Edward V. Thomas. Book review: Advanced Scientific Computing in BASIC with Applications in Chemistry, Biology, and Pharmacology by P. Valkó; S. Vajda. Technometrics, 32(4):464–465, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270149.

Thompson:1990:BRB

[Tho90b] Carol B. Thompson. Book review: Introductory Statistical Methods: An Integrated Approach Using Minitab by Richard A. Groeneveld. Technometrics, 32(3):348, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269121.

Thomas:1991:EVE

[Tho91] Edward V. Thomas. Errors-in-variables estimation in multivariate calibration. Technometrics, 33(4):405–413, November 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269413. See correction

[Tho92a].

Thomas:1992:CEV

[Tho92a] Edward V. Thomas. Correction: “Errors-in-Variables Estimation in Multivariate Calibration”. Technometrics, 34(3):372, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270074. See

[Tho91].

Thomson:1992:BRB

[Tho92b] Steve Thomson. Book review: How SAS Works by Paul Herzberg. Technometrics, 34(1):117, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269582.

Thomas:1993:BRB

[Tho93] Edward V. Thomas. Book review: Modelling Binary Data by D. Collett. Technometrics, 35(2):224, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269672.

Thomas:1994:EIS

[Tho94] Edward V. Thomas. Evaluating the ignition sensitivity of thermalbattery heat pellets. Technometrics, 36(3):273–282, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269371.

Thode:1999:BRB

[Tho99] Henry C. Thode, Jr. Book review: Applied Probability for Engineers and Scientists by Ephraim Suhir. Technometrics, 41(1): 76, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271002.

Tan:1999:MMC

[TI99] Charles Y. Tan and Boris Iglewicz. Measurement-methods comparisons and linear statistical relationship. Technometrics, 41 (3):192–201, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270564.

Timm:1995:BRB

[Tim95] Neil H. Timm. Book review: Elements of Multivariate Time Series Analysis by Gregory C. Reinsel. Technometrics, 37(1):118, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269165.

Timm:1996:BRB

[Tim96] Neil H. Timm. Book review: Nonlinear Models for Repeated Measurement Data by Marie Davidian; David M. Giltinan. Technometrics, 38(2):186–187, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270420.

Timm:1999:BRB

[Tim99] Neil H. Timm. Book review: Multivariate Reduced-Rank Regression by Gregory C. Reinsel; Raja P. Velu. Technometrics, 41 (4):369–370, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271356.

Tobias:1996:FDR

[Tob96] Randall D. Tobias. [Follow-Up Designs to Resolve Confounding in Multifactor Experiments]: Discussion. Technometrics, 38(4): 326–327, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271302. See [MSB96a].

Tobias:1999:BRB

[Tob99] Randall D. Tobias. Book review: Chemometrics: A Practical Guide by Kenneth R. Beebe; Randy J. Pell; Mary Beth Seasholtz. Technometrics, 41(4):375–376, November 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271366.

Toman:1995:BRB

[Tom95] Blaza Toman. Book review: Design and Analysis of Experiments, Volume 1: Introduction to Experimental Design by Klaus Hinkelman; Oscar Kempthorne. Technometrics, 37(3):348–349, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269927.

Toran:1994:UCS

[Tor94] Laura Toran. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Discussion. Technometrics, 36(2):150–152, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270225. See [Got94c].

Tortorella:1995:BRB

[Tor95] Michael Tortorella. Book review: Evaluating Control Systems Reliability: Techniques and Applications by William M. Goble. Technometrics, 37(3):344–345, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269922.

Tortorella:1999:BRB

[Tor99] Michael Tortorella. Book review: Advances in Stochastic Models for Reliability, Quality and Safety by Waltraud Kahle; Elart von Collaní; Jurgen Franz; Uwe Jensen. Technometrics, 41(3):268– 269, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270581.

Tribus:1990:CCB

[Tri90] Myron Tribus. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3): 271–272, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269102. See [HK90].

Trosset:1997:BRB

[Tro97] Michael W. Trosset. Book review: Biplots by J. C. Gower; D. J. Hand. Technometrics, 39(3):339–340, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271148.

Tsai:1991:BRB

[Tsa91] Chih-Ling Tsai. Book review: Analysis of Means in Some Non-Standard Situations by J. B. Dijkstra. Technometrics, 33(1):115– 116, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269024.

Tsokos:1992:BRB

[Tso92] Chris P. Tsokos. Book review: Time Series: A Biostatistical Introduction by Peter J. Diggle. Technometrics, 34(2):229–230, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269245.

Tucker:1992:SPM

[Tuc92] William T. Tucker. [Statistical Process Monitoring and Feedback Adjustment]: Discussion. Technometrics, 34(3):275–277, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270031. See

[BK92a].

Tucker:1996:ASM

[Tuc96] William T. Tucker. [Applications of Statistical Methods to Nondestructive Evaluation]: Discussion. Technometrics, 38(2):126–127, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270404. See

[OM96a].

Ture:1994:ORC

[Tür94] Tahsin Erkan Türe. Optimal row-column designs for multiple comparisons with a control: A complete catalog. Technometrics, 36 (3):292–299, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269373.

Tsai:1990:DTW

[TW90] Chih-Ling Tsai and Xizhi Wu. Diagnostics in transformation and weighted regression. Technometrics, 32(3):315–322, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269108.

Tsai:1992:TMD

[TW92] Chih-Ling Tsai and Xizhi Wu. Transformation-model diagnostics. Technometrics, 34(2):197–202, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269235.

Thomas:1996:SEE

[TW96] James P. Thomas and Robert P. Wei. Standard-error estimates for rates of change from indirect measurements. Technometrics, 38(1):59–68, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268903.

Tsung:1998:ERD

[TWN98] Fugee Tsung, Huaiqing Wu, and Vijayan N. Nair. On the efficiency and robustness of discrete proportional-integral control schemes. Technometrics, 40(3):214–222, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271177.

Umble:1999:DFB

[USM99] Elisabeth J. Umble, John W. Seaman, Jr., and Harry F. Martz. A distribution-free Bayesian approach for determining the joint probability of failure of materials subject to multiple proof loads. Technometrics, 41(3):183–191, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270563.

Vangel:1991:BRB

[Van91] Mark G. Vangel. Book review: Continuous Mining Reliability: Design and Operation of Mechanized Systems by Vladimir Pavlovic. Technometrics, 33(2):241, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269054.

VanBrackle:1992:BRB

[Van92a] Lewis N. VanBrackle. Book review: Probability for Applications by Paul Pfeiffer. Technometrics, 34(1):119–120, February 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269586.

Vangel:1992:NMO

[Van92b] Mark G. Vangel. New methods for one-sided tolerance limits for a one-way balanced random-effects ANOVA model. Technometrics, 34(2):176–185, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269233.

VanderWiel:1996:MPW

[Van96a] S. A. Vander Wiel. Monitoring processes that wander using integrated moving average models. Technometrics, 38(2):139–151, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270407.

Vangel:1996:BRB

[Van96b] Mark G. Vangel. Book review: System Reliability Theory: Models and Statistical Methods by Arnljot Høyland; Marvin Rausand. Technometrics, 38(1):79–80, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268910.

Vangel:1999:SNP

[Van99] Mark G. Vangel. [Simultaneous Nonparametric Prediction Limits]: Discussion. Technometrics, 41(2):108, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270725. See [DM99a].

Vardeman:1994:ER

[Var94a] Stephen B. Vardeman. Editor’s report. Technometrics, 36(1):1– 2, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269190.

Vardeman:1994:NE

[Var94b] Steve Vardeman. New Editor. Technometrics, 36(4):337, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269948.

Velilla:1994:DRU

[VB94a] Santiago Velilla and Juan A. Barrio. A discriminant rule under transformation. Technometrics, 36(4):348–353, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269950.

Viveros:1994:IEP

[VB94b] Román Viveros and N. Balakrishnan. Interval estimation of parameters of life from progressively censored data. Technometrics, 36(1):84–91, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269201.

Voss:1994:SIP

[VD94] Daniel T. Voss and Steven DeMartino. Spatial interpolation for photographic reproduction of color. Technometrics, 36(3):249– 259, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269369.

vanderVoet:1997:PMR

[vdVtBM97] Hilko van der Voet, Cajo J. F. ter Braak, and Colin Mallows. C_{P} and prediction with many regressors: Comments on Mallows (1995). Technometrics, 39(1):115–116, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270805.

Verducci:1990:BRB

[Ver90] Joseph S. Verducci. Book review: Topics in Applied Multivariate Analysis by D. M. Hawkins. Technometrics, 32(1):99–100, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269857.

Vining:1992:BRB

[Vin92] Geoff Vining. Book review: Data Analysis: A Model-Comparison Approach by Charles M. Judd; Gary H. McClelland. Technometrics, 34(1):107–108, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269570.

Viveros:1991:CSS

[Viv91] R. Viveros. Combining series system data to estimate component characteristics. Technometrics, 33(1):13–23, February 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269004.

Vining:1991:GAE

[VM91] G. Geoffrey Vining and Raymond H. Myers. A graphical approach for evaluating response surface designs in terms of the mean squared error of prediction. Technometrics, 33(3):315–326, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268783.

vanNostrand:1990:BRB

[vN90] R. Craig van Nostrand. Book review: Applied Regression Analysis: A Research Tool by John O. Rawlings. Technometrics, 32 (1):95–96, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269851.

Voelkel:1994:BRB

[Voe94] Joseph O. Voelkel. Book review: Understanding Variation: The Key to Managing Chaos by D. Wheeler. Technometrics, 36(2):225, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270246.

Venter:1996:HTA

[VS96] J. H. Venter and S. J. Steel. A hypothesis-testing approach toward identifying active contrasts. Technometrics, 38(2):161–169, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270409.

Venter:1998:IAC

[VS98] J. H. Venter and S. J. Steel. Identifying active contrasts by stepwise testing. Technometrics, 40(4):304–313, November 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270531.

vonTress:1991:BRB

[vT91] Mark von Tress. Book review: Analysis of Messy Data, Volume 2: Nonreplicated Experiments by George A. Milliken; Dallas E. Johnson. Technometrics, 33(1):114–115, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269023.

vonTress:1995:BRB

[vT95] Mark von Tress. Book review: Statistical Models for Ordinal Variables by Clifford C. Clogg; Edward S. Shihadeh. Technometrics, 37(1):118–119, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269166.

VanderWiel:1992:ASP

[VTFD92] Scott A. Vander Wiel, William T. Tucker, Frederick W. Faltin, and Necip Doganaksoy. Algorithmic statistical process control: Concepts and an application. Technometrics, 34(3):286–297, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270035.

Vukovinsky:1998:BRB

[Vuk98] Kim Vukovinsky. Book review: Statistical Process Control by Leonard A. Doty. Technometrics, 40(1):73–74, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271395.

VanderWiel:1992:SPM

[VV92] Scott A. Vander Wiel and Stephen B. Vardeman. [Statistical Process Monitoring and Feedback Adjustment]: Discussion. Technometrics, 34(3):278–281, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270032. See [BK92a].

VanderWiel:1994:DAN

[VV94] Scott A. Vander Wiel and Stephen B. Vardeman. A discussion of all-or-none inspection policies. Technometrics, 36(1):102–109, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269203.

Vardeman:1999:TWR

[VV99] Stephen B. Vardeman and Enid S. VanValkenburg. Two-way random-effects analyses and gauge R&R studies. Technometrics, 41(3):202–211, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270565.

Wade:1999:BRB

[Wad99] Mark R. Wade. Book review: Construction and Assessment of Classification Rules by David J. Hand. Technometrics, 41(3): 267, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270579.

Waller:1992:BRB

[Wal92a] Ray A. Waller. Book review: Probability and Statistics for the Engineering, Computing, and Physical Sciences by Edward R. Dougherty. Technometrics, 34(2):241–242, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269263.

Walters:1992:BRB

[Wal92b] Sherry K. Walters. Book review: Quality and Reliability Methods for Primary Batteries by P. Bro; S. C. Levy. Technometrics, 34(3):361, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270053.

Wangen:1990:CCB

[Wan90] Larry E. Wangen. Commentary: [Communications between Statisticians and Engineers/Physical Scientists]. Technometrics, 32(3):273–274, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269103. See [HK90].

Wang:1994:BRB

[Wan94] Min-Chiang Wang. Book review: Pyzdek’s Guide to SPC: Volume One, Fundamentals by Thomas Pyzdek. Technometrics, 36 (4):422, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269961.

Warren:1990:BRB

[War90] John Warren. Book review: Mathematical and Statistical Methods in Nuclear Safeguards by F. Argentesi; R. Avenhaus; M. Franklin; J. P. Shipley. Technometrics, 32(3):360–361, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269138.

Wasserman:1991:BRB

[Was91] Gary S. Wasserman. Book review: Probability and Statistics for Modern Engineering by Lawrence Lapin. Technometrics, 33(4): 490, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269438.

Wasserstein:1997:BRB

[Was97] Ronald L. Wasserstein. Book review: Monte Carlo: Concepts, Algorithms, and Applications by G. S. Fishman. Technometrics, 39(3):338, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271146.

Watt:1990:BRB

[Wat90] Terry L. Watt. Book review: Applied Statistics: A First Course by M. L. Berenson; D. M. Levine; D. Rindskopf. Technometrics, 32(2):224, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268871.

Watt:1996:BRB

[Wat96] Terry L. Watt. Book review: Numerical Analysis for the Geological Sciences by James R. Carr. Technometrics, 38(3):293–294, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270623.

Wendelberger:1994:UCS

[WB94] Joanne Wendelberger and Richard Beckman. [The Use of Conditional Simulation in Nuclear-Waste-Site Performance Assessment]: Discussion. Technometrics, 36(2):144–147, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270222. See [Got94c].

Welch:1992:SPC

[WBS^{+}92] William J. Welch, Robert. J. Buck, Jerome Sacks, Henry P. Wynn, Toby J. Mitchell, and Max D. Morris. Screening, predicting, and computer experiments. Technometrics, 34(1):15–25, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269548. See

comments [LLW^{+}96].

Wu:1992:GAM

[WC92] C. F. J. Wu and Youyi Chen. A graph-aided method for planning two-level experiments when certain interactions are important. Technometrics, 34(2):162–175, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269232.

Webster:1990:BRB

[Web90] J. T. Webster. Book review: Experimental Design and Analysis by Marvin Lentner; Thomas Bishop. Technometrics, 32(1):95, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269850.

Wehrly:1991:BRB

[Weh91] Thomas E. Wehrly. Book review: Engineering Applications of Stochastic Processes by Alexander Zayadezny; Daniel Tabak; Dov Wulich. Technometrics, 33(1):120, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269032.

Wehrly:1993:BRB

[Weh93] Thomas Wehrly. Book review: Complex Stochastic Systems by D. R. Cox; D. M. Titterington. Technometrics, 35(4):461–462, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270294.

Wehrly:1995:BRB

[Weh95] Thomas E. Wehrly. Book review: Statistical Analysis of Circular Data by N. I. Fisher. Technometrics, 37(2):229–230, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269626.

Weisberg:1990:BRB

[Wei90] Sanford Weisberg. Book review: Transformations and Weighting in Regression by R. J. Carroll; D. Ruppert. Technometrics, 32 (1):96–97, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1269852. See corrections [Ano90g].

Weisberg:1994:BRB

[Wei94] Sanford Weisberg. Book review: Handbook of Statistics 9: Computational Statistics by C. R. Rao. Technometrics, 36(3):328– 329, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269391.

Welch:1990:BRB

[Wel90] Michael E. Welch. Book review: Bayesian Analysis of Time Series and Dynamic Models by James C. Spall. Technometrics, 32 (3):354–355, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269130.

Welch:1993:BRB

[Wel93] Michael E. Welch. Book review: Bivariate Discrete Distributions by Subrahmaniam Kocherlakota; Kathleen Kocherlakota. Technometrics, 35(3):328–329, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269532.

Wendelberger:1993:BRB

[Wen93] Joanne R. Wendelberger. Book review: Adventures in Stochastic Processes by Sidney I. Resnick. Technometrics, 35(4):461, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270293.

Wendelberger:1994:BRB

[Wen94] Joanne R. Wendelberger. Book review: Statistics and Society, Data Collection and Interpretation by Walter T. Federer. Technometrics, 36(3):329–330, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269393.

Wesolowsky:1990:SAC

[Wes90] G. O. Wesolowsky. Simultaneous acceptance control charts for two correlated processes. Technometrics, 32(1):43–48, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269843.

Westfall:1992:BRB

[Wes92] Peter H. Westfall. Book review: Bayesian Forecasting and Dynamic Models by M. West; J. Harrison. Technometrics, 34(1): 115–116, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269581.

Wardrop:1992:SPM

[WG92] Daniel M. Wardrop and Carlos E. Garcia. [Statistical Process Monitoring and Feedback Adjustment]: Discussion. Technometrics, 34(3):281–282, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270033. See [BK92a].

Woo:1997:SIG

[WH97] T. C. Woo and Chung-Chi Hsieh. [Statistical Issues in Geometric Feature Inspection Using Coordinate Measuring Machines]: Discussion. Technometrics, 39(1):20–22, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270766. See [DGTZ97a].

Wharton:1995:BRB

[Wha95] Robert M. Wharton. Book review: Model-Free Curve Estimation by Michael E. Tarter; Michael D. Lock. Technometrics, 37 (4):460, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269743.

Wheeler:1993:BRB

[Whe93] Robert E. Wheeler. Book review: Optimum Experimental Designs by A. C. Atkinson; A. N. Donev. Technometrics, 35(4): 450, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270278.

Wang:1994:TID

[WI94] C. Ming Wang and Hari K. Iyer. Tolerance intervals for the distribution of true values in the presence of measurement errors. Technometrics, 36(2):162–170, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270228.

Willner:1990:BRB

[Wil90a] Ozzie Willner. Book review: How to Choose the Proper Sample Size by Gary Brush. Technometrics, 32(1):94–95, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269849.

Wilson:1990:BRB

[Wil90b] William J. Wilson. Book review: Elementary Statistics by Robert Johnson. Technometrics, 32(1):113, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269873.

Wilson:1998:BRB

[Wil98] William J. Wilson. Book review: Statistics: The Conceptual Approach by Gudmund R. Iversen; Mary Gergen. Technometrics, 40(1):77, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271401.

Wincek:1990:BRB

[Win90] Michael A. Wincek. Book review: Applied Statistical Time Series Analysis by Robert H. Shumway. Technometrics, 32(3):353– 354, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269129.

Wincek:1993:BRB

[Win93] Michael A. Wincek. Book review: Forecasting with Dynamic Regression Models by Alan Pankrantz. Technometrics, 35(1):87– 88, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269295.

Witt:1994:BRB

[Wit94] Thomas Witt. Book review: A MINITAB Companion with Macros by Peter W. Zehna. Technometrics, 36(1):124, February 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269222.

Weerahandi:1992:TRS

[WJ92] Samaradasa Weerahandi and Richard A. Johnson. Testing reliability in a stress-strength model when X and Y are normally distributed. Technometrics, 34(1):83–91, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269555. See [PGT95].

Wang:1995:CL

[WKWL95] P. C. Wang, Nouna Kettaneh-Wold, and Dennis K. J. Lin. Comments on Lin (1993). Technometrics, 37(3):358–359, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269944. See

[Lin93b].

Whitmore:1991:MSD

[WL91] G. Alex Whitmore and Mei Ling Ting Lee. A multivariate survival distribution generated by an inverse Gaussian mixture of exponentials. Technometrics, 33(1):39–50, February 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269006.

Wang:1997:MEM

[WL97] C. Ming Wang and C. Teresa Lam. A mixed-effects model for the analysis of circular measurements. Technometrics, 39(2):119–126, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270900.

Woodall:1990:EWM

[WM90] William H. Woodall and Hazem D. Maragah. [Exponentially Weighted Moving Average Control Schemes: Properties and Enhancements]: Discussion. Technometrics, 32(1):17–18, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269837. See

[LS90b].

Wardell:1994:RLDa

[WMP94a] Don G. Wardell, Herbert Moskowitz, and Robert D. Plante. Runlength distributions of special-cause control charts for correlated processes. Technometrics, 36(1):3–17, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269191. See discussion [Luc94,

AWS94, Fel94] and rejoinder [WMP94b].

Wardell:1994:RLDb

[WMP94b] Don G. Wardell, Herbert Moskowitz, and Robert D. Plante. [Run-Length Distributions of Special-Cause Control Charts for Correlated Processes]: Rejoinder. Technometrics, 36(1):23–27, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269195.

Wludyka:1997:AMT

[WN97] Peter S. Wludyka and Peter R. Nelson. An analysis-of-means-type test for variances from normal populations. Technometrics, 39(3):274–285, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271132.

Wolf:1991:BRB

[Wol91] Dennis A. Wolf. Book review: Algorithms for Chemists by Jure Zupan. Technometrics, 33(3):367–368, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268799.

Wolkan:1992:BRB

[Wol92] Joseph A. Wolkan. Book review: Introduction to Probability and Statistics by J. S. Milton; J. C. Arnold. Technometrics, 34(2):244, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269268.

Wold:1993:DPC

[Wol93] Svante Wold. Discussion: PLS in chemical practice. Technometrics, 35(2):136–139, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269657.

Wolstenholme:1995:NTW

[Wol95] Linda C. Wolstenholme. A nonparametric test of the weakestlink principle. Technometrics, 37(2):169–175, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269618.

Wolf:1996:BRB

[Wol96] Dennis A. Wolf. Book review: Recent Advances in Descriptive Multivariate Analysis by Wojtek J. Krzanowski. Technometrics, 38(3):291–292, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270621.

Wong:1992:ISP

[Won92] Augustine C. M. Wong. Inferences on the shape parameter of a gamma distribution: A conditional approach. Technometrics, 34(3):348–351, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270041.

Wong:1994:BRB

[Won94] Aldous Wong. Book review: Statistical Methods by R. J. Freund; W. J. Wilson. Technometrics, 36(2):222, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270242.

Wong:1998:BRB

[Won98] Aldous G. Wong. Book review: The Theory of Constraints: Applications in Quality and Manufacturing by Robert E. Stein. Technometrics, 40(3):257, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271183.

Woodall:1990:CCV

[Woo90] William H. Woodall. [CUSUM Charts with Variable Sampling Intervals]: Discussion. Technometrics, 32(4):389–391, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270117. See

[RAA90a].

Wood:1991:BRB

[Woo91] Jack Wood. Book review: News and Numbers: A Guide to Reporting Statistical Claims and Controversies in Health and Related Fields by Victor Cohn. Technometrics, 33(3):373–374, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268806.

Woodall:1992:BRB

[Woo92] William H. Woodall. Book review: SPC and Continuous Improvement by Mal Owen. Technometrics, 34(1):100–101, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269561.

Wozniak:1991:BRB

[Woz91] Patricia J. Wozniak. Book review: Applied Nonparametric Statistics by Wayne W. Daniel. Technometrics, 33(3):364–365, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268795.

Wolfinger:1998:JEL

[WT98] Russell D. Wolfinger and Randall D. Tobias. Joint estimation of location, dispersion, and random effects in robust design. Technometrics, 40(1):62–71, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271393.

Wu:1999:BRB

[Wu99] Yuhai Wu. Book review: Statistical Learning Theory by Vladimir N. Vapnik. Technometrics, 41(4):377–378, November 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271368.

Wang:1997:RRA

[WVYB97] C. Ming Wang, Dominic F. Vecchia, Matt Young, and Nathan A. Brilliant. Robust regression applied to optical-fiber dimensional quality control. Technometrics, 39(1):25–33, February 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270769.

Wang:1992:NOA

[WW92] J. C. Wang and C. F. J. Wu. Nearly orthogonal arrays with mixed levels and small runs. Technometrics, 34(4):409–422, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268940.

Wyckoff:1996:BRB

[Wyc96] James Wyckoff. Book review: Statistical Methods for Quality: With Applications to Engineering and Management by Irwin Miller; Marylees Miller. Technometrics, 38(2):183, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270414.

Yakowitz:1997:BRB

[Yak97] Sidney J. Yakowitz. Book review: An Introduction to Bayesian Networks by Finn V. Jensen. Technometrics, 39(3):336–337, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271143.

Yandell:1996:BRB

[Yan96] Brian S. Yandell. Book review: Kernel Smoothing by M. P. Wand; M. C. Jones. Technometrics, 38(1):75–76, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268906.

Yandell:1997:BRB

[Yan97] Brian S. Yandell. Book review: Smoothing Methods in Statistics by J. S. Simonoff. Technometrics, 39(3):338–339, August 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271147.

Yashchin:1990:LE

[Yas90] Emmanuel Yashchin. Letter to the Editor. Technometrics, 32 (4):469, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270155.

Yashchin:1992:ACO

[Yas92] Emmanuel Yashchin. Analysis of CUSUM and other Markov-type control schemes by using empirical distributions. Technometrics, 34(1):54–63, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269552.

Yashchin:1993:PCC

[Yas93] Emmanuel Yashchin. Performance of CUSUM control schemes for serially correlated observations. Technometrics, 35(1):37–52, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269288.

Yashchin:1994:MVC

[Yas94] Emmanuel Yashchin. Monitoring variance components. Technometrics, 36(4):379–393, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269953.

Yashchin:1995:ECM

[Yas95] Emmanuel Yashchin. Estimating the current mean of a process subject to abrupt changes. Technometrics, 37(3):311–323, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269915.

Yanagisawa:1990:CLA

[YDB90] Y. Yanagisawa, J. Disney, and A. Bendell. [A Critical Look at Accumulation Analysis and Related Methods]: Discussion. Technometrics, 32(2):153–157, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268858. See [HW90b].

Young:1994:EEQ

[YE94] Linda J. Young and Robert G. Easterling. Estimation of extreme quantiles based on sensitivity tests: A comparative study. Technometrics, 36(1):48–60, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269198.

Yeh:1993:BRB

[Yeh93] Chyon-Hwa Yeh. Book review: Design Statistics in Pharmacochemistry by Peter P. Mager. Technometrics, 35(3):331, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269536.

Ying:1991:BRB

[Yin91] Zhiliang Ying. Book review: Multi-Armed Bandit Allocation Indices by John C. Gittins. Technometrics, 33(4):488, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269435.

Yashchin:1995:CWM

[YWMP95] Emmanuel Yashchin, Don G. Wardell, Herbert Moskowitz, and Robert D. Plante. Comments on Wardell, Moskowitz, and Plante (1994). Technometrics, 37(2):243–245, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269653.

Zadeh:1995:DPT

[Zad95] Lotfi A. Zadeh. Discussion: Probability theory and fuzzy logic are complementary rather than competitive. Technometrics, 37 (3):271–276, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269908.

Zeitler:1994:BRB

[Zei94] David W. Zeitler. Book review: Introduction to Quality Control by Donald DelMar; George Sheldon. Technometrics, 36(1): 120–121, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269217.

Zelterman:1994:BRB

[Zel94] Daniel Zelterman. Book review: Statistical Models for Causal Analysis by Robert D. Retherford; Minja Kim Choe. Technometrics, 36(3):327–328, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269389.

Zhang:1992:BRB

[Zha92] Nien Fan Zhang. Book review: Forecasting and Time Series Analysis by D. C. Montgomery; L. A. Johnson; J. S. Gardiner.

Technometrics, 34(1):98–99, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269557.

Zhang:1994:MCA

[Zha94] Heping Zhang. Maximal correlation and adaptive splines. Technometrics, 36(2):196–201, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270231.

Zhang:1998:SCC

[Zha98] Nien Fan Zhang. A statistical control chart for stationary process data. Technometrics, 40(1):24–38, February 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271390. See comments

[Sch98b].

Ziegel:1990:BRBx

[Zie90a] Eric R. Ziegel. Book review: Applied Nonparametric Statistical Models by P. Sprent. Technometrics, 32(4):455–456, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270134.

Ziegel:1990:BRBv

[Zie90b] Eric R. Ziegel. Book review: Applied Regression Analysis in Econometrics by Howard E. Doran. Technometrics, 32(4):451, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270128.

Ziegel:1990:BRBw

[Zie90c] Eric R. Ziegel. Book review: Applied Time Series Analysis by Robert Shumway. Technometrics, 32(4):451–452, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270129.

Ziegel:1990:BRBo

[Zie90d] Eric R. Ziegel. Book review: COMPSTAT 1988 by D. Edwards; N. Raun. Technometrics, 32(2):239, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268892.

Ziegel:1990:BRBl

[Zie90e] Eric R. Ziegel. Book review: Contouring Geologic Surfaces with the Computer by T. Jones; D. Hamilton; C. Johnson. Technometrics, 32(2):232, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268882.

Ziegel:1990:BRBp

[Zie90f] Eric R. Ziegel. Book review: Data Analysis and Informatics 5 by E. Diday. Technometrics, 32(2):239–240, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268893.

Ziegel:1990:BRBz

[Zie90g] Eric R. Ziegel. Book review: Data, Expert Knowledge, and Decisions by W. Gaul; M. Schader. Technometrics, 32(4):465–466, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270151.

Ziegel:1990:BRBh

[Zie90h] Eric R. Ziegel. Book review: Encyclopedia of Statistical Sciences, Volume 9: Strata Chart to End with Cumulative Index by S. Kotz; N. Johnson; C. Read. Technometrics, 32(2):225, May 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268872.

Ziegel:1990:BRBd

[Zie90i] Eric R. Ziegel. Book review: Geostatistical Case Studies by G. Matheron; M. Armstrong. Technometrics, 32(1):105–106, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269864.

Ziegel:1990:BRBj

[Zie90j] Eric R. Ziegel. Book review: Geostatistics and Petroleum Geology by Michael Hohn. Technometrics, 32(2):231, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268880.

Ziegel:1990:BRBy

[Zie90k] Eric R. Ziegel. Book review: Handbook of Statistics, Volume 6: Sampling by P. R. Krishnaiah; C. R. Rao. Technometrics, 32 (4):460, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270142.

Ziegel:1990:BRBa

[Zie90l] Eric R. Ziegel. Book review: Juran’s Quality Control Handbook by J. M. Juran; F. M. Gryna. Technometrics, 32(1):97–98, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269854.

Ziegel:1990:BRBe

[Zie90m] Eric R. Ziegel. Book review: Mathematical Geology by N. Rock. Technometrics, 32(1):106, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269865.

Ziegel:1990:BRBq

[Zie90n] Eric R. Ziegel. Book review: Mathematical Modelling by J. N. Kapur. Technometrics, 32(2):240, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268894.

Ziegel:1990:BRBf

[Zie90o] Eric R. Ziegel. Book review: Model-Oriented Data Analysis by V. Federov; H. Lauter. Technometrics, 32(1):106, February 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269866.

Ziegel:1990:BRBaa

[Zie90p] Eric R. Ziegel. Book review: On Model Uncertainty and Its Statistical Implications by Theo Dijkstra. Technometrics, 32(4):466, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270152.

Ziegel:1990:BRBi

[Zie90q] Eric R. Ziegel. Book review: Optimal Design and Analysis of Experiments by Y. Dodge; V. Federov; H. Wynn. Technometrics, 32(2):229, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268877.

Ziegel:1990:BRBs

[Zie90r] Eric R. Ziegel. Book review: Principles of Environmental Sampling by Lawrence Keith. Technometrics, 32(3):357, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269133.

Ziegel:1990:BRBn

[Zie90s] Eric R. Ziegel. Book review: Probability and Statistics by Jaya Srivastava. Technometrics, 32(2):239, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268891.

Ziegel:1990:BRBr

[Zie90t] Eric R. Ziegel. Book review: Statistical Quality Control by E. Grant; R. Leavenworth. Technometrics, 32(3):343–344, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269115.

Ziegel:1990:BRBu

[Zie90u] Eric R. Ziegel. Book review: Statistical Tables and Formulae by S. Kokoska; C. Nevison. Technometrics, 32(3):361, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269140.

Ziegel:1990:BRBg

[Zie90v] Eric R. Ziegel. Book review: Statistical Theory and Data Analysis II by K. Matusita. Technometrics, 32(1):113–114, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269874.

Ziegel:1990:BRBt

[Zie90w] Eric R. Ziegel. Book review: Statistics: A Guide to the Unknown by J. Tanur; F. Mosteller; W. Kruskal; E. Lehman; R. Link; R. Pieters; G. Rising. Technometrics, 32(3):361, August 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269139.

Ziegel:1990:BRBb

[Zie90x] Eric R. Ziegel. Book review: Statistics in Research by Bernard Ostle; Linda Malone. Technometrics, 32(1):100–101, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269858.

Ziegel:1990:BRBm

[Zie90y] Eric R. Ziegel. Book review: The Evolving Role of Statistical Assessments as Evidence in the Courts by Stephen Fienberg. Technometrics, 32(2):238–239, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268890.

Ziegel:1990:BRBk

[Zie90z] Eric R. Ziegel. Book review: Use and Abuse of Statistical Methods in the Earth Sciences by William Size. Technometrics, 32(2): 231–232, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268881.

Ziegel:1990:BRBc

[Zie90-27] Eric R. Ziegel. Book reviews: Data Analysis for Managers by Harry Roberts; Minitab Handbook for Business and Economics by Robert Miller; Essential Business Statistics by K. Bond; J. Scott; Mathematical Modelling with Minitab by W. Fox; F. Giordano; S. Maddox; M. Weir. Technometrics, 32(1):102, February 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269860.

Ziegel:1990:BRBab

[Zie90-28] Eric R. Ziegel. Book reviews: Survey Errors and Survey Costs by Robert M. Groves; Telephone Survey Methodology by R. Groves; P. Biemer; L. Lyberg; J. Massey; W. Nicholls; J. Waksberg; Panel Surveys by D. Karprzyk; G. Duncan; G. Kalton; M. Singh. Technometrics, 32(4):466–467, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270153.

Ziegel:1991:BRBl

[Zie91a] Eric R. Ziegel. Book review: A Dictionary of Statistical Terms by F. H. C. Marriott. Technometrics, 33(2):247–248, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269067.

Ziegel:1991:BRBp

[Zie91b] Eric R. Ziegel. Book review: Assignments in Applied Statistics by Simon Conrad. Technometrics, 33(2):249, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269071.

Ziegel:1991:BRBm

[Zie91c] Eric R. Ziegel. Book review: Beta Mathematics Handbook by L. Rade; B. Westergren. Technometrics, 33(2):248, May 1991. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269068.

Ziegel:1991:BRBo

[Zie91d] Eric R. Ziegel. Book review: Classical and Modern Regression with Applications by Raymond Myers. Technometrics, 33(2):248, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269070.

Ziegel:1991:BRBk

[Zie91e] Eric R. Ziegel. Book review: Collected Works of John W. Tukey, Volume 6: More Mathematical 1938–1984 by Colin Mallows and John W. Tukey. Technometrics, 33(2):247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269066?.

Ziegel:1991:BRBj

[Zie91f] Eric R. Ziegel. Book review: Collected Works of John W. Tukey, Volume 6: More Mathematical 1938–1984 by Colin Mallows; John W. Tukey. Technometrics, 33(2):247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269066.

Ziegel:1991:BRBa

[Zie91g] Eric R. Ziegel. Book review: Concise Encyclopedia of Science and Technology by Sybil Parker. Technometrics, 33(1):121, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269034.

Ziegel:1991:BRBq

[Zie91h] Eric R. Ziegel. Book review: Data Analysis in the Chemical Industry: Volume 1. Basics by Roland Caulcott. Technometrics, 33(3):366, August 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268797.

Ziegel:1991:BRBc

[Zie91i] Eric R. Ziegel. Book review: Handbook of Nonlinear Regression Models by David Ratkowsky. Technometrics, 33(2):240–241, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269053.

Ziegel:1991:BRBn

[Zie91j] Eric R. Ziegel. Book review: Linear Statistical Models: An Applied Approach by Bruce Bowerman; Richard O’Connell. Technometrics, 33(2):248, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269069.

Ziegel:1991:BRBi

[Zie91k] Eric R. Ziegel. Book review: Multivariate Statistics and Probability: Essays in Honor of P. R. Krishnaiah by C. Rao; M. Rao. Technometrics, 33(2):247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269065.

Ziegel:1991:BRBh

[Zie91l] Eric R. Ziegel. Book review: Nonlinear Regression, Functional Relations, and Robust Methods by H. Bunke; O. Bunke. Technometrics, 33(2):247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269064.

Ziegel:1991:BRBd

[Zie91m] Eric R. Ziegel. Book review: Probability and Statistics in Geodesy and Geophysics by L. Kubackova; L. Kubacek; J. Kukuca. Technometrics, 33(2):241, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269055.

Ziegel:1991:BRBb

[Zie91n] Eric R. Ziegel. Book review: Risk Management and Assessment by Lester Lave. Technometrics, 33(1):121, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269035.

Ziegel:1991:BRBr

[Zie91o] Eric R. Ziegel. Book review: Statistical Design and Analysis of Industrial Experiments by Subir Ghosh. Technometrics, 33(4): 470, November 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269418.

Ziegel:1991:BRBg

[Zie91p] Eric R. Ziegel. Book review: Taguchi Methods (Special Issue of Quality and Reliability Engineering) by Norman Harris. Technometrics, 33(2):246–247, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269063.

Ziegel:1991:BRBe

[Zie91q] Eric R. Ziegel. Book reviews: Analysis of Binary Data by D. R. Cox; E. J. Snell; Design and Analysis of Cross-Over Trials by Byron Jones; Michael Kenward. Technometrics, 33(2):242–243, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269058.

Ziegel:1991:BRBf

[Zie91r] Eric R. Ziegel. Book reviews: Statistical Methodology in the Pharmaceutical Sciences by Donald L. Berry; Statistical Issues in Drug Research and Development by Karl E. Peace. Technometrics, 33 (2):243, May 1991. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269059.

Ziegel:1992:BRBd

[Zie92a] Eric R. Ziegel. Book review: A History of Probability and Statistics and Their Applications before 1750 by Anders Hald. Technometrics, 34(1):122, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269591.

Ziegel:1992:BRBa

[Zie92b] Eric R. Ziegel. Book review: Applied Linear Statistical Models by John Neter; William Wasserman; Michael Kutner. Technometrics, 34(1):121, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269588.

Ziegel:1992:BRBl

[Zie92c] Eric R. Ziegel. Book review: Chemometrics Tutorials by D. Massart; R. Brereton; R. Dessy; P. Hopke; C. Spiegelman; W. Wegscheider. Technometrics, 34(2):247, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269275.

Ziegel:1992:BRBh

[Zie92d] Eric R. Ziegel. Book review: Computation Methods in the Chemical Sciences by A. Carley; P. Morgan. Technometrics, 34(2): 245–246, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269271.

Ziegel:1992:BRBk

[Zie92e] Eric R. Ziegel. Book review: Computer-Aided Multivariable Analysis by A. Afifi; V. Clark. Technometrics, 34(2):246–247, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269274.

Ziegel:1992:BRBc

[Zie92f] Eric R. Ziegel. Book review: Data Analysis for Research Designs by G. Keppel; C. Zedick. Technometrics, 34(1):122, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269590.

Ziegel:1992:BRBg

[Zie92g] Eric R. Ziegel. Book review: Encyclopedia of Statistical Sciences, Supplement Volume by S. Kotz; N. Johnson; C. Read. Technometrics, 34(1):123, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269594.

Ziegel:1992:BRBf

[Zie92h] Eric R. Ziegel. Book review: Geostatistics by M. Armstrong. Technometrics, 34(1):123, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269593.

Ziegel:1992:BRBn

[Zie92i] Eric R. Ziegel. Book review: How to Run Mixture Experiments for Product Quality by John Cornell. Technometrics, 34(3):370, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270067.

Ziegel:1992:BRBb

[Zie92j] Eric R. Ziegel. Book review: Mastering Technical Writing by Joseph Mancuso. Technometrics, 34(1):121–122, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269589.

Ziegel:1992:BRBj

[Zie92k] Eric R. Ziegel. Book review: Multiway Data Analysis by R. Coppi; S. Bolasco. Technometrics, 34(2):246, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269273.

Ziegel:1992:BRBt

[Zie92l] Eric R. Ziegel. Book review: Probability and Statistics for Engineering and the Sciences by Jay L. Devore. Technometrics, 34(3): 372, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270073.

Ziegel:1992:BRBs

[Zie92m] Eric R. Ziegel. Book review: Sampling of Populations: Methods and Applications by P. Levy; S. Lemeshaw. Technometrics, 34(3):372, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270072.

Ziegel:1992:BRBu

[Zie92n] Eric R. Ziegel. Book review: SAS System for Linear Models by R. Littell; R. Freund; P. Spector. Technometrics, 34(4):500, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268965.

Ziegel:1992:BRBv

[Zie92o] Eric R. Ziegel. Book review: SAS System for Regression by R. Freund; R. Littell. Technometrics, 34(4):500–501, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268966.

Ziegel:1992:BRBr

[Zie92p] Eric R. Ziegel. Book review: State Space Modeling of Time Series by M. Aoki. Technometrics, 34(3):371–372, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270071.

Ziegel:1992:BRBi

[Zie92q] Eric R. Ziegel. Book review: Statistical Data Analysis and Inference by Yadolah Dodge. Technometrics, 34(2):246, May 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269272.

Ziegel:1992:BRBo

[Zie92r] Eric R. Ziegel. Book review: Statistical Process Control in Manufacturing by J. Keats; D. Montgomery. Technometrics, 34(3):370– 371, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270068.

Ziegel:1992:BRBx

[Zie92s] Eric R. Ziegel. Book review: Statistical Theory and Modelling by D. Hinkley; N. Reid; E. J. Snell. Technometrics, 34(4):501, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268968.

Ziegel:1992:BRBe

[Zie92t] Eric R. Ziegel. Book review: Statistics for Lawyers by M. Finkelstein; B. Levin. Technometrics, 34(1):122–123, February 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269592.

Ziegel:1992:BRBw

[Zie92u] Eric R. Ziegel. Book review: Statistics in Research and Development by R. Caulcott. Technometrics, 34(4):501, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268967.

Ziegel:1992:BRBq

[Zie92v] Eric R. Ziegel. Book review: Time Series: Theory and Methods by P. Brockwell; R. Davis. Technometrics, 34(3):371, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270070.

Ziegel:1992:BRBp

[Zie92w] Eric R. Ziegel. Book review: Total Quality Control by A. Fergenbaum. Technometrics, 34(3):371, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270069.

Ziegel:1992:BRBm

[Zie92x] Eric R. Ziegel. Book review: Understanding Industrial Experimentation by Donald J. Wheeler. Technometrics, 34(3):370, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270066.

Ziegel:1993:BRBm

[Zie93a] Eric R. Ziegel. Book review: A Leader’s Journey to Quality by Dana Cound. Technometrics, 35(1):102, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269322.

Ziegel:1993:BRBae

[Zie93b] Eric R. Ziegel. Book review: ANOVA: Repeated Measures by E. Girden. Technometrics, 35(4):464–465, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270301.

Ziegel:1993:BRBab

[Zie93c] Eric R. Ziegel. Book review: Applied Statistics for Engineers and Physical Scientists by R. Hogg; J. Ledolter. Technometrics, 35 (4):463, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270298.

Ziegel:1993:BRBd

[Zie93d] Eric R. Ziegel. Book review: Chemometrics: Experimental Design by E. Morgan. Technometrics, 35(1):99, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269313.

Ziegel:1993:BRBaa

[Zie93e] Eric R. Ziegel. Book review: Computing and Graphics in Statistics by A. Buja; P. Tukey. Technometrics, 35(4):463, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270297.

Ziegel:1993:BRBr

[Zie93f] Eric R. Ziegel. Book review: Control and Analysis of Noisy Processes by D. Koenig. Technometrics, 35(2):231–232, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269681.

Ziegel:1993:BRBq

[Zie93g] Eric R. Ziegel. Book review: Correspondence Analysis Handbook by J. P. Benzecri. Technometrics, 35(1):103, February 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269326.

Ziegel:1993:BRBc

[Zie93h] Eric R. Ziegel. Book review: Data Analysis for Managers by Harry Roberts. Technometrics, 35(1):99, February 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269312.

Ziegel:1993:BRBs

[Zie93i] Eric R. Ziegel. Book review: Deming’s Road to Continual Improvement by W. Scherkenbach. Technometrics, 35(2):233–234, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269685.

Ziegel:1993:BRBac

[Zie93j] Eric R. Ziegel. Book review: Design and Inference in Finite Population Sampling by A. Hedayat; B. Sinha. Technometrics, 35 (4):463–464, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270299.

Ziegel:1993:BRBp

[Zie93k] Eric R. Ziegel. Book review: Directions in Robust Statistics and Diagnostics (Parts 1 and 2) by W. Stahel; S. Wersberg. Technometrics, 35(1):103, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269325.

Ziegel:1993:BRBaj

[Zie93l] Eric R. Ziegel. Book review: Handbook of Quality Assurance for the Analytical Chemical Laboratory by J. Dux. Technometrics, 35 (4):466, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270306.

Ziegel:1993:BRBi

[Zie93m] Eric R. Ziegel. Book review: Introductory Statistics by Prem Mann. Technometrics, 35(1):100–101, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269318.

Ziegel:1993:BRBw

[Zie93n] Eric R. Ziegel. Book review: ISO 9000 by J. Lamprecht. Technometrics, 35(2):240, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269696.

Ziegel:1993:BRBah

[Zie93o] Eric R. Ziegel. Book review: Latent Variable Models by J. Loehlin. Technometrics, 35(4):465, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270304.

Ziegel:1993:BRBal

[Zie93p] Eric R. Ziegel. Book review: Managing Quality by R. Schuler; D. Harris. Technometrics, 35(4):466, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270308.

Ziegel:1993:BRBo

[Zie93q] Eric R. Ziegel. Book review: Managing Quality Dynamics by James Teboul. Technometrics, 35(1):102–103, February 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269324.

Ziegel:1993:BRBg

[Zie93r] Eric R. Ziegel. Book review: Measurement Errors in Surveys by P. Biemer; R. Groves; L. Lyberg; N. Mathiowetz; S. Sudman. Technometrics, 35(1):100, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269316.

Ziegel:1993:BRBt

[Zie93s] Eric R. Ziegel. Book review: Nonparametric Statistical Inference by J. Gibbons; S. Chakraborti. Technometrics, 35(2):239–240, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269693.

Ziegel:1993:BRBad

[Zie93t] Eric R. Ziegel. Book review: Nonsampling Error in Surveys by J. Lessler; W. Kalsbeek. Technometrics, 35(4):464, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270300.

Ziegel:1993:BRBn

[Zie93u] Eric R. Ziegel. Book review: Quality at Work in Research and Development by George Kidd. Technometrics, 35(1):102, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269323.

Ziegel:1993:BRBb

[Zie93v] Eric R. Ziegel. Book review: Quality Engineering Handbook by T. Pyzdek; R. Berger. Technometrics, 35(1):98–99, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269311.

Ziegel:1993:BRBai

[Zie93w] Eric R. Ziegel. Book review: Quality Through Engineering Design by W. Kuo; M. Pierson. Technometrics, 35(4):465–466, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270305.

Ziegel:1993:BRBe

[Zie93x] Eric R. Ziegel. Book review: Regression Diagnostics by John Fox; Newbury Park. Technometrics, 35(1):99–100, February 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269314.

Ziegel:1993:BRBz

[Zie93y] Eric R. Ziegel. Book review: Sampling by Steven Thompson. Technometrics, 35(3):324, August 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269527.

Ziegel:1993:BRBj

[Zie93z] Eric R. Ziegel. Book review: SAS System for Statistical Graphics by M. Friendly. Technometrics, 35(1):101, February 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269319.

Ziegel:1993:BRBk

[Zie93-27] Eric R. Ziegel. Book review: Statistical Methods for Survival Data Analysis by Elisa T. Lee. Technometrics, 35(1):101, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269320.

Ziegel:1993:BRBaf

[Zie93-28] Eric R. Ziegel. Book review: Statistical Methods in Biological and Medical Sciences by C. R. Rao; R. Chakraborty. Technometrics, 35(4):465, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270302.

Ziegel:1993:BRBh

[Zie93-29] Eric R. Ziegel. Book review: Statistical Methods: The Geometric Approach by D. Saville; G. Wood. Technometrics, 35(1):100, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269317.

Ziegel:1993:BRBa

[Zie93-30] Eric R. Ziegel. Book review: Statistical Theory and Modelling by D. Hinkley; N. Reid; E. J. Snell. Technometrics, 35(1):98, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269310.

Ziegel:1993:BRBag

[Zie93-31] Eric R. Ziegel. Book review: Statistics for Toxicologists by David Salsburg. Technometrics, 35(4):465, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270303.

Ziegel:1993:BRBu

[Zie93-32] Eric R. Ziegel. Book review: Survey Sampling by A. Chaudhuri; H. Stenger. Technometrics, 35(2):240, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269694.

Ziegel:1993:BRBf

[Zie93-33] Eric R. Ziegel. Book review: Survey Sampling Principles by E. K. Foreman. Technometrics, 35(1):100, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269315.

Ziegel:1993:BRBx

[Zie93-34] Eric R. Ziegel. Book review: The Team Approach to Quality by K. Shilliff; P. Motiska. Technometrics, 35(2):240–241, May 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269697.

Ziegel:1993:BRBy

[Zie93-35] Eric R. Ziegel. Book review: The Whats, Whys, and Hows of Quality Improvement by G. Miller; L. Krumm. Technometrics, 35 (2):241, May 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269698.

Ziegel:1993:BRBv

[Zie93-36] Eric R. Ziegel. Book review: Total Engineering Quality Management by R. Cottman. Technometrics, 35(2):240, May 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269695.

Ziegel:1993:BRBak

[Zie93-37] Eric R. Ziegel. Book review: Total Quality Management by A. Tenner; I. DeToro. Technometrics, 35(4):466, November 1993. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270307.

Ziegel:1993:BRBl

[Zie93-38] Eric R. Ziegel. Book review: Understanding Statistical Process Control by D. Wheeler; D. Chambers. Technometrics, 35(1):101– 102, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269321.

Ziegel:1994:BRBao

[Zie94a] Eric R. Ziegel. Book review: 100 Statistical Tests by G. K. Kanji. Technometrics, 36(4):436, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269982.

Ziegel:1994:BRBh

[Zie94b] Eric R. Ziegel. Book review: A Geostatistical Primer by Peter Brooker. Technometrics, 36(2):219–220, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270239.

Ziegel:1994:BRBu

[Zie94c] Eric R. Ziegel. Book review: Adequacy of Sample Size in Health Studies by S. Lemeshow; D. Hosmer; J. Klar; S. Luanga. Technometrics, 36(2):232, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270261.

Ziegel:1994:BRBag

[Zie94d] Eric R. Ziegel. Book review: Advanced Calculus with Applications in Statistics by Andre Khuri. Technometrics, 36(3):333, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269402.

Ziegel:1994:BRBab

[Zie94e] Eric R. Ziegel. Book review: Advances in GLIM and Statistical Modelling by L. Fahrmeir; B. Francis; R. Gilchrist; G. Tutz. Technometrics, 36(3):331, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269397.

Ziegel:1994:BRBad

[Zie94f] Eric R. Ziegel. Book review: An Introduction to Statistical Methods and Data Analysis by Lyman Ott. Technometrics, 36(3): 332, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269399.

Ziegel:1994:BRBe

[Zie94g] Eric R. Ziegel. Book review: Basic Statistics for Laboratories by W. Kelly; T. Ratliff; C. Nenadic. Technometrics, 36(1):126, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269226.

Ziegel:1994:BRBac

[Zie94h] Eric R. Ziegel. Book review: Basic Statistics: Tools for Continuous Improvement by M. Kiemele; S. Schmidt. Technometrics, 36(3):331, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269398.

Ziegel:1994:BRBs

[Zie94i] Eric R. Ziegel. Book review: Biostatistics by L. Fisher; G. van Belle. Technometrics, 36(2):231–232, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270259.

Ziegel:1994:BRBan

[Zie94j] Eric R. Ziegel. Book review: Boostrapping: A Nonparametric Approach to Statistical Inference by C. Mooney; R. Duvall. Technometrics, 36(4):435–436, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269981.

Ziegel:1994:BRBae

[Zie94k] Eric R. Ziegel. Book review: Chemometrics Tutorials II by R. Brereton; D. Scott; R. Massart; R. Dessy; P. Hopke; C. Spiegelman; W. Wegscheider. Technometrics, 36(3):332, August 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269400.

Ziegel:1994:BRBf

[Zie94l] Eric R. Ziegel. Book review: Continuous Improvement: Teams and Tools by R. Lynch; T. Werner. Technometrics, 36(1):126, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269227.

Ziegel:1994:BRBv

[Zie94m] Eric R. Ziegel. Book review: Cross-Over Experiments by D. Ratkowsky; M. Evans; J. Alldredge. Technometrics, 36(2):232– 233, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270262.

Ziegel:1994:BRBw

[Zie94n] Eric R. Ziegel. Book review: Cross-Over Trials in Clinical Research by Stephen Senn. Technometrics, 36(2):233, May 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270263.

Ziegel:1994:BRBa

[Zie94o] Eric R. Ziegel. Book review: Data Analysis for the Chemical Sciences by Richard Graham. Technometrics, 36(1):116, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269210.

Ziegel:1994:BRBt

[Zie94p] Eric R. Ziegel. Book review: Design and Analysis of Bioavailability and Bioequivalence Studies by S. Chow; J. Liu. Technometrics, 36(2):232, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270260.

Ziegel:1994:BRBy

[Zie94q] Eric R. Ziegel. Book review: Ecological Risk Estimation by S. Bartell; R. Gardner; R. O’Neill. Technometrics, 36(3):324– 325, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269386.

Ziegel:1994:BRBai

[Zie94r] Eric R. Ziegel. Book review: Essentials of SPC in the Process Industries by J. Pruett; H. Schneider. Technometrics, 36(4):421– 422, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269960.

Ziegel:1994:BRBi

[Zie94s] Eric R. Ziegel. Book review: Experimental Design: A Chemometric Approach by Stanley N. Deming; Stephen L. Morgan. Technometrics, 36(2):227–228, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270249.

Ziegel:1994:BRBaj

[Zie94t] Eric R. Ziegel. Book review: Forecasting and Time Series: An Applied Approach by Bruce Bowerman; Richard O’Connell. Technometrics, 36(4):434, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269977.

Ziegel:1994:BRBaa

[Zie94u] Eric R. Ziegel. Book review: Handbook of Sequential Analysis by B. Ghosh; P. Sen. Technometrics, 36(3):331, August 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269396.

Ziegel:1994:BRBap

[Zie94v] Eric R. Ziegel. Book review: Handbook of Statistics 11: Econometrics by G. Maddala; C. Rao; H. Vinod. Technometrics, 36 (4):436, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269983.

Ziegel:1994:BRBah

[Zie94w] Eric R. Ziegel. Book review: Handbook of Writing for the Mathematical Sciences by N. Higham. Technometrics, 36(3):333, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269403.

Ziegel:1994:BRBak

[Zie94x] Eric R. Ziegel. Book review: Introduction to Multiple Time Series Analysis by H. Lütkepohl. Technometrics, 36(4):434–435, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269978.

Ziegel:1994:BRBb

[Zie94y] Eric R. Ziegel. Book review: Introduction to the Practice of Statistics by D. Moore; G. McCabe. Technometrics, 36(1):124– 125, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269223.

Ziegel:1994:BRBr

[Zie94z] Eric R. Ziegel. Book review: Longitudinal Data with Serial Correlation: A State-Space Approach by R. Jones. Technometrics, 36 (2):231, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270258.

Ziegel:1994:BRBq

[Zie94-27] Eric R. Ziegel. Book review: Multiple Comparisons, Selection, and Applications in Biometry by Fred Hoppe. Technometrics, 36(2):230–231, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270257.

Ziegel:1994:BRBal

[Zie94-28] Eric R. Ziegel. Book review: Networks and Chaos: Statistical and Probabilistic Aspects by O. Barndorff-Nielson; J. L. Jensen; W. S. Kendall. Technometrics, 36(4):435, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269979.

Ziegel:1994:BRBam

[Zie94-29] Eric R. Ziegel. Book review: New Directions in Statistical Data Analysis and Robustness by S. Morgenthaler; E. Ronchetti; W.

Stahel. Technometrics, 36(4):435, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269980.

Ziegel:1994:BRBaq

[Zie94-30] Eric R. Ziegel. Book review: Principles of Biostatistics by M. Pagano; K. Gauvreau. Technometrics, 36(4):436–437, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269984.

Ziegel:1994:BRBc

[Zie94-31] Eric R. Ziegel. Book review: Probability and Statistics for Engineers and Scientists by R. Walpole; R. Myers. Technometrics, 36(1):125, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269224.

Ziegel:1994:BRBj

[Zie94-32] Eric R. Ziegel. Book review: Probability by J. Pitman. Technometrics, 36(2):228, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270250.

Ziegel:1994:BRBn

[Zie94-33] Eric R. Ziegel. Book review: Quality Control for Profit by R. Lester; N. Enrick; H. Mottley. Technometrics, 36(2):229, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270254.

Ziegel:1994:BRBx

[Zie94-34] Eric R. Ziegel. Book review: Real-Time Statistical Process Control by P. Badavas. Technometrics, 36(3):321–322, August 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269383.

Ziegel:1994:BRBz

[Zie94-35] Eric R. Ziegel. Book review: Regression with Dummy Variables by M. Hardy. Technometrics, 36(3):330–331, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269395.

Ziegel:1994:BRBo

[Zie94-36] Eric R. Ziegel. Book review: Root Cause Analysis by P. Wilson; L. Dell; G. Anderson. Technometrics, 36(2):230, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270255.

Ziegel:1994:BRBp

[Zie94-37] Eric R. Ziegel. Book review: Sequential Data in Biological Experiments by E. Roberts. Technometrics, 36(2):230, May 1994. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270256.

Ziegel:1994:BRBl

[Zie94-38] Eric R. Ziegel. Book review: Statistical Distributions by M. Evans; N. Hastings; B. Peacock. Technometrics, 36(2):229, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270252.

Ziegel:1994:BRBas

[Zie94-39] Eric R. Ziegel. Book review: Statistical Models in Engineering by G. Hahn; S. Shapiro. Technometrics, 36(4):437, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269986.

Ziegel:1994:BRBk

[Zie94-40] Eric R. Ziegel. Book review: Statistics for Engineering and the Sciences by W. Mendenhall; T. Sincich. Technometrics, 36(2): 228–229, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270251.

Ziegel:1994:BRBar

[Zie94-41] Eric R. Ziegel. Book review: Statistics for Spatial Data by Noel Cressie. Technometrics, 36(4):437, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269985.

Ziegel:1994:BRBaf

[Zie94-42] Eric R. Ziegel. Book review: Statistics for the Environment by V. Barnett; K. Turkman. Technometrics, 36(3):332–333, August 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269401.

Ziegel:1994:BRBm

[Zie94-43] Eric R. Ziegel. Book review: Statistics Quality Control for the Workplace by F. Kaminsky; R. Davis; R. Burke. Technometrics, 36(2):229, May 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270253.

Ziegel:1994:BRBg

[Zie94-44] Eric R. Ziegel. Book review: TQM: The Mechanics of Quality Processes by R. Johnson; L. Kazense. Technometrics, 36(1): 126, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269228.

Ziegel:1994:BRBd

[Zie94-45] Eric R. Ziegel. Book reviews: Breakthrough in Statistics: Volume I, Foundations and Basic Theory by S. Kotz; N. Johnson; Breakthrough in Statistics: Volume II, Methodology and Distribution by S. Kotz; N. Johnson. Technometrics, 36(1):125–126, February 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269225.

Ziegel:1995:BRBx

[Zie95a] Eric R. Ziegel. Book review: A Handbook of Small Data Sets by D. Hand; F. Daly; A. Lunn; K. McConway; E. Ostrowski. Technometrics, 37(2):242, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269650.

Ziegel:1995:BRBav

[Zie95b] Eric R. Ziegel. Book review: A Step-by-Step Approach to Using the SAS System for Univariate and Multivariate Statistics by L. Hatcher; E. Stepanski. Technometrics, 37(4):471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269762.

Ziegel:1995:BRBw

[Zie95c] Eric R. Ziegel. Book review: AI and Computer Power: The Impact of Statistics by D. J. Hand. Technometrics, 37(2):241–242, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269649.

Ziegel:1995:BRBaj

[Zie95d] Eric R. Ziegel. Book review: Analysis of Longitudinal Data by P. Diggle; K. Liang; S. Zeger. Technometrics, 37(3):356, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269941.

Ziegel:1995:BRBas

[Zie95e] Eric R. Ziegel. Book review: Applied Econometric Time Series by W. Enders. Technometrics, 37(4):469–470, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269759.

Ziegel:1995:BRBf

[Zie95f] Eric R. Ziegel. Book review: Artificial Intelligence Frontiers in Statistics by D. Hand. Technometrics, 37(1):127, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269180.

Ziegel:1995:BRBag

[Zie95g] Eric R. Ziegel. Book review: Aspects of Uncertainty by P. Freeman; A. Smith. Technometrics, 37(3):355, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269938.

Ziegel:1995:BRBaw

[Zie95h] Eric R. Ziegel. Book review: Biostatistics by Wayne Daniel. Technometrics, 37(4):471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269763.

Ziegel:1995:BRBal

[Zie95i] Eric R. Ziegel. Book review: Biostatistics for Epidemiologists by A. Ahlbom. Technometrics, 37(3):357, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269943.

Ziegel:1995:BRBi

[Zie95j] Eric R. Ziegel. Book review: Design and Analysis of Ecological Experiments by S. Scheiner; J. Gurevitch. Technometrics, 37(1):128, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269183.

Ziegel:1995:BRBv

[Zie95k] Eric R. Ziegel. Book review: Developing New Products with TQM by C. Cervetz. Technometrics, 37(2):241, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269648.

Ziegel:1995:BRBg

[Zie95l] Eric R. Ziegel. Book review: Developments in Time Series Analysis by T. S. Rao. Technometrics, 37(1):127–128, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269181.

Ziegel:1995:BRBq

[Zie95m] Eric R. Ziegel. Book review: Ecological Risk Assessment by Glenn Suter. Technometrics, 37(2):240, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269643.

Ziegel:1995:BRBh

[Zie95n] Eric R. Ziegel. Book review: Environmental Statistics, Assessment, and Forecasting by C. R. Cothern; N. P. Ross. Technometrics, 37(1):128, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269182.

Ziegel:1995:BRBaq

[Zie95o] Eric R. Ziegel. Book review: Extreme Value Theory and Applications by J. Galambos; J. Lechner; E. Simiu. Technometrics, 37 (4):469, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269757.

Ziegel:1995:BRBab

[Zie95p] Eric R. Ziegel. Book review: Four Days with Dr. Deming by W. Latzko; D. Saunders. Technometrics, 37(3):353–354, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269933.

Ziegel:1995:BRBa

[Zie95q] Eric R. Ziegel. Book review: Fourth Generation Management: The New Business Consciousness by Brian Joiner. Technometrics, 37(1):113, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269158.

Ziegel:1995:BRBae

[Zie95r] Eric R. Ziegel. Book review: Geostatistics Troia ’92 by A. Soares. Technometrics, 37(3):355, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269936.

Ziegel:1995:BRBc

[Zie95s] Eric R. Ziegel. Book review: GSLIB: Geostatistical Software Library and User’s Guide by C. Deutsch; A. Journel. Technometrics, 37(1):126, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269177.

Ziegel:1995:BRBam

[Zie95t] Eric R. Ziegel. Book review: Introduction to Life Prediction of Industrial Plant Materials by M. Kowaka. Technometrics, 37(4): 457, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269739.

Ziegel:1995:BRBad

[Zie95u] Eric R. Ziegel. Book review: Mastering the SAS System by Jay Jaffe. Technometrics, 37(3):354–355, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269935.

Ziegel:1995:BRBe

[Zie95v] Eric R. Ziegel. Book review: Mathematical Statistics and Data Analysis by John Rice. Technometrics, 37(1):127, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269179.

Ziegel:1995:BRBk

[Zie95w] Eric R. Ziegel. Book review: Measuring Customer Satisfaction by Bob Hayes. Technometrics, 37(1):128–129, February 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269185.

Ziegel:1995:BRBj

[Zie95x] Eric R. Ziegel. Book review: Modelling for Survival Data in Medical Research by D. Collett. Technometrics, 37(1):128, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269184.

Ziegel:1995:BRBd

[Zie95y] Eric R. Ziegel. Book review: Near Infra-Red Spectroscopy: Bridging the Gap between Data Analysis and NIR Applications by K. Hildrum; T. Isaksson; T. Naes; A. Tandberg. Technometrics, 37(1):126–127, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269178.

Ziegel:1995:BRBb

[Zie95z] Eric R. Ziegel. Book review: Pierre Gy’s Sampling Theory and Sampling Practice by Francis Pitard; Pierre Gy. Technometrics, 37(1):125–126, February 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1269176.

Ziegel:1995:BRBt

[Zie95-27] Eric R. Ziegel. Book review: Practical Biostatistical Methods by Steve Selvin. Technometrics, 37(2):241, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269646.

Ziegel:1995:BRBo

[Zie95-28] Eric R. Ziegel. Book review: Probability and Statistics for Engineers by R. Scheaffer; J. McClave. Technometrics, 37(2):239, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269641.

Ziegel:1995:BRBai

[Zie95-29] Eric R. Ziegel. Book review: Proceedings of the First U.S./Japan Conference on the Frontiers of Statistical Modeling: An Informational Approach by H. Bozdogan. Technometrics, 37(3):356, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269940.

Ziegel:1995:BRBac

[Zie95-30] Eric R. Ziegel. Book review: Quality by Experimental Design by Thomas Barker. Technometrics, 37(3):354, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269934.

Ziegel:1995:BRBap

[Zie95-31] Eric R. Ziegel. Book review: Quality Management: Tools and Methods for Improvement by H. Gitlow; A. Oppenheim; R. Oppenheim. Technometrics, 37(4):468–469, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269756.

Ziegel:1995:BRBau

[Zie95-32] Eric R. Ziegel. Book review: SAS Software Roadmaps by L. Burch; S. King. Technometrics, 37(4):470–471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269761.

Ziegel:1995:BRBaf

[Zie95-33] Eric R. Ziegel. Book review: Selecting Models from Data by P. Cheeseman; R. Oldford. Technometrics, 37(3):355, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269937.

Ziegel:1995:BRBu

[Zie95-34] Eric R. Ziegel. Book review: SPC in Extrusion by Chris Rauwendaal. Technometrics, 37(2):241, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269647.

Ziegel:1995:BRBan

[Zie95-35] Eric R. Ziegel. Book review: SPC Made Easy by Peter McReynolds. Technometrics, 37(4):462–463, November 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269747.

Ziegel:1995:BRBs

[Zie95-36] Eric R. Ziegel. Book review: Statistical Analysis of Epidemiological Data by Steve Selvin. Technometrics, 37(2):240–241, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269645.

Ziegel:1995:BRBl

[Zie95-37] Eric R. Ziegel. Book review: Statistical Methods for Groundwater Monitoring by Robert Gibbons. Technometrics, 37(2):231–232, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269628.

Ziegel:1995:BRBar

[Zie95-38] Eric R. Ziegel. Book review: Statistical Methods for Physical Science by J. Stanford; S. Vardeman. Technometrics, 37(4):469, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1269758.

Ziegel:1995:BRBr

[Zie95-39] Eric R. Ziegel. Book review: Statistical Models in Epidemiology by D. Clayton; M. Hills. Technometrics, 37(2):240, May 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269644.

Ziegel:1995:BRBax

[Zie95-40] Eric R. Ziegel. Book review: Statistics for Accountants by S. Letchford. Technometrics, 37(4):471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269764.

Ziegel:1995:BRBak

[Zie95-41] Eric R. Ziegel. Book review: Statistics in Toxicology by D. Krewski; C. Franklin. Technometrics, 37(3):356–357, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269942.

Ziegel:1995:BRBah

[Zie95-42] Eric R. Ziegel. Book review: Subjective Probability by G. Wright; P. Ayton. Technometrics, 37(3):355–356, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269939.

Ziegel:1995:BRBz

[Zie95-43] Eric R. Ziegel. Book review: The Collected Works of John Tukey: Volume VIII, Multiple Comparisons by H. Braun; John Tukey. Technometrics, 37(2):242, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269652.

Ziegel:1995:BRBy

[Zie95-44] Eric R. Ziegel. Book review: The Handbook of Research Synthesis by H. Cooper; L. Hedges. Technometrics, 37(2):242, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269651.

Ziegel:1995:BRBao

[Zie95-45] Eric R. Ziegel. Book review: The Quality Toolbox by Nancy Tague. Technometrics, 37(4):468, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269755.

Ziegel:1995:BRBn

[Zie95-46] Eric R. Ziegel. Book review: Time Series Analysis, Forecasting, and Control by G. Box; G. Jenkins; G. Reinsel. Technometrics, 37(2):238–239, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269640.

Ziegel:1995:BRBm

[Zie95-47] Eric R. Ziegel. Book review: Uncertainty by M. Morgan; M. Henrion. Technometrics, 37(2):237–238, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269639.

Ziegel:1995:BRBaa

[Zie95-48] Eric R. Ziegel. Book review: Weibull Analysis by Bryan Dodson. Technometrics, 37(3):347–348, August 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269925.

Ziegel:1995:BRBat

[Zie95-49] Eric R. Ziegel. Book reviews: Neural Networks in Computer Intelligence by L. Fu; Fundamentals of Neural Networks by L. Fausett. Technometrics, 37(4):470, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269760.

Ziegel:1995:BRBp

[Zie95-50] Eric R. Ziegel. Book reviews: Statistics for Business: Data Analysis and Modelling by J. Cryer; R. Miller; Introduction to Modern Business Statistics by G. Canavos; D. Miller; Introductory Business Statistics with Computer Applications by R. Shiffler; A. Adams; Statistics for Business and Economics by P. Mann. Technometrics, 37(2):239–240, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269642.

Ziegel:1996:BRBaz

[Zie96a] Eric R. Ziegel. Book review: Advances in Stochastic Modelling and Data Analysis by J. Janssen; C. Skiadas; C. Zopounidis. Technometrics, 38(4):414, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271338.

Ziegel:1996:BRBaf

[Zie96b] Eric R. Ziegel. Book review: Analysing Survival Data from Clinical Trials and Observational Studies by E. Marubini; M. Valsecchi. Technometrics, 38(3):299, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270634.

Ziegel:1996:BRBq

[Zie96c] Eric R. Ziegel. Book review: Applied Logistic Regression Analysis by Scott Menard. Technometrics, 38(2):192, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270433.

Ziegel:1996:BRBai

[Zie96d] Eric R. Ziegel. Book review: Applied Multivariate for Statistics with SAS(R) Software by R. Khattree; D. Naik. Technometrics, 38(3):300, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270637.

Ziegel:1996:BRBap

[Zie96e] Eric R. Ziegel. Book review: Bayesian Methods in Reliability by P. Sander; R. Badoux. Technometrics, 38(4):411, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271328.

Ziegel:1996:BRBu

[Zie96f] Eric R. Ziegel. Book review: Business Process Benchmarking by Robert Camp. Technometrics, 38(2):193, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270437.

Ziegel:1996:BRBh

[Zie96g] Eric R. Ziegel. Book review: Business Survey Methods by B. Cox; D. Binder; B. Chinnappa; A. Christianson; M. Colledge; P. Kott. Technometrics, 38(1):89, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268928.

Ziegel:1996:BRBaw

[Zie96h] Eric R. Ziegel. Book review: Case Studies in Bayesian Statistics, Volume II by C. Gatsonis; J. Hodges; R. Kass; N. Singpurwalla. Technometrics, 38(4):413, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271335.

Ziegel:1996:BRBav

[Zie96i] Eric R. Ziegel. Book review: Categorical Data Analysis Using the SAS System by M. Stokes; C. Davis; G. Koch. Technometrics, 38(4):413, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271334.

Ziegel:1996:BRBat

[Zie96j] Eric R. Ziegel. Book review: Chemometrics in Environmental Chemistry: Statistical Methods by J. Einax. Technometrics, 38 (4):412, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271332.

Ziegel:1996:BRBe

[Zie96k] Eric R. Ziegel. Book review: Customer-Focused Quality by N. Mowery; P. Reavis; S. Poling. Technometrics, 38(1):87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268925.

Ziegel:1996:BRBs

[Zie96l] Eric R. Ziegel. Book review: Data Analysis Handbook by Ildiko Frank; Roberto Todeschini. Technometrics, 38(2):193, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270435.

Ziegel:1996:BRBa

[Zie96m] Eric R. Ziegel. Book review: Data Analysis Using Microsoft Excel 5.0 by Michael Middleton. Technometrics, 38(1):78–79, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268909.

Ziegel:1996:BRBam

[Zie96n] Eric R. Ziegel. Book review: Decisions Involving Uncertainty by James Murtha. Technometrics, 38(4):408, November 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271322.

Ziegel:1996:BRBad

[Zie96o] Eric R. Ziegel. Book review: Encyclopedia of Quality Terms and Concepts by J. Cortada; J. Woods. Technometrics, 38(3):298, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270632.

Ziegel:1996:BRBm

[Zie96p] Eric R. Ziegel. Book review: Environmental Statistics by G. Patil; C. Rao. Technometrics, 38(2):190–191, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270429.

Ziegel:1996:BRBae

[Zie96q] Eric R. Ziegel. Book review: Fundamentals of Quality Auditing by B. Parsowith. Technometrics, 38(3):298–299, August 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270633.

Ziegel:1996:BRBo

[Zie96r] Eric R. Ziegel. Book review: Geostatistics for the Next Century by R. Dimitrakopoulos. Technometrics, 38(2):191–192, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270431.

Ziegel:1996:BRBp

[Zie96s] Eric R. Ziegel. Book review: Growth Curves by A. M. Kshirsagar; W. B. Smith. Technometrics, 38(2):192, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270432.

Ziegel:1996:BRBah

[Zie96t] Eric R. Ziegel. Book review: Image Analysis for the Biological Sciences by C. Glasbey; C. Horgan. Technometrics, 38(3):299, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270636.

Ziegel:1996:BRBr

[Zie96u] Eric R. Ziegel. Book review: Interpreting Probability Models by T. Lao. Technometrics, 38(2):192–193, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270434.

Ziegel:1996:BRBy

[Zie96v] Eric R. Ziegel. Book review: Introduction to Biostatistics by R. Forthofer; E. Lee. Technometrics, 38(2):195, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270441.

Ziegel:1996:BRBak

[Zie96w] Eric R. Ziegel. Book review: Introduction to Market Research Using the SAS(R) System by SAS Staff. Technometrics, 38(3):300– 301, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270639.

Ziegel:1996:BRBb

[Zie96x] Eric R. Ziegel. Book review: Logistic Regression Examples Using the SAS System by SAS Institute Staff. Technometrics, 38 (1):86–87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268922.

Ziegel:1996:BRBz

[Zie96y] Eric R. Ziegel. Book review: Managing Customer Value by Bradley Gale. Technometrics, 38(3):295–296, August 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270627.

Ziegel:1996:BRBv

[Zie96z] Eric R. Ziegel. Book review: Managing Quality by B. Dale. Technometrics, 38(2):194, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270438.

Ziegel:1996:BRBay

[Zie96-27] Eric R. Ziegel. Book review: Monte Carlo and Quasi-Monte Carlo Methods in Scientific Computing by H. Niederreiter; P. Shiue. Technometrics, 38(4):414, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271337.

Ziegel:1996:BRBag

[Zie96-28] Eric R. Ziegel. Book review: Observational Studies by Paul Rosenbaum. Technometrics, 38(3):299, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270635.

Ziegel:1996:BRBj

[Zie96-29] Eric R. Ziegel. Book review: Practical Methods for Design and Analysis of Complex Surveys by R. Lehtonen; E. Pahkinen. Technometrics, 38(1):89–90, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268930.

Ziegel:1996:BRBal

[Zie96-30] Eric R. Ziegel. Book review: Probability, Statistics and Optimization by F. Kelly. Technometrics, 38(3):301, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270640.

Ziegel:1996:BRBaq

[Zie96-31] Eric R. Ziegel. Book review: Problem Solving: A Statistician’s Guide by C. Chatfield. Technometrics, 38(4):411, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271329.

Ziegel:1996:BRBac

[Zie96-32] Eric R. Ziegel. Book review: Quality Function Deployment by L. Cohen. Technometrics, 38(3):298, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270631.

Ziegel:1996:BRBc

[Zie96-33] Eric R. Ziegel. Book review: Quick Results with SAS/Graph Software by A. Carpenter; C. Shipp. Technometrics, 38(1):87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268923.

Ziegel:1996:BRBab

[Zie96-34] Eric R. Ziegel. Book review: Recent Advances in Life Testing and Reliability by N. Balakrishnan. Technometrics, 38(3):297–298, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270630.

Ziegel:1996:BRBd

[Zie96-35] Eric R. Ziegel. Book review: Reliability-Centered Maintenance by John Moubray. Technometrics, 38(1):87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268924.

Ziegel:1996:BRBx

[Zie96-36] Eric R. Ziegel. Book review: SAS Programming by Example by R. Cody; R. Pass. Technometrics, 38(2):194–195, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270440.

Ziegel:1996:BRBan

[Zie96-37] Eric R. Ziegel. Book review: Statistical Applications in Process Control by J. B. Keats; D. C. Montgomery. Technometrics, 38 (4):410, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271326.

Ziegel:1996:BRBn

[Zie96-38] Eric R. Ziegel. Book review: Statistical Data Analysis for Ocean and Atmospheric Sciences by H. Jean Thiebaux. Technometrics, 38(2):191, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270430.

Ziegel:1996:BRBk

[Zie96-39] Eric R. Ziegel. Book review: Statistical Design and Analysis in Pharmaceutical Science by S. Chow; J. Liu. Technometrics, 38(1): 90, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268931.

Ziegel:1996:BRBax

[Zie96-40] Eric R. Ziegel. Book review: Statistical Modelling by G. Seeber; B. Francis; R. Hatzinger; G. Stechel-Berger. Technometrics, 38 (4):413–414, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271336.

Ziegel:1996:BRBao

[Zie96-41] Eric R. Ziegel. Book review: Statistical Process Control by C. Mamzic. Technometrics, 38(4):410–411, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271327.

Ziegel:1996:BRBl

[Zie96-42] Eric R. Ziegel. Book review: Statistics and the Evaluation of Evidence for Forensic Scientists by C. G. G. Aitken. Technometrics, 38(1):90–91, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268932.

Ziegel:1996:BRBau

[Zie96-43] Eric R. Ziegel. Book review: Statistics in Epidemiology by H. Sahai; A. Khurshid. Technometrics, 38(4):412–413, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271333.

Ziegel:1996:BRBar

[Zie96-44] Eric R. Ziegel. Book review: Stochastic Modeling and Geostatistics by J. Yarus; R. Chambers. Technometrics, 38(4):411–412, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271330.

Ziegel:1996:BRBaj

[Zie96-45] Eric R. Ziegel. Book review: The How-To Book for SAS/Graph Software by Thomas Miron. Technometrics, 38(3):300, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270638.

Ziegel:1996:BRBt

[Zie96-46] Eric R. Ziegel. Book review: The Malcolm Baldrige National Quality Award by M. Heaphy; G. Gruska. Technometrics, 38(2):193, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270436.

Ziegel:1996:BRBaa

[Zie96-47] Eric R. Ziegel. Book review: The Survey Kit (Partial: Volumes 2, 5, 6, 8) by Arlene Fink. Technometrics, 38(3):297, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270629.

Ziegel:1996:BRBi

[Zie96-48] Eric R. Ziegel. Book review: Understanding Industrial Designed Experiments by S. Schmidt; R. Launsby. Technometrics, 38(1): 89, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1268929.

Ziegel:1996:BRBw

[Zie96-49] Eric R. Ziegel. Book review: Understanding Neural Networks by M. Caudill; C. Butler. Technometrics, 38(2):194, May 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270439.

Ziegel:1996:BRBas

[Zie96-50] Eric R. Ziegel. Book review: Variowin: Software for Spatial Data Analysis by Y. Pannatier. Technometrics, 38(4):412, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271331.

Ziegel:1996:BRBf

[Zie96-51] Eric R. Ziegel. Book review: Winning with Quality by J. Wesner; J. Hiatt; D. Trimble. Technometrics, 38(1):87–88, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268926.

Ziegel:1996:BRBg

[Zie96-52] Eric R. Ziegel. Book reviews: Statistics for Business by J. Cryer; R. Miller; Statistics for Managers by U. Menzefricke; Business Statistics for Quality and Productivity by D. Levine; P. Ramsey; M. Berenson. Technometrics, 38(1):88, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268927.

Ziegel:1997:BRBr

[Zie97a] Eric R. Ziegel. Book review: A Toolbox for Quality Improvement and Problem Solving by D. Straker. Technometrics, 39(2):238, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270933.

Ziegel:1997:BRBaa

[Zie97b] Eric R. Ziegel. Book review: Basic Principles of Structural Equation Modelling by R. Mueller. Technometrics, 39(3):343, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271158.

Ziegel:1997:BRBg

[Zie97c] Eric R. Ziegel. Book review: Bayesian Analysis in Statistics and Econometrics by D. Berry; K. Chaloner; J. Geweke. Technometrics, 39(1):111, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270799.

Ziegel:1997:BRBh

[Zie97d] Eric R. Ziegel. Book review: Bayesian Biostatistics by D. Berry; D. Stangl. Technometrics, 39(1):111–112, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270800.

Ziegel:1997:BRBam

[Zie97e] Eric R. Ziegel. Book review: Bayesian Forecasting and Dynamic Models by M. West; J. Harrison. Technometrics, 39(4): 433, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271526.

Ziegel:1997:BRBao

[Zie97f] Eric R. Ziegel. Book review: Cases in Managerial Data Analysis by W. Carlson. Technometrics, 39(4):434, November 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271528.

Ziegel:1997:BRBp

[Zie97g] Eric R. Ziegel. Book review: Computational Modeling by C. Taber; R. Timpone. Technometrics, 39(2):238, May 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270931.

Ziegel:1997:BRBab

[Zie97h] Eric R. Ziegel. Book review: Computer-Aided Multivariate Analysis by A. Afifi; V. Clark. Technometrics, 39(3):343, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271156.

Ziegel:1997:BRBv

[Zie97i] Eric R. Ziegel. Book review: Customer Satisfaction Measurement and Management by E. Naumann; K. Giel. Technometrics, 39(2):239–240, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270937.

Ziegel:1997:BRBm

[Zie97j] Eric R. Ziegel. Book review: Data Analysis with Microsoft Excel 5.0 for Windows by K. Berk; P. Carey. Technometrics, 39(2):237, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270928.

Ziegel:1997:BRBu

[Zie97k] Eric R. Ziegel. Book review: Design and Management of Service Processes by R. Ramaswamy. Technometrics, 39(2):239, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270936.

Ziegel:1997:BRBal

[Zie97l] Eric R. Ziegel. Book review: Elements of Multivariate Time Series Analysis by Gregory Reinsel. Technometrics, 39(4):433, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271525.

Ziegel:1997:BRBaf

[Zie97m] Eric R. Ziegel. Book review: Forecasting Examples for Business and Economics Using the SAS(R) System by SAS Staff. Technometrics, 39(3):345, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271162.

Ziegel:1997:BRBz

[Zie97n] Eric R. Ziegel. Book review: Geostatistical Simulations by M. Armstrong; P. Dowd. Technometrics, 39(3):343, August 1997. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271157.

Ziegel:1997:BRBe

[Zie97o] Eric R. Ziegel. Book review: Glossary and Tables for Statistical Quality Control by ASQC Statistics Division. Technometrics, 39(1):111, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270797.

Ziegel:1997:BRBaq

[Zie97p] Eric R. Ziegel. Book review: Handbook of Statistical Analysis Using SAS by B. Everitt; G. Der. Technometrics, 39(4):434, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271530.

Ziegel:1997:BRBl

[Zie97q] Eric R. Ziegel. Book review: Improving Survey Questions by F. Fowler. Technometrics, 39(1):113, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270804.

Ziegel:1997:BRBai

[Zie97r] Eric R. Ziegel. Book review: Lifetime Data Models in Reliability and Survival Analysis by N. Jewell; A. Kimber; M. Lee; G. Whitmore. Technometrics, 39(4):432, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271522.

Ziegel:1997:BRBah

[Zie97s] Eric R. Ziegel. Book review: Measuring and Managing Customer Satisfaction by Sheila Kessler. Technometrics, 39(3):345–346, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271164.

Ziegel:1997:BRBi

[Zie97t] Eric R. Ziegel. Book review: Practical Longitudinal Data Analysis by D. Hand; M. Crowder. Technometrics, 39(1):112, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270801.

Ziegel:1997:BRBan

[Zie97u] Eric R. Ziegel. Book review: Process Redesign by A. Tenner; I. DeToro. Technometrics, 39(4):433–434, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271527.

Ziegel:1997:BRBf

[Zie97v] Eric R. Ziegel. Book review: Process Validation for Business Success by A. Elfe; J. Lahey; R. Launsby. Technometrics, 39(1): 111, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270798.

Ziegel:1997:BRBs

[Zie97w] Eric R. Ziegel. Book review: Quality Planning, Control, and Improvement in Research and Development by George Roberts. Technometrics, 39(2):238–239, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270934.

Ziegel:1997:BRBap

[Zie97x] Eric R. Ziegel. Book review: Quick Start to Data Analysis with SAS by F. Di Iorio; K. Hardy. Technometrics, 39(4):434, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271529.

Ziegel:1997:BRBw

[Zie97y] Eric R. Ziegel. Book review: Response Surfaces: Designs and Analyses by Andre Khuri; John Cornell. Technometrics, 39(3): 342, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271153.

Ziegel:1997:BRBo

[Zie97z] Eric R. Ziegel. Book review: Robust Statistics, Data Analysis, and Computer Intensive Methods by H. Rieder. Technometrics, 39 (2):238, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270930.

Ziegel:1997:BRBad

[Zie97-27] Eric R. Ziegel. Book review: SAS(R) System for Mixed Models by R. Littell; G. Milliken; W. Stroup; R. Wolfinger. Technometrics, 39(3):344, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271160.

Ziegel:1997:BRBak

[Zie97-28] Eric R. Ziegel. Book review: Simulation by S. M. Ross. Technometrics, 39(4):433, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271524.

Ziegel:1997:BRBag

[Zie97-29] Eric R. Ziegel. Book review: Simulation Modeling Using @Risk by Wayne Winston. Technometrics, 39(3):345, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271163.

Ziegel:1997:BRBq

[Zie97-30] Eric R. Ziegel. Book review: Spatial Analysis and GIS by S. Fotheringham; Peter Rogerson. Technometrics, 39(2):238, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270932.

Ziegel:1997:BRBt

[Zie97-31] Eric R. Ziegel. Book review: Specifications for the Chemical and Process Industries by ASQC Chemical Interest Committee. Technometrics, 39(2):239, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270935.

Ziegel:1997:BRBac

[Zie97-32] Eric R. Ziegel. Book review: Statistical Quality Control Using the SAS(R) System by Dennis King. Technometrics, 39(3):343– 344, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271159.

Ziegel:1997:BRBb

[Zie97-33] Eric R. Ziegel. Book review: Statistics and Data Analysis by Andrew Siegel; Charles Morgan. Technometrics, 39(1):110, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270794.

Ziegel:1997:BRBaj

[Zie97-34] Eric R. Ziegel. Book review: Statistics of Quality by S. Ghosh; W. Schucany; W. Smith. Technometrics, 39(4):432–433, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271523.

Ziegel:1997:BRBd

[Zie97-35] Eric R. Ziegel. Book review: Statistics: Theory and Methods by Donald Berry; Bernard Lindgren. Technometrics, 39(1):110– 111, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270796.

Ziegel:1997:BRBar

[Zie97-36] Eric R. Ziegel. Book review: Stock Market Analysis Using the SAS System: Technical Analysis by SAS Staff. Technometrics, 39(4):434–435, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271531.

Ziegel:1997:BRBj

[Zie97-37] Eric R. Ziegel. Book review: Survival Analysis by M. Parmar; D. Machin. Technometrics, 39(1):112–113, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270802.

Ziegel:1997:BRBae

[Zie97-38] Eric R. Ziegel. Book review: Survival Analysis Using the SAS System by Paul D. Allison. Technometrics, 39(3):344–345, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271161.

Ziegel:1997:BRBa

[Zie97-39] Eric R. Ziegel. Book review: Taguchi Techniques for Quality Engineering by Phillip Ross. Technometrics, 39(1):109–110, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270793.

Ziegel:1997:BRBy

[Zie97-40] Eric R. Ziegel. Book review: The Analysis of Time Series by Chris Chatfield. Technometrics, 39(3):342–343, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271155.

Ziegel:1997:BRBn

[Zie97-41] Eric R. Ziegel. Book review: The Elements of Graphing Data by William S. Cleveland. Technometrics, 39(2):237–238, May 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270929.

Ziegel:1997:BRBc

[Zie97-42] Eric R. Ziegel. Book review: Time Series Models by D. R. Cox; D. V. Hinkley; O. E. Barndorff-Nielsen. Technometrics, 39(1): 110, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270795.

Ziegel:1997:BRBk

[Zie97-43] Eric R. Ziegel. Book review: Wavelets and Statistics by A. Antoniadis; G. Oppenheim. Technometrics, 39(1):113, February 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270803.

Ziegel:1997:BRBx

[Zie97-44] Eric R. Ziegel. Book reviews: Applied Linear Statistical Models by J. Neter; M. Kutner; C. Nachtsheim; W. Wasserman; Applied Linear Regression Models by J. Neter; M. Kutner; C. Nachtsheim; W. Wasserman. Technometrics, 39(3):342, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271154.

Ziegel:1998:BRBaj

[Zie98a] Eric R. Ziegel. Book review: A First Course in Probability by Sheldon Ross. Technometrics, 40(3):268, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271207.

Ziegel:1998:BRBx

[Zie98b] Eric R. Ziegel. Book review: A Probabilistic Analysis of the Sacco and Vanzetti Evidence by Joseph B. Kadane; David A. Schum. Technometrics, 40(2):166, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270674.

Ziegel:1998:BRBe

[Zie98c] Eric R. Ziegel. Book review: Advances in Knowledge Discovery and Data Mining by Usama M. Fayyad; Gregory Piatetski-Shapiro; Padhraic Smyth; Ramasamy Uthurusamy. Technometrics, 40(1):83, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271414.

Ziegel:1998:BRBt

[Zie98d] Eric R. Ziegel. Book review: Advances in Statistical Decision Theory and Applications by S. Panchapakesan; N. Balakrishnan. Technometrics, 40(2):165, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270670.

Ziegel:1998:BRBu

[Zie98e] Eric R. Ziegel. Book review: Advances in the Theory and Practice of Statistics by Norman L. Johnson; N. Balakrishnan. Technometrics, 40(2):165, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270671.

Ziegel:1998:BRBf

[Zie98f] Eric R. Ziegel. Book review: Applications of Computer Aided Time Series Modeling by Masanao Aoki; Arthur M. Havenner. Technometrics, 40(1):83, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271415.

Ziegel:1998:BRBn

[Zie98g] Eric R. Ziegel. Book review: Applications of Time Series Analysis in Astronomy and Meteorology by T. Subba Rao; M. B. Priestly; O. Lessi. Technometrics, 40(2):163, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270664.

Ziegel:1998:BRBae

[Zie98h] Eric R. Ziegel. Book review: Applied Statistics and the SAS Programming Language by Ronald P. Cody; Jeffrey K. Smith. Technometrics, 40(3):266–267, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271202.

Ziegel:1998:BRBq

[Zie98i] Eric R. Ziegel. Book review: Bayesian Computation Using Minitab by James H. Albert. Technometrics, 40(2):164, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270667.

Ziegel:1998:BRBao

[Zie98j] Eric R. Ziegel. Book review: Bayesian Statistics: An Introduction by Peter M. Lee. Technometrics, 40(4):358, November 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270551.

Ziegel:1998:BRBv

[Zie98k] Eric R. Ziegel. Book review: Breakthroughs in Statistics, Volume III by Samuel Kotz; Norman L. Johnson. Technometrics, 40 (2):165, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270672.

Ziegel:1998:BRBh

[Zie98l] Eric R. Ziegel. Book review: Case Studies in Bayesian Statistics, Volume III by Constantine Gatsonis; James S. Hodges; Robert E. Kass; Robert McCulloch; Peter Rossi; Nozer D. Singpurwalla. Technometrics, 40(1):84, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271417.

Ziegel:1998:BRBi

[Zie98m] Eric R. Ziegel. Book review: Chemometrics in Environmental Analysis by J. W. Einax; H. W. Zwanziger; S. Geiss. Technometrics, 40(1):84, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271418.

Ziegel:1998:BRBap

[Zie98n] Eric R. Ziegel. Book review: Data, Statistics, and Decisions Models with Excel by Donald L. Harnett; James F. Horrell. Technometrics, 40(4):358, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270552.

Ziegel:1998:BRBa

[Zie98o] Eric R. Ziegel. Book review: Encyclopedia of Statistical Sciences, Update Volume 1 by S. Kotz; C. Read; D. Banks. Technometrics, 40(1):81–82, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271410.

Ziegel:1998:BRBas

[Zie98p] Eric R. Ziegel. Book review: Encyclopedia of Statistical Sciences: Update Volume 2 by S. Kotz; C. Read; D. Banks. Technometrics, 40(4):359, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270555.

Ziegel:1998:BRBar

[Zie98q] Eric R. Ziegel. Book review: Evolutionary Operation: A Statistical Method for Process Improvement by George Box; Norman R. Draper. Technometrics, 40(4):359, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270554.

Ziegel:1998:BRBr

[Zie98r] Eric R. Ziegel. Book review: Extending SAS(R) Survival Analysis Techniques for Medical Research by Alan Cantor. Technometrics, 40(2):164, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270668.

Ziegel:1998:BRBam

[Zie98s] Eric R. Ziegel. Book review: GeoENV1: Geostatistics for Environmental Applications by A. Soares; J. Gomez-Hernandez; R. Froidevaux. Technometrics, 40(4):357–358, November 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270549.

Ziegel:1998:BRBal

[Zie98t] Eric R. Ziegel. Book review: Geostatistical Software Library and User’s Guide by Clayton V. Deutsch; Andre G. Journel. Technometrics, 40(4):357, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270548.

Ziegel:1998:BRBz

[Zie98u] Eric R. Ziegel. Book review: Geostatistics Wollongong ’96 by E. Y. Baafi; N. A. Schofield. Technometrics, 40(3):265, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271197.

Ziegel:1998:BRBav

[Zie98v] Eric R. Ziegel. Book review: Getting Started with the Market Research Application by Donna M. Sawyer; Maura Stokes. Technometrics, 40(4):360, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270558.

Ziegel:1998:BRBy

[Zie98w] Eric R. Ziegel. Book review: Handbook of Chemometrics and Qualimetrics: Part A by D. Massart; B. Vandeginote; L. Buydens; S. DeJong; P. Lewi; J. Smeyers-Verbeke. Technometrics, 40(3):264–265, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271196.

Ziegel:1998:BRBs

[Zie98x] Eric R. Ziegel. Book review: Industrial Statistics by C. Kitsos; L. Edler. Technometrics, 40(2):164–165, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270669.

Ziegel:1998:BRBan

[Zie98y] Eric R. Ziegel. Book review: Introduction to Chemical Exposure and Risk Assessment by W. Brock Neely. Technometrics, 40(4): 358, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270550.

Ziegel:1998:BRBw

[Zie98z] Eric R. Ziegel. Book review: Leading Personalities in Statistical Science by Norman L. Johnson; Samuel Kotz. Technometrics, 40(2):165–166, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270673.

Ziegel:1998:BRBai

[Zie98-27] Eric R. Ziegel. Book review: Log-Linear Models and Logistic Regression by Ronald Christensen. Technometrics, 40(3):267– 268, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271206.

Ziegel:1998:BRBl

[Zie98-28] Eric R. Ziegel. Book review: Log-Linear Models for Event Histories by J. Vermunt. Technometrics, 40(1):85, February 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271421.

Ziegel:1998:BRBat

[Zie98-29] Eric R. Ziegel. Book review: Loss Models by S. Klugman; H. Panju; G. Willmot. Technometrics, 40(4):359, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270556.

Ziegel:1998:BRBp

[Zie98-30] Eric R. Ziegel. Book review: Matrix Algebra from a Statistician’s Perspective by David A. Harville. Technometrics, 40(2):164, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270666.

Ziegel:1998:BRBb

[Zie98-31] Eric R. Ziegel. Book review: Measuring Customer Satisfaction by Bob Hayes. Technometrics, 40(1):82, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271411.

Ziegel:1998:BRBah

[Zie98-32] Eric R. Ziegel. Book review: Monte Carlo Simulation by C. Mooney. Technometrics, 40(3):267, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271205.

Ziegel:1998:BRBc

[Zie98-33] Eric R. Ziegel. Book review: Practical Handbook of Spatial Statistics by S. Arlinghaus. Technometrics, 40(1):82, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271412.

Ziegel:1998:BRBaf

[Zie98-34] Eric R. Ziegel. Book review: Proceedings in Computational Statistics by Albert Prat. Technometrics, 40(3):267, August 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271203.

Ziegel:1998:BRBaq

[Zie98-35] Eric R. Ziegel. Book review: Quality Improvement through Statistical Methods by Bovas Abraham. Technometrics, 40(4):358– 359, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270553.

Ziegel:1998:BRBj

[Zie98-36] Eric R. Ziegel. Book review: Randomization, Bootstrap and Monte Carlo Methods in Biology by B. Manly. Technometrics, 40(1): 84, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271419.

Ziegel:1998:BRBo

[Zie98-37] Eric R. Ziegel. Book review: Randomization Tests by E. Edgington. Technometrics, 40(2):163, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270665.

Ziegel:1998:BRBak

[Zie98-38] Eric R. Ziegel. Book review: S+ Spatial Stats by S. Kaluzny; S. Vega; T. Cardoso; A. Shelley. Technometrics, 40(4):357, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270547.

Ziegel:1998:BRBad

[Zie98-39] Eric R. Ziegel. Book review: Spreadsheet Modeling and Decision Analysis by C. Ragsdale. Technometrics, 40(3):266, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271201.

Ziegel:1998:BRBm

[Zie98-40] Eric R. Ziegel. Book review: Statistical Case Studies for Industrial Process Improvement by Veronica Czitrom; Patrick D. Spagon. Technometrics, 40(2):163, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270663.

Ziegel:1998:BRBag

[Zie98-41] Eric R. Ziegel. Book review: Statistical Theory and Computational Aspects of Smoothing by W. Härdle; M. G. Schimek. Technometrics, 40(3):267, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271204.

Ziegel:1998:BRBaa

[Zie98-42] Eric R. Ziegel. Book review: Statistics for Environmental Biology and Toxicology by Walter W. Piegorsch; A. John Bailar. Technometrics, 40(3):265–266, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271198.

Ziegel:1998:BRBg

[Zie98-43] Eric R. Ziegel. Book review: Survey Measurement and Process Quality by L. Lyberg; P. Biemer; M. Collins; E. de Leeuw; C. Dippo; N. Schwartz; D. Trewin. Technometrics, 40(1):83–84, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271416.

Ziegel:1998:BRBk

[Zie98-44] Eric R. Ziegel. Book review: Survival Analysis with Long-Term Survivors by R. Maller; X. Zhou. Technometrics, 40(1):85, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271420.

Ziegel:1998:BRBab

[Zie98-45] Eric R. Ziegel. Book review: The Balanced Scorecard by R. Kaplan; D. Norton. Technometrics, 40(3):266, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271199.

Ziegel:1998:BRBd

[Zie98-46] Eric R. Ziegel. Book review: The Basics of S and S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 40(1):82–83, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271413.

Ziegel:1998:BRBau

[Zie98-47] Eric R. Ziegel. Book review: The Robust Organization by William Stinson. Technometrics, 40(4):359–360, November 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270557.

Ziegel:1998:BRBac

[Zie98-48] Eric R. Ziegel. Book review: Value Leadership by Michael Harris. Technometrics, 40(3):266, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271200.

Ziegel:1999:BRBl

[Zie99a] Eric R. Ziegel. Book review: 101 Good Ideas: How to Improve Just about Any Process by Karen Bemowski; Brad Stratton. Technometrics, 41(1):84–85, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271021.

Ziegel:1999:BRBm

[Zie99b] Eric R. Ziegel. Book review: A History of Mathematical Statistics from 1750 to 1930 by Anders Hald. Technometrics, 41(1): 85, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271022.

Ziegel:1999:BRBb

[Zie99c] Eric R. Ziegel. Book review: Applied Multivariate Statistical Analysis by Richard A. Johnson; Dean W. Wichern. Technometrics, 41(1):81–82, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271011.

Ziegel:1999:BRBd

[Zie99d] Eric R. Ziegel. Book review: Applied Regression Analysis: A Research Tool by John O. Rawlings; Sastry G. Pantula; David A. Dickey. Technometrics, 41(1):82, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271013.

Ziegel:1999:BRBav

[Zie99e] Eric R. Ziegel. Book review: ARCH Models and Financial Applications by Christian Gourieroux. Technometrics, 41(4):382–383, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271379.

Ziegel:1999:BRBah

[Zie99f] Eric R. Ziegel. Book review: Case Studies in Environmental Statistics by Douglas Nychka; Walter W. Piegorsch; Lawrence H. Cox. Technometrics, 41(3):273, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270593.

Ziegel:1999:BRBs

[Zie99g] Eric R. Ziegel. Book review: Common Statistical Methods for Clinical Research with SAS Examples by Glenn A. Walker. Technometrics, 41(2):179–180, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270753.

Ziegel:1999:BRBu

[Zie99h] Eric R. Ziegel. Book review: Communication for Engineers by Richard Ellis. Technometrics, 41(2):180, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270755.

Ziegel:1999:BRBao

[Zie99i] Eric R. Ziegel. Book review: Data Analysis with Microsoft(R) Excel by Kenneth N. Berk; Patrick Carey. Technometrics, 41 (4):380, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271372.

Ziegel:1999:BRBax

[Zie99j] Eric R. Ziegel. Book review: Discretization and MCMC Convergence Assessment by Christian P. Robert. Technometrics, 41 (4):383, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271381.

Ziegel:1999:BRBac

[Zie99k] Eric R. Ziegel. Book review: Doing Statistics with Excel^{T M} 97 by Marilyn Pelosi; Theresa M. Sandifer; Jersy J. Letkowski. Technometrics, 41(3):271–272, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270588.

Ziegel:1999:BRBaa

[Zie99l] Eric R. Ziegel. Book review: Engineering Statistics by Douglas C. Montgomery; George C. Runger; Norma F. Hubele. Technometrics, 41(3):271, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270586.

Ziegel:1999:BRBr

[Zie99m] Eric R. Ziegel. Book review: Environmental Stats for S-Plus by Steven P. Millard. Technometrics, 41(2):179, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270752.

Ziegel:1999:BRBp

[Zie99n] Eric R. Ziegel. Book review: Frontiers in Statistical Quality Control 5 by H. J. Lenz; P. T. Wilrich. Technometrics, 41(2):179, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270750.

Ziegel:1999:BRBn

[Zie99o] Eric R. Ziegel. Book review: Fundamentals of Industrial Quality Control by Lawrence S. Aft. Technometrics, 41(2):178, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270748.

Ziegel:1999:BRBf

[Zie99p] Eric R. Ziegel. Book review: Handbook of Statistical Methods for Engineers and Scientists by Harrison M. Wadsworth. Technometrics, 41(1):83, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271015.

Ziegel:1999:BRBg

[Zie99q] Eric R. Ziegel. Book review: How to Conduct Organizational Surveys by Jack E. Edwards; Marie D. Thomas; Paul Rosenfeld; Stephanie Booth-Kewley. Technometrics, 41(1):83, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271016.

Ziegel:1999:BRBi

[Zie99r] Eric R. Ziegel. Book review: JMP Start Statistics by John Sall; Ann Lehman (SAS Institute). Technometrics, 41(1):84, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271018.

Ziegel:1999:BRBaz

[Zie99s] Eric R. Ziegel. Book review: Linear Models by S. R. Searle. Technometrics, 41(4):383–384, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271383.

Ziegel:1999:BRBe

[Zie99t] Eric R. Ziegel. Book review: Modelling Longitudinal and Spatially Correlated Data by Timothy G. Gregoire; David R. Brillinger; Peter J. Diggle; Estelle Russek-Cohen; William G. Warren; Russell D. Wolfinger. Technometrics, 41(1):82–83, February 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271014.

Ziegel:1999:BRBan

[Zie99u] Eric R. Ziegel. Book review: Nonparametric Statistical Methods by Myles Hollander; Douglas A. Wolfe. Technometrics, 41 (4):380, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271371.

Ziegel:1999:BRBar

[Zie99v] Eric R. Ziegel. Book review: Nonresponse in Household Interview Surveys by Robert M. Groves; Mick P. Couper. Technometrics, 41(4):381, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271375.

Ziegel:1999:BRBag

[Zie99w] Eric R. Ziegel. Book review: Numerical Linear Algebra for Applications in Statistics by James E. Gentle. Technometrics, 41 (3):272–273, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270592.

Ziegel:1999:BRBz

[Zie99x] Eric R. Ziegel. Book review: Odds Ratios in the Analysis of Contingency Tables by T. Rudas. Technometrics, 41(2):181–182, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270760.

Ziegel:1999:BRBo

[Zie99y] Eric R. Ziegel. Book review: Optimization in Quality Control by K. Alsultan; M. A. Rahim. Technometrics, 41(2):178–179, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270749.

Ziegel:1999:BRBak

[Zie99z] Eric R. Ziegel. Book review: Practical Management Science by Wayne L. Winston; S. Christian Albright. Technometrics, 41(3): 274, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270596.

Ziegel:1999:BRBa

[Zie99-27] Eric R. Ziegel. Book review: Practical Nonparametric and Semiparametric Bayesian Statistics by D. D. Dey; P. Müller. Technometrics, 41(1):81, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271010.

Ziegel:1999:BRBap

[Zie99-28] Eric R. Ziegel. Book review: Practical Risk Assessment for Project Management by Stephen Grey. Technometrics, 41(4):380–381, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271373.

Ziegel:1999:BRBas

[Zie99-29] Eric R. Ziegel. Book review: Principles of Total Quality by Jill A. Swift; Joel E. Ross; Vincent K. Omachonu. Technometrics, 41(4):381–382, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271376.

Ziegel:1999:BRBv

[Zie99-30] Eric R. Ziegel. Book review: Probability and Statistics for Engineers and Scientists by Ronald E. Walpole; Raymond H. Meyers; Sharon L. Meyers. Technometrics, 41(2):180–181, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270756.

Ziegel:1999:BRBae

[Zie99-31] Eric R. Ziegel. Book review: Programming Techniques for Object-Based Statistical Analysis with SAS Software by Tanya Kolosova; Samuel Berestizhevsky. Technometrics, 41(3):272, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270590.

Ziegel:1999:BRBx

[Zie99-32] Eric R. Ziegel. Book review: Prove It with Figures by H. Zeisel; D. Kaye. Technometrics, 41(2):181, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270758.

Ziegel:1999:BRBat

[Zie99-33] Eric R. Ziegel. Book review: Quality Assurance for the Chemical and Process Industries by American Society for Quality. Technometrics, 41(4):382, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271377.

Ziegel:1999:BRBaq

[Zie99-34] Eric R. Ziegel. Book review: Quantitative Risk Analysis by David Vose. Technometrics, 41(4):381, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271374.

Ziegel:1999:BRBj

[Zie99-35] Eric R. Ziegel. Book review: SAS(R) System for Elementary Statistical Analysis by Sandra D. Schlotzhauer; Ramon C. Littell. Technometrics, 41(1):84, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271019.

Ziegel:1999:BRBay

[Zie99-36] Eric R. Ziegel. Book review: Selected Papers of Hirotugu Akaike by Hirotugu Akaike; Emanuel Parzen; Kunio Tanabe; Genshiro Kitagawa. Technometrics, 41(4):383, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271382.

Ziegel:1999:BRBau

[Zie99-37] Eric R. Ziegel. Book review: Statistical Case Studies by Roxy Peck; Larry D. Haugh; Arnold Goodman. Technometrics, 41(4):

382, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271378.

Ziegel:1999:BRBad

[Zie99-38] Eric R. Ziegel. Book review: Statistical Design and Analysis of Experiments by Peter John. Technometrics, 41(3):272, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270589.

Ziegel:1999:BRBaw

[Zie99-39] Eric R. Ziegel. Book review: Statistical Issues in Drug Development by Stephen Senn. Technometrics, 41(4):383, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271380.

Ziegel:1999:BRBy

[Zie99-40] Eric R. Ziegel. Book review: Statistical Significance by Siu Chow. Technometrics, 41(2):181, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270759.

Ziegel:1999:BRBw

[Zie99-41] Eric R. Ziegel. Book review: Statistical Thinking for Managers by David Hildebrand; Lyman Ott. Technometrics, 41(2):181, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270757.

Ziegel:1999:BRBab

[Zie99-42] Eric R. Ziegel. Book review: Statistics: An Introduction by Robert D. Mason; Douglas A. Lind; William G. Marchal. Technometrics, 41(3):271, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270587.

Ziegel:1999:BRBam

[Zie99-43] Eric R. Ziegel. Book review: Statistics for Petroleum Engineers and Geoscientists by Jerry L. Jensen; Larry W. Lake; Patrick W. M. Corbett; David J. Goggin. Technometrics, 41(4):379–380, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271370.

Ziegel:1999:BRBq

[Zie99-44] Eric R. Ziegel. Book review: Statistics for the Environment 2 by Vic Barnett; K. Feridun Turkman. Technometrics, 41(2):179, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270751.

Ziegel:1999:BRBai

[Zie99-45] Eric R. Ziegel. Book review: Statistics for the Environment 3 by V. Barnett; K. Feridun Turkman. Technometrics, 41(3):273–274, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270594.

Ziegel:1999:BRBaj

[Zie99-46] Eric R. Ziegel. Book review: Stochastic Models in Geosystems by Stanislav A. Molchanov; Wojbor A. Woyczynski. Technometrics, 41(3):274, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270595.

Ziegel:1999:BRBaf

[Zie99-47] Eric R. Ziegel. Book review: Strategic Data Warehousing Principles Using SAS Software by Peter R. Welbrock. Technometrics, 41(3):272, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270591.

Ziegel:1999:BRBt

[Zie99-48] Eric R. Ziegel. Book review: Survey Research Using the SAS System by Archer Gravely. Technometrics, 41(2):180, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270754.

Ziegel:1999:BRBal

[Zie99-49] Eric R. Ziegel. Book review: Theory of Point Estimation by E. L. Lehmann; George Casella. Technometrics, 41(3):274, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270597.

Ziegel:1999:BRBk

[Zie99-50] Eric R. Ziegel. Book review: Total Quality Tools by PQ Systems, Inc. Technometrics, 41(1):84, February 1999. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271020.

Ziegel:1999:BRBc

[Zie99-51] Eric R. Ziegel. Book review: Univariate and Multivariate General Linear Models by Neil H. Timm; Tammy A. Mieczkowski. Technometrics, 41(1):82, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271012.

Ziegel:1999:BRBh

[Zie99-52] Eric R. Ziegel. Book reviews: The Survey Research Handbook by Pamela L. Alreck; Robert B. Settle; How to Conduct Surveys: A Step-by-Step Guide by Arlene Fink; Jacqueline Kosecoff. Technometrics, 41(1):83–84, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271017.

Zinkgraf:1993:BRB

[Zin93] Stephen A. Zinkgraf. Book review: Computer-Interactive Data Analysis by A. D. Lunn; D. R. McNeil. Technometrics, 35(4): 459, November 1993. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270290.

Zhou:1994:STV

[ZM94] Leping Zhou and Thomas Mathew. Some tests for variance components using generalized p values. Technometrics, 36(4):394–402, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1269954. See comment [BZM96].

Zhang:1994:AAD

[ZP94] Nien Fan Zhang and James F. Pollard. Analysis of autocorrelations in dynamic processes. Technometrics, 36(4):354–368, November 1994. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269951.

Zimmerman:1991:CSS

[ZZ91] Dale L. Zimmerman and M. Bridget Zimmerman. A comparison of spatial semivariogram estimators and corresponding ordinary kriging predictors. Technometrics, 33(1):77–91, February 1991. CO-REFERENCES 315

DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269009.